A Bibliography of Papers in Lecture Notes in Computer Science (2002) (Part 2 of 4)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

02 May 2020 Version 1.11

Title word cross-reference

(3, 3) [1732]. 0 [426]. 1 [426, 1647]. 1/f [1260]. 168 [1729]. 2 [1740, 943, 1677, 420, 50, 1732]. 3 [146, 154, 18, 1094, 1033, 1664, 196, 992, 1020, 1065, 84, 29, 1640, 1075, 1093, 1662, 1023, 142, 1011, 1019, 1637, 30, 219, 1364, 1107, 1430]. 3 × 3 [536]. 4 [1060]. 8 · 168 [1729]. ^{(G)} [659]. 2 [1056]. ^{st} [208]. ^{T M} [596]. _{2} [213]. ax^{7} + bx + c [1729]. ax^{8} + bx + c [1729]. Δ [1694]. Dy^{2} = x^{3} [1735]. k [1701, 434, 1677, 408, 1691]. k^{th} [1711]. MIN ERVA [292]. N EMESIS [358]. √

[1754]. n [301]. N = pq [1752]. P [164, 1720]. Ψ [285]. q [1705]. x [1735]. y′′ = f(x, y) [164].

-adic [1754, 1720]. -Center [1691]. -D [1664, 84, 1060, 1019, 1637, 1647, 1364]. -Gram [1705]. -Level [50]. -List [1694]. -nearest [408]. -Partition [434]. -SAT [420]. -Split [1732]. -Stability [164]. -Stage [1662]. -Tier [1430].

.NET [88].

/Geom/c [659].

1003.1q [1336]. 128 [1540]. 1980-88 [19]. 1989-1997 [20].

2 [637]. 21 [208]. 2nd [1247].

3G [649, 102, 640]. 3GPP [673]. 3rd [709].

90 [186]. 95 [1328, 1329].

a-Si [216]. Abelian [1731, 1726]. Aborted [503]. Absent [495]. Abstract [296, 356, 501, 1290]. Accelerated [1727]. Accelerating [158]. Accelerator [155, 153]. Accelerators [154, 152]. Access [617, 679, 385, 1180, 696, 787, 746, 1482, 770, 98, 1229, 771, 668]. Accesses [762]. Accessibility [784]. Accessory [139]. According [1283]. Accuracy [1495]. Accurate [985, 1056, 369, 1130, 994]. Achieve [702]. Achieving [1049]. Acid [538]. Acoustic [1026]. Acquisition [1094, 1026, 901, 730, 865, 1438, 897, 1557, 1602]. ACT [1410]. ACT-R [1410]. Action [1734, 960]. Actional [1437]. Actions [299]. Active [1367, 1113, 60, 666, 61, 1110, 1349, 1390, 1127, 1308]. Activities [891, 895]. Activity [1610]. Acts [297, 460]. Ad [628, 630, 641, 629, 696, 698]. Ad-Hoc [630, 641, 696]. Ada [1328, 1329, 1326, 1337, 1327, 1333, 1336, 1325, 1322, 1338, 1331, 1318]. Ada95 [1320]. AdaBoost [1496]. Adaptability [723, 742]. Adaptable [1649]. Adaptation [774, 780, 756, 727, 729, 1085, 732, 806, 1188, 744, 47, 834, 90, 802, 961]. Adapted [922]. Adapting [793, 796, 772]. Adaption [1277]. Adaptive [752, 720, 721, 621, 754, 1398, 777, 778, 779, 724, 755, 725, 630, 726, 249, 728, 757, 804, 759, 1474, 784, 785, 730, 805, 643, 731, 1068, 761, 733, 734, 608, 84, 666, 788, 1397, 737, 1388, 1205, 1678, 400, 716, 1600, 1383, 792, 763, 764, 701, 186, 795, 746, 1508, 1446, 1259, 915, 797, 1390, 798, 926, 1484, 1486, 747, 800, 749, 801, 1229, 807, 320, 717, 652, 803]. Adaptively [196]. Adaptivity [723, 783, 767]. Added [390]. Adding [1470, 1327, 3, 381]. Addition [1176, 1195]. Additive [1754]. Addressed [1560]. Addressing [859]. Adequate [1488]. adic [1754, 1720, 114]. Adjoint [107]. Adjusted [654]. Adjustment [1144]. Admission [624, 664, 649, 612, 703, 602, 707, 611, 613]. Admission-Controlled [624]. Adoption [590]. Advance [613]. Advanced [588]. Adversary [254]. Advertisement [443]. AERO [450]. Aerodynamic [86]. Aeronautical [1426]. AES [252]. Affection [1376]. Affective [1466, 1479, 1260]. Affine [1052, 927, 932]. Affording [1377]. Against [1540, 1162, 1589, 1567, 277, 1172, 1529]. Age [1351]. Agent [1362, 295, 857, 1200, 1372, 1220, 1433, 468, 833, 805, 284, 1468, 760, 1588, 285, 315, 292, 1469, 1230, 301, 597, 1461, 311, 1546, 1385, 1198, 286, 877, 1262, 290, 1201, 467, 450, 449, 291, 288, 483, 776]. Agent-Based [1362, 857, 1230, 597, 1262, 1201]. Agent-Mediated [1546]. Agents [313, 1633, 296, 1372, 306, 1367, 887, 1384, 1474, 308, 309, 312, 1375, 1345, 1381, 293, 1426, 1467, 1369, 1382, 1463, 799, 639, 289, 1222, 480, 1386, 1199, 451, 486, 1376]. AgentSpeak [293]. Aggregated [682]. Aggregates [1568, 699, 1603]. Aggregation [462]. Aggressive [1497]. Agreement [277, 247]. AHA [783, 758]. AHAM [758]. AI [1240]. AI-Planning [1240]. Aided [1425]. AIMD [329]. Air [661, 1656]. Aircraft [112, 1198]. ALCOR [1625]. Algebra [1451, 1393, 1415]. Algebraic [285, 43, 1310, 1719, 1743]. Algebraically [1753]. Algebras [338]. Algorithm [614, 1634, 1649, 380, 394, 1208, 1713, 201, 1740, 1266, 1620, 1207, 402, 1738, 5, 1690, 476, 1065, 1744, 230, 415, 84, 1710, 1711, 473, 534, 362, 1751, 1165, 172, 379, 1203, 919, 1307, 1692, 977, 1681, 532, 1209, 1194, 1196, 1143, 432, 674, 1622, 1647, 417, 451, 486, 1727]. Algorithmic [259, 1625, 536, 601]. Algorithms [1641, 1684, 1458, 1699, 193, 1087, 442, 1197, 385, 123, 1571, 1095, 63, 411, 1689, 444, 174, 676, 1691, 1679, 1678, 1176, 1119, 1237, 1195, 463, 1170, 365, 1704, 1658, 1202, 430, 426, 1581, 1235, 1234, 1720, 1693, 683, 150, 1117, 416, 873, 951, 375, 341, 631]. Alignment [1550, 1270]. Alignments [1550]. All-Norm [1699]. All-Optical [1666]. Allocation [324, 442, 634, 664, 666, 604, 300, 1682, 1622, 317]. Also [1488]. Alternating [289, 1533]. Alternating-Time [289]. Alternative [1458, 1550, 846]. Ambiguity [976]. Ambre [1449]. AMICO [1382]. AMnet [708]. among [393, 902]. Amorphous [216]. Amortized [490, 1714]. Analogies [1403]. Analyser [1310]. Analysing [1304]. Analysis [1323, 1167, 638, 330, 724, 1328, 257, 963, 854, 1293, 1329, 1136, 616, 661, 358, 625, 1435, 1249, 1358, 1180, 944, 1477, 404, 1289, 1301, 1074, 1016, 627, 1600, 1333, 1501, 1247, 256, 179, 1379, 1514, 1023, 1144, 1697, 677, 1511, 948, 658, 957, 659, 1260, 361, 662, 1201, 631]. Analytic [1735, 1243]. Analytical [646, 921, 360]. Analytics [596]. Analyze [846]. Anchored [609]. Angiographic [142]. Animated [1467]. Animation [19, 20]. Animations [460]. Anisotropic [925]. Annotated [1576]. Annotation [587]. Annular [1131]. Annulus [121]. Anonymity [501]. Anonymous [1221, 1183]. Answer [1429]. Answering [1560, 1232, 471]. Ant [1245]. Antisocial [306]. Antitag [531]. Antithetic [40]. Apart [1357]. Aperture [156]. Aplusix [1393]. Aplusix-Editor [1393]. Appearance [1280, 1140, 952, 923]. Applicability [497]. Application [147, 1633, 935, 727, 642, 1724, 1689, 1475, 1507, 176, 174, 458, 1663, 896, 1186, 1165, 1481, 79, 69, 1128, 1322, 1223, 639, 1194, 1308, 1199, 1248, 1504, 227, 479, 535, 830, 1217, 1286]. Application-Defined [1322]. Application-Specific [1663]. Applications [88, 614, 1339, 1321, 92, 778, 36, 279, 688, 16, 137, 761, 101, 986, 608, 94, 737, 999, 835, 816, 1334, 1742, 1682, 1637, 66, 921, 280, 497, 751]. Applied [161, 1245, 1179, 1270, 1284, 423].

Apply [1586]. Applying [890, 1211].

Apportionment [1228]. Appraisal [1465]. Approach [1573, 433, 1624, 1518, 1471, 823, 843, 859, 388, 450, 1560, 756, 322, 726, 1329, 387, 2, 1599, 987, 943, 625, 1435, 973, 1283, 1120, 476, 38, 570, 401, 651, 1582, 866, 974, 474, 441, 75, 1243, 815, 1423, 1121, 763, 920, 300, 816, 1247, 210, 340, 828, 1024, 597, 3, 869, 1170, 1191, 1569, 1428, 601, 1039, 1549, 749, 872, 423, 1262, 1107, 1131, 1201]. Approaches [1458, 212, 22, 1063, 1437, 211]. Approximability [1700]. Approximate [1034]. Approximating [443]. Approximation [1673, 1699, 442, 437, 1687, 1689, 438, 1691, 1679, 1680, 430, 432]. Approximations [1698]. Aqueous [536, 538]. Arbitrarily [469]. Arbitrarily-Nested [469]. Arbitrary [1050]. Arc [458, 1682, 1309]. Architectural [1384, 1313, 27, 983, 1008]. Architecture [617, 1619, 670, 686, 1624, 103, 562, 691, 805, 864, 608, 1645, 292, 835, 191, 68, 663, 1484, 1662, 1486, 483, 800, 776, 820, 1391, 1618, 1430]. Architectures [843, 291, 1640, 1627, 602, 844]. Area [8, 1024, 69, 4]. ARGOS [1487]. Argumentation [1431, 1481]. Arguments [1431]. Arithmetic [346, 548]. Arithmetically [1721]. Armed [1692]. Arms [303]. array [1616]. Arrival [603]. Arrovian [429]. Art [636, 205]. Artificial [1187, 1186, 1578, 1248]. Artin [1740]. ASIMIL [1422]. Asimov [303]. ASM [345]. ASP [891]. Aspects [636, 1725, 494, 770, 708]. assembly [547]. Assess [1448]. Assessing [1441, 561]. Assessment [1452, 731, 505, 1408, 1426, 47]. Assessments [897]. Asset [1246]. Assignment [440, 476, 400, 631]. Assistance [718, 479]. Assistant [1433, 1405, 1564, 1416]. Assistants [1198, 808]. Assisted [1129]. Association [364, 461, 1592, 389, 1591, 363, 370, 477, 390, 362, 365, 391, 373]. Association-Based [477]. Association-Rule [365]. Associations [371, 1184]. Associative [540]. Assorted [1135]. Assume [341]. Assume-Guarantee [341]. Assumptions [273, 355, 266]. Assurance [891]. Assured [604]. Asymmetric [677]. Asymptotics [606, 611]. Asynchronous [1665]. Atlas [1391]. ATM [626]. Attack [248, 1149, 1161, 1165, 1540, 1529]. Attacks [1162, 1523, 1171, 259, 1538, 1530, 1158]. Attention [1188]. Attractive [888]. Attribute [1593, 1243]. Auction [516, 1230]. Auctions [517, 306]. Audio [967, 1255]. Audio-Video [967]. Audio-Visual [1255]. Audiovisual [1071]. Auditing [847]. Augmentation [1563]. Augmented [41, 134]. Augmenting [1360]. Authentication [270, 1156, 506, 500]. Author [116, 245, 1371, 1489, 1522, 1544, 1609, 1669, 1715, 1755, 281, 316, 335, 357, 414, 448, 489, 518, 555, 598, 714, 809, 879, 918, 972, 1032, 1092, 1148, 1177, 1185, 1265, 1285, 1311, 1340, 1461]. Authoring [1459, 1458, 1402, 1414, 728, 1472, 790, 794, 792, 764]. Authorizations [1566]. Auto [1219]. Auto-downloading [1219]. Autoimmune [1599]. Automata [1571, 1299, 1645]. Automated [1470, 1620, 1664, 1605, 321, 1588, 1115, 242, 454, 750, 1580, 1008]. Automatic [719, 112, 386, 1134, 1210, 1104, 565, 950, 1052, 1435, 107, 114, 1192, 1018, 1130, 1413, 1455, 1559, 1186, 766, 109, 90, 953, 111, 113, 1190, 341]. Automatically [1257]. Automation [1243, 1487, 917]. Automorphism [1732]. Autonomous [1222]. AutoTutor [1453]. Availability [468]. Average [1035]. Avionics [1324]. AVL [1714]. AVL-Trees [1714]. AVS [91]. AVS/Express [91]. Aware [630, 690, 604, 648, 706]. Awareness [235]. Axis [1020, 954, 1017]. Axisymmetric [148].

B [355, 32, 342, 9, 343]. B-Spline [32, 9].

Babel [830]. Baby [1751]. Back [1490, 1186]. Back-Propagation [1186]. Background [1068]. Backward [1499]. Bagging [1495]. Bait [1149]. Balance [1665]. Balanced [119, 992]. Balancing [621, 1660, 1663, 230, 227]. Ball [128, 139]. Bandwidth [324, 665, 666, 26]. Barrier [722]. Base [1679, 1484]. Based [190, 774, 1362, 566, 1118, 857, 1431, 514, 296, 1398, 937, 472, 1220, 388, 689, 1224, 380, 194, 1005, 979, 690, 1208, 1601, 171, 1620, 565, 1395, 987, 950, 943, 642, 137, 152, 5, 641, 411, 833, 246, 1405, 217, 336, 1661, 760, 732, 1449, 1109, 1213, 1228, 1358, 1258, 276, 120, 1608, 612, 1112, 666, 477, 789, 1193, 1254, 1276, 1368, 291, 260, 1244, 1141, 372, 1645, 1450, 740, 75, 337, 1243, 1301, 1192, 836, 247, 838, 695, 847, 395, 366, 815, 835, 1423, 1121, 1413, 794, 763, 764, 1230, 1333, 1403, 1437, 1227, 795, 920, 1404, 816]. Based [1247, 1392, 1241, 1656, 766, 891, 172, 767, 187, 619, 703, 191, 597, 533, 962, 994, 471, 1387, 1390, 1353, 939, 1421, 1073, 799, 1569, 601, 1566, 1483, 639, 15, 1277, 602, 1516, 894, 1080, 611, 1183, 483, 632, 747, 1194, 1196, 1549, 989, 769, 749, 872, 220, 1412, 913, 98, 78, 1268, 508, 317, 149, 1262, 1107, 1406, 1257, 603, 1260, 367, 347, 452, 1267, 713, 830, 1239, 1009, 222, 878, 1201, 653, 548]. Baseline [923]. Baselines [998]. Bases [1579]. Basic [1238, 539, 426, 1565, 1274, 340]. Basic-REAL [340]. Basis [1690, 176, 1739, 1045, 175, 1387, 9]. Batch [436]. Batching [1236]. Bayes [400, 367]. Bayesian [1441, 368, 1029, 1283, 1558, 369, 1130, 991, 1607, 1064, 479, 487]. BDD [260]. BDD-Based [260]. BDI [1376]. Be [1105, 1037, 787, 495, 503, 1488]. Beads [535]. Beam [154, 149]. Beams [148, 153]. Bed [688]. Been [915]. BeepBeep [1535]. Behavior [1435, 949, 378, 384, 318]. Behavioral [762, 406]. Behaviors [1199]. Behaviour [723]. Behavioural [831].

Belief [1062, 1006]. Beltrami [941]. Benchmarking [75, 567, 484]. Bending [1584]. Benzene [213]. between [143, 166, 909, 354, 716, 1041, 523, 1485, 1011, 1215]. Beyond [1541, 1436, 1429, 214, 844]. Bézier [8]. Bi [959]. Bi-centric [959]. Bias [1166, 1513, 1511]. Bias-Variance [1511]. Bid [1230]. Bidirectional [1146]. Bids [299]. Bin [437]. Binding [1336]. Bioelectric [55]. Biological [138, 1599, 398, 519, 1235]. Biometric [1283, 1284]. Biometrics [1273, 1268, 1274]. Biomolecular [529]. Birthday [1541]. Bit [532]. Blending [10]. Blind [514]. Block [270, 1644, 1151, 1174]. Block-Cipher [270]. Blockcipher [1169]. Blocked [1641]. Blocking [1639, 681, 1647]. Blocks [1236, 1718]. Bluetooth [687, 636, 261, 631]. Blurring [928]. Board [727]. Boasting [1513]. Bodies [201, 52]. Body [955, 1076]. Boltzmann [1668]. Boolean [1258, 1154]. Boomerang [1523]. Boost [368, 403]. Boosted [1493]. Boosting [1556, 1608, 1497, 1515, 1495, 1491]. Boosting-Based [1608]. Boson [174]. Both [209]. Bottlenecks [320]. Bound [1087, 425]. Boundaries [1047]. Boundary [1059, 525, 1128]. Bounded [436, 1688]. Bounds [262, 1659, 1532, 947]. Box [1662]. Boxes [1342]. BP [1213]. Braid [246, 247]. Brain [1130, 1517]. Branch [1087, 425, 1211]. Branch-and-Cut [425]. Branching [1235]. Breakthrough [886]. Bridges [1246]. Bridging [1315]. Brief [334]. Bringing [577]. Broadband [625, 1473]. Broadcast [670, 444, 1161]. Broadcasts [1653]. Broom [1462]. Browser [761]. Browsing [597]. Buchmann [1727]. Buffer [1594, 658]. Buffet [1187]. Bug [905]. Bugs [915]. Build [826]. Building [856, 445, 1236, 1029, 864, 125, 1718, 1368, 561, 568, 983, 1331, 1143, 956, 1654]. Built [783]. Bulk [1714, 447]. Bundling [1703].

Bureaucrat [1674]. Burst [686, 668]. Burst-Switched [686, 668]. Bus [695]. Business [885, 884, 510, 839, 593, 413, 1286]. Buy [447]. Buy-at-Bulk [447]. Buying [1564, 451, 486]. BVMs [161]. Bystander [1453]. Byzantine [1314, 277].

C [1450, 10, 659]. Cable [607, 83]. CAC [664]. Cache [322, 698, 1181, 1594, 1229]. Caching [615, 706]. Cactus [423]. CAD [16, 141, 130]. Cadence [346]. Calculate [1650]. Calculating [673]. Calculation [52]. Calculations [212, 216]. Calculus [623, 1288, 285]. Calibration [1104, 1109, 1001, 1101, 991, 997, 1103]. Call [1701, 651, 649, 612, 303, 707, 807]. Call-Level [651]. Calls [652]. Camellia [1532, 1529, 1172]. Camera [1104, 992, 991, 959, 1103, 1038]. Camfield [1349]. Can [1105, 1037, 826]. Cannot [495, 1462]. Capabilities [581]. Capability [1214, 540, 886]. Capacitated [1698]. Capacity [324, 675, 620, 669, 613, 631]. Capture [1400, 953]. Car [1186]. Card [507, 509, 508]. Card-Based [508]. Cardinality [435]. Care [482, 106]. Carlo [63, 1660, 1, 64, 62, 1039, 66, 40]. Carriers [333]. Carrying [869]. Cars [727]. Cascadable [523]. Cascaded [221]. Case [377, 1685, 1339, 1431, 724, 1449, 789, 1244, 1305, 1355, 21, 482, 360, 1030, 872]. Case-Based [1431, 1449, 789]. Cases [1283, 899, 1330]. Cash [490]. Cast [1052, 1291]. Catadioptric [1108, 982]. Catalogs [834]. Categorical [382]. Category [769]. Category-Based [769]. Causal [387]. Causality [477]. Causation [412]. Cavity [147, 158]. CAWICOMS [571]. Cayley [1748]. CBC [1541, 280]. CBC-MAC [1541]. cbCPT [487]. CDMA [649]. CEA [159]. Cell [152, 678]. Cellular [476, 1645, 707, 652, 713]. Center [1691, 683]. Centered [1401]. Centers [807]. Central [1108]. Centre [194, 1109].

Centric [834, 959]. Century [208]. Certificate [504, 505]. Chain [1195, 1039]. Chains [1678, 1654, 1580]. Challenge [810, 1289, 332]. Challenges [359]. Change [104, 845, 1140]. Changes [621]. Changing [1245, 738, 1363]. Channel [257, 599, 1361, 644, 476, 650, 652, 713, 1365]. Channelization [610]. Channels [267]. Chaos [237]. Chaotic [238]. Character [1237, 1253, 1506]. Characteristic [1740, 393, 1532]. Characteristics [806, 1444, 1467, 69, 893, 1274, 1260]. Characterization [1132, 428, 1096, 465]. Characterize [679]. Characterizing [1298]. Characters [1706]. Charge [148, 149]. Charts [342]. Chaum [514]. Checker [1252]. Checking [595, 832, 345]. Checks [1326]. Chemical [217]. Children [1461, 1463]. Chinese [1743]. Choice [1294]. Cholesky [35]. Chosen [249]. Ciliates [544]. Cinematographic [460]. Cipher [270, 1166, 1151, 261, 1537, 1174]. Ciphertext [249]. Circle [1109]. Circuits [162]. Circulation [230]. Cities [1360, 1361, 1348, 1343, 1359, 1344, 1346, 1341]. Citizen [569]. Citizens [1346]. City [1361, 1358, 1345]. Civil [1347]. ClaiMaker [592]. CLARISSE [1443]. Class [305, 980, 1598, 44, 1736, 676, 337, 1510, 1726, 603, 1301]. Class-Specific [980]. Classes [296, 425, 678]. Classical [1214]. Classification [584, 1279, 402, 949, 409, 1115, 408, 1547, 1515, 1508, 1126, 1492, 1516, 328]. Classifier [1505, 1507, 1606, 400, 366, 1517, 1503, 1521, 1512, 1502, 1506, 1495, 1284, 1504]. Classifiers [1518, 368, 401, 1510, 369, 1514, 1607, 1138]. Classifying [407, 398, 907, 1049]. Classroom [900]. Clause [534, 547]. Client [864, 28, 1183]. Client-Side [1183]. Climate [80, 78]. CLOS [1662]. Closed [1753]. Closing [1323]. Closure [424, 1634]. Clue [454]. Cluster [65, 322, 411, 1180, 94, 1656, 379, 1658, 1693]. Cluster-Based [411, 1656]. Clustered [1619]. Clustering [1118, 462, 380, 232, 1069, 1052, 1689, 1070, 382, 29, 701, 1555, 1692, 977, 381, 399, 361]. Clustering-Type [1689]. Clusters [226, 1654]. Cluttered [920]. CM [1734]. CM3 [840]. CMM [911, 135]. CMPP [625]. Co [1616]. Co-array [1616]. Coach [764, 1391]. Coalition [305, 467]. Coarse [1277, 1010]. Coarsening [1178]. Coatings [52]. Coaxial [679]. Cockpit [1421]. Code [1323, 610, 661, 152, 1665, 107, 100, 186, 1627, 220, 913]. Code-Division [610]. Code-Points [220]. Codes [1655, 159, 1630, 99]. CODEWORD [1700]. Coding [1208, 524, 762]. Coefficient [183]. Cognitive [1372, 1451, 1465, 1423, 1426, 749]. Cognitively [383]. Cohesion [472]. Collaboration [67, 909, 1462, 1379]. Collaborative [73, 775, 1414, 1439, 1433, 730, 1477, 740, 1399, 231, 791, 1381, 1427, 573, 1613]. Collage [1364]. Collection [1631]. Collective [239, 150, 1261]. Collider [157]. Collision [499, 1153, 127]. Collisions [1152]. Colony [1245]. Color [1254, 1122, 962, 1277, 1239]. Color-Based [962]. Colored [4]. Coloring [1686, 1694]. Coloured [1303]. Comb [422]. Combination [1517]. Combinatorial [34, 1258]. Combine [23]. Combined [1267]. Combiners [1514]. Combining [1518, 692, 349, 353, 1510, 791, 1083, 1304, 1671, 1503, 430, 923]. Come [1376]. COMET [818]. Comment [759]. Commerce [481, 584, 785, 805, 1588, 1230, 225, 1380]. Commercial [512, 1355]. Commitment [298]. Commitments [299]. Common [716, 1421]. Commonsense [741]. Communicating [296, 820, 1290]. Communication [313, 1417, 1630, 1350, 667, 1626, 1345, 71]. Communication/Computation [1626]. Communications [68, 632]. Communicative [460, 1437]. Communities [1360, 1347]. Community [775, 779, 1347, 1361, 1350, 1352, 578, 825, 1349, 1354]. Commutator [247]. Compact [964]. Companies [902, 893]. Companion [1487]. Comparative [682, 1228, 46, 871, 1055]. Compare [1458]. Comparing [1736, 1122, 55, 850]. Comparison [147, 1087, 166, 459, 627, 95, 911, 77, 1521, 1512, 1189]. Comparisons [143]. Compass [1124]. Competition [902, 1382]. Competitive [1696, 1222]. Compiler [87, 1657]. Compiler-Controlled [1657]. Compiling [1318]. Complete [336]. Completion [436, 939]. Complex [1605, 308, 945, 1585, 175, 1270, 1617, 1411, 135]. Complexes [213]. Complexity [545, 866, 1355, 1754, 1714, 290]. Component [1398, 1136, 908, 944, 75, 838, 1074, 1240, 191, 90, 1091, 1144, 814]. Component-Based [75, 838, 191]. Component-Oriented [1240]. Components [1395, 1468, 525, 1666, 826, 1333]. Composable [267]. Composite [34, 704, 1044]. Composition [1716, 279, 814]. Compositional [469]. Comprehension [1408]. Compressing [1710]. Compression [1543, 29, 1209]. Computable [258]. Computation [1634, 240, 112, 1717, 63, 277, 529, 548, 1626, 546, 1726, 526, 81, 547, 1727]. Computational [207, 1611, 48, 1612, 196, 1725, 1188, 51, 210, 77, 33, 69, 204, 1373, 99, 255, 290]. Computationally [383]. Computations [159, 1730]. Computer [1451, 934, 1599, 1435, 19, 20, 1425, 208, 1455, 1227, 1415, 1453, 483, 1622, 1632, 1406]. Computer-Based [483].

Computer-Mediated [1435]. Computers [617, 1564, 1647]. Computing [73, 18, 545, 551, 955, 202, 1709, 67, 140, 1612, 1207, 1424, 1652, 1296, 512, 1744, 519, 1705, 520, 1572, 208, 1176, 550, 77, 1125, 1753, 113, 530, 538, 535, 222, 522, 1613]. Concentration [537]. Concept [93, 394, 1625, 1238, 455, 130]. Conception [1437, 1481]. Concepts [1431, 1403, 844]. Conceptual [823, 863, 842, 1555, 855, 812, 751]. Concise [372]. Concurrency [1335, 72]. Concurrent [580, 1328, 1289, 350]. Condition [926, 1217]. Conditional [466]. Conditionalized [656]. Conditions [783, 986, 46]. Confidence [1606, 1256, 371]. Confidence-Lift [371]. Configuration [571, 904, 1212, 677]. Configurations [1076]. Confirmation [539]. Conflict [1305]. Conflict-Free [1305]. Conflicting [295]. Conflicts [880, 1434]. Congestion [633, 645, 632]. ConGolog [833]. Conics [954]. Connecting [1359]. Connection [885, 1208, 664, 618, 653]. Connection-Based [1208]. Connectionist [474]. Connections [323, 1349]. Connectivity [1692]. Conquer [903]. Cons [735]. Consanguinity [1361]. Consensus [924]. Conservation [161]. Conservative [1497]. Consideration [896]. Considering [797]. Consistency [595, 458, 932]. Constant [275, 447]. Constellation [664]. Constrained [439, 935, 435, 438, 1704]. Constraint [1109, 395, 1413, 794, 300, 115, 1412, 878]. Constraint-Based [395, 1413, 794, 1412]. Constraints [829, 1236, 1591, 115, 1128, 983, 1107, 361]. Construct [192]. Constructing [888, 1678, 1045]. Construction [1224, 521, 409, 1541, 27, 795, 540, 1394, 71, 423, 466]. Constructions [1169, 1578, 21, 1154]. Constructive [1272, 1397, 474]. Constructivist [1484]. Contact [201, 1249, 203, 83]. Content [756, 732, 761, 498, 1137, 1516, 1125]. Content-Based [1516]. Content-Plots [1125]. Contention [607]. Context [481, 1178, 299, 738, 815, 1122, 1076, 1492, 824]. Context-Dependent [824]. Contextual [464]. Contextware [1315]. Continued [1746]. Continuous [1074, 463, 1215, 10]. Contour [919, 939]. Contours [433, 131]. Contract [1327, 480]. Contractual [480]. Contrast [1146]. Contributions [342]. Control [1701, 76, 664, 649, 1264, 612, 666, 888, 1205, 180, 695, 1302, 1469, 319, 902, 1320, 703, 1335, 602, 707, 611, 632, 613, 286, 1262, 537, 1201]. Controllability [1294, 735]. Controllable [1675, 1294]. Controlled [624, 470, 1657, 1308]. Controller [650, 1231, 111]. Controls [320]. Convection [84, 45]. Convection-Diffusion [84]. Conventional [1206, 77]. Conversational [1468, 1467]. Conversations [1435]. Convex [1678, 417]. Cooperate [903]. Cooperation [1605, 902, 1382, 1009]. Cooperative [811, 813, 792, 768, 1222]. Cooperatively [397]. Coordinate [1011]. Coordinatewise [417]. Coordination [1362]. Copy [322, 498]. Copy-Protected [498]. CORBA [189]. Core [327, 317]. Core-Stateless [317]. Corporate [888]. Corporation [1227]. Correction [50, 1004]. Correlation [259, 1005, 1165, 1271]. Correspondence [936, 1078, 1071, 1131]. Correspondences [1734, 1003]. Cost [1708, 1702, 1630, 1188, 898]. Cosymmetry [45]. COTS [909, 907, 826]. Counteract [1355]. Counting [1722, 1739, 1751, 95, 1720]. Coupled [203, 51]. Coupling [114, 79]. Couresware [1459]. Courses [755, 1396, 765]. Courseware [1402]. Cover [1508, 466]. Coverage [913]. Covering [1556, 1702]. CPTs [487]. Crack [1667]. Crash [1314].

Created [1459, 495]. Creating [87, 1361, 1415, 1349]. Creation [908, 608, 910]. Credit [614, 1228, 507, 509]. Crew [1198]. Crisp [1512]. Critical [1002, 928]. Cross [15]. Crossover [1236]. Crossroads [1717]. Cruise [1205]. Cryptanalysis [1163, 1150, 279, 1151, 1159, 1164, 246, 265, 264, 261, 1175, 1542, 1173, 1160, 1172, 1531, 260, 1527, 1526, 1528, 1525]. Cryptoanalysis [224]. Cryptographic [1742, 1170, 1534]. Cryptography [501, 255]. Cryptosystems [250, 1718, 497]. Crystal [155]. Crystalline [48]. Crystallographic [57]. CSCL [1389]. CSE [206]. CSP [348, 352, 350, 74]. CSPs [1241]. Cube [1180]. Cubes [395]. Cue [1114, 1063]. Cues [1060]. CUMULVS [91]. Curricula [210]. Curvature [137]. Curve [1752, 1739, 9, 1058]. Curved [135]. Curves [24, 1728, 1072, 990, 1717, 1740, 1735, 120, 1002, 1751, 258, 1719, 17, 1720, 1732, 1737, 951]. Customer [1179, 760, 903, 360]. Customer-Supplier [903]. Customizable [1212]. Customizing [481]. Cut [425, 428, 419, 420]. Cuts [424, 1038, 1037]. Cyber [1368]. Cycle [336, 1690, 457, 872, 1201]. Cycles [1650, 1747]. Cyclic [101, 1290]. Cyclopentene [215]. Cylinder [15].

D [146, 154, 18, 1094, 1033, 1664, 196, 943, 992, 1020, 1677, 1065, 84, 29, 1640, 1075, 1060, 1093, 1023, 142, 1011, 1019, 1637, 30, 219, 1647, 1364, 1107]. DAG [129, 465]. DAG-Isomorphic [465]. Damage [1225]. DAML [580, 582]. DAML-S [582, 580]. Dangerous [1686]. Danish [903]. Data [377, 1685, 1638, 1619, 1595, 584, 965, 829, 380, 860, 1293, 1329, 1604, 1620, 861, 818, 862, 1113, 358, 730, 863, 1365, 1592, 865, 203, 349, 1663, 1608, 382, 408, 827, 183, 1242, 836, 395, 1621, 1671, 1515, 1455, 1501, 1508, 495, 1480, 379, 1013, 650, 1627, 869, 993, 1594, 1636, 910, 456, 1019, 770, 359, 824, 1654, 220, 873, 219, 1184, 98, 399, 1248, 1215, 837, 396, 373, 361, 397, 930]. Data-Intensive [1627]. Data-Parallel [1663]. Data-Semantics [1604]. Database [841, 843, 1337, 469, 1510, 835, 27, 1737, 1412]. Databases [1567, 912, 1560, 1575, 540, 1569, 1080]. Datacubes [1572]. DataFAN [1329]. DDS [1357]. Deadline [614]. Deadlock [1200]. Dealing [411]. Debug [1450]. Debugger [1624]. Debugging [96, 1257]. Decay [237]. Decimation [5]. Decision [1590, 1293, 1520, 1243, 1225, 611, 1199, 1580, 1488]. Decision-Making [1243, 1199]. Decision-Support [1225]. Decisions [1454]. Declarative [1]. Decoding [1165]. Decomposition [1081, 1709, 965, 1644, 194, 201, 949, 84, 50]. Decompositions [1708]. Decoupling [146]. Dedicated [1472, 692, 256]. Deduction [296]. Defect [904]. Deferred [867]. Defined [1249, 1322]. Defining [863]. Definite [1638]. Definition [13, 1272, 785]. Defocus [1021, 974]. Deformable [1062, 1056, 1019, 1061, 1055]. Deformation [136, 1019, 25]. Deformations [1020, 23]. Deforming [1035]. DEFORMOTION [1035]. Degree [1721, 279, 553]. Delaunay [129]. Delay [698, 600, 655]. Delegatable [1566]. Deletion [552]. Delimited [376]. DELISP [376]. Delivering [886]. Delivery [134]. Demand [614, 633, 233, 446]. Democratic [1347]. Deniability [268]. Dense [966, 1700, 1642, 963, 987, 981]. Density [934, 215, 410]. Deontic [288]. Dependence [699]. Dependency [1328, 1502, 465]. Dependent [1236, 824]. Deployment [822, 821]. Depth [1695, 1046, 984]. Derivation [1235]. Derivatives [115]. Deriving [1438, 1307]. Descent [248]. Describing [820]. Description [582, 551, 1033, 51, 1551]. Descriptions [1044]. Descriptor [1015].

Design [1323, 1414, 1439, 352, 112, 1197, 628, 252, 1417, 1702, 845, 60, 1452, 284, 1476, 906, 1258, 1244, 1327, 1205, 1243, 1666, 27, 1469, 186, 1212, 340, 1369, 1231, 622, 430, 798, 111, 530, 849, 72, 153, 1239, 192]. Designated [1230]. Designer [16]. Designing [1347, 82, 104, 866, 815]. Designs [339]. Detail [1178, 3]. Detected [915]. Detecting [942, 1270, 233]. Detection [1100, 966, 1200, 1094, 1069, 1314, 950, 127, 1475, 1635, 1608, 1112, 1254, 1192, 1097, 83, 1636, 1277, 1275, 1190, 406, 341, 1464]. Detections [410]. Detector [927]. Determination [1078, 6]. Determining [1584]. Determinism [237]. Deterministic [240, 644, 551]. Deutsche [243]. Developing [34, 725, 104, 208, 1386, 530, 78]. Development [1456, 588, 1599, 1625, 1337, 784, 904, 336, 905, 1319, 821, 353, 896, 339, 1237, 816, 868, 1446, 1320, 1324, 893, 910, 916, 573, 1354, 1389]. Developmental [1392]. Developments [898, 117, 158]. Device [313, 687, 761, 806, 864]. Devices [314, 1321, 312, 506, 494, 315]. DI-CUT [420]. Diabetic [1126]. Diabetic-Related [1126]. Diagnose [1423, 1257]. Diagnostic [1392, 1580]. Diagonalization [160]. Diagrams [831, 1598, 1293, 337, 222]. Dialectic [1431]. Dialectical [1354]. Dialogs [799]. Dialogue [1409, 1427, 311, 1430]. Dialogues [295, 311]. Dictionary [1557]. Didactic [1431]. Diff [703]. Diff-Serv [703]. Difference [190, 45, 941, 199]. Different [462, 166, 1551, 869]. Differential [1163, 279, 165, 163, 1164, 1175, 1173, 1172, 43, 1176, 1526, 211, 1532, 168, 1525]. Differential-Algebraic [43]. Differentials [1524]. Differentiated [664, 642, 649, 672, 604]. Differentiation [112, 114, 109, 702, 111, 655, 113]. Difficult [594]. Difficulties [1331]. Diffie [266].

DiffServ [688, 602, 603, 633, 1231, 318]. Diffuse [1046]. Diffuse-Specular [1046]. Diffusion [935, 945, 84, 1066]. Digital [1360, 493, 1361, 1348, 1343, 1352, 1358, 1359, 494, 1345, 1363, 133, 1355, 1353, 1461, 496, 1516, 1346, 219, 497, 1341]. Digitizing [1368]. Digraph [1692]. Digraphs [1748]. Dilation [1691]. Dimension [797, 817]. Dimensional [53, 1360, 437, 152, 223, 1103]. Dimensionality [1573]. Dimensioning [665, 644, 620]. Dimensions [860]. Direct [1033, 1664, 148, 1]. Directed [1706]. Directing [1456]. Direction [1106]. Directions [559]. Directory [1350]. Dirichlet [46, 1019]. Discipline [666]. Discourse [1427]. Discovering [1585, 1591, 362, 572]. Discovery [462, 687, 609, 1439, 387, 1554, 389, 404, 376, 1671, 392, 705, 1623, 1580]. Discrete [212, 1747, 1748, 198, 179, 659, 47]. Discrete-Time [179, 659]. Discriminant [1722, 1501]. Discrimination [539, 1083]. Discriminators [1083]. Discussion [1220, 1503, 478]. Discussions [1436]. Disease [1126]. Dishonest [277]. Disjunction [372, 1586, 575]. Disjunction-Free [372, 1586]. Disjunctive [1577]. Disk [132]. Dispatching [1672]. Dispersion [1684]. Displacement [23]. Displaying [1455]. Disposable [509]. Dissection [223]. Dissemination [1365]. Distance [1458, 1422, 1276, 836, 1455, 1137, 1024, 1379, 1486, 1386]. Distance-Based [1276]. Distances [118]. Distinguishing [1538]. Distortion [985, 932]. Distributed [614, 1200, 1633, 610, 144, 1649, 1398, 92, 1494, 1433, 322, 1207, 76, 904, 512, 667, 101, 229, 353, 453, 739, 553, 1332, 1333, 300, 1404, 574, 77, 340, 1657, 1307, 1553, 669, 1386, 1654, 71, 1389]. Distribution [440, 1134, 599, 1213, 404, 498]. Distributional [1070]. Distributions [739, 1041]. Distributive [1576]. District [482, 1262]. Divergence [1041]. Diverse [851]. Diversity [1497, 1495]. Divide [1355]. Division [610, 305]. Divisions [301]. DNA [521, 519, 548, 524, 520, 534, 542, 523, 540, 541, 530, 547, 527, 535, 522]. DNA-based [548]. DNASequenceGenerator [521]. Do [271, 787, 946]. Document [1589, 823, 1600, 766, 1308]. Documents [780, 782, 407, 393, 1563, 743, 392, 870, 478]. Does [1031, 1410, 1407]. Dog [949]. Doing [206]. Domain [162, 194, 201, 84, 1256, 1450, 50, 1403, 561, 1602, 417, 874, 688]. Domain-Specific [1256, 874]. Domains [305, 753, 1232, 851, 769]. Dominance [374]. Dominant [1028]. Dominated [148, 149]. Dominating [1688, 1685]. Done [929]. DÓRIS [1385]. Down [980, 391]. downloading [1219]. Downsizing [1227]. DOWNSIZINGX [1227]. dproc [94]. Drago [1335]. Dramatic [1462, 989]. DREAM [1056]. Driven [780, 726, 884, 108, 128, 1056, 351, 453, 816, 812]. Drop [657]. Dropping [642]. Drowning [1112]. DSQL [870]. DTDs [1589]. Duality [132]. DUCO [740]. Dumb [1459]. Duration [613]. During [1435]. DVB [663]. DVB-RCS [663]. Dynamic [1098, 1573, 723, 22, 266, 1000, 853, 1574, 623, 321, 1689, 1660, 176, 101, 229, 1399, 292, 156, 300, 236, 1241, 703, 834, 825, 1692, 1400, 1581, 799, 9, 1278, 769, 1030, 220, 674, 677, 98, 406, 713, 707]. Dynamical [986]. Dynamically [1245, 1321, 1363]. Dynamics [243, 619, 239, 213, 1454]. ’Dynamism [949]. Dyslexia [1392].

E-CARE [106]. E-Cash [490]. E-Catalogs [834]. e-Commerce [584, 785, 805, 481, 1380]. e-Documents [780]. e-Learning [1480, 776, 73, 1374]. E-mail [1221]. Earth [664]. Easy [325, 23]. ECN [648]. ECN-Aware [648]. Economic [1353]. Eden [338]. Edge [18, 445, 944, 612, 1691, 941, 926, 57]. Edge-Based [612]. Edge-Dilation [1691].

Edge-Failure [445]. Edge-Preserving [926]. Edges [940]. Edit [1458, 1086]. Edit-Union [1086]. Editor [1393]. Edmonds [416]. Education [207, 752, 1405, 1363, 1386]. Educational [1398, 778, 725, 1465, 728, 730, 731, 1396, 1403, 1400, 802, 803]. Educational-Game [725]. Effect [1051]. Effective [1459, 1441, 781, 1233, 762, 214, 767, 413]. Effectively [895]. Effectiveness [410]. Effects [906, 1587, 745, 1495]. Efficacy [1467]. Efficiency [645, 365, 669]. Efficient [1685, 1672, 12, 34, 467, 388, 394, 162, 628, 226, 402, 469, 1650, 1635, 1670, 1537, 107, 363, 184, 1588, 681, 1193, 1192, 1242, 836, 395, 376, 1176, 1726, 272, 1130, 1195, 181, 611, 1183, 396, 375, 221, 383]. Efficiently [1325, 258, 399]. Efficiently-Computable [258]. Effort [1456]. EHL [1661]. Eigen [949]. Eigen-Decomposition [949]. EigenSegments [1081]. Eigensolvers [147]. Eigenvalue [42]. EJB [835]. Elastic [201, 665]. Elasticity [203]. eLearning [726]. Electrical [1610]. Electromagnetic [138]. Electronic [551, 103, 162, 1218, 510, 1588, 790, 1515, 1230, 480, 451, 486]. Element [143, 1206, 84, 1711, 195, 189, 192]. Elements [155, 196]. Elevation [219]. Elimination [113]. Ellipses [1051]. Elliptic [1717, 1752, 1739, 1719, 1720, 1737, 1746]. EM-Based [1620]. EM-ICP [1120]. Email [1554]. Embedded [1321, 1312, 1316, 1535, 312]. Embedding [1086]. Emei [218]. Emerging [402]. Emotional [1470, 1474]. Emotions [1465, 284]. Emphasizing [536]. Empirical [912, 735, 1408, 749, 1330, 1464]. Empowering [578]. Emptiness [1040]. Empty [121, 1300]. Enabled [67, 82, 556]. Enabling [1610, 1613]. ENCAL [1476]. Encapsulating [1314]. Encoded [143]. Encryption [251, 1157, 249, 1535, 1156, 276]. End [1601, 710]. End-To-End [710]. End-Users [1601]. Ended [795]. Endomorphisms [258]. Energy [628, 1037, 1617, 1618]. Engine [741]. Engineering [1414, 449, 1405, 283, 890, 351, 851, 204, 852, 674, 329, 487]. Engines [565]. Enhance [468]. Enhanced [1642, 476, 550, 328]. Enhancement [1181]. Enhancing [321, 1663, 941, 410, 711]. Ennis [1351]. Ensemble [1081, 1211]. Ensembles [1507, 1607, 1511, 948, 1493, 1498]. Enterococcus [543]. Enterprise [1352, 827, 593, 218]. Entrepreneurial [1218]. Entropies [244, 933]. Entropy [370, 1273]. Entry [722]. Environment [207, 1414, 756, 89, 1337, 833, 730, 512, 1476, 739, 740, 1399, 1478, 1381, 1332, 792, 1403, 1437, 561, 1457, 1363, 1334, 15, 233, 1386, 749, 801, 709, 1389, 1380]. Environmental [1363, 1201]. Environments [18, 1619, 22, 1441, 725, 887, 1356, 1605, 643, 1455, 77, 704, 1483, 71, 1488]. Envy [301]. Envy-Free [301]. Enzymes [236]. Ephemeral [746]. Epistemic [1375, 287]. Equality [439]. Equation [171, 167, 1639, 400]. Equations [200, 165, 43, 1420, 64, 62, 109, 211, 37]. Equi [1631]. Equi-Join [1631]. Equilibration [101]. Equivalence [1304]. Equivalent [1721]. Era [883]. Erlangen [204]. Erlangen-Nuremberg [204]. ERP [218]. Error [1475, 1186, 1093, 1011]. Essay [1391]. Essential [402]. Estates [1349]. Estates-MIT [1349]. Estimates [947]. Estimating [1076, 66]. Estimation [966, 1111, 990, 1281, 934, 1029, 229, 898, 400, 1234, 984, 951, 924, 1050, 1064, 1106]. Estimators [947]. ETH [206]. Ethical [1468]. Ethically [208]. Ethics [208]. Ethnomethodology [1481]. Ethylene [215]. ETL [876]. euclidean [1061, 1002]. Eulerian [1056, 81]. Europe [588, 1374, 588]. European [1170]. Evaluate [1170]. Evaluating [752, 1360, 889, 1316, 1095, 1396, 840, 745, 639].

Evaluation [614, 1633, 723, 688, 783, 398, 641, 1445, 906, 1425, 1054, 793, 1469, 319, 228, 1457, 650, 484, 1483, 1442, 917, 640]. Evaluationof [1260]. Evaluator [1560]. Event [355, 599, 128, 1305]. Event-Driven [128]. Events [1585, 931]. Ever [826]. EVMS [896]. Evolution [551, 843, 1496, 901, 1313, 928, 1344]. Evolutionary [1649, 401, 1205, 362, 744, 1581, 532, 1209, 1196]. Evolving [1360, 1236, 845, 842, 401, 286]. Exact [421, 1679]. Example [218]. Examples [56, 1557]. Exception [450]. Exchange [268, 266, 267, 578, 515, 574]. Excitation [139]. Exclusion [1357]. Exclusion-Processes [1357]. Execution [580, 1648, 1321]. Exemplar [1141]. Exemplar-Based [1141]. Exemplification [797]. Exhaustive [1617]. Existent [1632]. Existing [691]. Expansion [1629]. Experience [899, 1350, 1319, 894]. Experienced [1432]. Experiences [900, 1632]. experiment [587]. Experimental [688, 641, 519, 539, 540, 1521, 1512, 602, 1434]. Experiments [1468, 64, 1410, 1353]. Expert [1223, 1226, 1519]. Expertise [819]. Experts [1513, 1520]. Explanation [1409]. Explanations [1431]. Explicit [339, 645, 951]. Exploiting [1181, 771, 1089]. Exploration [1441, 1617]. Explorativeness [861]. Exploratory [395, 795, 808]. Exploring [819, 1418]. Exponential [425]. Exponentiation [1195]. Exposing [1326]. Exposition [797]. Exposition-Exemplification [797]. Express [91]. Expressed [23]. Expressing [1294]. Expression [399]. Extended [501, 913, 875]. Extending [1326, 32, 248, 130]. Extensible [87, 94, 192]. Extension [1067, 1740, 1232, 1460]. Extensional [869]. Extensions [319]. External [1085, 838]. Extract [857]. Extracted [1401].

Extracting [412, 131, 393]. Extraction [1597, 719, 475, 386, 407, 262, 1085, 1193, 1237, 994, 1492, 1055]. Extreme [118, 1556, 839, 911]. Extruded [653]. Eye [1126, 1004].

Face

[1279, 1266, 1254, 1276, 1141, 1097, 1018, 1269, 1093, 1277, 1278, 1088, 1268, 1267]. Faces [1192, 1275]. Faceted [875]. Facilitating [795]. Facilitation [1220]. Facilities [1224]. Facility [1468, 430, 432]. Factor [1680]. Factor-2 [1680]. Factorial [1053, 1501]. Factoring [1752]. Factorisation [977]. Factorization [147]. Factorization-Free [147]. faecalis [543]. Failure [445, 1314]. Failures [1314, 341]. Fair [666, 502, 515, 702, 327, 620]. Fairness [355, 325, 648]. Families [892]. Family [1159, 1020]. Fast [259, 196, 5, 925, 1156, 1120, 1188, 1739, 139, 942, 941, 1165, 1653, 1209, 1116, 149, 328]. Faster [415]. FastMap [379]. Fault [685, 257]. Fault-Tolerant [685]. FCFS [684]. FDD [700]. Feasibility [1352, 287, 1202]. Feasible [305, 958]. Feature [841, 1003, 1303, 1237, 1508, 1492, 1271, 989, 1017, 1607]. Feature-Preserving [1017]. Features [937, 943, 784, 1303, 1408, 1215]. FEC [616]. Feedback [643, 1068, 1463]. Feeling [1260]. FEM [198]. FFT [57, 1647]. Fiat [273]. Fiber [679]. Fibre [644]. Fidelity [531]. Fiedler [66]. Field [1749, 55, 1743]. Fields [1721, 1736, 1249, 1053, 1745, 1726, 1751, 1754, 1720, 81, 1753, 1727, 1746]. Fighter [1198]. Figure [1063]. Figure-Ground [1063]. Fill [1709]. Fill-in [1709]. Filling [120, 145]. Film [1190]. Filter [1057]. Filtering [1072, 925, 455, 231, 791, 743, 769]. Filters [1051, 943, 945, 1705, 922, 1270]. Filtration [45, 199]. Filtration-Convection [45]. Finance [53]. Find [388, 1690]. Finding [421, 385, 1062, 1713, 1676, 1091, 452]. Fine [212, 1377, 1277]. Fingerprint [1272, 1270, 1271]. Finite [190, 122, 143, 1571, 1296, 84, 195, 49, 1745, 189, 1751, 1305, 1720, 86, 192]. Finland [900]. FIPA [314, 315]. FIPA-OS [315]. Firewalls [227]. First [1635, 303, 1737]. Fit [1019]. Fitness [1234]. Fitting [1620, 954, 1093]. Five [899]. Fixed [346, 162, 1099, 1688, 1702, 1747, 1512]. Fixed-Point [346]. Flaws [1175, 280]. FLD [1267]. Flexibility [692]. Flexible [82, 184, 1073, 877]. Flight [1422]. Flight-Training [1422]. Flow [966, 1329, 418, 428, 419, 941, 198, 109, 1654, 1668, 958, 1286]. Flows [935, 634, 1702, 132, 419, 673, 86]. Fluctuation [1260]. Fluid [963, 145, 42]. Fluid-Structure [145, 42]. Fluids [239]. Fluxes [81]. Fly [1085, 1636]. Focusing [915]. Fool [254]. Footprint [315]. Forbidden [1686]. Force [1012]. Forced [238, 170]. Forces [1085, 1355]. Fore [1692]. Fore-Armed [1692]. Forecasting [1189]. Foreign [1567]. Forensic [1283]. Forewarned [1692]. Forklift [1707]. Form [136, 1223, 1019]. Formal [563, 551, 1598, 304, 570, 297, 353, 1306, 30]. Formalization [346, 1595]. Formalizing [307]. Formally [347]. Format [1638]. Formation [305, 467, 1344]. Formats [1551, 869, 28]. Forms [1641]. Formula [1724]. Formulation [1638, 825]. Fortran [186, 1616]. Fortran90 [163]. Forums [478]. Forward [273, 1499, 234, 272]. Forward-Secure [272]. Forward-Security [273]. Forwarding [638]. Fostering [1408]. Fosters [1448]. Foundation [1288, 848, 1427]. Four [1563, 590, 6]. Four-Valued [1563]. Fourier [1644]. Foveated [1133]. FP [461, 391]. FP-Growth [391]. FP-Tree [461]. FPGA [1194]. FPGA-Based [1194]. Fractions [1746]. Fracture [198]. Frame [1119].

Frames [1011]. Framework [1597, 88, 584, 13, 778, 1598, 936, 727, 690, 1575, 1123, 1643, 865, 1068, 344, 195, 1696, 61, 743, 704, 480, 1183, 1386, 850, 1654, 508, 317, 830]. Frameworks [822]. Free [147, 1153, 372, 1586, 136, 1121, 1305, 301, 1128, 1019]. Freeze [1697]. Freeze-Tag [1697]. Frequency [637, 700]. Frequent [372, 1586, 392, 375]. Friction [180]. Frictionless [203]. Friendly [905]. Frieze [1016]. Front [1095]. Frontal [1275]. Frontier [1622]. FSMs [1196]. Full [138, 1572, 940, 1616]. Fully [152, 35]. Function [1159, 1542, 1749, 1745, 366, 471, 1146, 417, 1746]. Function-Based [366, 471]. Functional [473, 215]. Functionality [318]. Functions [279, 935, 1724, 415, 46, 1568, 1037, 256, 1024, 1093, 429, 1154, 1533]. Fundamental [982]. Fusion [967, 1113, 1606, 1520, 1269, 1514, 1521, 1512, 1117, 1504]. Future [1652, 1490, 1614, 1363, 671]. Futures [243]. Fuzzy [1243, 180, 1057, 763, 1231, 179, 747, 1518]. Fuzzy-Based [763].

G [658, 659]. GA [1584]. Gait

[1281, 1280, 1016]. Galois [1729]. Game [725, 453, 1462, 1291]. Game-Tree [453]. Games [1465, 308, 310]. Gamma [1041]. Gap [421, 447]. GARP [901]. Gas [1244, 721]. Gate [543]. Gateway [663, 227]. Gauss [1716, 925, 43]. Gaussian [177, 1214, 1068, 1739, 928, 605]. Gaze [1004]. GDT [397]. GDT-RS [397]. GEC [401]. GEMS [1655]. Gene [399]. Genealogy [1377]. General [262, 1027, 606, 56, 1504, 1440]. Generality [1706]. Generalization [416, 1215]. Generalizations [1716, 1561]. Generalized [1641, 200, 461, 933, 458, 986, 372, 1586, 400, 376]. Generate [1533]. Generating [1597, 1395, 1507, 299, 80]. Generation [1634, 22, 565, 759, 107, 651, 697, 1413, 1559, 765, 1400, 28, 799, 507, 709, 959, 10]. Generative [957]. Generator [246, 1536]. Generators [269, 260, 372, 1586, 627]. Generic [154, 562, 724, 472, 853, 262, 272, 855, 851, 1191, 931]. Genes [544]. Genetic [1197, 1208, 173, 1207, 476, 174, 224, 1276, 366, 1195, 1203, 463, 601, 1235, 1234, 1194, 1622, 225]. Genus [1728, 1688, 1732]. Geo [1480]. Geo-Referenced [1480]. Geographic [223]. Geom [659]. Geomechanics [202]. Geometric [1684, 1108, 1087, 1000, 6, 21, 1683, 1107]. Geometry [22, 986, 954, 33, 1617, 1487, 1050]. Getting [1447]. GHS [248]. GI [658]. GI-G-1 [658]. Giant [1751]. GIS [124, 1246, 130]. GIS-Integrated [1246]. Given [1436, 6]. Global [904, 578, 498, 1587, 1024, 549]. Global-to-Local [1024]. Glocalization [1342]. GMRES [41]. GNAT [1337, 1325, 1324]. Go [1407]. Goal [859, 291, 815, 741]. Goal-Based [291]. Goal-Oriented [741]. Goals [1428]. Goes [1407]. Going [1436]. Golden [496]. Good [890, 1488]. GOOSE [741]. Got [271]. Governing [1252]. GPRS [678]. Gradient [1110, 1066]. Grain [212]. GrAL [193]. Gram [1705]. Grammar [475, 754, 1252]. Grand [810]. Granularity [615]. Graph [1686, 554, 1377, 1038, 1037, 739, 337, 1137, 1693, 66]. Graphical [967, 352, 1417, 353]. Graphics [19, 20, 28]. Graphs [1684, 122, 558, 1687, 1688, 1650, 525, 413, 423, 975, 846]. Gravity [380]. Greedy [1556]. Grid [59, 53, 193, 1612, 82, 321, 60, 1661, 184, 70, 78, 1618, 452, 58, 1613, 217, 1660, 557]. Grid-Based [452]. Gridding [1131]. gridlib [184]. Grids [1651, 61, 77, 181]. Gross [167]. Ground [1063, 1508]. Group [721, 266, 1220, 165, 166, 693, 299, 500, 68, 1378, 1732, 1434, 1533]. Group-Oriented [1378]. Group-Related [299]. Grouping [1060]. Groups [1729, 262, 246, 247, 703, 1379]. Growth [376, 1667, 391]. GSM [1663, 678]. GSM/GPRS [678]. Guarantee [341]. Guaranteed [326]. Guaranteeing [651]. Guarantees [600]. Guard [88, 1676, 713]. Guidance [1418]. Guide [890]. Guided [756, 924]. Guideline [910].

H [1162, 553, 216]. H.263 [641]. H.263-Based [641]. H323 [691]. Haemodynamics [1611]. Hairstyle [14]. Half [1074]. Half-Spaces [1074]. Halonium [215]. Hamiltonian [161, 171, 169]. Hand [312]. Hand-Held [312]. Handheld [313, 314]. Handicapped [742]. Handling [450, 633, 550, 1553, 320]. Handoff [638, 330, 656, 652]. Handover [713]. Handwriting [1507, 1253]. Handwritten [1237, 1191, 1223, 1506]. Hard [439, 1496, 1333, 431]. Hardness [1687]. Hardware [491, 1204, 692, 494, 95, 1209, 1318]. Harmonic [41, 46]. Harmonious [1260]. Harness [76]. Hash [249, 1542, 256]. Hausdorff [1094, 1276]. Having [1551, 869]. HAWAII [638]. Health [103, 104, 482]. Heart [1610]. Heating [1262]. Height [1281]. Held [312]. Helicopter [1421]. Helios [670]. Hellman [266]. Helmholtz [1089]. Help [1450]. Helping [16, 1447]. Helps [1461]. Hessenberg [1641]. Hessenberg-Triangular [1641]. Heterogeneous [856, 1651, 843, 599, 861, 1555, 1378, 1569]. Heuristics [1696, 124, 1697, 1202]. Hidden [1676, 1743]. Hierarchical [475, 936, 380, 232, 276, 1640, 1243, 1018, 701, 1308, 1442, 222]. Hierarchy [1421]. Higgs [174]. High [324, 1339, 207, 18, 144, 564, 53, 202, 1625, 1612, 159, 692, 1652, 182, 635, 325, 1068, 188, 1621, 210, 1656, 47, 209, 1324, 333, 150, 1487, 331, 334, 913, 332, 1618, 837, 603]. High-Dimensional [53]. High-Integrity [1324]. High-Level [1068, 188, 209]. High-Performance [202, 1652, 837]. High-Precision [564]. High-School [1487]. High-Speed [325, 331, 334, 332]. Higher [279, 1723]. Higher-Rank [1723]. Highlight [1014]. Highlighting [1496]. Highly [279]. Hilbert [1719]. Hiperlan [637]. Hiperlan/2 [637]. HIRLAM [1665]. Histograms [933, 98]. Hitachi [891]. HLR [234]. Hoc [628, 630, 641, 629, 696, 698]. HOL [346, 347]. Hold [1614]. Holiday [946]. Homage [500]. Home [482, 709]. homeomorphic [7]. Homogeneous [1664]. Homotopy [162]. Horn [547]. HPC [1614]. HTML [777]. HTML/XML [777]. HTTP [322, 618]. Hulls [1678]. Human [1470, 1372, 1610, 950, 1468, 1432, 1436, 1273, 1076, 1453, 970, 960, 1199, 1077]. Human-Agent [1372]. Human-Provided [1470]. Humanoid [1255]. Humor [1373]. Hungarian [70]. HUTI [1643]. Hybrid [13, 679, 1499, 474, 1256, 791, 198, 178, 1387, 1428, 717, 803]. Hybridisation [523]. Hydro [157]. Hydro- [157]. Hydrodynamics [1250]. Hydrogenated [216]. Hyperbolic [47]. Hyperdynamics [969]. Hyperelliptic [1751, 258]. Hypermedia [720, 777, 778, 779, 783, 726, 728, 804, 729, 731, 733, 734, 737, 1388, 763, 743, 744, 1403, 868, 1390, 747, 802, 803]. Hyperspace [1397]. Hypertext [723, 757, 800]. Hypertexts [754]. Hypothesis [953]. hypre [182]. Hysteresis [1301].

i* [833]. I-P-O [906]. I-P-O/Multilateral [906]. I-Search [1562]. I/O [695]. IBM [1655]. ICA [1267]. Iceberg [1572]. Icing [242]. Iconic [1268]. ICP [1120]. Identification [273, 1283, 619, 1093]. Identity [276, 1520, 1521]. Identity-Based [276]. ie [1473]. ie-TV [1473]. IEEE [1336]. IEROCRYPT [1162]. IFT [692]. II [20, 1023]. ILE [1448, 1449]. Illuminant [1051, 1106]. Illuminants [1043, 1050]. Illumination [1045, 1049, 932]. Illusion [810]. I’m [1462]. Image [1104, 1133, 1099, 1095, 943, 1059, 992, 1070, 1056, 1193, 929, 1368, 1510, 981, 1046, 1045, 939, 1128, 1073, 1209, 946, 921, 952, 1268, 948, 1129, 1096, 1050, 932, 1260, 1066]. Image-Based [1368, 1268]. Imagery [963, 133]. Images [1072, 1081, 1108, 364, 1147, 1105, 1600, 996, 1042, 1219, 1270, 1140, 940, 142, 1143, 1035, 1055, 1058, 1049, 1064]. Imaginary [1736]. Imaging [146, 18, 1135]. IMAP [790]. Imbalanced [1514]. Impact [615]. Implementation [551, 12, 1417, 1228, 1238, 43, 1631, 110, 1646, 256, 181, 1194, 1239, 221]. Implementations [591]. Implementing [1741, 134, 492]. Implements [1560]. Implication [1561]. Implicit [504, 789, 339, 310, 970]. Importance [969]. Impossible [1173, 1526]. Impoverished [456]. Improve [235, 325, 1225, 1584]. Improved [1162, 442, 123, 387, 1153, 498, 1527, 420, 700, 1751, 115, 1693, 1532, 769, 432, 961]. Improvement [890, 1282, 895, 133, 883, 894]. Improvements [906, 327, 989]. Improving [1665, 1547, 301, 892, 1581, 1223, 776, 98]. In-Place [1711]. Inapproximability [1676]. Inbound [320]. Incidence [1179]. Including [1424, 1381, 1557]. Income [1349]. Incomplete [309, 1706, 1248]. Inconsistencies [1553]. Incorporating [1432, 1444, 1201]. Increases [1470]. Increasing [968]. Incremental [965, 1157, 380, 394, 407, 177, 1266, 1574, 461, 867, 1548, 375]. Indefinite [1067]. Independence [464, 1049, 938]. Independencies [466]. Independent [781, 1292, 865, 1288, 761, 1657]. Indeterminism [240]. Index [190, 116, 245, 1371, 1489, 1522, 1544, 1609, 1669, 1715, 1755, 281, 316, 335, 357, 414, 448, 489, 518, 555, 598, 714, 809, 879, 918, 972, 1032, 1092, 1148, 1177, 1185, 1265, 1285, 1311, 1340, 43, 1189]. Indexing [1123, 1080]. Indirect [1184]. Indistinguishability [560, 253]. Individual [821]. Individualism [1342]. Induced [132, 280]. Inductive [455]. Industrial [1194]. Industry [1584, 218]. Inequalities [422, 426]. Inference [464, 565, 847, 1576, 526, 583]. Infinite [1302, 1304]. Infinite-State [1304]. Influence [1323]. Information [1597, 1362, 718, 719, 857, 810, 112, 1596, 386, 235, 1629, 1570, 1562, 1554, 104, 309, 1059, 569, 786, 1663, 1233, 1543, 231, 1344, 302, 787, 1351, 746, 828, 1551, 1545, 1253, 824, 381, 1366, 771, 543, 596, 837, 223, 218]. Informational [1251]. Informative [1071]. Informing [1464]. Infrastructure [313, 1546, 78]. Inheritance [858, 1577]. Inheriting [1581]. Initial [782, 700, 897]. Initialize [1443]. Injection [257]. Inner [202]. Inner/Outer [202]. Input [203]. Inquiry [1440]. Insertion [552]. Insertion-Deletion [552]. Inspectable [1418]. Inspection [917, 135]. Instability [239]. Instance [1548, 874]. Instances [1700, 1547]. Institutions [307]. Instruction [764]. Instructional [1414, 1433, 1410, 1446, 1454]. Instrument [1549]. Insulated [250]. Integer [439, 441, 1205, 426, 429]. Integers [1719]. Integral [1531, 1731, 1317]. Integrality [421, 447]. Integrate [767]. Integrated [679, 861, 337, 340, 1324, 430, 1246, 220, 399]. Integrating [831, 1596, 1474, 1254, 1392, 572, 91]. Integration [566, 584, 829, 1114, 1059, 1575, 1063, 102, 1455, 211, 869, 1545, 350, 984, 218, 830]. Integrative [760]. Integrator [909]. Integrity [1339, 829, 1324]. Intellectual [1377]. Intelligence [175, 1578, 1261, 1552]. Intelligent [1220, 1250, 1570, 1562, 1395, 1452, 1432, 1588, 1582, 1444, 1466, 790, 1564, 1450, 1416, 1408, 1404, 1457, 1219, 1415, 1387, 1421, 799, 1594, 1484, 1246, 1386, 1442, 1385, 225, 877]. Intensional [869]. Intensities [159]. Intensity [1122, 150, 1106]. Intensive [1627]. Inter [467, 1403, 844]. Inter-agent [467]. Inter-Domain [1403]. Inter-organizational [844]. Interacting [1216, 660]. Interaction [1356, 1477, 1455, 1255, 802, 145, 42, 1418]. Interactions [139, 1303, 214, 1485]. Interactive [144, 22, 1417, 1085, 409, 100, 14, 27, 799, 11, 153, 25, 517]. Interactivity [776]. Interconnection [1659, 228, 1662]. Interest [942, 927]. Interesting [371, 488, 373, 397]. Interestingness [374]. Interests [738]. Interface [82, 661, 14, 27, 1400, 1322, 1463]. Interfaces [864, 505, 796, 1338, 717]. Interference [676]. Intermolecular [214]. Internet [753, 679, 694, 1233, 498, 1564, 739, 231, 699, 627, 619, 326, 1306, 750, 640, 712, 28]. Internet-Domains [753]. Interoperability [1546]. Interoperable [103]. Interpolating [1013]. Interpolation [54, 133]. Intersection [1684, 424, 1074, 416, 1733]. Intersession [1282]. Interval [847, 1241, 115, 227]. Interval-Based [847, 1241]. Intervention [1218]. Interworking [691, 671]. Intramodal [1520]. Intrinsic [1137]. Introducing [710]. Introduction [904, 282, 1374, 1341]. Intrusion [1406, 406]. Intuitive [797]. Invariant [943, 1052, 927, 1281]. Invariants [1736, 1122]. Invasive [1411, 768]. Inverse [1, 1497]. Inversion [138, 1638, 514]. Investment [1622]. Involving [1347]. Ions [215]. IP [626, 330, 689, 690, 693, 612, 673, 654, 655, 710]. IPSEC [280]. IPv6 [656]. Irregular [1628, 53, 1630, 25]. Irrelevant [883]. Islands [599]. Isn’t [905]. Isodistance [17]. Isogeny [1738]. Isomorphic [465]. Issue [1287]. Issues [694, 1226, 113, 1318, 1201, 1341]. ITA [774]. Itemset [388]. Iteration [56]. Iterative [202, 97, 36, 44, 1643, 1202]. ITSs [1459].

Jack [608]. Jack-in-the-Net [608]. Jacobian [1730, 1732]. Jacobians [113]. JADE [1386]. Janus} [68]. Java [1321, 92, 89, 565, 75, 1386, 74, 72]. Java/SQL [565]. JCSP.net [74]. JESA [705]. Job [400]. Join [1631]. Joins [1567]. Joint [251, 1035]. Jumping [402]. JumpStart [686]. Just [686]. Just-in-Time [686].

K-Order [221]. KASUMI [1171]. KDD [1601, 374, 1549]. KDD-Based [1549]. Kedlaya [1740]. KERMIT [1412]. Kernel [1685, 1094, 934, 1214, 1491, 1321]. Kernels [1067]. Key [268, 1171, 1567, 266, 267, 249, 1539, 250, 1718, 247, 1119, 254, 497]. Key-Insulated [250]. Keystream [260, 1536]. Keyterms [407]. Keyword [1256]. Kind [1393]. Knapsack [435, 533]. Knapsacks [439]. Knights [536]. Know [1447]. Knowledge [1414, 412, 1250, 1224, 1579, 1395, 283, 578, 733, 811, 813, 1563, 795, 1392, 1578, 1553, 1353, 1569, 1394, 1545, 1484, 1557, 872, 1623, 1401, 772, 487]. Knowledge-Based [1224, 1395, 1353, 1569]. Known [1105, 1158, 1050]. Kolmogorov [172]. Korean [1252]. Kronecker [1724]. KSACI [313]. Kullback [1041]. Kutta [164]. Kyoto [1358].

L [293, 1235]. L-system [1235]. Lab [1481]. Label [675]. Labeling [1680, 1681]. Labels [1680]. Lags [1591]. LAN [641]. Land [1508]. Landscapes [1617]. Language [412, 307, 469, 1232, 285, 1327, 1437, 591, 1247, 574, 1369, 1182, 1428, 583, 1318, 373]. Languages [554, 524, 560, 209, 850, 871]. LANs [635, 696, 680]. Large [154, 1642, 385, 599, 1659, 382, 1558, 96, 43, 1750, 395, 1374, 895, 540, 1080, 1732, 219, 710, 286].

Large-Scale [1642, 96, 43, 1374, 895, 710, 286]. Largest [388, 121, 1091]. Laser [151]. Last [322]. Last-Copy [322]. Latency [431]. Latent [1140]. Lattice [1668]. Lattices [1576]. Laundering [503]. Law [303]. Layer [139, 849, 681]. Layered [177, 642, 964, 643]. Layering [559]. Layers [585, 1064]. Layout [1224, 183, 1600]. Lazy [1674, 367]. LDA [1269]. Leakage [1543]. LEAP [314]. Learn [1450, 527]. Learned [1356, 890]. Learner [1417, 1435, 793, 1483, 1418, 1593]. Learners [734, 1447, 772]. Learning [1098, 1100, 1443, 1422, 1573, 1471, 841, 1263, 1439, 1550, 754, 1441, 1433, 756, 177, 1214, 1266, 1448, 1099, 1571, 1605, 1021, 1468, 1476, 1419, 1377, 1477, 788, 1397, 1444, 1375, 738, 739, 740, 1399, 1388, 455, 1097, 366, 1211, 1381, 1447, 1410, 1548, 1455, 1403, 1437, 745, 795, 1480, 1457, 178, 172, 1137, 767, 1393, 1140, 1382, 1379, 1139, 1394, 1483, 1486, 1144, 1386, 776, 456, 749, 1487, 801, 1199, 451, 486, 1488, 1454, 1142, 1440, 367, 773, 1380, 73, 1374]. Leash [949]. Left [145]. Leibler [1041]. Lemma [427]. Length [425, 1033, 539, 1119, 57]. Length-only [539]. Lens [985]. LEO [756, 664, 661, 662]. Less [488]. Lessons [1356, 890, 1480]. Level [1224, 1625, 283, 1068, 188, 651, 1520, 50, 911, 209, 978, 186]. Levenshtein [1137]. LEVIATHAN [1166]. Lexical [472, 473]. Lexicon [1099]. Libor [243]. Libraries [1612]. Library [1628, 193, 182]. License [1186]. Lid [1617]. Lie [166]. Lie-Group [166]. Life [336, 456, 872, 1201]. Life-Cycle [336, 1201]. Life-Long [456]. Lift [371]. Lifted [426]. Light [2, 1106]. Lighting [1]. Lightpath [677]. Lightweight [321]. Like [1031]. LILI [1540]. LILI-128 [1540]. Limit [1724, 1541]. Limited [1712, 1508, 507]. Limited-Use [507]. Linacs [159]. Line [1108, 1677, 1304, 507, 1253, 1282]. Linear [1685, 161, 1111, 1673, 197, 1642, 1150, 166, 261, 39, 1513, 64, 124, 62, 1514, 1093, 993, 37, 1528, 1532, 1694, 40, 1525, 995]. Linearity [1539]. Lines [154, 1111, 1213, 947]. Linguistic [747]. Link [620]. Linked [1476]. Linking [716, 347]. Links [324, 123, 711]. Linux [692]. Liquid [1197]. Liquids [236]. Lissajous [1051]. List [1165, 1694]. Listener [1338]. Listens [1470, 1407]. Listing [1052]. Literal [459]. Lithium [89]. Little [1342]. Lived [323]. Liveness [355, 1298]. Living [1372]. LKC [1394]. Load [621, 1660, 1665, 101, 1663, 230, 1264, 227]. Loading [1539]. Loads [1187]. Local [1550, 937, 1350, 944, 404, 1606, 1587, 1024, 1023, 1617, 956, 1010, 286, 930]. Localisation [1126]. Locality [98]. Localization [1695, 1113, 1276, 1018]. Localized [932]. Location [235, 630, 129, 234, 430, 683, 432]. Location-Aware [630]. LocationMessaging [105]. Loci [22]. Locus [1469]. Log [1358, 762, 1181, 1182]. Log-Coding [762]. Logarithm [1747, 1748]. Logic [774, 831, 1577, 1575, 459, 1243, 292, 1416, 288, 1566, 289, 1565]. Logic-Based [1243, 1566]. Logic-ITA [774]. Logical [1595, 848, 622, 1290, 876]. Logics [1375, 1576, 1551]. Logistic [1491]. Logistical [59]. Logistics [430]. LOGML [1182]. Logs [385]. Long [96, 699, 1088, 456]. Long-Range [699]. Long-Running [96]. Long-Term [1088]. Look [1031]. Lookahead [1692]. Looking [1109]. Lookup [692]. Loop [1323, 1648]. Loopy [1062]. Lorenz [374]. Loss [657]. Low [664, 506, 209, 1349, 410]. Low- [1349]. Low-Level [209]. Low-Power [506]. Lower [262, 425]. Lowering [722]. Loyalty [760]. LP [447]. LWFA [151]. Lyapunov [239].

M [516, 379, 417]. M-convex [417]. M-FastMap [379]. M2Tracker [920]. MAC [630, 661, 635, 1541]. Machine [1443, 841, 296, 1263, 1099, 1288, 1211, 1675, 816].

Machine-Independent [1288]. Machines [1713, 403, 1190, 298, 1491]. Machining [176]. Macronet [1206]. Made [777]. Magnetic [146, 1517, 142]. Magnetohydrodynamic [157]. mail [1221]. Maintaining [1446]. Maintains [760]. Maintenance [1246]. Majority [1560, 526]. Make [1376]. Makespan [1674]. Making [755, 114, 1243, 1199, 1286]. Malicious [1381]. Mammography [137, 1128]. Management [1251, 689, 690, 633, 904, 730, 693, 184, 811, 813, 840, 666, 827, 494, 234, 1317, 593, 618, 825, 704, 639, 1594, 1246, 496, 360, 916, 872, 613, 671, 318, 497, 603, 1286]. Managers [1406]. Managing [866, 1430]. Manhattan [1104]. Manifold [973]. Manifolds [1101]. Many [1728, 996, 1247, 983, 611]. Many-Sorted [1247]. Map [1358, 1646, 1247, 983]. Map-Based [1358]. Mapping [122, 473, 1555]. Mappings [1034]. Maps [563, 572, 23]. MArCo [1434]. Maria [1310]. Mark [243]. Market [1218, 511, 880, 1189]. Marketplaces [451, 486]. Markets [908]. Marking [1300, 654]. Markov [1053, 1039]. Markup [1182]. Mass [49, 81]. Master [1651]. Master-Slave [1651]. Mastery [1410]. Matching [841, 1589, 558, 1036, 987, 1242, 819, 988, 1076, 1259, 581, 1024, 1271, 1091, 946, 1116, 446, 1082, 416, 1006, 1007, 923, 1086, 932, 1239, 1009]. Material [1666]. Materialized [867, 396]. Materials [48, 51, 1615, 1049]. Mathematica [168]. Mathematical [626, 1207, 141, 205]. Matrices [1638, 197, 160, 1616]. Matrix [1641, 935, 171, 165, 63, 982, 54, 1639, 1667, 977, 168]. Matrix-Valued [935]. Matroid [1690, 416]. Matroids [428]. Maturity [886]. Max [428, 419, 420]. Max-Flow [428, 419]. MaxCSP [458]. Maximal [1154]. Maximally [356]. Maximizing [988]. Maximum [1698, 370, 1680]. MaxK [388]. MCMC [1029]. MD4 [1153]. MDG [347]. Me [1447]. Mean [1495].

Meaningful [719, 131]. Measure [374, 1256, 1041, 1502]. Measurement [1213, 827, 619, 703, 713, 1022]. Measurement-Based [619, 703, 713]. Measurements [143, 1235]. Measures [863, 484]. Measuring [65, 1446, 1617]. Mechanics [99]. Mechanism [378, 603]. Mechanisms [669, 1418]. Medial [1020, 1017]. Mediated [1435, 1546, 528]. Medical [140, 1580]. Medicine [102]. Medium [1417]. Meet [338]. Meeting [878]. Meets [758, 130]. Membrane [551, 550, 83]. Membranes [550]. Memory [226, 1635, 100, 1640, 819, 1317, 1548, 51, 109, 1540, 1647, 98]. Meningitis [397]. Mental [1383, 1222, 1274]. Mercy [1151]. Merging [1560]. Mersenne [1723]. Mesh [194, 5, 29, 669, 99]. Mesh-Based [194]. Meshes [2, 33, 1055]. Message [270, 278, 1249, 230, 319]. Meta [1593, 1568, 1202]. Meta-heuristics [1202]. Meta-learner [1593]. Metacognitive [1447]. Metadata [726, 578, 811, 813, 1400]. Metal [1584]. Metamodel [818]. Metamorphosis [7]. Metastable [237]. Meteorological [81]. Method [857, 53, 1644, 162, 1208, 1601, 171, 368, 403, 1752, 39, 1056, 1, 45, 474, 43, 455, 49, 1304, 1130, 62, 706, 1636, 852, 1617, 456, 1602, 42, 957, 86, 227, 222, 1672]. Methodological [216]. Methodologies [821]. Methodology [725, 449, 1504]. Methods [197, 462, 143, 36, 165, 166, 1449, 906, 167, 64, 1429, 868, 164, 851, 4, 169, 113]. METIORE [779]. Metric [1079, 432, 30]. Metrics [912, 910, 1061]. Metropolitan [69]. Micro [1321, 330]. Micro-kernel [1321]. Micro-mobility [330]. MICROBE [75]. MicroMint [492]. Micronuclear [544]. Micropayments [491]. Microreactor [523]. Microreactors [522]. Middleware [866, 189]. Might [1614]. Migrating [88]. Migration [98]. Mill [128]. Millennium [1210]. Min [428, 419]. Min-Cut [428, 419].

Miner [477]. Mini [587].

Mini-experiment [587]. Minima [956]. Minimal [995, 1090, 8, 1561, 110, 1195]. Minimally [356, 1411]. Minimization [1056, 417]. Minimized [1037]. Minimizing [273, 436, 444, 1674, 415, 698, 1011]. Minimum [119, 1709, 1676, 1689, 1690, 676, 836, 1307, 431]. Minimum-Interference [676]. Mining [377, 1619, 1178, 576, 364, 1574, 461, 402, 358, 1592, 363, 1582, 1558, 395, 371, 378, 1621, 1671, 1501, 1181, 379, 413, 365, 1182, 1183, 359, 1184, 399, 1215, 391, 373, 383, 397]. Minorities [277]. Minutiae [1272]. Mirror [1380]. Misclassification [1215]. Mislabeled [1547]. Misleading [782]. Missing [965, 1036, 411, 993]. MISTY [1168]. MISTY1 [1527]. Mixed [647, 63, 195, 198, 1469, 426]. Mixed-Hybrid [198]. Mixing [1567]. Mixture [1068]. Mixture-of-Gaussian [1068]. MML [387]. Mobile [685, 617, 1200, 1633, 314, 887, 689, 235, 661, 312, 656, 1365, 476, 629, 506, 180, 234, 1303, 1480, 69, 639, 233, 482, 331, 881, 1366, 837]. mobility [330]. modal [761]. Mode [1155, 270, 108, 700, 109]. Model [1098, 626, 563, 832, 1642, 1439, 1134, 853, 884, 1287, 1509, 1026, 901, 730, 863, 336, 1179, 805, 176, 1419, 1432, 1258, 1180, 351, 546, 1254, 827, 1276, 673, 737, 678, 1450, 1388, 579, 1192, 1587, 1121, 763, 198, 79, 940, 533, 994, 133, 939, 1127, 1093, 768, 852, 1693, 220, 219, 913, 1061, 1580, 771, 886, 345, 659, 255, 1257, 803, 726]. Model-Based [336, 1258, 1450, 1121, 994, 913, 1257]. Model-Driven [884]. Model-Free [1121]. Modeling [377, 1590, 143, 823, 1250, 1465, 48, 449, 647, 1620, 82, 152, 833, 323, 625, 284, 865, 733, 19, 20, 827, 928, 1450, 1301, 234, 1018, 1671, 14, 1333, 228, 619, 855, 1667, 15, 1412, 78, 807, 876, 812, 205, 1048, 751]. Modelisation [1487]. Modelling [1134, 1417, 804, 100, 1, 342, 1303, 1423, 1247, 1615, 828, 1378, 1390, 1379, 850, 802, 145, 1198, 1107, 1473]. Models [481, 1443, 1633, 1595, 967, 143, 607, 854, 387, 1033, 1207, 1071, 1068, 459, 944, 1054, 344, 736, 29, 931, 136, 141, 51, 745, 1135, 851, 622, 893, 1073, 983, 1235, 1485, 81, 970, 1019, 956, 655, 921, 874, 961, 465, 406, 1189, 1418, 25]. Moderate [1349]. Moderate-Income [1349]. Modern [36]. Modes [1156, 95]. Modification [574]. Modifications [646]. Modified [1542, 129, 648, 379]. Modular [1709, 1734, 1724, 1730, 1645, 1310, 1195, 828, 1324, 1733]. Module [1466]. Modules [1216, 523]. Molecular [217, 519, 214, 236, 213]. Molecules [547]. Money [503]. Monitoring [93, 92, 94, 95, 877, 1262, 773, 1217, 58]. Monotone [513]. Montages [1140]. Monte [63, 1660, 1, 64, 62, 1039, 66, 40]. Morphable [1093]. Morphological [1132]. Morphology [125]. Mosaic [1118, 4]. Mosaic-Based [1118]. Most [605]. Motion [146, 995, 990, 955, 1000, 985, 963, 987, 950, 992, 964, 1054, 1045, 1060, 117, 1028, 953, 970, 1035, 984, 1030, 952, 924, 957]. Motions [238, 949, 998]. Motivating [1452]. Motivational [1464]. Mount [218]. Movie [1519]. Movies [1052]. Moving [991, 1030]. MPEG [1472, 104, 1637]. MPEG-4 [1472]. MPEG-7 [104]. MPI [1665, 1631, 187, 1653]. MPIT [1626]. MPLS [675, 1231, 674, 329]. MPQS [1750]. MR [143, 137]. MR-Mammography [137]. MR-Measurements [143]. MRI [1130]. MUGI [1536]. Multi [1672, 688, 679, 450, 726, 1519, 449, 1509, 468, 277, 127, 643, 1365, 1120, 864, 761, 681, 1038, 291, 1205, 1243, 1301, 1097, 288, 85, 186, 920, 1212, 1135, 1615, 1203, 191, 1704, 939, 993, 601, 946, 483, 776, 1198, 286, 11]. Multi-Agent [1198, 450, 449, 468, 291, 288, 483, 776, 286]. Multi-attribute [1243]. Multi-camera [1038]. Multi-channel [1365]. Multi-class [1301]. Multi-constrained [1704]. Multi-device [864]. Multi-domain [688]. Multi-expert [1519]. Multi-layer [681]. Multi-layered [643]. Multi-level [186]. Multi-method [1672]. Multi-modal [761]. Multi-model [726]. Multi-objective [1205, 601]. Multi-part [1212]. Multi-particle [127]. Multi-party [277]. Multi-physics [85, 1615, 191]. Multi-purposed [679]. Multi-response [1509]. Multi-robot [1203]. Multi-sampled [1135]. Multi-scale [1120, 1615, 939]. Multi-view [1097, 920, 946]. Multi-volume [11]. Multiagent [287, 832]. Multicarrier [672]. Multicast [685, 684, 682, 643, 693, 681, 61, 600, 703, 601, 632, 683]. Multicasting [234, 1653]. Multichannel [630]. Multiclass [323, 1493]. Multiclassifier [1490]. Multicolored [118]. Multicommodity [418, 419]. Multidimensional [824]. Multifractal [606]. Multifrontal [35]. Multigrid [183, 1640]. Multihop [630]. Multilateral [906]. Multilayer [1188, 199, 52]. Multilinear [948]. Multilingual [407]. Multimedia [856, 644, 673, 790, 1302, 1392, 1594, 707, 709]. Multimodal [1521, 930]. Multinet [1500]. Multipeer [71]. Multiphase [99]. Multiple [1505, 386, 860, 1605, 1026, 1507, 1445, 1591, 1476, 986, 696, 678, 1043, 1501, 1492, 953, 1463, 1502, 1557, 1528, 1082, 1117, 1506, 320, 1050, 1064, 227, 1106]. Multiplication [1645, 258]. Multiplicative [1524]. Multipoint [629]. Multipole [38]. Multiresolution [12, 13, 933, 1254]. Multiresolution-Based [1254]. Multiset [1711]. Multistage [1659, 228, 1498]. Multisymplectic [167]. Multithreaded [93]. Multivariate [1069]. Multiview [1011]. MUMCUT [1330]. Mumford [979]. Munich [737]. Muon [157]. Musical [1549]. Must [1488]. Mutagenesis [546]. Mutual [506, 139]. My [1462, 946]. Myocardial [146].

Naïve [368, 367]. Nanoactuator [541]. Napster [495]. Natural [412, 469, 1232, 940, 3]. Natural-Language [469]. Navier [200, 107]. Navigating [597]. Navigation [804, 734, 788, 739, 740, 1469, 748, 26, 717]. Navigational [384]. Navigations [1364]. Navigator [748]. Naviz [384]. NCA [1713]. Nearest [1677, 1495, 408, 1700]. Needs [1236, 1447, 482]. Negation [469]. Negotiating [308]. Negotiation [304, 309, 760, 310, 311]. Neighbor [408, 221]. Neighborhood [173, 367]. Neighbors [1677, 1683]. NESSIE [1170]. Nested [862, 469, 1635]. Net [1293, 1296, 608]. NetAPI [586]. Nets [1295, 1298, 1299, 354, 1301, 1303, 1310, 744, 1305, 1308, 1309, 1286]. Network [377, 1499, 1214, 243, 623, 1702, 358, 1348, 1350, 1352, 667, 1663, 608, 1425, 1558, 1192, 791, 378, 198, 650, 704, 430, 1379, 639, 1229, 671, 710, 1354, 1498, 1217, 58]. Networked [1342]. Networking [636, 233, 334, 74]. Networks [686, 687, 753, 607, 621, 675, 1441, 679, 445, 628, 630, 177, 690, 599, 1347, 32, 1659, 833, 323, 476, 325, 176, 174, 1213, 1258, 651, 649, 1626, 1264, 612, 474, 697, 234, 1187, 698, 1186, 228, 175, 178, 703, 69, 663, 620, 1073, 669, 601, 1662, 333, 707, 655, 331, 334, 677, 709, 332, 710, 640, 1248, 329, 1261, 479, 652, 668, 1189, 662, 713, 487]. Neumann [46]. Neural [753, 1204, 1216, 243, 32, 176, 174, 1264, 474, 1192, 791, 1187, 1186, 175, 1229, 1248, 1498, 1189, 1217]. Neuro [1518, 1057, 707]. Neuro-dynamic [707]. Neuro-Fuzzy [1057, 1518]. Neuron [1206]. News [405, 750]. Next [651, 697]. Next-Generation [651]. Nimbus [1164]. NLG [1601]. NLG-Based [1601]. Node [1655]. Nodes [1587]. Noise [455, 1017, 680]. Noisy [1072, 1019]. Nokia [1303]. Non [551, 995, 517, 1294, 7, 1534, 931, 768, 1061, 932, 466]. Non-controllable [1294]. Non-cryptographic [1534]. Non-deterministic [551]. Non-euclidean [1061]. Non-generic [931]. Non-homeomorphic [7]. Non-interactive [517]. Non-invasive [768]. Non-linear [995]. Non-redundant [466]. Non-uniform [932]. Noncooperative [696]. Nonlinear [138, 279, 979, 244, 180, 111]. Nonlinearity [1154]. Nonmonotonic [585]. Nonmonotonicity [505]. Nonuniform [228]. Norm [1699]. Normal [29, 1739]. Normalized [1550, 1066]. Norms [307]. Nose [1515]. Notation [190, 352]. Note [554]. Notification [645]. Notions [267, 560, 274]. Novel [635]. Novice [1478]. NP [431]. NP-Hard [431]. NTRU [265]. Nucleic [538]. NUCOMP [1725]. Number [1721, 269, 1726, 272, 170, 1754]. Numbers [1726, 507, 509, 1743]. Numeric [362]. Numerical [1610, 165, 1664, 166, 1612, 64, 47, 187, 170, 157]. NURBS [24]. Nuremberg [204]. Nyström [1067, 164].

O [695]. O/Multilateral [906]. Object [1100, 1339, 1094, 1312, 348, 912, 1316, 854, 1099, 1084, 519, 822, 1115, 337, 1289, 1083, 27, 1247, 1045, 350, 1142, 345]. Object-Oriented [1316, 854, 822, 27]. Object-Relational [912]. Object-Z [348, 350, 345]. objective [1205, 601]. Objectives [413]. Objects [138, 1633, 853, 7, 1113, 1193, 1140, 1103]. Observation [961]. Observations [1088, 820]. Observer [180]. Occlusion [1003, 964]. Occlusions [966, 996]. Occupancy [1040]. Occurring [519]. Ocean [599, 100]. Octree [13]. Odd [1735]. ODEs [166]. Off [1587, 507, 692, 1215]. Off-Line [507]. Offline [491]. Offset [700]. OI [1561]. OI-Implication [1561]. Oil [1244]. OLAP [817]. OMC [1317]. OMC-INTEGRAL [1317].

Omnidirectional [1133]. On-Line [1253, 1282]. On-the-Fly [1636, 1085]. One [44, 1020, 569, 1510, 1719, 1103]. One-Class [1510]. One-Dimensional [1103]. One-Parameter [1020]. Ones [6]. Online [207, 1220, 676, 1696, 674]. only [539]. onto [1627]. OntoEdit [573]. Ontological [294, 1402, 1414]. Ontologies [595, 1553, 875, 1552]. Ontology [566, 585, 756, 1554, 827, 593, 561, 871, 573, 1401, 812]. Ontology-Based [566]. Ontology-Centered [1401]. Ontology-Driven [812]. Ontology-Guided [756]. Open [720, 1441, 1417, 323, 733, 353, 745, 795, 1306, 916, 878, 1341]. Open-Ended [795]. Open-Source [916]. Opening [1419]. OpenMP [187]. Operation [1155, 270, 541]. Operations [884, 363, 1264]. Operator [1263, 1132]. Operators [187, 181, 1581]. Opinion [1361]. OPO [?]denBurger:2002:TMT. Opportunities [557, 359]. Optical [966, 670, 635, 1666, 953, 669, 668, 958]. Optics [155]. Optimal [675, 385, 263, 140, 123, 1713, 309, 643, 651, 110, 1745, 988, 1259, 1627, 620, 1681]. Optimality [301]. Optimisation [1245, 1204, 1208, 1033, 1207]. Optimistic [515]. Optimization [1633, 995, 140, 127, 128, 520, 1276, 1640, 698, 1119, 1203, 1627, 1658, 1246, 632, 1096]. Optimization-Based [632]. Optimizations [183]. Optimized [1155, 174]. Optimizing [601, 532]. Optimum [1676]. Oracles [514]. Orbit [664]. Order [1729, 279, 165, 374, 1247, 926, 1514, 221]. Ordered [1592]. Orders [1742]. Ordinal [389]. ORESPICS [210]. Organization [1102, 895, 872]. Organizational [449, 1377, 291, 819, 844]. Organize [946]. Organizer [756]. Organizing [1646]. Orientation [1339, 1312, 1289, 1190].

Oriented [305, 859, 1316, 854, 269, 519, 822, 741, 27, 1240, 1378, 72]. Original [2]. Orthogonal [165, 163, 39, 15, 168]. Orthography [988]. Oscillatory [166]. OSCONVR [27]. OSIRIS [152]. Other [1698, 263]. Outdoor [1098, 1042]. Outer [202]. Outlier [1608, 410]. Output [684, 1665]. Outputs [1505, 1512]. Outsourced [450]. Outstanding [868]. Over-Simplification [1423]. Overcoming [456]. Overhearing [294]. Overlay [712]. Overturn [1664]. Overview [1422, 1618, 318]. Own [1447].

P [906, 551, 364, 408, 552, 549]. P-trees [364, 408]. Pacemaker [139]. Packaging [743]. Packed [1638]. Packet [1703, 650, 654, 241, 328, 603]. Packets [657]. Packing [1698, 437]. Packings [126]. Padding [280]. Page [226, 1358, 98]. Pages [386, 773]. Paging [1712]. Pair [1641]. Pairing [1741, 1733]. Pairings [1718]. Pairwise [214, 977]. PANAMA [1152]. Panoramic [1143]. Paper [505, 4]. Papers [592]. Paradigm [249, 1626, 1338]. Paradigms [499]. Paradox [1541]. Parallacs [1655]. Parallel [1641, 1624, 1634, 1648, 202, 97, 36, 1611, 1629, 89, 1625, 63, 1650, 1661, 107, 548, 1663, 35, 1264, 84, 1639, 96, 1645, 1631, 110, 1646, 1666, 1696, 62, 85, 186, 83, 256, 1657, 209, 191, 199, 211, 1658, 150, 873, 1622, 1647, 99, 86]. Parallelepiped [1107]. Parallelisation [100]. Parallelism [212, 1635, 1657, 1623]. Parallelism-Independent [1657]. Parallelisms [1637]. Parallelizable [270]. Parallelization [1628, 1668]. Parallelizing [1630, 1632]. Parameter [1620, 1688, 1020, 1231, 947]. Parameterized [1136, 930]. Parameters [995, 955, 1213]. Parametric [990, 1070, 17]. Parametrically [238]. PARDISO [37]. Parents [1581]. Pareto [1095]. Parse [1139]. Parsing [1252, 474, 1058]. Part [1208, 1212, 19, 20, 1023]. Partial [1114, 1252, 1548, 115, 773]. Participation [1347, 594, 1349]. Particle [1204, 152, 153, 127]. Particles [155]. Partition [119, 434, 1571]. Partition-Refining [1571]. Partitioning [1067, 1501, 1137, 99]. Partitions [842, 120, 1572]. Partner [902]. Parts [1036]. party [277]. Passenger [1365]. Passive [635]. Past [1652]. Pasting [1494]. Patches [1664, 1010]. Path [609, 1085, 46, 788, 605, 669, 537, 452]. Path-Finding [452]. Paths [31, 1686, 675, 139]. Patient [1611]. Patient-Specific [1611]. Pattern [1242, 376, 1671, 1122, 228, 1259, 1203]. Pattern-Growth [376]. Patterns [1496, 1384, 544, 402, 372, 1586, 376, 1016, 392, 410, 192]. Payment [832, 509]. Payoff [305]. PC [65, 695]. PC-Based [695]. PCA [1269, 1143]. PCR [520, 528]. PDE [185, 187, 191, 1131]. Pedagogical [1384, 1475, 1467, 1385]. Pedigrees [302]. Peer [67, 76, 916, 332]. Peer-to-Peer [67, 76]. Pencil [947]. Pencils [1075]. Pendulum [238, 170]. People [742, 920, 994, 1139, 1117]. Peptide [545, 538]. Per-Pixel [1068]. Perceived [1467]. Perception [1695, 457]. Perceptron [1188]. Perceptual [1060]. Perdu [1407]. Perfectly [278]. Performance [614, 1655, 18, 65, 144, 202, 97, 638, 330, 859, 679, 889, 1316, 1612, 692, 1652, 182, 321, 641, 229, 629, 1710, 1640, 234, 1621, 85, 319, 648, 55, 95, 1656, 228, 256, 650, 1627, 706, 484, 639, 769, 677, 113, 1584, 640, 711, 1096, 658, 318, 837, 603, 662]. Performing [897]. Perfusion [146]. Periodic [949, 1746]. Periods [272]. Permutable [1698]. Permutation [1534]. Permutations [1683]. Persistence [1470]. Persistent [634, 746]. Person [1274]. Personae [787]. Personal [900, 899, 1520, 1521, 105, 808]. Personalised [787]. Personality [284, 381].

Personalization

[715, 781, 727, 786, 746, 750]. Personalized [722, 726, 806, 765, 1401, 751]. Personalizing [731]. Personis [736]. Persons [1249]. Perspective [687, 908, 291, 1345, 996, 30]. Pesticides [1518]. Petri [1299, 354, 1301, 744, 1308, 1293, 1295, 1296, 1298, 1303, 1305, 1309, 1286]. PETSc [114]. Phantom [143]. Phase [937, 1213]. Phase-Based [937]. Phase-to-Phase [1213]. Phenomena [51, 241]. Phenomenon [495]. Pheromone [543]. Phones [1303, 1366]. Photo [1364]. Photographs [1008]. Photometric [976]. Photorefractive [1666]. Photos [716, 1516]. Phrases [475]. Phylogenies [1706]. Physical [955, 36, 1315, 1359]. Physics [1420, 1481, 1391, 1411, 1618, 85, 1615, 191]. Phytoplankton [136]. Pictures [1139]. Piecewise [9]. Pilot [1409, 1370]. Pilot-Testing [1409]. PIT [1628]. Pitaevskii [167]. Pixel [1068, 1007]. Pixels [1135]. PKC’98 [1542]. PKI [271]. Place [1218, 1711, 1308]. Placed [1102]. Placement [1208, 443, 1679]. Plaintext [1543, 1158]. Plan [460, 1376]. Planar [1685, 433, 1686, 998]. Plane [122, 22, 1109]. Plane-Based [1109]. PlanEdit [798]. Planes [1075, 1028, 1030]. Planning [1649, 1596, 140, 1238, 46, 1240, 117, 1602, 1434, 218, 1430]. Plans [287, 793, 1383]. Plasma [152]. Plate [1034, 1186]. Platform [314, 1424, 315, 1440]. Plausibility [1551]. Plausible [1565]. Plight [748]. Plots [1125]. PMI [498]. PNA [528]. PNA-mediated [528]. Point [346, 197, 162, 259, 1069, 118, 1079, 1001, 1031, 1739, 1751, 927, 258, 1681, 1720]. Point-Labeling [1681]. Pointer [1713, 234]. Points [1111, 1728, 1734, 973, 1744, 1747, 1731, 928, 942, 1270, 1680, 1030, 220]. Poisson [149]. Poker [527]. Polar [17].

Polemic [457]. Policy [689, 690, 693]. Policy-Based [690]. Polygonal [1678]. Polygons [1676, 124, 1022]. Polyhedral [433, 435]. Polynomial [1690, 54, 438, 422]. Polynomial-Time [438, 422]. Polynomials [1754]. Polytope [434]. Pool [651]. Popularity [888, 604]. Population [526, 1454]. Populations [1485]. Porphyry [1583]. Portable [93, 103]. Portal [217]. Portals [748]. Pose [1111, 1078]. Posing [1457]. Position [1001, 505]. Positioning [1711]. Positive [1638]. POSIX [1336, 1322]. POSIX-Ada [1322]. Possibilistic [1567]. Post [1505, 1436]. Post-practice [1436]. Post-processing [1505]. Postings [488]. Postponed [363]. Postviewer [1397]. Potential [1252, 247, 149]. Potentials [214, 926]. Power [617, 65, 514, 1316, 162, 506, 9, 1217]. Powers [1684]. Practical [857, 1329, 1161, 1636, 1226, 1632, 492]. Practice [269, 899, 890, 820, 1436]. Practice-Oriented [269]. Practices [902]. Precedence [1236]. Precision [564, 156]. Preconditioner [64]. Preconditioners [182, 38]. Preconditioning [197, 39]. Predictability [244, 1343]. Prediction [97, 243, 405, 762, 242, 739, 1187, 484]. Predictive [377, 235]. Predictor [1229]. Predictors [1211]. Preemptive [438, 589, 658]. Preference [1592]. Preference-Ordered [1592]. Preferences [295]. Prefetching [235]. Prefix [1296]. Preprocess [1325]. Preprocessor [1584]. Presence [657, 700, 680]. Present [1652]. Presentation [1601, 729, 1401]. Presentations [856, 782, 783, 1461]. Preservation [45]. Preserving [18, 926, 359, 1017]. Pretty [1331]. Pretty-Printer [1331]. Preventing [782]. Previewing [788]. Price [516]. Pricing [667]. Prime [57]. Primes [1750]. Primitive [1534]. Prince [1462]. Principal [1001, 944, 1074]. Principle [48, 370, 1421, 932]. Principles [539, 1438]. Printer [1331]. Prior [299, 738, 772]. Prior-Learning [738]. Priori [217]. Prioritization [1330, 713]. Priority [642, 655, 658, 1261]. Priors [978]. Privacy [511, 1343, 770, 359]. Private [517]. Probabalistic [1071]. Probabilistic [967, 1040, 464, 1123, 1063, 1084, 1671, 654, 962, 939, 1603, 970, 1077]. Probabilities [657, 681, 369]. Probability [940, 1532]. Probable [605]. Probing [602]. Probing-Based [602]. Problem [1685, 1245, 1698, 440, 1700, 200, 437, 1687, 121, 1688, 1702, 1723, 435, 418, 443, 1689, 476, 1674, 203, 840, 45, 1436, 400, 1413, 1675, 1240, 55, 1457, 199, 1719, 622, 1704, 798, 1681, 446, 431, 1194, 748, 432, 1697, 1411, 42]. Problems [161, 1698, 1628, 53, 1649, 421, 1596, 163, 1702, 1661, 536, 84, 1691, 1679, 420, 463, 559, 533, 1191, 216, 1493, 537, 535, 492]. Procedural [1316]. Procedures [202]. Process [900, 858, 194, 901, 905, 899, 176, 906, 890, 1243, 338, 816, 1481, 895, 883, 894, 850, 1584, 72, 1286]. Processes [1294, 244, 865, 519, 349, 1383, 1357, 157, 910, 150, 1290, 1058, 876, 844]. Processing [1593, 1250, 929, 1675, 1392, 1615, 873, 1505]. Processor [1651]. Processors [1316]. Produce [2, 883]. Producing [1152, 4]. Product [885, 584, 892]. Production [1448, 1201]. Products [907, 1212]. Professional [104]. Profile [108, 786]. Profile-Driven [108]. Profiles [1570, 739]. Profiling [1399, 406]. Program [722, 1625, 521, 206]. Programmed [546]. Programming [1633, 1316, 89, 1288, 188, 441, 195, 1205, 292, 366, 1332, 1240, 911, 209, 426, 707, 429, 878]. Programs [93, 1328, 1577, 1635, 459, 96, 1450, 204, 1632, 1257]. Progressive [771]. Project [571, 106, 1349, 910, 358]. Projection [1503, 42]. Projective [995]. Projects [897, 916]. Prominent [1270].

Promote [895]. Proof [355, 428]. Proofs [263, 249, 425]. Propagation [1062, 139, 1186, 151, 115, 1006, 241, 1239]. Properties [161, 355, 1108, 1539, 982, 524, 1176, 236]. Property [24, 428]. Proportionally [620]. Proposal [485, 1621, 1618, 535]. Proposals [782]. Proposed [1542]. Pros [735]. Prospect [1377]. Prostheses [1372]. Protected [498]. Protecting [493, 770]. Protocol [630, 661, 635, 1297, 500, 520, 516, 247, 1302, 600, 1306, 705, 90, 71, 640]. Protocols [832, 330, 866, 629, 334, 668]. Prototypes [1507]. Prototyping [831]. Provably [504]. Provided [1470]. Providing [710, 849]. Proving [454]. Provisioning [675]. Proxies [322]. Proximal [1446]. Proxy [618, 711]. ProxyCommunicative [297]. Pruning [322, 370, 328]. PSE [1661]. Pseudo [1079, 1273]. Pseudo-entropy [1273]. Pseudo-Metric [1079]. Pseudorandom [269, 1159, 246, 1534]. Pseudorandomness [1169, 1168]. PSS [263]. PTAS [436, 1675]. Public [249, 250, 1361, 1365, 1718, 770, 497]. Public-Key [249]. Publications [1583, 779, 746]. Publishing [790]. Pulse [151, 1203]. PUNCH [532]. Punctured [1731]. Pupils [1487]. Puppet [1461]. Purchase [1179]. Pure [1713]. Purified [541]. purposed [679]. Pursuit [383]. Putting [194]. Pwr4 [1655]. Python [188].

Q [1163]. QoS [688, 690, 656, 644, 651, 649, 612, 706, 601, 710]. QoS-Based [601]. QoS-Conditionalized [656]. Quadratic [440, 1736, 1284, 1727]. Quadrature [943]. Quadtrees [222]. Qualitative [1426, 1485, 1082, 1391]. Quality [207, 885, 1250, 863, 643, 906, 120, 888, 827, 1656, 891, 915, 902, 892, 880, 916, 881, 886]. Quality-Adaptive [915]. Quantification [469, 142]. Quantifiers [575]. Quantitative [143]. Quantum [268, 471]. QUARE [1162].

Quasi [1292, 203, 1003, 981, 62, 1622]. Quasi-Coupled [203]. Quasi-Dense [981]. Quasi-Monte [62]. Quasi-Parallel [1622]. Quasi-Sparse [1003]. Quasi-Static [1292]. Quasidictionary [1670]. Queries [1567, 1560, 469]. Query [472, 1560, 1629, 861, 591, 488, 583, 873]. Query-Based [472]. Query-Less [488]. Querying [562, 860, 862, 570, 1569]. Question [1232, 1429, 471]. Questions [1559]. Queue [666, 606, 603]. Queued [684]. Queueing [678, 605, 327, 655, 659]. Queues [1301, 660]. Queuing [889, 323]. Quickest [418]. Quite [1749].

R [1410]. RA [664]. Race [1635, 1636]. Racing [759]. Radial [942]. Radical [126]. Radio [775, 694, 650, 704]. Radiometric [1105]. RAM [1712, 1196]. RAM-Based [1196]. Random [914, 657, 1748, 1053, 253, 1682, 1693]. Randomized [1541, 1711]. Randomly [1102]. Range [1026, 1677, 699]. Rank [1735, 1723, 986, 1719]. Ranking [888]. Rapid [831, 10]. Rapid-Prototyping [831]. Raster [220, 222]. Rate [666, 275, 913, 317, 603]. Rating [789, 766]. Ratio [419]. Rational [1620, 42, 1380]. Rationale [820]. Ravenscar [1334]. Ray [2]. RBAC [765]. RBF [177, 1214, 178]. RC4 [1161]. RC6 [1528]. RCS [663]. RDF [562, 558, 565, 579, 819, 567, 575, 591, 574, 586]. Re [1662]. Re-routing [1662]. Reachability [1310]. Reactive [1292, 1264, 339, 211]. Reactivity [217]. Reading [1470, 723, 754, 1407, 767]. Real [614, 688, 1114, 1295, 1535, 759, 1085, 176, 1288, 519, 229, 1399, 1726, 1333, 79, 350, 1727, 286, 340]. Real-Time [688, 1114, 1295, 1535, 1085, 1288, 229, 1399, 1333, 350, 286]. Realisation [828]. Realistic [19, 20]. Reality [27, 1351]. Realization [49, 1363, 199, 1622, 543, 538]. Realm [221]. Rearrangeable [1662]. Reasoning [1598, 1449, 1003, 1450, 716, 1420, 1241, 1551, 1566, 1565, 1487]. Recherche [1407]. Recipient [1221]. Reciprocity [1089]. Recognition [1281, 1279, 1210, 1266, 1625, 1099, 1507, 1084, 1141, 1075, 1280, 141, 1186, 1122, 1237, 175, 1191, 1223, 1278, 1088, 1253, 1506, 1549, 1284, 938, 1215, 1138, 1267, 1077]. Recognizing [1585, 960]. Recommendation [722, 1249, 1588, 1582]. Recommendations [825]. Recommender [1179, 789, 791, 769]. Recommending [753]. Recomputation [108]. Reconstruction [31, 433, 144, 32, 1065, 1002, 1038, 981, 1110, 993, 1023, 30, 1008, 975, 1089]. Records [103]. Recovering [973, 1012]. Recovery [985, 992, 1046, 1043, 515, 1028]. Rectangle [1523]. Rectilinearity [1022]. Recursive [1638, 35, 1639, 922, 1501]. Recycling [457]. RED [325]. REDEEM [1459]. Reduce [976]. Reduced [1162, 1150, 1171, 1530, 1526, 1528, 1158, 1525]. Reduced-Round [1162, 1525]. Reducing [1630, 109]. Reduction [1641, 1685, 1573, 1642, 1513, 1188, 1304, 26]. Redundancy [854]. Redundant [657, 370, 466]. Reengineering [1319, 835, 892]. Reference [779, 737]. Referenced [1480]. Refined [196]. Refinement [348, 1288, 763, 1307, 289, 1308]. Refining [1571]. Reflectances [1043]. Reflective [599, 1015]. Refresh [867]. Region [1036, 1059, 920, 1058]. Region-Based [920]. Regional [1362]. Regions [1056]. Registration [146, 1133, 137, 1026, 1120, 1024, 1128, 1011, 1035, 1010, 1129]. Regression [1499, 1587, 1491]. Regular [1641, 238, 1690, 1305, 187]. Regularities [1344]. Regularization [18, 935, 926]. Regularized [945, 1073]. Regulated [1448]. Regulator [1727]. Rehearsal [456]. Reification [589]. Reinforcement [1471]. Rejections [1701]. Relabeling [1547].

Related [1171, 299, 1126]. Relation [888]. Relational [1573, 1590, 912, 1710, 773]. Relations [1604]. Relationship [1252, 903, 855, 413, 360]. Relationships [467]. Relativistic [152]. Relaxation [434]. Relaxed [485]. Relaying [629]. Relevance [1251, 931, 825]. Relevant [1144]. Reliability [1659, 468]. Reliable [685, 1321, 203, 61, 600, 1653, 71]. Remaindering [1743]. Remarks [165]. Remote [1642, 26]. Remotely [364]. Removal [1014, 1017]. Removing [1147, 1547, 106]. Rendering [1000, 19, 20]. Reno [647]. Reparation [1579]. Repeated [310]. Replacement [615, 226]. Replicated [1334]. Replication [468]. Report [1350, 1319, 1737]. Repositories [574]. Representation [1100, 950, 1123, 459, 964, 372, 1586, 855, 1545, 9, 1082, 1442]. Representational [1387, 1454]. Representations [1476, 1140, 978, 1143, 930]. Representing [862, 944, 1563, 575, 572, 824]. Reproducibility [1300]. Required [893]. Requirement [109]. Requirements [859, 890, 898, 838, 839, 868]. Rescue [1369]. Research [592, 1429, 334, 1341, 1613]. Reservation [651, 652]. Reservation-Pool [651]. Reservations [613]. Residuals [1134]. Resilient [445, 1154]. Resolution [607, 933, 47, 968, 958]. Resolve [880]. Resonance [146, 1517, 142]. Resonators [147]. Resource [442, 633, 634, 94, 438, 110, 604, 300, 618, 704, 645, 1622, 717, 318, 218, 1473]. Resource-Adaptive [717]. Resource-Optimal [110]. Resource/Connection [618]. Resources [566, 59, 664, 1396, 1712]. Response [444, 1105, 303, 1509]. Responsibility [208]. Responsible [208]. Rest [926]. Restarted [41]. Restoration [1202, 669]. Restoring [1012]. Restrict [363]. Restricted [914]. Restricting [221]. Restructuring [1232].

Results [1523, 723, 1114, 1676, 1169, 1223, 1434]. Resume [658]. Retaining [1197]. Rethinking [271]. Retransmissions [319]. Retrenchments [356]. Retrieval [1583, 564, 823, 1629, 1570, 1554, 363, 1233, 1510, 488, 1594, 1096, 478]. Retrospective [334]. Returns [1189]. Reusability [104]. Reusable [1395]. Reuse [814, 1473]. Reversal [110]. Reverse [108, 1230, 109]. Review [159]. Reviews [916]. Revised [265, 1297]. Revising [1396]. Revisited [1577, 278, 56]. Revisiting [999, 303]. Revocation [505]. Rewriting [831]. RF [158]. RFCT [477]. Rhapsody [336]. Rhetorically [1461]. Rhetorically-Structured [1461]. Right [321, 206, 929]. Right-Sizing [321]. Rights [493, 494, 496, 497]. Rigid [955]. Rigid-Body [955]. Rigidity [988]. RIJNDAEL [1533]. Ring [677, 668]. Rings [156, 1719, 150]. RIPEMD [1153]. Risk [505, 897]. Ritz [41]. RMD [318]. Road [242]. Roadmap [1375]. Roadmaps [956]. Roads [1246]. RoboCup [1199]. Robot [1695, 180, 1696, 1203]. RobotBased [1255]. Robots [1320]. Robust [934, 1136, 1120, 1282, 1073, 1127, 1428, 1278, 1117, 1143, 951]. Robustness [1294]. Rocket [882]. Role [464, 853, 825, 852, 938]. Roles [289, 1552]. Root [262]. Rooted [1351]. Roots [1754]. Rotation [1051, 943, 170]. Rough [1580]. Round [1162, 1150, 1171, 1530, 1168, 1526, 619, 1528, 1533, 1525]. Round-Trip [619]. Rounding [420]. Rounds [224]. Route [1238, 698]. Route-Cache [698]. Router [692]. Routers [695, 603]. Routine [163]. Routing [609, 621, 676, 866, 629, 698, 1704, 601, 674, 1261, 1662]. RS [397]. RSA [514]. RSA-Based [514]. RSBR [397]. RSIP [653]. RSVP [319]. Rule [585, 364, 1228, 363, 1227, 365, 747, 1580, 391]. Rule-Based [1228, 1227, 747]. RuleML [565]. Rules [449, 461, 1592, 389, 409, 1591, 370, 404, 839, 390, 362, 549, 1387, 1390, 1521, 1512, 611, 373, 397]. Run [1326, 94]. Run-Time [1326, 94]. Runge [164]. Running [96, 293].

S [580, 582, 1162, 1620, 1056]. S&P500 [243]. S-Parameter [1620]. Saclay [159]. Saddle [197, 1069]. Sail [759]. Salesman [421, 1194]. Sample [973, 1495]. sampled [1135]. Samples [1604]. Sampling [95, 969, 924]. Saskatoon [482]. SAT [420, 534, 535]. Satellite [664, 661, 663, 133, 711, 662]. Satisfaction [300]. Satisfiability [459]. SaTPEP [711]. Saturation [1530, 1149, 680]. Saving [617, 691]. SBLH [1160]. SC2000 [1174, 1525]. Scaffolded [734]. Scaffolding [1470, 1447]. Scalability [361]. Scalable [321, 701, 1667]. Scale [138, 154, 1642, 599, 96, 931, 43, 922, 1374, 895, 540, 710, 286, 1120, 1615, 939]. Scale-Space [922]. Scaled [988]. Scaled-Orthography [988]. Scales [1082]. Scaling [415, 417]. Scan [31, 402]. Scarf [427]. Scattering [211]. Scatternet [687]. Scatternets [631]. Scenario [859, 647, 815]. Scenario-Oriented [859]. Scenarios [759, 1389]. Scene [1118, 1065, 1031, 1038, 920, 134, 1042]. Scenes [1098, 986, 1011, 1030]. Schedulability [1333]. Schedule [1236]. Scheduled [1627]. Schedules [110, 597]. Scheduling [614, 1673, 1651, 610, 436, 1292, 444, 667, 438, 453, 672, 1675, 1657, 1322, 603, 878]. Schema [841, 843, 562, 565, 1555]. Schemas [842, 567]. Scheme [514, 34, 638, 437, 265, 1228, 762, 366, 50, 47, 650, 618, 1753, 652, 713]. Schemes [190, 1673, 504, 44, 263, 262, 616, 643, 438, 1679, 941, 199, 274, 1506]. Scholarly [1583, 746]. School [210, 1487]. Schreier [1740]. Schur [1641]. Science [207, 210, 882, 205]. Scientific [88, 87, 82, 1630, 188, 1481]. Scissors [542]. Scout [722]. Scratch [277]. Scream [1537]. Screw [1101]. Screw-Transform [1101]. Scriptable [87]. Scrutable [757]. SEA [1738]. SEAL [1159]. Search [1087, 470, 1562, 861, 173, 1554, 174, 1084, 1358, 453, 836, 741, 1202, 540, 1127]. Searchers [1478]. Searching [754, 781, 1677, 1696]. Second [165, 926]. Secure [1321, 103, 504, 267, 249, 278, 277, 512, 272, 274]. SecureClick [509]. Security [273, 251, 514, 887, 263, 693, 1168, 500, 1541, 1172, 494, 274, 770, 508, 280, 1529]. Segment [987, 1677]. Segmentation [980, 1519, 979, 1095, 1059, 1065, 1063, 1070, 1084, 1056, 1054, 1193, 1130, 1073, 15, 1035, 1129, 1066, 1009]. Segmentation-Based [1009]. Segmenting [920]. Selection [841, 1499, 1134, 16, 933, 1054, 1119, 523, 1492, 1428, 532, 1607]. Selections [1567]. Selective [1188, 961, 367]. Selectors [777]. Self [377, 1409, 1155, 624, 1649, 1448, 623, 1102, 1646, 1101, 991, 547, 997]. Self-Adaptable [1649]. Self-assembly [547]. Self-Calibration [991]. Self-Explanation [1409]. Self-Organization [1102]. Self-Organizing [1646]. Self-Regulated [1448]. Self-Similar [624, 623]. Self-Similarity [377]. Self-Synchronizing [1155]. Selling [760, 451, 486]. Semantic [857, 823, 571, 744, 581, 825, 1553, 587, 849, 582, 595, 576, 588, 556, 729, 594, 570, 560, 557, 592, 567, 590, 568, 559, 597, 577, 583, 573]. Semantically [719]. Semantically-Meaningful [719]. Semantics [580, 1583, 1604, 469, 297, 337, 1378, 1565]. Semi [47, 181, 30, 873]. Semi-discrete [47]. Semi-metric [30]. Semi-structured [873]. Semi-unstructured [181]. Semidefinite [434]. Semiotic [1345]. Semistructured [1595, 862, 393, 392, 824]. Sender [646].

Sender-Side [646]. Sense [716]. Sensed [364]. Sensing [1133, 797]. Sensing-Intuitive [797]. Sensitive [405]. Sensitivity [112, 1001, 648, 1201]. Sensor [967]. Sensors [1102]. Sentences [1325]. Separable [169]. Separation [422, 1046]. Sequence [1236, 981, 1016, 530, 997]. Sequence-Dependent [1236]. Sequence-to-Sequence [997]. Sequences [1178, 398, 521, 1132, 992, 1585, 1591, 762, 1046, 1119, 1555, 311, 952, 375]. Sequencing [797]. Sequential [985, 1591, 376]. Serendipity [1343]. Series [244, 1242, 836, 1189]. Serpent [1150]. Serv [703]. Served [1380]. Server [718, 781, 226, 736, 817]. Servers [856, 618, 320]. Service [582, 564, 235, 678, 1302, 326, 1306, 704, 705, 613, 671, 844, 814]. Serviceflow [844]. Services [1362, 1642, 1367, 726, 727, 571, 556, 642, 1179, 557, 827, 697, 604, 581, 577, 1353, 1546, 881, 709, 710]. Servoing [1121]. Sesame [562]. Session [604]. Session-Aware [604]. Sessions [385, 1180]. Set [1589, 1688, 978, 532, 219, 710, 1685, 388]. SET-Based [388]. SETM* [388]. SETM*-MaxK [388]. Sets [118, 1676, 1079, 1558, 946, 1580, 1330]. Settings [1222]. Setup [1236, 1482]. Several [1560]. Severe [633]. Seymour [428]. SFLASH [264]. SFS [133]. Shading [1025, 1014]. Shadow [1057, 975]. Shadows [2, 1147]. Shah [979]. Shaker [128]. Shamir [273]. Shape [990, 979, 1033, 1029, 1021, 1047, 974, 999, 1018, 51, 1076, 1024, 1025, 1014, 1127, 1093, 978, 1080, 1082, 1035]. Shape-from-Shading [1014]. Shape-to-Area [1024]. Shapes [1062, 1075, 4]. Shaping [623]. Share [225]. Shared [1635, 100, 1481, 683, 1647]. Shared-Memory [1647]. Sharing [1366]. Sheet [777, 1584]. Shelf [692]. Shipboard [175]. Shock [945, 1080]. Shock-Based [1080]. Shopping [479]. Short [243, 323, 428, 1429, 254, 478]. Short-Lived [323]. Short-Term [243]. Shortest [537]. Shot [1118, 232]. Should [503]. Si [216]. Side [646, 1183]. Sided [1639]. Siegel [1724]. SIETTE [1445]. Sieve [1749]. Sigma [886]. Signaling [686, 708]. Signalling [673]. Signature [251, 514, 263, 262, 265, 1282, 515, 274]. Signatures [273, 272, 513]. Significant [1550]. SIGUE [755]. Silhouette [994]. Silhouettes [1075]. Silicon [216]. SIM AGENT [293]. Similar [624, 623]. Similarity [377, 1273, 836, 1041]. SIMPA [1424]. Simple [777, 616, 308, 635, 548, 673, 422, 1083, 591, 236, 800, 1622, 1607]. Simpler [41]. Simplex [1677]. Simplification [5, 1423]. Simplified [326]. Simplifying [80]. SimStudents [1410]. Simulated [1456, 144]. Simulation [324, 144, 162, 194, 1664, 833, 128, 217, 176, 519, 184, 1663, 681, 1666, 1423, 214, 85, 1656, 191, 157, 1454, 149, 668]. Simulation-Based [1423]. Simulations [238, 148, 1660, 80, 706]. Simulator [1425]. Simultaneous [1078, 1445]. Single [1104, 402, 230, 954, 999, 1675, 133, 1662, 1622, 447, 1050]. Single-Box [1662]. Single-Scan [402]. Single-Sink [447]. Single-View [999]. Singular [965, 386]. Singularities [24]. Sink [447]. SIP [691, 673]. SIP-H323 [691]. Siphons [1298]. Site [889, 1224, 786]. Sites [784, 742]. Situation [175]. Situations [1299]. Six [886]. Size [1495]. Sizing [321]. Sketch [1048]. Sketches [30]. Sketching [14]. Skill [1263]. Skills [1425]. Skin [1254, 1192, 1277]. Skin-Color [1254]. Skin-Tone [1192]. Skipjack [1175, 1530]. Slave [1651]. Sliding [1680]. SLM [891]. Slopes [131]. Slovenia [697]. Small [1708, 1721, 421, 1494, 315, 1745, 893, 998]. Small-Footprint [315]. Smooth [1075, 1742]. Smoothing [1129]. SMP [1626, 1639, 1658]. Snaps [946]. SNet [833].

SNNB [367]. SNoW [1138]. SOBER [1539, 1538]. SOBER-t16 [1538]. SOBER-t32} [1539]. Social [1356, 1435, 1345, 740, 208, 815, 1255, 1379, 429, 1380]. Soft [614, 2]. Soft-Real-Time [614]. SoftPOSIT [1078]. Software [190, 900, 885, 185, 87, 1610, 194, 904, 901, 283, 905, 899, 1537, 908, 1319, 896, 195, 695, 189, 1393, 895, 882, 892, 893, 883, 880, 897, 820, 916, 881, 1199, 886, 341]. Software-Efficient [1537]. Solid [882]. Solution [691, 163, 16, 1650, 203, 84, 495, 1656, 332, 653]. Solutions [536, 1025, 537]. Solver [107, 1668, 40, 149]. Solvers [34, 97, 1643, 1639, 55, 187]. Solving [53, 1596, 201, 166, 43, 64, 62, 109, 533, 798, 37, 748, 168, 1411, 1493]. Some [1728, 723, 607, 782, 425, 165, 1169]. Sonar [1203]. Sorted [1247]. Sorting [1707]. Sound [1549]. Source [643, 47, 916]. Source-Adaptive [643]. Sources [861, 569, 302, 1551, 869, 611, 1106]. Space [1672, 185, 1649, 148, 1635, 120, 1368, 931, 1388, 1243, 922, 1302, 1503, 239, 1271, 1278, 968, 971, 149]. Space-Efficient [1635]. Space-Filling [120]. Space-Time [968, 971]. Spaces [394, 1074, 1508, 828]. SPAI [64]. Spanning [119, 1307]. Sparse [1100, 35, 1003, 187, 1667, 1616, 37]. Spatial [964, 408, 1247, 33, 69, 1364, 1366, 218]. Spatio [856, 1081, 1123, 922]. Spatio-Temporal [856, 1081, 1123, 922]. Speaker [1188]. Special [1749, 1681]. Specific [1597, 980, 1611, 1663, 1256, 523, 874]. Specification [355, 349, 371, 1302, 288, 1404, 340, 1306, 1389]. Specifications [845, 344, 339, 287, 1309]. Specifying [350]. SPECT [18]. Spectra [1517]. Spectral [1067, 936, 167, 45, 81]. Spectral-Difference [45]. Specular [1046, 1023]. Specularities [976, 952]. Speech [1210, 1392, 1191]. Speed [324, 635, 325, 767, 333, 331, 334, 332].

Speeding [1608, 258]. Speeds [1692]. Speedup [684]. Sphere [122]. Spheres [126, 6]. SPI [893]. Spin [579]. Splicing [554]. Spline [1034, 32, 9]. Split [424, 1732]. Splits [1501]. Splitting [44, 241]. Sporadic [1013]. Spotting [1256]. SPW [346]. SQL [565, 870]. Square [1529]. Squaring [1645]. Squashing [1608]. SquishQL [591]. SSC2 [1167]. SSL [280]. st [516]. st-Price [516]. Stability [1313, 164, 179]. Stable [305, 459]. Stacking [1509]. Stage [437, 1662]. Stakeholders [820]. Standard [266, 1336]. Standards [578]. Star [1698]. Starts [527]. State [636, 1571, 337, 342, 1302, 1304, 816, 52, 1464]. State-Charts [342]. Statecharts [343]. Stateless [327, 317]. States [1222, 237]. Static [1328, 1292, 1692, 406]. Statically [1627]. Station [1424, 1679]. Stations [696]. Statistical [1069, 1033, 1132, 1097, 1048]. Statistics [979, 939, 1514]. Steering [69]. Steiner [1687, 1702]. Step [1751, 164]. Steps [590]. Stereo [1090, 976, 920, 1091, 989, 984, 1006, 1007, 1009]. Stereopsis [1005, 1089]. Stereovision [1039, 1004]. Sticker [533]. Stiffness [47]. Stochastic [48, 1065, 1301, 338, 236]. Stock [511, 405, 1189]. Stokes [200, 107]. Storage [1638, 59]. Stories [1460, 1463]. Storing [562]. StoryML [1460]. Strategies [1651, 1087, 1605, 309, 83, 1349, 135]. Strategy [194, 252, 1556, 1403, 1110, 915, 1581, 1234, 221]. Stratified [1664, 1101]. Stream [1166, 261, 1537]. Streaming [712]. Streams [642, 681, 408, 1275]. Strengthening [1326]. Stride [1281]. String [534]. Strips [12]. Stroke [1253]. Strong [434, 68]. Structural [831, 1135, 465]. Structure [1197, 1000, 985, 177, 987, 992, 1168, 460, 1558, 996, 1028, 1030, 145, 998, 42, 837]. Structure-from-Motion [987]. Structured [89, 1554, 1563, 1646, 1461, 870, 1602, 873].

Structures

[185, 393, 125, 83, 1305, 1551, 1235, 142, 930]. Student [1443, 1470, 1465, 1419, 745, 768, 1482, 1442, 1464]. Students [1456, 1450]. Studies [215]. Study [324, 377, 1339, 912, 724, 682, 435, 1365, 1352, 1228, 735, 46, 646, 151, 1355, 213, 602, 21, 482, 871, 872, 1434, 1055, 668, 1330, 1464]. Studying [1360, 1610]. Style [777, 1363, 749]. Sub [1007]. Sub-Pixel [1007]. Subadditive [441]. Subdivision [4]. Subgraphs [1686]. Subject [1249]. Subjective [495]. Subjects [1468]. Submarine [1225]. Submodular [415]. Subnets [653]. Subspace [1279]. Subspaces [1071, 1074]. Substitute [1206]. Substrate [814]. Subsystems [673]. Subtypes [580]. Sugeno [179]. Suitability [1204]. Suite [90]. Summaries [1118, 747]. Summarization [472, 232, 470]. Super [1168, 958]. Super-Pseudorandomness [1168]. Super-Resolution [958]. Supercomputing [70]. Superconducting [158]. Supernodal [35]. Supplier [903]. Supply [1698]. Support [190, 685, 1402, 1625, 403, 1365, 964, 1436, 1225, 371, 1478, 530, 1488, 1190, 423, 487, 1418, 1491]. Supported [352, 1353, 1602, 877]. Supporting [207, 1601, 1419, 822, 1477, 482, 817, 844]. Supports [1409]. Surface [433, 32, 1120, 139, 1110, 1055, 1010, 975, 25, 1089]. Surfaces [31, 1090, 8, 1730, 1002, 1012, 1731, 17, 1023, 135, 10]. Surveillance [1115]. Survey [1722, 1059, 868, 903, 334]. Survivability [1225]. Survivable [826]. Surviving [1028]. Suspension [1668]. SVM [1214, 1511, 1138]. SWAD [588]. SWAD-Europe [588]. Swarm [1204]. Sweden [1355]. Sweep [1304]. Sweep-Line [1304]. Switch [684]. Switchback [108]. Switched [686, 675, 668]. Switches [692, 603]. Switching [650]. Sylvester [1639]. Sylvester-Type [1639]. Symbolic [22]. Symbols [1733]. Symmetric [161, 1638, 160, 1007]. Symmetrical [966]. Symmetries [1010]. Symmetry [1015, 942, 1270]. SymOntoX [593]. Symplectic [171, 169]. Synchronisation [700]. Synchronised [1295]. Synchronization [637]. Synchronizing [1155]. Synchronous [1435]. Synergically [869]. Synthesis [1003, 970, 1196, 957]. Synthetic [3]. System [1422, 1409, 721, 723, 1398, 92, 1220, 1250, 1224, 1519, 171, 1562, 641, 904, 785, 127, 644, 1452, 336, 1475, 1432, 1588, 1582, 1112, 1282, 789, 242, 1466, 1244, 1221, 1450, 339, 1243, 231, 1225, 180, 791, 1416, 64, 1621, 1310, 553, 27, 794, 1408, 744, 1237, 1227, 795, 1392, 134, 175, 1203, 619, 1320, 1415, 199, 471, 1387, 1390, 1421, 531, 1223, 1246, 1226, 483, 509, 482, 1500, 1487, 1364, 530, 1442, 1389, 543, 877, 1406, 347, 218, 1239, 1217, 878, 192, 1235]. Systematic [51, 892]. Systems [1597, 614, 1505, 685, 585, 752, 1200, 197, 551, 355, 1471, 1642, 1312, 1451, 724, 450, 1292, 449, 1599, 1395, 104, 661, 468, 1490, 1365, 731, 39, 1228, 101, 229, 821, 353, 1639, 354, 552, 43, 1289, 1205, 838, 287, 1420, 826, 815, 288, 1304, 672, 763, 62, 253, 1479, 1333, 1227, 1404, 236, 1446, 47, 340, 1657, 605, 1324, 549, 179, 1553, 1492, 1521, 1502, 1484, 37, 483, 747, 776, 350, 169, 168, 769, 671, 1385, 508, 1488, 26, 1198, 286, 717, 1262, 837, 223].

T [58]. t16 [1538]. t32 [1538, 1539]. Table [621, 453]. Tables [862, 1233, 1710]. Tackle [1649]. Tackling [1302]. Tag [392, 531, 1697]. Tag-Antitag [531]. Tail [606]. Tailorable [1332]. Tailoring [910]. Takagi [179]. Takagi-Sugeno [179]. Taking [579]. Tale [1138]. Tame [1745]. Tangent [1075, 239, 6]. Target [157]. Targeted [1475]. Task [1597, 305, 80, 456]. Tasking [1651]. Tasks [1627]. Tasksfor [1292]. Tate [1741, 1718]. TBR [109]. TCP [324, 647, 323, 325, 646, 648, 711]. TD [1655]. TEA [224, 1526]. Teach [1449]. Teachers [1472]. Teaching [1431, 1416, 209, 801, 205]. Team [904]. Teams [902]. Teamwork [1372]. Technique [1648, 257, 321, 896, 766, 1128, 916, 149]. Techniques [398, 128, 349, 19, 20, 888, 789, 420, 605, 1627, 1667, 133, 66, 1623, 399, 26, 1008, 1217]. Technologies [1610, 557, 102]. Technology [715, 810, 636, 1660, 376, 1392]. TeleCities [1348]. Teleconferencing [1004]. Teleoperated [1320]. TeLoDe} [1415]. Template [1116, 149]. Temporal [856, 1081, 860, 818, 1123, 404, 922, 1241, 289]. Temps [1407]. Tensor [935, 1060]. TensorFaces [948]. Tensors [999]. Tenth [1719]. Term [243, 1088]. Terminode [609]. Terms [1298, 47]. Terrain [131, 3, 133]. Terrains [1676]. Tessellations [126]. Test [1634, 688, 529, 1453, 526, 1330]. Test-Bed [688]. Testbed [679, 454]. Testing [1409, 914, 887, 110, 915, 913]. Text [462, 475, 472, 470, 1574, 716]. Textons [1145]. Texts [412]. Textual [354]. Texture [1051, 1047, 1041, 1146, 1093, 1048, 1044]. Textured [957]. Textures [1000]. Thai [1237]. Their [1360, 139, 1447, 1122, 921]. Theorem [172, 454]. Theoretical [1287]. Theories [304]. Theory [1040, 238, 889, 380, 1717, 579, 999, 56, 426, 1427, 660, 1260]. Thermal [52]. Thermo [203]. Thermo-Elasticity [203]. These [880]. Thickened [525]. Thickness [1131]. Thin [1034, 864]. Thin-Client [864]. Thing [206]. Three [1360, 177, 152, 1497, 1750, 141, 591, 223]. Three-Dimensional [1360, 152, 223]. Three-Layered [177]. Threshold [1705, 1301]. Threshold-Based [1301]. throughout [1374]. Throughput [680]. Tier [1430]. TigerSharc [1627]. Time [614, 1672, 1673, 686, 160, 637, 436, 688, 162, 1114, 1326, 1295, 1535, 244, 759, 661, 1296, 444, 1085, 569, 1690, 176, 1591, 1288, 349, 229, 94, 438, 1399, 1242, 836, 110, 422, 922, 124, 272, 553, 1333, 1407, 619, 179, 880, 289, 1540, 968, 1694, 350, 971, 286, 659, 1189]. Time-Code [661]. Time-Domain [162]. Time-Memory [1540]. Time-Minimal [110]. Time-Recursive [922]. Time-to-Market [880]. Time-Varying [553]. Timed [352, 1309]. Timed-Arc [1309]. Times [1236, 1675, 496]. Timing [257]. TIN [130]. Tissues [1131]. Today [787]. Together [577, 1376, 1201]. Token [230, 1308, 1291]. Token-Controlled [1308]. Tolerant [685]. Tomographic [1504]. Tone [1192]. Too [594]. Tool [88, 1443, 1459, 352, 1472, 1625, 1232, 905, 865, 1212, 151, 798, 90, 1602, 58]. Tool-Independent [865]. Tool-Supported [352]. Toolkit [100, 75, 79]. Tools [1451, 790, 1455, 1378, 871, 917]. Top [585, 980, 391]. Top-Down [980]. Topic [563, 462]. Topics [1445]. Topology [627, 622, 923, 1142, 58]. Total [18, 436]. Tourism [218]. Tourist [718, 727]. Towers [1745]. Town [1368, 1351]. Toxicity [1518]. TPAL [1309]. Traceback [654]. Tracing [2, 275, 502, 1336]. Tracking [155, 1000, 1664, 950, 904, 905, 1115, 738, 156, 920, 1255, 919, 962, 994, 953, 1183, 970, 1117, 960, 951, 971, 961]. Tractable [1688]. Tradeoff [1540]. Trading [1587, 1306, 225, 1215]. Trading-Off [1587]. Traditional [214]. Traffic [624, 623, 665, 625, 699, 672, 606, 228, 674, 329]. Trail [252]. Trails [753]. Trainable [1506]. Trained [1512]. Trainee [1425]. Training [1422, 1204, 1424, 1405, 1423, 1426, 1404, 1421, 1557, 749, 1248, 1406]. Traitor [275]. Trajectories [140]. Trajectory [1027]. Transaction [1297]. Transactional [1335]. Transfer [1436, 495, 523]. Transform [273, 942, 1101]. Transformation [843, 337, 9, 852, 583]. Transformations [1122]. Transformed [1140]. Transformer [1264, 1217]. Transient [634]. Transition [551, 354, 1307]. Transitions [1020]. Transitive [1634]. Translating [344, 343, 1309]. Translation [1099]. Translations [354]. Transmission [610, 278, 641, 275, 650]. Transparent [515, 1334]. Transport [81]. Transportable [1733]. Transportation [1698, 1365]. Transposition [453]. Trapezoidation [124]. Travel [1596]. Traveling [421, 1194]. Treatment [140, 269, 1468, 1392, 170]. Tree [185, 1590, 1708, 1687, 461, 453, 1646, 392, 1307, 1493]. Tree-Structured [1646]. Trees [1098, 119, 1686, 1509, 1587, 683, 431, 1714, 1086, 364, 408]. Treewidth [1709]. Trend [405]. Triangle [12]. Triangular [1641, 5]. Triangulation [1027, 129]. Triangulations [1678]. Trinomials [1729]. Trip [619]. TRIPLE [583]. Trivia [1559]. Truly [114]. Truncated [1172]. Trust [467, 499, 271]. Trusted [491]. Trustees [502]. Trusting [569]. TrustME [90]. Trustworthiness [302]. Truth [1508]. Trying [1376]. TSP [423]. TTL [698]. Tube [526]. Tubes [529]. TUC [1355]. Tumours [1517]. Tuning [1710, 655]. Tuple [328]. Turbulence [1664]. Turing [1453]. Tutor [1470, 1432, 1444, 1429, 1407, 1412]. Tutorial [1409, 1427, 1430]. Tutoring [1471, 1395, 1432, 1438, 1466, 1450, 1383, 1408, 1479, 1387, 1421, 1428, 1484, 483, 1442, 1385, 1411]. Tutors [1424, 1436, 1413]. TV [722, 1473]. TWIMC [1221]. Twin [1360]. Two [437, 899, 224, 1747, 459, 1639, 164, 1485, 1138]. Two-Cycles [1747]. Two-Dimensional [437]. Two-Literal [459]. Two-Sided [1639]. Two-Stage [437]. Twofish [1149]. Type [1689, 1168, 1639]. Types [855, 1500].

Ubiquitous [718, 712]. ULTRAflow} [845]. UML [831, 1598, 854, 842, 337, 898, 342, 1320, 874]. UMTS [700]. Unambiguous [1091].

Unbounded [272, 254, 255]. Uncalibrated [976, 992, 1028]. Uncapacitated [432]. Uncertain [965, 203, 613]. Uncertainty [1251]. Unconditional [277]. Unconditionally [274]. Undecidability [1719]. Understand [1263, 1447]. Understanding [854, 1604, 732, 908, 928, 1428, 1268, 596]. Unforgeable [1157]. Unification [485]. Unified [1121, 1247]. uniform [932]. Unilateral [203]. Uninitialized [1326]. Union [1086]. Universal [249, 546, 553]. Universality [545]. Universally [267]. Universities [1374]. University [204]. Unknown [1593]. Unlicensed [672]. Unordered [946]. Unsteady [233]. unstructured [181]. Unsuitable [1312]. Unsymmetric [37]. Update [394, 375, 221]. Updates [1714, 817, 773]. Uploaded [1321]. Upper [1532]. Urban [1370]. Usability [784, 888]. Usable [561]. Usage [1178, 730, 1582, 695, 834, 1231, 1182, 1183]. Usage-Centric [834]. Use [481, 1471, 1361, 1449, 1436, 494, 1187, 1325, 1462, 507]. Used [1270, 1379]. User [481, 718, 780, 385, 783, 1570, 730, 786, 905, 864, 762, 736, 738, 505, 763, 825, 770, 406, 1473]. User-Driven [780]. Users [689, 1601, 578, 208, 807]. Using [1458, 1339, 831, 1431, 462, 475, 610, 995, 1067, 753, 841, 1321, 1263, 467, 1451, 1441, 859, 472, 887, 235, 1000, 1062, 48, 1208, 1570, 1033, 1114, 1448, 1029, 364, 1207, 623, 104, 1095, 468, 862, 786, 1365, 933, 230, 100, 1264, 866, 370, 1003, 1639, 1252, 460, 408, 789, 1606, 29, 1739, 1497, 1388, 1572, 1301, 180, 1631, 234, 839, 716, 366, 189, 1016, 1303, 14, 50, 141, 1041, 744, 1110, 920, 765, 95, 1076, 1480, 1195, 1137, 767, 1060, 1203, 619, 1320, 209, 1324, 3, 1719, 1387, 1014, 133, 1355, 1553, 90, 983, 28, 1093, 1271, 1235, 15, 1278]. Using [707, 910, 1253, 1019, 655, 1006, 1622, 113, 98, 951, 1229, 1055, 1580, 320, 543, 328, 1107, 345, 135, 773, 406, 397, 192]. Utilising [1476]. Utility [616, 317]. Utilization [396].

Validating [863]. Validation

[1287, 1579, 838, 379, 399, 361]. Validity [810]. Valleys [1617]. Value [965, 908, 390, 495]. Valued [935, 1563]. Values [440, 965, 1593, 411, 1557]. Variability [1282]. Variable [324, 212, 183, 1119]. Variables [1326, 1558]. Variance [1511]. Variants [1153, 1497, 550, 172]. Variation [18, 932]. Variational [48, 171, 973, 974, 1024]. Varieties [311]. Various [1418]. Varying [160, 1043, 553]. Vascular [144, 142]. Vector [1698, 403, 43, 1388, 66, 220, 1190, 1066, 1491]. Vector-Space [1388]. Vectors [41, 29]. Vegas [647]. Vehicle [1673]. Velocity [143, 922]. Velocity-Adapted [922]. Vendors [909]. Ventricular [145]. Verification [832, 1297, 519, 1282, 1520, 1269, 340, 1521, 1226, 290]. Verificationfor [1258]. Verified [347]. Version [1542]. versus [22, 1587, 187, 880]. Vertex [5, 1689, 1694, 113]. Very [385, 540, 1116]. via [273, 832, 1496, 889, 387, 667, 944, 1038, 1037, 288, 362, 1555, 1463, 399]. Vickrey [306]. Video [614, 1118, 615, 967, 955, 232, 642, 641, 643, 1123, 1056, 1112, 1193, 1124, 1141, 1137, 1275, 1125, 968, 1004, 1077]. Video-Based [1112]. Video-on-Demand [614]. Video-Teleconferencing [1004]. Vietnamese [1191]. View [1281, 259, 950, 986, 1003, 999, 993, 1290, 1354, 959, 1097, 920, 946]. View-Based [950]. View-invariant [1281]. Viewpoint [1049]. Views [1590, 1026, 867, 1043, 983, 1142, 396]. Virtual [1356, 1315, 529, 1405, 1352, 27, 1363, 1462, 1382, 71, 709]. Virtuality [1351]. Viscosity [1025]. Vision [934, 1113, 1613]. Visual [1113, 732, 1121, 379, 1255, 956, 938, 961, 596, 1138]. Visualization [154, 1398, 184, 142, 1637, 219, 11, 153, 1217].

Visualizations [80]. Visualizer [384]. Visualizing [1178, 69]. Visually [742]. Vivo [1517]. Vlasov [148]. VLSI [1634]. Vocabularies [572]. Vocabulary [1099, 572]. VoIP [616]. Volcanoes [1738]. Volterra [1072]. Volume [86, 11]. Volumetric [1055]. Voronoi [126, 222]. VORPAL [151]. Votes [1494]. Voting [510, 1063, 1060]. Vs [1316, 1431, 1214, 578, 1138]. VxWorks [1324].

Want [787]. WAP [1297, 640]. War [503]. Warehouse [396]. Warehouses [856, 818]. Warehousing [865]. Warping [1278]. Water [1250, 19, 20]. Watersheds [1193, 1254]. Wave [1213, 241]. Wavelet [1005, 1193, 3]. Wavelet-Based [1005]. Wavepacket [212]. Waves [139]. Way [1109, 23, 846]. WDM [686, 621, 677]. Weak [464]. Weaknesses [247]. Weather [1665]. Weaving [592]. Web [1406, 1596, 774, 856, 566, 617, 582, 719, 857, 1402, 595, 1178, 576, 588, 1398, 1367, 755, 781, 1220, 386, 889, 887, 385, 407, 226, 1562, 571, 556, 784, 862, 729, 1405, 786, 1468, 760, 864, 732, 594, 1396, 1358, 570, 1180, 1582, 560, 762, 788, 557, 827, 1450, 592, 815, 1478, 742, 567, 794, 792, 764, 590, 787, 795, 392, 746, 1404, 816, 765, 1181, 796, 767, 581, 568, 559, 597, 577, 1387, 1182, 1483, 1183, 509, 583, 749, 770, 824, 573, 1184, 881, 1229, 1401, 320, 596, 773, 814, 383, 1552, 751, 1380]. Web-Applications [816]. Web-Based [1406, 774, 1398, 1220, 1405, 760, 1450, 815, 794, 764, 795, 767, 1387, 1483]. Web-Served [1380]. Web-Services [1367]. WebForum [488]. WebFrame [383]. Website [384]. Websites [481, 1350, 888]. WebTheme [596]. WEDS [1482]. Weierstrass [1744]. Weight [1307]. Weighted [123, 1079, 431, 227]. Weights [1496]. Weil [248, 1718]. Welfare [429]. Well [1244]. Westfield [1350]. Westwood [325].

WETAS [794]. Where [1407, 915]. Which [6, 1560]. Whiplash [528]. Who [787]. Why2 [1391]. Why2-Atlas [1391]. Wide [781, 252, 923]. Widespread [590]. WINDS [801]. Wireless [685, 628, 641, 1297, 651, 696, 678, 69, 68, 601, 748, 331, 671, 640, 652]. Wires [106]. WIS [849]. WISECON [1564]. Witch [1462]. within [1249, 819]. without [491, 684, 1695, 1047, 502, 501]. Wizard [1238]. Word [484]. Words [393]. Work [1424, 453, 1410, 1481, 1286]. Workers [482]. Workflow [1294, 845, 1308, 877, 844, 846]. Working [68]. Workstations [1626]. World [1367, 781]. Worlds [1360, 1315, 1357]. Woven [1460]. Writing [1391]. WTLS [280].

XCBC [1156]. XECB [1156]. XML

[566, 1589, 777, 782, 104, 1460, 867, 848, 766, 568, 28, 830, 192]. XML-Based [830]. XSL [104]. XTEA [1526].

Years [899]. You’re [1462].

Z [348, 350, 345]. ZIP [1158]. Zodiac [1173]. Zone [1446]. Zurich [206].

References

Jolivet:2002:IDL

[1] V. Jolivet, D. Plemenos, and P. Poulingeas. Inverse direct lighting with a Monte Carlo method and declarative modelling. Lecture Notes in Computer Science, 2330:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                              link.springer-ny.com/link/service/
                                              series/0558/bibs/2330/23300003.
                                              htm;   http://link.springer-ny.
                                              com/link/service/series/0558/papers/
                                              2330/23300003.pdf.

Debelov:2002:LMO

[2] V. A. Debelov and I. M. Sevastyanov. Light meshes — original approach to produce soft shadows in ray tracing. Lecture Notes in Computer Science, 2330:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300013.
pdf.

Perez:2002:ASD

[3] M. Perez, M. Fernandez, and M. Lozano. Adding synthetic detail to natural terrain using a wavelet approach. Lecture Notes in Computer Science, 2330:22– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300022.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300022.
pdf.

Seo:2002:NAS

[4] S. H. Seo, D. W. Kang, Y. S. Park, and K. H. Yoon. The new area subdivision methods for producing shapes of colored paper mosaic. Lecture Notes in Computer Science, 2330:32– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300032.
pdf.

Franc:2002:FAT

[5] M. Franc and V. Skala. Fast algorithm for triangular mesh simplification based on vertex decimation. Lecture Notes in Computer Science, 2330: 42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300042.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300042.
                                        pdf.

Roanes-Macias:2002:GDS

[6] E. Roanes-Macías and E. Roanes-Lozano. Geometric determination of the spheres which are tangent to four given ones. Lecture Notes in Computer Science, 2330:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300052.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300052.
                                        pdf.

Elkouhen:2002:MNH

[7] M. Elkouhen and D. Bechmann. Metamorphosis of non-homeomorphic objects. Lecture Notes in Computer Science, 2330:62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2330/23300062.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2330/23300062.pdf.

Cosin:2002:BSM

[8] C. Cosín and J. Monterde. Bézier surfaces of minimal area. Lecture Notes in Computer Science, 2330:72– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300072.
pdf.

Ryu:2002:TDB

[9] J. Ryu, Y. Cho, and D.-S. Kim. Transformation of a dynamic Bspline curve into piecewise power basis representation. Lecture Notes in Computer Science, 2330:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300082.
pdf.

Zhang:2002:RGC

[10] J. J. Zhang and L. You. Rapid generation of C continuous blending surfaces. Lecture Notes in Computer Science, 2330:92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2330/23300092.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2330/23300092.
pdf.

Wilson:2002:IMV

[11] B. Wilson, E. B. Lum, and K.-L. Ma. Interactive multi-volume visualization. Lecture Notes in Computer Science, 2330:102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2330/23300102.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2330/23300102.
pdf.

Belmonte:2002:EIM

[12] Ó. Belmonte, I. Remolar, J. Ribelles, M. Chover, and M. Fernández. Efficient implementation of multiresolution triangle strips. Lecture Notes in Computer Science, 2330:111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300111.
pdf.

Boada:2002:HOT

[13] I. Boada and I. Navazo. The hybrid octree: Towards the definition of a multiresolution hybrid framework. Lecture Notes in Computer Science, 2330:121– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300121.
pdf.

Mao:2002:IHM

[14] X. Mao, K. Kashio, H. Kato, and A. Imamiya. Interactive hairstyle modeling using a sketching interface. Lecture Notes in Computer Science, 2330: 131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300131.
pdf.

Ryoo:2002:OCC

[15] S. T. Ryoo and K. H. Yoon. Orthogonal cross cylinder using segmentation based environment modeling. Lecture Notes in Computer Science, 2330:141– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300141.
pdf.

Essert-Villard:2002:HDS

[16] C. Essert-Villard. Helping the designer in solution selection: Applications in CAD. Lecture Notes in Computer Science, 2330:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300151.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300151.
pdf.

Puig-Pey:2002:PIC

[17] J. Puig-Pey, A. Gálvez, and A. Iglesias. Polar isodistance curves on parametric surfaces. Lecture Notes in Computer Science, 2330:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300161.
pdf.

Antonelli:2002:TVR

[18] L. Antonelli, L. Carracciuolo, M. Ceccarelli, L. D’Amore, and A. Murli. Total variation regularization for edge preserving 3D SPECT imaging in high performance computing environments. Lecture Notes in Computer Science, 2330:171–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300171.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300171.
pdf.

Iglesias:2002:CGTa

[19] A. Iglesias. Computer graphics techniques for realistic modeling, rendering, and animation of water. Part I: 1980-88. Lecture Notes in Computer Science, 2330:181–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300181.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2330/23300181.
pdf.

Iglesias:2002:CGTb

[20] A. Iglesias. Computer graphics techniques for realistic modeling, rendering and animation of water. Part II: 1989-1997. Lecture Notes in Computer Science, 2330:191–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300191.
pdf.

Schramm:2002:CSG

[21] É. Schramm and P. Schreck. A case study in geometric constructions. Lecture Notes in Computer Science, 2330: 201–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300201.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300201.
pdf.

Botana:2002:IVS

[22] F. Botana. Interactive versus symbolic approaches to plane loci generation in dynamic geometry environments. Lecture Notes in Computer Science, 2330: 211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300211.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2330/23300211.
pdf.

Peyre:2002:DED

[23] H. Peyré and D. Bechmann. Deformations expressed as displacement maps: An easy way to combine deformations. Lecture Notes in Computer Science, 2330:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300219.
pdf.

Arnal:2002:PSN

[24] A. Arnal, A. Lluch, and J. Monterde. A property on singularities of NURBS curves. Lecture Notes in Computer Science, 2330:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300229.
pdf.

Zheng:2002:IDI

[25] J. J. Zheng and J. J. Zhang. Interactive deformation of irregular surface models. Lecture Notes in Computer Science, 2330:239–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300239.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2330/23300239.
pdf.

Vazquez:2002:BRT

[26] P.-P. Vázquez and M. Sbert. Bandwidth reduction techniques for remote navigation systems. Lecture Notes in Computer Science, 2330:249– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300249.
pdf.

Marir:2002:OIV

[27] F. Marir, K. Ouazzane, and K. Zerzour. OSCONVR: An interactive virtual reality interface to an objectoriented database system for construction architectural design. Lecture Notes in Computer Science, 2330:258– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300258.
pdf.

Rodeiro:2002:ICG

[28] J. Rodeiro and G. Pérez. Internet client graphics generation using XML formats. Lecture Notes in Computer Science, 2330:268–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300268.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300268.
                                        pdf.

Kim:2002:CNV

[29] D.-S. Kim, Y. Cho, and D. Kim. The compression of the normal vectors of 3D mesh models using clustering. Lecture Notes in Computer Science, 2330: 275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300275.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300275.
                                        pdf.

Sosnov:2002:SMF

[30] A. Sosnov, P. Macé, and G. Hégron. Semi-metric formal 3D reconstruction from perspective sketches. Lecture Notes in Computer Science, 2330:285– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300285.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300285.
                                        pdf.

Alberts:2002:RSS

[31] C.-P. Alberts. Reconstruction of surfaces from scan paths. Lecture Notes in Computer Science, 2330:295–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300295.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2330/23300295.
pdf.

Echevarria:2002:ENN

[32] G. Echevarría, A. Iglesias, and A. Gálvez. Extending neural networks for B-spline surface reconstruction. Lecture Notes in Computer Science, 2330:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2330/23300305.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2330/23300305.
pdf.

Otero:2002:CGS

[33] C. Otero and R. Togores. Computational geometry and spatial meshes. Lecture Notes in Computer Science, 2330:315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2330/23300315.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2330/23300315.
pdf.

Bhowmick:2002:CSD

[34] S. Bhowmick, P. Raghavan, and K. Teranishi. A combinatorial scheme for developing efficient composite solvers. Lecture Notes in Computer Science, 2330:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2330/23300325.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2330/23300325.pdf.

Irony:2002:PFR

[35] D. Irony, G. Shklarski, and S. Toledo. Parallel and fully recursive multifrontal supernodal sparse cholesky. Lecture Notes in Computer Science, 2330:335– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300335.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300335.
                                        pdf.

Cai:2002:PIM

[36] X. Cai, Y. Saad, and M. Sosonkina. Parallel iterative methods in modern physical applications. Lecture Notes in Computer Science, 2330:345– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300345.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300345.
                                        pdf.

Schenk:2002:SUS

[37] O. Schenk and K. Gärtner. Solving unsymmetric sparse systems of linear equations with PARDISO. Lecture Notes in Computer Science, 2330:355– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300355.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300355.
                                        pdf.

Guillaume:2002:MAP

[38] P. Guillaume, A. Huard, and C. Le Calvez. A multipole approach for preconditioners. Lecture Notes in Computer Science, 2330:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300364.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300364.
pdf.

Herrero:2002:OML

[39] H. Herrero, E. Castillo, and R. E. Pruneda. Orthogonal method for linear systems. preconditioning. Lecture Notes in Computer Science, 2330:374– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300374.
pdf.

Tan:2002:AMC

[40] C. J. K. Tan. Antithetic Monte Carlo linear solver. Lecture Notes in Computer Science, 2330:383–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300383.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300383.
pdf.

Boojhawon:2002:RSG

[41] R. Boojhawon and M. Bhuruth. Restarted simpler GMRES augmented with harmonic Ritz vectors. Lecture Notes in Computer Science, 2330:393– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300393.
pdf.

Voss:2002:PMR

[42] H. Voss. A projection method for a rational eigenvalue problem in fluidstructure interaction. Lecture Notes in Computer Science, 2330:403–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300403.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300403.
pdf.

Kulikov:2002:IVG

[43] G. Y. Kulikov and G. Y. Benderskaya. On implementation of vector Gauss method for solving largescale systems of index 1 differentialalgebraic equations. Lecture Notes in Computer Science, 2330:412–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300412.
pdf.

Ciegis:2002:OCS

[44] R. Ciegis and V. Pakalnyte. One class of splitting iterative schemes. Lecture Notes in Computer Science, 2330:422– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300422.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300422.
pdf.

Kantur:2002:FCP

[45] O. Kantur and V. Tsybulin. Filtrationconvection problem: Spectral-difference method and preservation of cosymmetry. Lecture Notes in Computer Science, 2330:432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2330/23300432.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2330/23300432.pdf.

Karnik:2002:CSD

[46] M. Karnik, B. Dasgupta, and V. Eswaran. A comparative study of Dirichlet and Neumann conditions for path planning through harmonic functions. Lecture Notes in Computer Science, 2330:442–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300442.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300442.
pdf.

Naidoo:2002:AAH

[47] R. Naidoo and S. Baboolal. Adaptation and assessment of a high resolution semi-discrete numerical scheme for hyperbolic systems with source terms and stiffness. Lecture Notes in Computer Science, 2330:452–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300452.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300452.
pdf.

Cox:2002:CMC

[48] D. Cox, P. Klouček, and D. R. Reynolds. The computational modeling of crystalline materials using a stochastic variational principle. Lecture Notes in Computer Science, 2330: 461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300461.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300461.
pdf.

Leinen:2002:RFM

[49] P. Leinen. Realization of the finite mass method. Lecture Notes in Computer Science, 2330:470–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300470.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300470.
pdf.

Marsden:2002:DDU

[50] R. H. Marsden, T. N. Croft, and C.- H. Lai. Domain decomposition using a 2-level correction scheme. Lecture Notes in Computer Science, 2330:480– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300480.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300480.
pdf.

Melnik:2002:CMM

[51] R. V. N. Melnik and A. J. Roberts. Computational models for materials with shape memory: Towards a systematic description of coupled phenomena. Lecture Notes in Computer Science, 2330:490–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300490.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300490.
pdf.

Shevchuk:2002:CTS

[52] V. A. Shevchuk. Calculation of thermal state of bodies with multilayer coatings. Lecture Notes in Computer Science, 2330:500–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300500.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300500.
pdf.

Berridge:2002:IGM

[53] S. Berridge and H. Schumacher. An irregular grid method for solving highdimensional problems in finance. Lecture Notes in Computer Science, 2330: 510–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300510.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300510.
                                        pdf.

Husek:2002:PPM

[54] P. Hušek and R. Pytelková. On polynomial and polynomial matrix interpolation. Lecture Notes in Computer Science, 2330:520–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300520.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300520.
                                        pdf.

Mohr:2002:CPS

[55] M. Mohr and B. Vanrumste. Comparing the performance of solvers for a bioelectric field problem. Lecture Notes in Computer Science, 2330:528– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300528.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300528.
                                        pdf.

Pedersen:2002:IRE

[56] P. W. Pedersen. Iteration revisited examples from a general theory. Lecture Notes in Computer Science, 2330:538– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300538.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300538.
pdf.

Seguel:2002:NPE

[57] J. Seguel, D. Bollman, and E. Orozco. A new prime edge length crystallographic FFT. Lecture Notes in Computer Science, 2330:548–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300548.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300548.
pdf.

denBurger:2002:TMT

[58] M. den Burger, T. Kielmann, and H. E. Bal. TOPOMON : a monitoring tool for grid network topology. Lecture Notes in Computer Science, 2330:558– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300558.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300558.
pdf.

Bassi:2002:LSR

[59] A. Bassi, M. Beck, E. Fuentes, T. Moore, and J. S. Plank. Logistical storage resources for the grid. Lecture Notes in Computer Science, 2330: 568–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300568.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300568.
                                        pdf.

Gelas:2002:TDA

[60] J.-P. Gelas and L. Lefèvre. Towards the design of an active grid. Lecture Notes in Computer Science, 2330:578– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300578.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300578.
                                        pdf.

Maimour:2002:ARM

[61] M. Maimour and C. Pham. An active reliable multicast framework for the grids. Lecture Notes in Computer Science, 2330:588–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300588.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300588.
                                        pdf.

Mascagni:2002:PQM

[62] M. Mascagni and A. Karaivanova. A parallel quasi-Monte Carlo method for solving systems of linear equations. Lecture Notes in Computer Science, 2330:598–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2330/23300598.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2330/23300598.
pdf.

Fathi:2002:MMC

[63] B. Fathi, B. Liu, and V. Alexandrov. Mixed Monte Carlo parallel algorithms for matrix computation. Lecture Notes in Computer Science, 2330: 609–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300609.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300609.
pdf.

Liu:2002:NEM

[64] B. Liu, B. Fathi, and V. Alexandrov. Numerical experiments with Monte Carlo methods and SPAI preconditioner for solving system of linear equations. Lecture Notes in Computer Science, 2330:619–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300619.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300619.
pdf.

Atanassov:2002:MPP

[65] E. I. Atanassov. Measuring the performance of a power PC cluster. Lecture Notes in Computer Science, 2330: 628–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300628.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300628.
                                        pdf.

Srinivasan:2002:MCT

[66] A. Srinivasan and M. Mascagni. Monte Carlo techniques for estimating the Fiedler vector in graph applications. Lecture Notes in Computer Science, 2330:635–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2330/23300635.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300635.
                                        pdf.

Curley:2002:PPC

[67] M. G. Curley. Peer-to-peer computing enabled collaboration. Lecture Notes in Computer Science, 2330:646– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300646.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300646.
                                        pdf.

Pascoe:2002:WTS

[68] J. S. Pascoe, V. S. Sunderam, and R. J. Loader. Working towards strong wireless group communications: The Janus architecture. Lecture Notes in Computer Science, 2330:655–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300655.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300655.
pdf.

Pascoe:2002:TMC

[69] J. S. Pascoe, V. S. Sunderam, R. J. Loader, and G. Sibley. Towards mobile computational application steering: Visualizing the spatial characteristics of metropolitan area wireless networks. Lecture Notes in Computer Science, 2330:665–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300665.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300665.
pdf.

Kacsuk:2002:HSG

[70] P. Kacsuk. Hungarian supercomputing grid. Lecture Notes in Computer Science, 2330:671–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300671.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300671.
pdf.

Stuer:2002:CRM

[71] G. Stuer, F. Arickx, and J. Broeckhove. The construction of a reliable multipeer communication protocol for distributed virtual environments. Lecture Notes in Computer Science, 2330: 679–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300679.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300679.
                                        pdf.

Welch:2002:POD

[72] P. H. Welch. Process oriented design for Java: Concurrency for all. Lecture Notes in Computer Science, 2330:687– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300687.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300687.
                                        pdf.

Alexandrov:2002:CCL

[73] N. Alexandrov, J. S. Pascoe, and V. Alexandrov. Collaborative computing and E-learning. Lecture Notes in Computer Science, 2330:688–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300688.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300688.
                                        pdf.

Welch:2002:CNJ

[74] P. H. Welch, J. R. Aldous, and J. Foster. CSP networking for Java (JCSP.net). Lecture Notes in Computer Science, 2330:695–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300695.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300695.
pdf.

Kurzyniec:2002:MBT

[75] D. Kurzyniec and V. Sunderam. The MICROBE benchmarking toolkit for Java: a component-based approach. Lecture Notes in Computer Science, 2330:709–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2330/23300709.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2330/23300709.
pdf.

Engelmann:2002:DPP

[76] C. Engelmann, S. L. Scott, and G. A. Geist. Distributed peer-to-peer control in harness. Lecture Notes in Computer Science, 2330:720– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300720.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300720.
pdf.

Nemeth:2002:CCD

[77] Z. Németh and V. Sunderam. A comparison of conventional distributed computing environments and computational grids. Lecture Notes in Computer Science, 2330:729–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300729.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300729.
pdf.

Taylor:2002:DGB

[78] J. Taylor, M. Dvorak, and S. A. Mickelson. Developing grid based infrastructure for climate modeling. Lecture Notes in Computer Science, 2330:739– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300739.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300739.
pdf.

Ong:2002:RAM

[79] E. T. Ong, J. W. Larson, and R. L. Jacob. A real application of the model coupling toolkit. Lecture Notes in Computer Science, 2330:748–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300748.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300748.
pdf.

Mickelson:2002:STG

[80] S. A. Mickelson, J. A. Taylor, and M. Dvorak. Simplifying the task of generating climate simulations and visualizations. Lecture Notes in Computer Science, 2330:758–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300758.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300758.
pdf.

Segers:2002:CMF

[81] A. Segers, P. van Velthoven, B. Bregman, and M. Krol. On the computation of mass fluxes for Eulerian transport models from spectral meteorological fields. Lecture Notes in Computer Science, 2330:767–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300767.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300767.
pdf.

Dvorak:2002:DFG

[82] M. Dvorak, J. Taylor, and S. A. Mickelson. Designing a flexible grid enabled scientific modeling interface. Lecture Notes in Computer Science, 2330:777– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300777.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2330/23300777.
                                        pdf.

Muylle:2002:PCD

[83] J. Muylle and B. H. V. Topping. Parallel contact detection strategies for cable and membrane structures. Lecture Notes in Computer Science, 2330:787– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300787.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300787.
                                        pdf.

Jimack:2002:PDD

[84] P. K. Jimack and S. A. Nadeem. A parallel domain decomposition algorithm for the adaptive finite element solution of 3-D convectiondiffusion problems. Lecture Notes in Computer Science, 2330:797–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300797.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300797.
                                        pdf.

McManus:2002:PPM

[85] K. McManus, M. Cross, C., Walshaw, N. Croft, and A. Williams. Parallel performance in multi-physics simulation. Lecture Notes in Computer Science, 2330:806–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2330/23300806.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2330/23300806.
pdf.

Weatherill:2002:PFV

[86] N. Weatherill, K. Sørensen, O. Hassan, and K. Morgan. A parallel finite volume method for aerodynamic flows. Lecture Notes in Computer Science, 2330:816–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300816.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300816.
pdf.

Beazley:2002:ECC

[87] D. M. Beazley. An extensible compiler for creating scriptable scientific software. Lecture Notes in Computer Science, 2330:824–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300824.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300824.
pdf.

Abramson:2002:GTM

[88] D. Abramson, G. Watson, and L. P. Dung. Guard: a tool for migrating scientific applications to the .NET framework. Lecture Notes in Computer Science, 2330:834–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2330/23300834.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300834.
                                        pdf.

Danelutto:2002:LSP

[89] M. Danelutto and P. Teti. Lithium: a structured parallel programming environment in Java. Lecture Notes in Computer Science, 2330:844–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300844.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300844.
                                        pdf.

Reussner:2002:UTT

[90] R. Reussner, I. Poernomo, and H. W. Schmidt. Using the TrustME tool suite for automatic component protocol adaptation. Lecture Notes in Computer Science, 2330:854–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300854.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300854.
                                        pdf.

Wilde:2002:ICA

[91] T. Wilde, J. A. Kohl, and R. E. Flanery, Jr. Integrating CUMULVS into AVS/Express. Lecture Notes in Computer Science, 2330:864–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2330/23300864.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300864.
pdf.

Bubak:2002:MSD

[92] M. Bubak, W. Funika, P. Metel, R. Orłowski, and R. Wismüller. Monitoring system for distributed Java applications. Lecture Notes in Computer Science, 2330:874–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300874.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300874.
pdf.

Balis:2002:CPM

[93] B. Balis, M. Bubak, W. Funika, and R. Wismüller. A concept of portable monitoring of multithreaded programs. Lecture Notes in Computer Science, 2330:884–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300884.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300884.
pdf.

Jancic:2002:DER

[94] J. Jancic, C. Poellabauer, K. Schwan, M. Wolf, and N. Bright. dproc — extensible run-time resource monitoring for cluster applications. Lecture Notes in Computer Science, 2330:894– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300894.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300894.
                                        pdf.

Moore:2002:CCS

[95] S. V. Moore. A comparison of counting and sampling modes of using performance monitoring hardware. Lecture Notes in Computer Science, 2330: 904–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300904.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300904.
                                        pdf.

Kranzlmuller:2002:DLS

[96] D. Kranzlmüller, N. Thoai, and J. Volkert. Debugging large-scale, longrunning parallel programs. Lecture Notes in Computer Science, 2330:913– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300913.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300913.
                                        pdf.

Blanco:2002:PPP

[97] V. Blanco, P. González, J. C. Cabaleiro, D. B. Heras, T. F. Pena, J. J. Pombo, and F. F. Rivera. Performance prediction for parallel iterative solvers. Lecture Notes in Computer Science, 2330:923–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300923.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300923.
pdf.

Tao:2002:IDL

[98] J. Tao, M. Schulz, and W. Karl. Improving data locality using dynamic page migration based on memory access histograms. Lecture Notes in Computer Science, 2330:933–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300933.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300933.
pdf.

Walshaw:2002:MMP

[99] C. Walshaw, M. Cross, and K. Mc-Manus. Multiphase mesh partitioning for parallel computational mechanics codes. Lecture Notes in Computer Science, 2330:943–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300943.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300943.
pdf.

Ierotheou:2002:SMP

[100] C. S. Ierotheou, S. Johnson, P. Leggett, and M. Cross. The shared memory parallelisation of an ocean modelling code using an interactive parallelisation toolkit. Lecture Notes in Computer Science, 2330:953–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300953.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300953.
                                        pdf.

Hofinger:2002:DLE

[101] S. Höfinger. Dynamic load equilibration for cyclic applications in distributed systems. Lecture Notes in Computer Science, 2330:963–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300963.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300963.
                                        pdf.

Marsh:2002:MIT

[102] A. Marsh. 3G medicine — the integration of technologies. Lecture Notes in Computer Science, 2330:972– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23300972.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23300972.
                                        pdf.

Blobel:2002:ASP

[103] B. Blobel. Architecture of secure portable and interoperable electronic health records. Lecture Notes in Computer Science, 2330:982–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300982.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300982.
pdf.

Emmen:2002:DCR

[104] A. Emmen. Designing for change and reusability — using XML, XSL, and MPEG-7 for developing professional health information systems. Lecture Notes in Computer Science, 2330:995– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23300995.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23300995.
pdf.

Saarelainen:2002:PL

[105] M. Saarelainen. Personal LocationMessaging. Lecture Notes in Computer Science, 2330:1003–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23301003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23301003.
pdf.

Marsh:2002:CPR

[106] A. Marsh. The E-CARE project — removing the wires. Lecture Notes in Computer Science, 2330:1012–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23301012.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23301012.
                                        pdf.

Heimbach:2002:AGE

[107] P. Heimbach, C. Hill, and R. Giering. Automatic generation of efficient adjoint code for a parallel Navier–Stokes solver. Lecture Notes in Computer Science, 2330:1019–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23301019.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23301019.
                                        pdf.

Fagan:2002:SPD

[108] M. Fagan and A. Carle. Switchback: Profile-driven recomputation for reverse mode. Lecture Notes in Computer Science, 2330:1029–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23301029.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23301029.
                                        pdf.

Naumann:2002:RMR

[109] U. Naumann. Reducing the memory requirement in reverse mode automatic differentiation by solving TBR flow equations. Lecture Notes in Computer Science, 2330:1039–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23301039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23301039.
pdf.

Lehmann:2002:ITT

[110] U. Lehmann and A. Walther. The implementation and testing of timeminimal and resource-optimal parallel reversal schedules. Lecture Notes in Computer Science, 2330:1049–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23301049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23301049.
pdf.

Robenack:2002:ADN

[111] K. Röbenack. Automatic differentiation for nonlinear controller design. Lecture Notes in Computer Science, 2330:1059–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2330/23301059.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2330/23301059.
pdf.

Bucker:2002:CSI

[112] H. M. Bücker, B. Lang, A. Rasch, and C. H. Bischof. Computation of sensitivity information for aircraft design by automatic differentiation. Lecture Notes in Computer Science, 2330:1069– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23301069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23301069.
                                        pdf.

Tadjouddine:2002:PIV

[113] M. Tadjouddine, S. A. Forth, J. D. Pryce, and J. K. Reid. Performance issues for vertex elimination methods in computing Jacobians using automatic differentiation. Lecture Notes in Computer Science, 2330:1077–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2330/23301077.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2330/23301077.
                                        pdf.

Hovland:2002:MAD

[114] P. Hovland, B. Norris, and B. Smith. Making automatic differentiation truly automatic: Coupling PETSc with ADIC. Lecture Notes in Computer Science, 2330:1087–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2330/23301087.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2330/23301087.pdf.

Petrov:2002:IIC

[115] E. Petrov and F. Benhamou. Improved interval constraint propagation for constraints on partial derivatives. Lecture Notes in Computer Science, 2330:1097– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2330/23301097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2330/23301097.
pdf.

Anonymous:2002:AIa

[116] Anonymous. Author index. Lecture Notes in Computer Science, 2330:1107– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2330/2330auth.pdf.

Overmars:2002:RDM

[117] M. H. Overmars. Recent developments in motion planning. Lecture Notes in Computer Science, 2331:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310003.
pdf.

Dumitrescu:2002:EDM

[118] A. Dumitrescu and S. Guha. Extreme distances in multicolored point sets. Lecture Notes in Computer Science, 2331:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2331/23310014.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2331/23310014.
pdf.

Andersson:2002:BPM

[119] M. Andersson, J. Gudmundsson, C. Levcopoulos, and G. Narasimhan. Balanced partition of minimum spanning trees. Lecture Notes in Computer Science, 2331:26–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2331/23310026.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2331/23310026.pdf.

Hungershofer:2002:QPB

[120] J. Hungershöfer and J.-M. Wierum. On the quality of partitions based on space-filling curves. Lecture Notes in Computer Science, 2331:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310036.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310036.
                                        pdf.

Diaz-Banez:2002:LEA

[121] J. M. Díaz-Báñez, F. Hurtado, H. Meijer, D. Rappaport, and T. Sellares. The largest empty annulus problem. Lecture Notes in Computer Science, 2331: 46–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310046.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310046.
                                        pdf.

Bekker:2002:MGS

[122] H. Bekker and K. De Raedt. Mapping graphs on the sphere to the finite plane. Lecture Notes in Computer Science, 2331:55–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310055.
pdf.

Daescu:2002:IOW

[123] O. Daescu. Improved optimal weighted links algorithms. Lecture Notes in Computer Science, 2331:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310065.
pdf.

Lorenzetto:2002:LTH

[124] G. P. Lorenzetto and A. Datta. A linear time heuristics for trapezoidation of GIS polygons. Lecture Notes in Computer Science, 2331:75– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310075.
pdf.

Huybers:2002:MBS

[125] P. Huybers. The morphology of building structures. Lecture Notes in Computer Science, 2331:85–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310085.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310085.
                                        pdf.

Gervois:2002:VRT

[126] A. Gervois, L. Oger, P. Richard, and J. P. Troadec. Voronoi and radical tessellations of packings of spheres. Lecture Notes in Computer Science, 2331: 95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310095.
                                        pdf.

Gavrilova:2002:CDO

[127] M. L. Gavrilova and J. Rokne. Collision detection optimization in a multi-particle system. Lecture Notes in Computer Science, 2331:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310105.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310105.
                                        pdf.

Gavrilova:2002:OTE

[128] M. Gavrilova, J. Rokne, D. Gavrilov, and O. Vinogradov. Optimization techniques in an event-driven simulation of a shaker ball mill. Lecture Notes in Computer Science, 2331:115– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310115.
pdf.

Kolingerova:2002:MDL

[129] I. Kolingerová. Modified DAG location for Delaunay triangulation. Lecture Notes in Computer Science, 2331: 125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310125.
pdf.

Tse:2002:TMC

[130] R. O. C. Tse and C. Gold. TIN meets CAD — extending the TIN concept in GIS. Lecture Notes in Computer Science, 2331:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310135.
pdf.

Dakowicz:2002:EMS

[131] M. Dakowicz and C. Gold. Extracting meaningful slopes from terrain contours. Lecture Notes in Computer Science, 2331:144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310144.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310144.
                                        pdf.

Giesen:2002:DDI

[132] J. Giesen and M. John. Duality in disk induced flows. Lecture Notes in Computer Science, 2331:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310154.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310154.
                                        pdf.

Rajabi:2002:IDT

[133] M. A. Rajabi and J. A. R. Blais. Improvement of digital terrain model interpolation using SFS techniques with single satellite imagery. Lecture Notes in Computer Science, 2331:164– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310164.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310164.
                                        pdf.

Mower:2002:IAS

[134] J. E. Mower. Implementing an augmented scene delivery system. Lecture Notes in Computer Science, 2331:174– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310174.
pdf.

Wirza:2002:ISC

[135] R. Wirza, M. S. Bloor, and J. Fisher. Inspection strategies for complex curved surfaces using CMM. Lecture Notes in Computer Science, 2331:184– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310184.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310184.
pdf.

Lyakh:2002:FFD

[136] A. Lyakh. The free form deformation of phytoplankton models. Lecture Notes in Computer Science, 2331:194– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310194.
pdf.

Fischer:2002:CBR

[137] B. Fischer and J. Modersitzki. Curvature based registration with applications to MR-mammography. Lecture Notes in Computer Science, 2331:202– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310202.
pdf.

Abubakar:2002:FSN

[138] A. Abubakar and P. M. van den Berg. Full scale nonlinear electromagnetic inversion for biological objects. Lecture Notes in Computer Science, 2331:207– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310207.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310207.
pdf.

Kroc:2002:PEW

[139] J. Kroc. Propagation of excitation waves and their mutual interactions in the surface layer of the ball with fast accessory paths and the pacemaker. Lecture Notes in Computer Science, 2331: 217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310217.
pdf.

Daescu:2002:COT

[140] O. Daescu and A. Bhatia. Computing optimal trajectories for medical treatment planning and optimization. Lecture Notes in Computer Science, 2331: 227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310227.
pdf.

Martyniak:2002:CRU

[141] J. Martyniak, K. Stanisz-Wallis, and L. Walczycka. CAD recognition using three mathematical models. Lecture Notes in Computer Science, 2331: 234–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310234.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310234.
pdf.

Schaap:2002:QVV

[142] J. A. Schaap, P. J. H. de Koning, J. P. Janssen, J. J. M. Westenberg, R. J. van der Geest, and J. H. C. Reiber. 3D quantification visualization of vascular structures in magnetic resonance angiographic images. Lecture Notes in Computer Science, 2331:242– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310242.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310242.
pdf.

Box:2002:QMC

[143] F. M. A. Box, M. C. M. Rutten, M. A. van Buchem, J. Doornbos, R. J. van der Geest, P. J. H. de Koning, J. Schaap, F. N. van de Vosse, and J. H. C. Reiber. Quantitative methods for comparisons between velocity encoded MRmeasurements and finite element modeling in phantom models. Lecture Notes in Computer Science, 2331:255– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310255.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310255.
                                        pdf.

Bellemann:2002:HPD

[144] R. G. Bellemann and R. Shulakov. High performance distributed simulation for interactive simulated vascular reconstruction. Lecture Notes in Computer Science, 2331:265–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310265.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310265.
                                        pdf.

Verdonck:2002:FSI

[145] P. R. Verdonck and J. A. Vierendeels. Fluid-structure interaction modelling of left ventricular filling. Lecture Notes in Computer Science, 2331: 275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2331/23310275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310275.
pdf.

Ablitt:2002:MDR

[146] N. Ablitt, J. Gao, P. Gatehouse, and G.-Z. Yang. Motion decoupling and registration for 3D magnetic resonance myocardial perfusion imaging. Lecture Notes in Computer Science, 2331:285– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310285.
pdf.

Arbenz:2002:CFF

[147] P. Arbenz. A comparison of factorization-free eigensolvers with application to cavity resonators. Lecture Notes in Computer Science, 2331:295– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310295.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310295.
pdf.

Filbet:2002:DAV

[148] F. Filbet, J.-L. Lemaire, and E. Sonnendrücker. Direct axisymmetric Vlasov simulations of space charge dominated beams. Lecture Notes in Computer Science, 2331:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310305.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310305.
                                        pdf.

Vorobiev:2002:FPS

[149] L. G. Vorobiev and R. C. York. Fast Poisson solver for space charge dominated beam simulation based on the template potential technique. Lecture Notes in Computer Science, 2331:315– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310315.
                                        pdf.

Shishlo:2002:PAC

[150] A. Shishlo, J. Holmes, and V. Danilov. Parallel algorithms for collective processes in high intensity rings. Lecture Notes in Computer Science, 2331:325– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310325.
                                        pdf.

Nieter:2002:VTS

[151] C. Nieter and J. R. Cary. VOR-PAL as a tool for the study of laser pulse propagation in LWFA. Lecture Notes in Computer Science, 2331:334– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2331/23310334.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310334.
pdf.

Fonseca:2002:OTD

[152] R. A. Fonseca, L. O. Silva, F. S. Tsung, V. K. Decyk, W. Lu, C. Ren, W. B. Mori, S. Deng, S. Lee, T. Katsouleas, and J. C. Adam. OSIRIS: a three-dimensional, fully relativistic particle in cell code for modeling plasma based accelerators. Lecture Notes in Computer Science, 2331:342– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310342.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310342.
pdf.

Wilson:2002:IVP

[153] B. Wilson, K.-L. Ma, J. Qiang, and R. Ryne. Interactive visualization of particle beams for accelerator design. Lecture Notes in Computer Science, 2331:352–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310352.
pdf.

Adelmann:2002:GLS

[154] A. Adelmann and D. Feichtinger. Generic large scale 3D visualization of accelerators and beam lines. Lecture Notes in Computer Science, 2331:362– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310362.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310362.
                                        pdf.

Biryukov:2002:TPA

[155] V. Biryukov, A. Drees, R. P. Fliller, N. Malitsky, and D. Trbojevic. Tracking particles in accelerator optics with crystal elements. Lecture Notes in Computer Science, 2331:372–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310372.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310372.
                                        pdf.

Meot:2002:PDA

[156] F. Méot. Precision dynamic aperture tracking in rings. Lecture Notes in Computer Science, 2331:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310381.
                                        pdf.

Samulyak:2002:NSH

[157] R. Samulyak. Numerical simulation of hydro- and magnetohydrodynamic processes in the muon collider target. Lecture Notes in Computer Science, 2331:

391–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310391.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310391.
pdf.

Zaplatin:2002:SRA

[158] E. Zaplatin. Superconducting RF accelerating cavity developments. Lecture Notes in Computer Science, 2331: 401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310401.
pdf.

Duperrier:2002:CSC

[159] R. Duperrier, N. Pichoff, and D. Uriot. CEA saclay codes review for high intensities linacs computations. Lecture Notes in Computer Science, 2331:411– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310411.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310411.
pdf.

Baumann:2002:DTV

[160] M. Baumann and U. Helmke. Diagonalization of time varying symmetric matrices. Lecture Notes in Computer Science, 2331:419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310419.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310419.
                                        pdf.

Amodio:2002:CPS

[161] P. Amodio, F. Iavernaro, and D. Trigiante. Conservation properties of symmetric BVMs applied to linear Hamiltonian problems. Lecture Notes in Computer Science, 2331:429–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310429.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310429.
                                        pdf.

Chiarantoni:2002:FPH

[162] E. Chiarantoni, G. Fornarelli, S. Vergura, and T. Politi. A fixed point homotopy method for efficient time-domain simulation of power electronic circuits. Lecture Notes in Computer Science, 2331:439–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310439.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310439.
                                        pdf.

Diele:2002:FRS

[163] F. Diele, T. Politi, and I. Sgura. A Fortran90 routine for the solution of orthogonal differential problems. Lecture Notes in Computer Science, 2331:

449–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310449.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310449.
pdf.

Paternoster:2002:TSR

[164] B. Paternoster. Two step Runge– Kutta–Nyström methods for y′′ = f(x, y) and P -stability. Lecture Notes in Computer Science, 2331:459–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310459.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310459.
pdf.

DelBuono:2002:SRN

[165] N. Del Buono and C. Elia. Some remarks on numerical methods for second order differential equations on the orthogonal matrix group. Lecture Notes in Computer Science, 2331:467– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310467.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310467.
pdf.

Diele:2002:NCB

[166] F. Diele and S. Ragni. Numerical comparison between different liegroup methods for solving linear oscillatory ODEs. Lecture Notes in Computer Science, 2331:476–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310476.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310476.
                                        pdf.

Islas:2002:MSM

[167] A. L. Islas and C. M. Schober. Multisymplectic spectral methods for the Gross–Pitaevskii equation. Lecture Notes in Computer Science, 2331:486– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310486.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310486.
                                        pdf.

Sofroniou:2002:SOM

[168] M. Sofroniou and G. Spaletta. Solving orthogonal matrix differential systems in mathematica. Lecture Notes in Computer Science, 2331:496–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310496.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310496.
                                        pdf.

Sofroniou:2002:SMS

[169] M. Sofroniou and G. Spaletta. Symplectic methods for separable Hamiltonian systems. Lecture Notes in Computer Science, 2331:506–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310506.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310506.
pdf.

Pavani:2002:NTR

[170] R. Pavani. Numerical treatment of the rotation number for the forced pendulum. Lecture Notes in Computer Science, 2331:516–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310516.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310516.
pdf.

DelBuono:2002:SMB

[171] N. Del Buono, C. Elia, and L. Lopez. Symplectic method based on the matrix variational equation for Hamiltonian system. Lecture Notes in Computer Science, 2331:526–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310526.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310526.
pdf.

Neruda:2002:VLA

[172] R. Neruda, A. Štědrý, and J. Drkošová. Variants of learning algorithm based on Kolmogorov theorem. Lecture Notes in Computer Science, 2331:536– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310536.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310536.
                                        pdf.

Dominguez:2002:GNS

[173] J. J. Domínguez, S. Lozano, and M. Calle. Genetic neighborhood search. Lecture Notes in Computer Science, 2331:544–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2331/23310544.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310544.
                                        pdf.

Hakl:2002:ANN

[174] F. Hakl, M. Hlaváček, and R. Kalous. Application of neural networks optimized by genetic algorithms to Higgs boson search. Lecture Notes in Computer Science, 2331:554–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310554.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310554.
                                        pdf.

Nechaev:2002:CSR

[175] Y. Nechaev, A. Degtyarev, and I. Kiryukhin. Complex situation recognition on the basis of neural networks in shipboard intelligence system. Lecture Notes in Computer Science, 2331:

564–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310564.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310564.
pdf.

Haber:2002:DMM

[176] R. E. Haber, R. H. Haber, A. Alique, S. Ros, and J. R. Alique. Dynamic model of the machining process on the basis of neural networks: From simulation to real time application. Lecture Notes in Computer Science, 2331: 574–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310574.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310574.
pdf.

Coufal:2002:ISL

[177] D. Coufal. Incremental structure learning of three-layered Gaussian RBF networks. Lecture Notes in Computer Science, 2331:584–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310584.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310584.
pdf.

Neruda:2002:HLR

[178] R. Neruda and P. Kudová. Hybrid learning of RBF networks. Lecture Notes in Computer Science, 2331:594– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310594.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310594.
                                        pdf.

Pytelkova:2002:SAD

[179] R. Pytelková and P. Hušek. Stability analysis of discrete-time Takagi-Sugeno fuzzy systems. Lecture Notes in Computer Science, 2331:604–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310604.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310604.
                                        pdf.

Lee:2002:FCS

[180] W.-Y. Lee, I.-S. Lim, and U.-Y. Huh. Fuzzy control system using nonlinear friction observer for the mobile robot. Lecture Notes in Computer Science, 2331:613–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310613.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310613.
                                        pdf.

Pflaum:2002:EIO

[181] C. Pflaum and D. Seider. Efficient implementation of operators on semiunstructured grids. Lecture Notes in Computer Science, 2331:622–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310622.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310622.
pdf.

Falgout:2002:HLH

[182] R. D. Falgout and U. Meier Yang. hypre: a library of high performance preconditioners. Lecture Notes in Computer Science, 2331:632–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310632.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310632.
pdf.

Kowarschik:2002:DLO

[183] M. Kowarschik, U. Rüde, and C. Weiß. Data layout optimizations for variable coefficient multigrid. Lecture Notes in Computer Science, 2331:642– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310642.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310642.
pdf.

Hulsemann:2002:GFE

[184] F. Hülsemann, P. Kipfer, U. Rüde, and G. Greiner. gridlib: Flexible and efficient grid management for simulation and visualization. Lecture Notes in Computer Science, 2331:652– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310652.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310652.
                                        pdf.

Bader:2002:STS

[185] M. Bader, H.-J. Bungartz, A. Frank, and R. Mundani. Space tree structures for PDE software. Lecture Notes in Computer Science, 2331:662– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310662.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310662.
                                        pdf.

Mitchell:2002:DPA

[186] W. F. Mitchell. The design of a parallel adaptive multi-level code in Fortran 90. Lecture Notes in Computer Science, 2331:672–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310672.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310672.
                                        pdf.

Norden:2002:OVM

[187] M. Nordén, S. Holmgren, and M. Thuné. OpenMP versus MPI for PDE solvers based on regular sparse numerical operators. Lecture Notes in Computer Science, 2331:681– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310681.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310681.
pdf.

Hinsen:2002:HLS

[188] K. Hinsen. High-level scientific programming with python. Lecture Notes in Computer Science, 2331:691–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310691.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310691.
pdf.

Lindemann:2002:UCM

[189] J. Lindemann, O. Dahlblom, and G. Sandberg. Using CORBA middleware in finite element software. Lecture Notes in Computer Science, 2331: 701–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310701.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310701.
pdf.

Aahlander:2002:SSF

[190] K. Åhlander and K. Otto. On software support for finite difference schemes based on index notation. Lecture Notes in Computer Science, 2331:711– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310711.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310711.
                                        pdf.

Parker:2002:CBA

[191] S. G. Parker. A component-based architecture for parallel multi-physics PDE simulation. Lecture Notes in Computer Science, 2331:719–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310719.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310719.
                                        pdf.

vonOehsen:2002:UDP

[192] J. Barr von Oehsen, C. L. Cox, E. C. Cyr, and B. A. Malloy. Using design patterns and XML to construct an extensible finite element system. Lecture Notes in Computer Science, 2331:735– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310735.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310735.
                                        pdf.

Berti:2002:GGA

[193] G. Berti. GrAL — the grid algorithms library. Lecture Notes in Computer Science, 2331:745–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2331/23310745.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310745.
pdf.

Chow:2002:SST

[194] P. Chow and C. Addison. A software strategy towards putting domain decomposition at the centre of a meshbased simulation process. Lecture Notes in Computer Science, 2331:755– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310755.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310755.
pdf.

Langtangen:2002:SFM

[195] H. P. Langtangen and K.-A. Mardal. A software framework for mixed finite element programming. Lecture Notes in Computer Science, 2331:764– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310764.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310764.
pdf.

Douglas:2002:FAR

[196] C. C. Douglas, J. Hu, J. Ray, D. Thorne, and R. Tuminaro. Fast, adaptively refined computational elements in 3D. Lecture Notes in Computer Science, 2331:774–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310774.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310774.
                                        pdf.

Axelsson:2002:PML

[197] O. Axelsson and M. Neytcheva. Preconditioning methods for linear systems with saddle point matrices. Lecture Notes in Computer Science, 2331: 784–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310784.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310784.
                                        pdf.

Maryska:2002:MHF

[198] J. Maryška, O. Severýn, and M. Vohralík. Mixed-hybrid FEM discrete fracture network model of the fracture flow. Lecture Notes in Computer Science, 2331: 794–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310794.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310794.
                                        pdf.

Pavlus:2002:PRD

[199] M. Pavluš and E. Hayryan. Parallel realization of difference schemes of filtration problem in a multilayer system. Lecture Notes in Computer Science, 2331:804–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310804.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310804.
pdf.

Bourchtein:2002:SPG

[200] A. Bourchtein and L. Bourchtein. Stokes problem for the generalized Navier–Stokes equations. Lecture Notes in Computer Science, 2331:813–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310813.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310813.
pdf.

Danek:2002:DDA

[201] J. Daněk. Domain decomposition algorithm for solving contact of elastic bodies. Lecture Notes in Computer Science, 2331:820–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310820.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310820.
pdf.

Blaheta:2002:PHP

[202] R. Blaheta, O. Jakl, and J. Starý. Parallel high-performance computing in geomechanics with inner/outer iterative procedures. Lecture Notes in Computer Science, 2331:830–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310830.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310830.
                                        pdf.

Hlavacek:2002:RSU

[203] I. Hlaváček and J. Nedoma. Reliable solution of a unilateral frictionless contact problem in quasicoupled thermo-elasticity with uncertain input data. Lecture Notes in Computer Science, 2331:840–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310840.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310840.
                                        pdf.

Ruede:2002:CEP

[204] U. Ruede. Computational engineering programs at the University of Erlangen-Nuremberg. Lecture Notes in Computer Science, 2331:852–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310852.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310852.
                                        pdf.

Wiechert:2002:TMM

[205] W. Wiechert. Teaching mathematical modeling: Art or science? Lecture Notes in Computer Science, 2331:858– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310858.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310858.
pdf.

Jeltsch:2002:CPE

[206] R. Jeltsch and K. Nipp. CSE program at ETH zurich: Are we doing the right thing? Lecture Notes in Computer Science, 2331:863–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310863.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310863.
pdf.

Anido:2002:OES

[207] L. Anido, J. Santos, M. Fernández, M. Caeiro, J. Rodríguez, and M. Llamas. An online environment supporting high quality education in computational science. Lecture Notes in Computer Science, 2331:872–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310872.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310872.
pdf.

Lintner:2002:CES

[208] M. D. Lintner. Computing, ethics and social responsibility: Developing ethically responsible computer users for the 21^{st} Century. Lecture Notes in Computer Science, 2331:882–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310882.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310882.
                                        pdf.

Pan:2002:TPP

[209] Y. Pan. Teaching parallel programming using both high-level and low-level languages. Lecture Notes in Computer Science, 2331:888–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310888.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310888.
                                        pdf.

Mori:2002:CSH

[210] P. Mori and L. Ricci. Computational science in high school curricula: The ORESPICS approach. Lecture Notes in Computer Science, 2331: 898–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310898.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310898.
                                        pdf.

Piermarini:2002:PAI

[211] V. Piermarini, L. Pacifici, S. Crocchianti, and A. Laganà. Parallel approaches to the integration of the differential equations for reactive scattering. Lecture Notes in Computer Science, 2331:908–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310908.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310908.
pdf.

Bellucci:2002:FGP

[212] D. Bellucci, S. Tasso, and A. Laganà. Fine grain parallelism for discrete variable approaches to wavepacket calculations. Lecture Notes in Computer Science, 2331:918–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2331/23310918.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2331/23310918.pdf.

Riganelli:2002:MDS

[213] A. Riganelli, M. Memelli, and A. Laganà. A molecular dynamics study of the benzene… ar_{2} complexes. Lecture Notes in Computer Science, 2331:926– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310926.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310926.
pdf.

Marcelli:2002:BTE

[214] G. Marcelli, B. D. Todd, and R. J. Sadus. Beyond traditional effective intermolecular potentials and pairwise interactions in molecular simulation. Lecture Notes in Computer Science, 2331:932–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310932.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310932.
                                        pdf.

Sigalas:2002:DFS

[215] M. P. Sigalas and V. I. Teberekidis. Density functional studies of halonium ions of ethylene and cyclopentene. Lecture Notes in Computer Science, 2331: 942–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310942.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310942.
                                        pdf.

Sax:2002:MPC

[216] A. F. Sax and T. Krüger. Methodological problems in the calculations on amorphous hydrogenated silicon, a-Si:H. Lecture Notes in Computer Science, 2331:950–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310950.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310950.
                                        pdf.

Gervasi:2002:TGB

[217] O. Gervasi, A. Laganà, and M. Lobbiani. Towards a GRID based portal for an a priori molecular simulation of chemical reactivity. Lecture Notes in Computer Science, 2331:956– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310956.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310956.
pdf.

Yan:2002:ERP

[218] L. Yan, J. b. Wang, Y. a. Ma, and J. Dou. The enterprise resource planning (ERP) system and spatial information integration in tourism industry – mount emei for example. Lecture Notes in Computer Science, 2331: 966–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310966.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310966.
pdf.

Sun:2002:VLD

[219] M. Sun, Y. Xue, A.-N. Ma, and S.-J. Mao. 3D visualization of large digital elevation model (DEM) data set. Lecture Notes in Computer Science, 2331: 975–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23310975.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23310975.
pdf.

Sun:2002:DVR

[220] M. Sun, Y. Xue, A.-N. Ma, and S.-J. Mao. Dynamic vector and raster integrated data model based on code-points. Lecture Notes in Computer Science, 2331:984–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310984.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310984.
                                        pdf.

Zhang:2002:KON

[221] Yong Zhang, Lizhu Zhou, Jun Chen, and RenLiang Zhao. K-order neighbor: The efficient implementation strategy for restricting cascaded update in realm. Lecture Notes in Computer Science, 2331:994–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23310994.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23310994.
                                        pdf.

Zhao:2002:HRM

[222] R. Zhao, Z. Li, J. Chen, C. M. Gold, and Y. Zhang. A hierarchical raster method for computing Voronoi diagrams based on quadtrees. Lecture Notes in Computer Science, 2331:1004– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23311004.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2331/23311004.
pdf.

Xue:2002:DTD

[223] Y. Xue, M. Sun, Y. Zhang, and R. Zhao. The dissection of threedimensional geographic information systems. Lecture Notes in Computer Science, 2331:1014–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2331/23311014.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2331/23311014.pdf.

Hernandez:2002:GCT

[224] J. C. Hernández, J. M. Sierra, P. Isasi, and A. Ribagorda. Genetic cryptoanalysis of two rounds TEA. Lecture Notes in Computer Science, 2331:1024– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23311024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23311024.
pdf.

Vassell:2002:GCI

[225] C. Vassell. Genetic commerce — intelligent share trading. Lecture Notes in Computer Science, 2331:1032–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23311032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23311032.
pdf.

Chung:2002:EMP

[226] J. Y. Chung and S. Kim. Efficient memory page replacement on Web server clusters. Lecture Notes in Computer Science, 2331:1042–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23311042.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23311042.
                                        pdf.

Woo:2002:IWL

[227] B. K. Woo, D. S. Kim, S. S. Hong, K. H. Kim, and T. M. Chung. Interval weighted load balancing method for multiple application gateway firewalls. Lecture Notes in Computer Science, 2331:1051–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23311051.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23311051.
                                        pdf.

Mun:2002:MPE

[228] Y. Mun and H. Choo. Modeling and performance evaluation of multistage interconnection networks with nonuniform traffic pattern. Lecture Notes in Computer Science, 2331:1061– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23311061.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2331/23311061.
pdf.

Huh:2002:RTP

[229] E.-N. Huh, L. R. Welch, and Y. Mun. Real-time performance estimation for dynamic, distributed real-time systems. Lecture Notes in Computer Science, 2331:1071–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23311071.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23311071.
pdf.

Hwang:2002:LBA

[230] J. Hwang, W. J. Lee, B. G. Lee, and Y. S. Kim. A load balancing algorithm using the circulation of A single message token. Lecture Notes in Computer Science, 2331:1080– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23311080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23311080.
pdf.

Lee:2002:CFS

[231] D. Lee and H. Choi. A collaborative filtering system of information on the internet. Lecture Notes in Computer Science, 2331:1090–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23311090.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23311090.
                                        pdf.

Choi:2002:HSC

[232] Y. Choi, S. J. Kim, and S. Lee. Hierarchical shot clustering for video summarization. Lecture Notes in Computer Science, 2331:1100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23311100.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23311100.
                                        pdf.

Shakhov:2002:DUD

[233] V. V. Shakhov, H. Choo, and H. Y. Youn. On detecting unsteady demand in mobile networking environment. Lecture Notes in Computer Science, 2331:1108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23311108.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23311108.
                                        pdf.

Lee:2002:PML

[234] D. C. Lee, S.-K. Han, and Y. S. Mun. Performance modeling of location management using multicasting HLR with forward pointer in mobile networks. Lecture Notes in Computer Science, 2331:1118–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2331/23311118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23311118.
pdf.

Cho:2002:UPP

[235] G. Cho. Using predictive prefetching to improve location awareness of mobile information service. Lecture Notes in Computer Science, 2331:1128– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23311128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23311128.
pdf.

Morozov:2002:DSP

[236] I. V. Morozov, G. E. Norman, and V. V. Stegailov. Dynamic and stochastic properties of molecular systems: From simple liquids to enzymes. Lecture Notes in Computer Science, 2331:1137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23311137.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23311137.
pdf.

Stegailov:2002:DCD

[237] V. V. Stegailov. Determinism and chaos in decay of metastable states. Lecture Notes in Computer Science, 2331:1147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2331/23311147.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23311147.
                                        pdf.

Butikov:2002:RCM

[238] E. I. Butikov. Regular and chaotic motions of the parametrically forced pendulum: Theory and simulations. Lecture Notes in Computer Science, 2331:1154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2331/23311154.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23311154.
                                        pdf.

Posch:2002:LIC

[239] H. A. Posch and C. Forster. Lyapunov instability and collective tangent space dynamics of fluids. Lecture Notes in Computer Science, 2331:1170– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23311170.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23311170.
                                        pdf.

Bogdanov:2002:DCT

[240] A. V. Bogdanov, A. S. Gevorkyan, E. N. Stankova, and M. I. Pavlova. Deterministic computation towards indeterminism. Lecture Notes in Computer Science, 2331:1176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2331/23311176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23311176.
pdf.

Valuev:2002:SPW

[241] I. A. Valuev and B. Esser. Splitting phenomena in wave packet propagation. Lecture Notes in Computer Science, 2331:1184–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2331/23311184.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2331/23311184.
pdf.

Korotenko:2002:ASP

[242] K. Korotenko. An automated system for prediction of icing on the road. Lecture Notes in Computer Science, 2331:1193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2331/23311193.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2331/23311193.
pdf.

Dmitrieva:2002:NNP

[243] L. Dmitrieva, Y. Kuperin, and I. Soroka. Neural network prediction of short-term dynamics of futures on deutsche mark, libor, and S&P500. Lecture Notes in Computer Science, 2331:1201–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2331/23311201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23311201.
                                        pdf.

Ebeling:2002:EPN

[244] W. Ebeling. Entropies and predictability of nonlinear processes and time series. Lecture Notes in Computer Science, 2331:1209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2331/23311209.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2331/23311209.
                                        pdf.

Anonymous:2002:AIb

[245] Anonymous. Author index. Lecture Notes in Computer Science, 2331:1219– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2331/2331auth.pdf.

Gennaro:2002:CPG

[246] Rosario Gennaro and Daniele Micciancio. Cryptanalysis of a pseudorandom generator based on braid groups. Lecture Notes in Computer Science, 2332: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2332/23320001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320001.
                                        pdf.

Lee:2002:PWC

[247] Sang Jin Lee and Eonkyung Lee. Potential weaknesses of the commutator key agreement protocol based on braid groups. Lecture Notes in Computer Science, 2332:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320014.
pdf.

Galbraith:2002:EGW

[248] Steven D. Galbraith, Florian Hess, and Nigel P. Smart. Extending the GHS Weil descent attack. Lecture Notes in Computer Science, 2332:29– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320029.
pdf.

Cramer:2002:UHP

[249] Ronald Cramer and Victor Shoup. Universal hash proofs and a paradigm for adaptive chosen ciphertext secure public-key encryption. Lecture Notes in Computer Science, 2332:45– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320045.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320045.
pdf.

Dodis:2002:KIP

[250] Yevgeniy Dodis, Jonathan Katz, Shouhuai Xu, and Moti Yung. Keyinsulated public key cryptosystems. Lecture Notes in Computer Science, 2332:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2332/23320065.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320065.
                                        pdf.

An:2002:SJS

[251] Jee Hea An, Yevgeniy Dodis, and Tal Rabin. On the security of joint signature and encryption. Lecture Notes in Computer Science, 2332:83– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2332/23320083.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320083.
                                        pdf.

Daemen:2002:AWT

[252] Joan Daemen and Vincent Rijmen. AES and the wide trail design strategy. Lecture Notes in Computer Science, 2332:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2332/23320108.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320108.
                                        pdf.

Maurer:2002:IRS

[253] Ueli Maurer. Indistinguishability of random systems. Lecture Notes in Computer Science, 2332:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320110.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320110.
pdf.

Russell:2002:HFU

[254] Alexander Russell and Hong Wang. How to fool an unbounded adversary with a short key. Lecture Notes in Computer Science, 2332:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320133.
pdf.

Woodruff:2002:CUC

[255] David P. Woodruff and Marten van Dijk. Cryptography in an unbounded computational model. Lecture Notes in Computer Science, 2332:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320149.
pdf.

Nakajima:2002:PAP

[256] Junko Nakajima and Mitsuru Matsui. Performance analysis and parallel implementation of dedicated hash functions. Lecture Notes in Computer Science, 2332:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320165.
pdf.

Clark:2002:FIT

[257] John A. Clark and Jeremy L. Jacob. Fault injection and a timing channel on an analysis technique. Lecture Notes in Computer Science, 2332:181– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320181.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320181.
pdf.

Park:2002:SPM

[258] Young-Ho Park, Sangtae Jeong, and Jongin Lim. Speeding up point multiplication on hyperelliptic curves with efficiently-computable endomorphisms. Lecture Notes in Computer Science, 2332:197–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320197.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2332/23320197.
pdf.

Chose:2002:FCA

[259] Philippe Chose, Antoine Joux, and Michel Mitton. Fast correlation attacks: An algorithmic point of view. Lecture Notes in Computer Science, 2332:209–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2332/23320209.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2332/23320209.
pdf.

Krause:2002:BBC

[260] Matthias Krause. BDD-based cryptanalysis of keystream generators. Lecture Notes in Computer Science, 2332: 222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320222.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320222.
pdf.

Golic:2002:LCB

[261] Jovan Dj. Golić, Vittorio Bagini, and Guglielmo Morgari. Linear cryptanalysis of Bluetooth stream cipher. Lecture Notes in Computer Science, 2332: 238–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320238.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320238.
pdf.

Damgaard:2002:GLB

[262] Ivan Damgård and Maciej Koprowski. Generic lower bounds for root extraction and signature schemes in general groups. Lecture Notes in Computer Science, 2332:256–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2332/23320256.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320256.
                                        pdf.

Coron:2002:OSP

[263] Jean-Sébastien Coron. Optimal security proofs for PSS and other signature schemes. Lecture Notes in Computer Science, 2332:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2332/23320272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320272.
                                        pdf.

Gilbert:2002:CS

[264] Henri Gilbert and Marine Minier. Cryptanalysis of SFLASH. Lecture Notes in Computer Science, 2332:288– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2332/23320288.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320288.
                                        pdf.

Gentry:2002:CRN

[265] Craig Gentry and Mike Szydlo. Cryptanalysis of the revised NTRU signature scheme. Lecture Notes in Computer Science, 2332:299–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320299.
pdf.

Bresson:2002:DGD

[266] Emmanuel Bresson, Olivier Chevassut, and David Pointcheval. Dynamic group Diffie–Hellman key exchange under standard assumptions. Lecture Notes in Computer Science, 2332:321– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320321.
pdf.

Canetti:2002:UCN

[267] Ran Canetti and Hugo Krawczyk. Universally composable notions of key exchange and secure channels. Lecture Notes in Computer Science, 2332:337– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320337.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320337.
pdf.

Beaver:2002:DQK

[268] Donald Beaver. On deniability in quantum key exchange. Lecture Notes in Computer Science, 2332:352–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2332/23320352.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320352.
                                        pdf.

Desai:2002:POT

[269] Anand Desai, Alejandro Hevia, and Yiqun Lisa Yin. A practiceoriented treatment of pseudorandom number generators. Lecture Notes in Computer Science, 2332:368–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2332/23320368.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320368.
                                        pdf.

Black:2002:BCM

[270] John Black and Phillip Rogaway. A block-cipher mode of operation for parallelizable message authentication. Lecture Notes in Computer Science, 2332: 384–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2332/23320384.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320384.
                                        pdf.

Kent:2002:RPW

[271] Stephen Kent. Rethinking PKI: What’s trust got to do with it? Lecture Notes in Computer Science, 2332:398– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320398.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320398.
pdf.

Malkin:2002:EGF

[272] Tal Malkin, Daniele Micciancio, and Sara Miner. Efficient generic forwardsecure signatures with an unbounded number of time periods. Lecture Notes in Computer Science, 2332:400– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320400.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320400.
pdf.

Abdalla:2002:ISF

[273] Michel Abdalla, Jee Hea An, Mihir Bellare, and Chanathip Namprempre. From identification to signatures via the Fiat–Shamir transform: Minimizing assumptions for security and forward-security. Lecture Notes in Computer Science, 2332:418– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320418.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2332/23320418.
                                        pdf.

Shikata:2002:SNU

[274] Junji Shikata, Goichiro Hanaoka, Yuliang Zheng, and Hideki Imai. Security notions for unconditionally secure signature schemes. Lecture Notes in Computer Science, 2332:434–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2332/23320434.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320434.
                                        pdf.

Kiayias:2002:TTC

[275] Aggelos Kiayias and Moti Yung. Traitor tracing with constant transmission rate. Lecture Notes in Computer Science, 2332:450–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2332/23320450.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2332/23320450.pdf.

Horwitz:2002:THI

[276] Jeremy Horwitz and Ben Lynn. Toward hierarchical identity-based encryption. Lecture Notes in Computer Science, 2332:466–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2332/23320466.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320466.
                                        pdf.

Fitzi:2002:UBA

[277] Matthias Fitzi, Nicolas Gisin, Ueli Maurer, and Oliver von Rotz. Unconditional Byzantine agreement and multiparty computation secure against dishonest minorities from scratch. Lecture Notes in Computer Science, 2332: 482–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320482.
pdf.

Desmedt:2002:PSM

[278] Yvo Desmedt and Yongge Wang. Perfectly secure message transmission revisited. Lecture Notes in Computer Science, 2332:502–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2332/23320502.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2332/23320502.pdf.

Canteaut:2002:DCH

[279] Anne Canteaut and Marion Videau. Degree of composition of highly nonlinear functions and applications to higher order differential cryptanalysis. Lecture Notes in Computer Science, 2332: 518–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2332/23320518.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2332/23320518.
pdf.

Vaudenay:2002:SFI

[280] Serge Vaudenay. Security flaws induced by CBC padding — applications to SSL, IPSEC, WTLS, …. Lecture Notes in Computer Science, 2332:534– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2332/23320534.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2332/23320534.
                                        pdf.

Anonymous:2002:AIc

[281] Anonymous. Author index. Lecture Notes in Computer Science, 2332:547– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2332/2332auth.pdf.

Meyer:2002:I

[282] John-Jules Ch. Meyer and Milind Tambe. Introduction. Lecture Notes in Computer Science, 2333:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2333/23330001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330001.
                                        pdf.

Giunchiglia:2002:KLS

[283] Fausto Giunchiglia, Anna Perini, and Fabrizio Sannicolò. Knowledge level software engineering. Lecture Notes in Computer Science, 2333:6– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330006.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330006.
pdf.

Gmytrasiewicz:2002:EPA

[284] Piotr J. Gmytrasiewicz and Christine L. Lisetti. Emotions and personality in agent design and modeling. Lecture Notes in Computer Science, 2333: 21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330021.
pdf.

Kinny:2002:BCA

[285] David Kinny. The Ψ calculus: An algebraic agent language. Lecture Notes in Computer Science, 2333:32– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330032.
pdf.

Wagner:2002:ERT

[286] Thomas Wagner and Victor Lesser. Evolving real-time local agent control for large-scale multi-agent systems. Lecture Notes in Computer Science, 2333:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2333/23330051.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330051.
                                        pdf.

Lesperance:2002:EFP

[287] Yves Lespérance. On the epistemic feasibility of plans in multiagent systems specifications. Lecture Notes in Computer Science, 2333:69– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2333/23330069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330069.
                                        pdf.

Lomuscio:2002:MAS

[288] Alessio Lomuscio and Marek Sergot. On multi-agent systems specification via deontic logic. Lecture Notes in Computer Science, 2333:86– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2333/23330086.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330086.
                                        pdf.

Ryan:2002:ARR

[289] Mark Ryan and Pierre-Yves Schobbens. Agents and roles: Refinement in alternating-time temporal logic. Lecture Notes in Computer Science, 2333: 100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2333/23330100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330100.
pdf.

Wooldridge:2002:CCA

[290] Michael Wooldridge and Paul E. Dunne. The computational complexity of agent verification. Lecture Notes in Computer Science, 2333:115– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330115.
pdf.

Kolp:2002:GBO

[291] Manuel Kolp, Paolo Giorgini, and John Mylopoulos. A goal-based organizational perspective on multiagent architectures. Lecture Notes in Computer Science, 2333:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330128.
pdf.

Leite:2002:DLP

[292] João Alexandre Leite, José Júlio Alferes, and Luís Moniz Pereira. MIN ERVA — A dynamic logic programming agent architecture. Lecture Notes in Computer Science, 2333:141– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2333/23330141.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330141.
                                        pdf.

Machado:2002:RAA

[293] Rodrigo Machado and Rafael H. Bordini. Running AgentSpeak(L) agents on SIM AGENT. Lecture Notes in Computer Science, 2333:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2333/23330158.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330158.
                                        pdf.

Aiello:2002:OO

[294] Marco Aiello, Paolo Busetta, Antonia Donà, and Luciano Serafini. Ontological overhearing. Lecture Notes in Computer Science, 2333:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2333/23330175.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330175.
                                        pdf.

Amgoud:2002:ADC

[295] Leila Amgoud and Simon Parsons. Agent dialogues with conflicting preferences. Lecture Notes in Computer Science, 2333:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2333/23330190.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330190.
pdf.

Bonzon:2002:AMC

[296] Pierre Bonzon. An abstract machine for classes of communicating agents based on deduction. Lecture Notes in Computer Science, 2333:206– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330206.
pdf.

Huber:2002:FSP

[297] Marcus J. Huber, Sanjeev Kumar, Philip R. Cohen, and David R. McGee. A formal semantics for ProxyCommunicative acts. Lecture Notes in Computer Science, 2333:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330221.
pdf.

Yolum:2002:CM

[298] Pinar Yolum and Munindar P. Singh. Commitment machines. Lecture Notes in Computer Science, 2333:235–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330235.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330235.
                                        pdf.

Hunsberger:2002:GBG

[299] Luke Hunsberger. Generating bids for group-related actions in the context of prior commitments. Lecture Notes in Computer Science, 2333:248– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2333/23330248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330248.
                                        pdf.

Modi:2002:DDR

[300] Pragnesh Jay Modi, Hyuckchul Jung, Milind Tambe, Wei-Min Shen, and Shriniwas Kulkarni. Dynamic distributed resource allocation: a distributed constraint satisfaction approach. Lecture Notes in Computer Science, 2333:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2333/23330264.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2333/23330264.pdf.

Nuchia:2002:IOA

[301] Stephen W. Nuchia and Sandip Sen. Improving optimality of n agent envy-free divisions. Lecture Notes in Computer Science, 2333:277–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2333/23330277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330277.
pdf.

Liu:2002:TISa

[302] Wei Liu and Mary-Anne Williams. Trustworthiness of information sources and information pedigrees. Lecture Notes in Computer Science, 2333:290– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330290.
pdf.

Pynadath:2002:RAF

[303] David V. Pynadath and Milind Tambe. Revisiting Asimov’s First Law: a response to the call to arms. Lecture Notes in Computer Science, 2333:307– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330307.
pdf.

Dignum:2002:FTN

[304] Frank Dignum. Formal theories of negotiation. Lecture Notes in Computer Science, 2333:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330321.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2333/23330321.
                                        pdf.

Belmonte:2002:SFP

[305] Maria Victoria Belmonte, Ricardo Conejo, José Luis Pérez de-la Cruz, and Francisco Triguero. A stable and feasible payoff division for coalition formation in a class of task oriented domains. Lecture Notes in Computer Science, 2333:324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2333/23330324.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330324.
                                        pdf.

Brandt:2002:AAV

[306] Felix Brandt and Gerhard Weiß. Antisocial agents and Vickrey auctions. Lecture Notes in Computer Science, 2333:335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2333/23330335.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330335.
                                        pdf.

Esteva:2002:FLI

[307] Marc Esteva, Julian Padget, and Carles Sierra. Formalizing a language for institutions and norms. Lecture Notes in Computer Science, 2333:348– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2333/23330348.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2333/23330348.
pdf.

Faratin:2002:SNA

[308] Peyman Faratin, Mark Klein, Hiroki Sayama, and Yaneer Bar-Yam. Simple negotiating agents in complex games. Lecture Notes in Computer Science, 2333:367–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330367.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330367.
pdf.

Fatima:2002:ONS

[309] S. Shaheen Fatima, Michael Wooldridge, and Nicholas R. Jennings. Optimal negotiation strategies for agents with incomplete information. Lecture Notes in Computer Science, 2333:377– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330377.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330377.
pdf.

Littman:2002:INR

[310] Michael L. Littman and Peter Stone. Implicit negotiation in repeated games. Lecture Notes in Computer Science, 2333:393–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2333/23330393.htm;   http:

                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330393.
                                        pdf.

Sadri:2002:DNA

[311] Fariba Sadri, Francesca Toni, and Paolo Torroni. Dialogues for negotiation: Agent varieties and dialogue sequences. Lecture Notes in Computer Science, 2333:405–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2333/23330405.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330405.
                                        pdf.

Finin:2002:AHH

[312] Tim Finin. Agents for hand-held, mobile, or embedded devices. Lecture Notes in Computer Science, 2333:422– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2333/23330422.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2333/23330422.
                                        pdf.

Albuquerque:2002:KHD

[313] Ryan L. Albuquerque, Jomi F. Hübner, Gustavo E. de Paula, Jaime S. Sichman, and Geber L. Ramalho. KSACI: a handheld device infrastructure for agents communication. Lecture Notes in Computer Science, 2333:423–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2333/23330423.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330423.
pdf.

Bergenti:2002:LFP

[314] Federico Bergenti and Agostino Poggi. LEAP: a FIPA platform for handheld and mobile devices. Lecture Notes in Computer Science, 2333:436– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330436.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330436.
pdf.

Laukkanen:2002:FAP

[315] Mikko Laukkanen, Sasu Tarkoma, and Jani Leinonen. FIPA-OS agent platform for small-footprint devices. Lecture Notes in Computer Science, 2333: 447–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2333/23330447.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2333/23330447.
pdf.

Anonymous:2002:AId

[316] Anonymous. Author index. Lecture Notes in Computer Science, 2333:461– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2333/2333auth.pdf.

Venkitaraman:2002:CSU

[317] Narayanan Venkitaraman, Jayanth P. Mysore, and Mike Needham. A core-stateless utility based rate allocation framework. Lecture Notes in Computer Science, 2334:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2334/23340001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2334/23340001.
pdf.

Westberg:2002:RMD

[318] Lars Westberg, András Császár, Georgios Karagiannis, Ádám Marquetant, David Partain, Octavian Pop, Vlora Rexhepi, Róbert Szabó, and Attila Takács. Resource management in Diffserv (RMD): a functionality and performance behavior overview. Lecture Notes in Computer Science, 2334:17– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2334/23340017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2334/23340017.
pdf.

Menth:2002:PEE

[319] Michael Menth and Rüdiger Martin. Performance evaluation of the extensions for control message retransmissions in RSVP. Lecture Notes in Computer Science, 2334:35– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2334/23340035.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2334/23340035.
pdf.

Voigt:2002:HMB

[320] Thiemo Voigt and Per Gunningberg. Handling multiple bottlenecks in Web servers using adaptive inbound controls. Lecture Notes in Computer Science, 2334:50–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2334/23340050.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2334/23340050.
pdf.

Feng:2002:DRS

[321] Wu chun Feng, Mike Fisk, Mark Gardner, and Eric Weigle. Dynamic right-sizing: An automated, lightweight, and scalable technique for enhancing grid performance. Lecture Notes in Computer Science, 2334:69– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2334/23340069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2334/23340069.
pdf.

Cohen:2002:LCA

[322] Reuven Cohen and Itai Dabran. The “Last-Copy” approach for distributed cache pruning in a cluster of HTTP proxies. Lecture Notes in Computer Science, 2334:84–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2334/23340084.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2334/23340084.
                                        pdf.

Garetto:2002:MSL

[323] M. Garetto, R. Lo Cigno, M. Meo, E. Alessio, and M. Ajmone Marsan. Modeling short-lived TCP connections with open multiclass queuing networks. Lecture Notes in Computer Science, 2334:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2334/23340100.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2334/23340100.
                                        pdf.

Abrahamsson:2002:THS

[324] Henrik Abrahamsson, Olof Hagsand, and Ian Marsh. TCP over high speed variable capacity links: a simulation study for bandwidth allocation. Lecture Notes in Computer Science, 2334: 117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2334/23340117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2334/23340117.
                                        pdf.

Grieco:2002:TWE

[325] Luigi Alfredo Grieco and Saverio Mascolo. TCP westwood and easy RED to improve fairness in highspeed networks. Lecture Notes in Computer Science, 2334:130–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2334/23340130.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2334/23340130.
pdf.

Ossipov:2002:SGS

[326] Evgueni Ossipov and Gunnar Karlsson. A simplified guaranteed service for the internet. Lecture Notes in Computer Science, 2334:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2334/23340147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2334/23340147.
pdf.

Pelsser:2002:ICS

[327] Cristel Pelsser and Stefaan De Cnodder. Improvements to core stateless fair queueing. Lecture Notes in Computer Science, 2334:164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2334/23340164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2334/23340164.
pdf.

Wang:2002:FPC

[328] Pi-Chung Wang, Chia-Tai Chan, Wei-Chun Tseng, and Yaw-Chung Chen. A fast packet classification by using enhanced tuple pruning. Lecture Notes in Computer Science, 2334:180– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2334/23340180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2334/23340180.
                                        pdf.

Wang:2002:TEA

[329] Jianping Wang, Stephen Patek, Haiyong Wang, and Jörg Liebeherr. Traffic engineering with AIMD in MPLS networks. Lecture Notes in Computer Science, 2334:192–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2334/23340192.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2334/23340192.
                                        pdf.

Blondia:2002:PAI

[330] Chris Blondia, Olga Casals, Peter De Cleyn, and Gert Willems. Performance analysis of IP micro-mobility handoff protocols. Lecture Notes in Computer Science, 2334:211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2334/23340211.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2334/23340211.
                                        pdf.

Sterbenz:2002:HSM

[331] James P. G. Sterbenz. High-speed mobile and wireless networks. Lecture Notes in Computer Science, 2334:227– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2334/23340227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2334/23340227.
pdf.

Touch:2002:PNH

[332] Joseph D. Touch. Peer networks — high-speed solution or challenge? Lecture Notes in Computer Science, 2334: 228–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2334/23340228.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2334/23340228.
pdf.

Schrodi:2002:HSN

[333] Karl J. Schrodi. High speed networks for carriers. Lecture Notes in Computer Science, 2334:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2334/23340229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2334/23340229.
pdf.

Sterbenz:2002:PHS

[334] James P. G. Sterbenz. Protocols for high-speed networks: a brief retrospective survey of high-speed networking research. Lecture Notes in Computer Science, 2334:243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2334/23340243.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2334/23340243.
                                        pdf.

Anonymous:2002:AIe

[335] Anonymous. Author index. Lecture Notes in Computer Science, 2334:267– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2334/2334auth.pdf.

Gery:2002:RCL

[336] Eran Gery, David Harel, and Eldad Palachi. Rhapsody: a complete life-cycle model-based development system. Lecture Notes in Computer Science, 2335:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2335/23350001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2335/23350001.
                                        pdf.

Kuske:2002:ISU

[337] Sabine Kuske, Martin Gogolla, Ralf Kollmann, and Hans-Jörg Kreowski. An integrated semantics for UML class, object and state diagrams based on graph transformation. Lecture Notes in Computer Science, 2335:11– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2335/23350011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2335/23350011.
pdf.

Lopez:2002:SPA

[338] Natalia López, Manuel Núñez, and Fernando Rubio. Stochastic process algebras meet eden. Lecture Notes in Computer Science, 2335:29– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2335/23350029.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2335/23350029.
pdf.

Lano:2002:ISE

[339] K. Lano, D. Clark, and K. Androutsopoulos. From implicit specifications to explicit designs in reactive system development. Lecture Notes in Computer Science, 2335:49– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2335/23350049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2335/23350049.
pdf.

Nepomniaschy:2002:BRI

[340] V. A. Nepomniaschy, N. V. Shilov, E. V. Bodin, and V. E. Kozura. Basic-REAL: Integrated approach for design, specification and verification of distributed systems. Lecture Notes in Computer Science, 2335:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2335/23350069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2335/23350069.
                                        pdf.

Zulkernine:2002:AGA

[341] Mohammad Zulkernine and Rudolph E. Seviora. Assume-guarantee algorithms for automatic detection of software failures. Lecture Notes in Computer Science, 2335:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2335/23350089.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2335/23350089.
                                        pdf.

Ledang:2002:CMU

[342] Hung Ledang and Jeanine Souquières. Contributions for modelling UML state-charts in B. Lecture Notes in Computer Science, 2335:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2335/23350109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2335/23350109.
                                        pdf.

Sekerinski:2002:TSB

[343] Emil Sekerinski and Rafik Zurob. Translating statecharts to B. Lecture Notes in Computer Science, 2335:128– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2335/23350128.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2335/23350128.
pdf.

Katz:2002:FTM

[344] Shmuel Katz and Orna Grumberg. A framework for translating models and specifications. Lecture Notes in Computer Science, 2335:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2335/23350145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2335/23350145.
pdf.

Winter:2002:MCO

[345] Kirsten Winter and Roger Duke. Model checking object-Z using ASM. Lecture Notes in Computer Science, 2335:165– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2335/23350165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2335/23350165.
pdf.

Akbarpour:2002:FCS

[346] Behzad Akbarpour, Abdelkader Dekdouk, and Sofiène Tahar. Formalization of cadence SPW fixed-point arithmetic in HOL. Lecture Notes in Computer Science, 2335:185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2335/23350185.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2335/23350185.
                                        pdf.

Xiong:2002:FLM

[347] Haiyan Xiong, Paul Curzon, Sofiène Tahar, and Ann Blandford. Formally linking MDG and HOL based on a verified MDG system. Lecture Notes in Computer Science, 2335:205– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2335/23350205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2335/23350205.
                                        pdf.

Bolton:2002:ROZ

[348] Christie Bolton and Jim Davies. Refinement in object-Z and CSP. Lecture Notes in Computer Science, 2335: 225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2335/23350225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2335/23350225.
                                        pdf.

Hoenicke:2002:CST

[349] Jochen Hoenicke and Ernst-Rüdiger Olderog. Combining specification techniques for processes, data and time. Lecture Notes in Computer Science, 2335:245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2335/23350245.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2335/23350245.
pdf.

Smith:2002:IRT

[350] Graeme Smith. An integration of realtime object-Z and CSP for specifying concurrent real-time systems. Lecture Notes in Computer Science, 2335:267– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2335/23350267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2335/23350267.
pdf.

Kent:2002:MDE

[351] Stuart Kent. Model driven engineering. Lecture Notes in Computer Science, 2335:286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2335/23350286.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2335/23350286.
pdf.

Brooke:2002:DTS

[352] Phillip J. Brooke and Richard F. Paige. The design of a tool-supported graphical notation for timed CSP. Lecture Notes in Computer Science, 2335:299– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2335/23350299.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2335/23350299.
pdf.

Johnsen:2002:CGF

[353] Einar B. Johnsen, Wenhui Zhang, Olaf Owe, and Demissie B. Aredo. Combining graphical and formal development of open distributed systems. Lecture Notes in Computer Science, 2335:319– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2335/23350319.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2335/23350319.
                                        pdf.

Korenblat:2002:TBT

[354] Katerina Korenblat, Orna Grumberg, and Shmuel Katz. Translations between textual transition systems and petri nets. Lecture Notes in Computer Science, 2335:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2335/23350339.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2335/23350339.
                                        pdf.

Barradas:2002:SPL

[355] Héctor Ruíz Barradas and Didier Bert. Specification and proof of liveness properties under fairness assumptions in B event systems. Lecture Notes in Computer Science, 2335:360– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2335/23350360.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2335/23350360.
pdf.

Jeske:2002:MMA

[356] C. Jeske and R. Banach. Minimally and maximally abstract retrenchments. Lecture Notes in Computer Science, 2335:380–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2335/23350380.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2335/23350380.
pdf.

Anonymous:2002:AIf

[357] Anonymous. Author index. Lecture Notes in Computer Science, 2335:401– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2335/2335auth.pdf.

Garofalakis:2002:NDM

[358] Minos Garofalakis and Rajeev Rastogi. Network data mining and analysis: The N EMESIS Project. Lecture Notes in Computer Science, 2336: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360001.
pdf.

Srikant:2002:PPD

[359] Ramakrishnan Srikant. Privacy preserving data mining: Challenges and opportunities. Lecture Notes in Computer Science, 2336:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360013.
pdf.

Srivastava:2002:CAC

[360] Jaideep Srivastava, Jau-Hwang Wang, Ee-Peng Lim, and San-Yih Hwang. A case for analytical customer relationship management. Lecture Notes in Computer Science, 2336:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360014.
pdf.

Zaiane:2002:DCA

[361] Osmar R. Zaïane, Andrew Foss, Chi-Hoon Lee, and Weinan Wang. On data clustering analysis: Scalability, constraints, and validation. Lecture Notes in Computer Science, 2336:28– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360028.
pdf.

Mata:2002:DNA

[362] Jacinto Mata, José-Luis Alvarez, and José-Cristobal Riquelme. Discovering numeric association rules via evolutionary algorithm. Lecture Notes in Computer Science, 2336:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360040.
pdf.

Hipp:2002:ERR

[363] Jochen Hipp, Christoph Mangold, Ulrich Güntzer, and Gholamreza Nakhaeizadeh. Efficient rule retrieval and postponed restrict operations for association rule mining. Lecture Notes in Computer Science, 2336:52– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360052.
pdf.

Ding:2002:ARM

[364] Qin Ding, Qiang Ding, and William Perrizo. Association rule mining on remotely sensed images using P-trees. Lecture Notes in Computer Science, 2336:66–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2336/23360066.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2336/23360066.
pdf.

Pudi:2002:EAR

[365] Vikram Pudi and Jayant R. Haritsa. On the efficiency of associationrule mining algorithms. Lecture Notes in Computer Science, 2336:80– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360080.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360080.
                                        pdf.

Lin:2002:FBC

[366] Jung-Yi Lin, Been-Chian Chien, and Tzung-Pei Hong. A function-based classifier learning scheme using genetic programming. Lecture Notes in Computer Science, 2336:92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360092.
                                        pdf.

Xie:2002:SSN

[367] Zhipeng Xie, Wynne Hsu, Zongtian Liu, and Mong Li Lee. SNNB: a selective neighborhood based naïve Bayes for lazy learning. Lecture Notes in Computer Science, 2336:104– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360104.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360104.
                                        pdf.

Diao:2002:MBN

[368] Lili Diao, Keyun Hu, Yuchang Lu, and Chunyi Shi. A method to boost naïve Bayesian classifiers. Lecture Notes in Computer Science, 2336:115– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360115.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360115.
pdf.

Ling:2002:TBC

[369] Charles X. Ling and Huajie Zhang. Toward Bayesian classifiers with accurate probabilities. Lecture Notes in Computer Science, 2336:123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360123.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360123.
pdf.

Jaroszewicz:2002:PRA

[370] Szymon Jaroszewicz and Dan A. Simovici. Pruning redundant association rules using maximum entropy principle. Lecture Notes in Computer Science, 2336:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360135.
pdf.

Lin:2002:CLS

[371] Wen-Yang Lin, Ming-Cheng Tseng, and Ja-Hwung Su. A confidencelift support specification for interesting associations mining. Lecture Notes in Computer Science, 2336:148– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360148.
pdf.

Kryszkiewicz:2002:CRF

[372] Marzena Kryszkiewicz and Marcin Gajek. Concise representation of frequent patterns based on generalized disjunction-free generators. Lecture Notes in Computer Science, 2336:159– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360159.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360159.
pdf.

Yen:2002:MIA

[373] Show-Jane Yen and Yue-Shi Lee. Mining interesting association rules: a data mining language. Lecture Notes in Computer Science, 2336:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360172.
pdf.

Hilderman:2002:LDO

[374] Robert J. Hilderman. The Lorenz dominance order as a measure of interestingness in KDD. Lecture Notes in Computer Science, 2336:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360177.
pdf.

Zhang:2002:EAI

[375] Minghua Zhang, Ben Kao, David Cheung, and Chi-Lap Yip. Efficient algorithms for incremental update of frequent sequences. Lecture Notes in Computer Science, 2336:186– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360186.
pdf.

Lin:2002:DED

[376] Ming-Yen Lin, Suh-Yin Lee, and Sheng-Shun Wang. DELISP: Efficient discovery of generalized sequential patterns by delimited patterngrowth technology. Lecture Notes in Computer Science, 2336:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360198.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2336/23360198.
                                        pdf.

Adibi:2002:SSD

[377] Jafar Adibi, Wei-Min Shen, and Eaman Noorbakhsh. Self-similarity for data mining and predictive modeling — A case study for network data. Lecture Notes in Computer Science, 2336: 210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360210.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360210.
                                        pdf.

Lin:2002:NMM

[378] Shun-Chieh Lin, Shian-Shyong Tseng, and Yao-Tsung Lin. A new mechanism of mining network behavior. Lecture Notes in Computer Science, 2336:218– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360218.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360218.
                                        pdf.

Ng:2002:MFM

[379] Michael Ng and Joshua Huang. M-FastMap: a modified FastMap algorithm for visual cluster validation in data mining. Lecture Notes in Computer Science, 2336:224–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360224.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2336/23360224.
pdf.

Chen:2002:IHD

[380] Chien-Yu Chen, Shien-Ching Hwang, and Yen-Jen Oyang. An incremental hierarchical data clustering algorithm based on gravity theory. Lecture Notes in Computer Science, 2336: 237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360237.
pdf.

Tan:2002:API

[381] Ah-Hwee Tan and Hong Pan. Adding personality to information clustering. Lecture Notes in Computer Science, 2336:251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2336/23360251.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2336/23360251.
pdf.

Jollois:2002:CLC

[382] François-Xavier Jollois and Mohamed Nadif. Clustering large categorical data. Lecture Notes in Computer Science, 2336:257–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360257.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2336/23360257.
                                        pdf.

Zheng:2002:WPC

[383] Tong Zheng, Yonghe Niu, and Randy Goebel. WebFrame: In pursuit of computationally and cognitively efficient Web mining. Lecture Notes in Computer Science, 2336:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360264.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360264.
                                        pdf.

Prasetyo:2002:NWN

[384] Bowo Prasetyo, Iko Pramudiono, Katsumi Takahashi, and Masaru Kitsuregawa. Naviz: Website navigational behavior visualizer. Lecture Notes in Computer Science, 2336:276– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360276.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360276.
                                        pdf.

Chen:2002:OAF

[385] Zhixiang Chen, Ada Wai-Chee Fu, and Frank Chi-Hung Tong. Optimal algorithms for finding user access sessions from very large Web logs. Lecture Notes in Computer Science, 2336: 290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2336/23360290.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360290.
pdf.

Chang:2002:AIE

[386] Chia-Hui Chang, Shih-Chien Kuo, Kuo-Yu Hwang, Tsung-Hsin Ho, and Chih-Lung Lin. Automatic information extraction for multiple singular Web pages. Lecture Notes in Computer Science, 2336:297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360297.
pdf.

Dai:2002:IAD

[387] Honghua Dai and Gang Li. An improved approach for the discovery of causal models via MML. Lecture Notes in Computer Science, 2336:304– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360304.
pdf.

Chang:2002:SME

[388] Ye-In Chang and Yu-Ming Hsieh. SETM*-MaxK: An efficient SET-based approach to find the largest itemset. Lecture Notes in Computer Science, 2336:316–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2336/23360316.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360316.
                                        pdf.

Guillaume:2002:DOA

[389] Sylvie Guillaume. Discovery of ordinal association rules. Lecture Notes in Computer Science, 2336:322–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360322.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360322.
                                        pdf.

Lin:2002:VAA

[390] T. Y. Lin, Y. Y. Yao, and E. Louie. Value added association rules. Lecture Notes in Computer Science, 2336:328– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360328.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360328.
                                        pdf.

Wang:2002:TFG

[391] Ke Wang, Liu Tang, Jiawei Han, and Junqiang Liu. Top down FP-growth for association rule mining. Lecture Notes in Computer Science, 2336:334– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360334.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2336/23360334.
pdf.

Miyahara:2002:DFT

[392] Tetsuhiro Miyahara, Yusuke Suzuki, Takayoshi Shoudai, Tomoyuki Uchida, Kenichi Takahashi, and Hiroaki Ueda. Discovery of frequent tag tree patterns in semistructured Web documents. Lecture Notes in Computer Science, 2336: 341–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360341.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360341.
pdf.

Furukawa:2002:ECS

[393] Kazuyoshi Furukawa, Tomoyuki Uchida, Kazuya Yamada, Tetsuhiro Miyahara, Takayoshi Shoudai, and Yasuaki Nakamura. Extracting characteristic structures among words in semistructured documents. Lecture Notes in Computer Science, 2336:356– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360356.
pdf.

Cheung:2002:EAI

[394] Felix Cheung, Ben Kao, David Cheung, and Chi-Yuen Ng. An efficient algorithm for incremental update of concept spaces. Lecture Notes in Computer Science, 2336:368–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360368.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360368.
                                        pdf.

Li:2002:ECB

[395] Cuiping Li, Shengen Li, Shan Wang, and Xiaoyong Du. Efficient constraint-based exploratory mining on large data cubes. Lecture Notes in Computer Science, 2336:381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360381.
                                        pdf.

Yang:2002:EUM

[396] Don-Lin Yang, Man-Lin Huang, and Ming-Chuan Hung. Efficient utilization of materialized views in a data warehouse. Lecture Notes in Computer Science, 2336:393–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360393.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360393.
                                        pdf.

Zhong:2002:MIR

[397] Ning Zhong and Juzhen Dong. Mining interesting rules in meningitis data by cooperatively using GDT-RS and RSBR. Lecture Notes in Computer Science, 2336:405–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360405.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360405.
pdf.

Deshpande:2002:ETC

[398] Mukund Deshpande and George Karypis. Evaluation of techniques for classifying biological sequences. Lecture Notes in Computer Science, 2336: 417–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360417.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360417.
pdf.

Tseng:2002:EMG

[399] Vincent S. M. Tseng and Ching-Pin Kao. Efficiently mining gene expression data via integrated clustering and validation techniques. Lecture Notes in Computer Science, 2336:432– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360432.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360432.
pdf.

Liang:2002:AGE

[400] Yulan Liang, King-Ip Lin, and Arpad Kelemen. Adaptive generalized estimation equation with Bayes classifier for the job assignment problem. Lecture Notes in Computer Science, 2336:438– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360438.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360438.
                                        pdf.

Hsu:2002:GEA

[401] William W. Hsu and Ching-Chi Hsu. GEC: An evolutionary approach for evolving classifiers. Lecture Notes in Computer Science, 2336:450–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360450.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360450.
                                        pdf.

Fan:2002:ESS

[402] Hongjian Fan and Ramamohanarao Kotagiri. An efficient single-scan algorithm for mining essential jumping emerging patterns for classification. Lecture Notes in Computer Science, 2336:456–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2336/23360456.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360456.
                                        pdf.

Diao:2002:MBS

[403] Lili Diao, Keyun Hu, Yuchang Lu, and Chunyi Shi. A method to boost support vector machines. Lecture Notes in Computer Science, 2336:463– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360463.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360463.
pdf.

Jin:2002:DDL

[404] Xiaoming Jin, Yuchang Lu, and Chunyi Shi. Distribution discovery: Local analysis of temporal rules. Lecture Notes in Computer Science, 2336:469– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360469.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360469.
pdf.

Fung:2002:NSS

[405] Gabriel Pui Cheong Fung, Jeffrey Xu Yu, and Wai Lam. News sensitive stock trend prediction. Lecture Notes in Computer Science, 2336:481– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360481.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360481.
pdf.

Yeung:2002:UPI

[406] Dit-Yan Yeung and Yuxin Ding. User profiling for intrusion detection using dynamic and static behavioral models. Lecture Notes in Computer Science, 2336:494–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360494.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360494.
                                        pdf.

Chien:2002:IEK

[407] Lee-Feng Chien, Chien-Kang Huang, Hsin-Chen Chiao, and Shih-Jui Lin. Incremental extraction of keyterms for classifying multilingual documents in the Web. Lecture Notes in Computer Science, 2336:506–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360506.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360506.
                                        pdf.

Khan:2002:NNC

[408] Maleq Khan, Qin Ding, and William Perrizo. k-nearest neighbor classification on spatial data streams using P-trees. Lecture Notes in Computer Science, 2336:517–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2336/23360517.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2336/23360517.pdf.

Han:2002:ICC

[409] Jianchao Han and Nick Cercone.

Interactive construction of classification rules. Lecture Notes in Computer Science, 2336:529–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360529.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360529.
pdf.

Tang:2002:EEO

[410] Jian Tang, Zhixiang Chen, Ada Wai chee Fu, and David W. Cheung. Enhancing effectiveness of outlier detections for low density patterns. Lecture Notes in Computer Science, 2336:535– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360535.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360535.
pdf.

Fujikawa:2002:CBA

[411] Yoshikazu Fujikawa and TuBao Ho. Cluster-based algorithms for dealing with missing values. Lecture Notes in Computer Science, 2336:549–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2336/23360549.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2336/23360549.
pdf.

Chan:2002:ECK

[412] Ki Chan, Boon-Toh Low, Wai Lam, and Kai-Pui Lam. Extracting causation knowledge from natural language texts. Lecture Notes in Computer Science, 2336:555–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360555.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360555.
                                        pdf.

Ong:2002:MRG

[413] Kok-Leong Ong, Wee-Keong Ng, and Ee-Peng Lim. Mining relationship graphs for effective business objectives. Lecture Notes in Computer Science, 2336:561–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2336/23360561.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2336/23360561.
                                        pdf.

Anonymous:2002:AIg

[414] Anonymous. Author index. Lecture Notes in Computer Science, 2336:567– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2336/2336indx.pdf.

Iwata:2002:FSA

[415] Satoru Iwata. A faster scaling algorithm for minimizing submodular functions. Lecture Notes in Computer Science, 2337:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2337/23370001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370001.
pdf.

Spille:2002:GEM

[416] Bianca Spille and Robert Weismantel. A generalization of Edmonds’ matching and matroid intersection algorithms. Lecture Notes in Computer Science, 2337:9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2337/23370009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370009.
pdf.

Tamura:2002:CDS

[417] Akihisa Tamura. A coordinatewise domain scaling algorithm for Mconvex function minimization. Lecture Notes in Computer Science, 2337:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2337/23370021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370021.
pdf.

Fleischer:2002:QMF

[418] Lisa Fleischer and Martin Skutella. The quickest multicommodity flow problem. Lecture Notes in Computer Science, 2337:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2337/23370036.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2337/23370036.pdf.

Gunluk:2002:NMC

[419] Oktay Günlük. A new min-cut maxflow ratio for multicommodity flows. Lecture Notes in Computer Science, 2337:54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2337/23370054.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370054.
                                        pdf.

Lewin:2002:IRT

[420] Michael Lewin, Dror Livnat, and Uri Zwick. Improved rounding techniques for the MAX 2-SAT and MAX DI-CUT problems. Lecture Notes in Computer Science, 2337:67– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2337/23370067.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370067.
                                        pdf.

Boyd:2002:FEI

[421] Sylvia Boyd and Geneviève Labonté. Finding the exact integrality gap for small traveling salesman problems. Lecture Notes in Computer Science, 2337:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2337/23370083.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2337/23370083.
pdf.

Letchford:2002:PTS

[422] Adam N. Letchford and Andrea Lodi. Polynomial-time separation of simple comb inequalities. Lecture Notes in Computer Science, 2337:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2337/23370093.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370093.
pdf.

Wenger:2002:NAC

[423] Klaus M. Wenger. A new approach to cactus construction applied to TSP support graphs. Lecture Notes in Computer Science, 2337:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2337/23370109.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370109.
pdf.

Andersen:2002:SCI

[424] Kent Andersen, Gérard Cornuéjols, and Yanjun Li. Split closure and intersection cuts. Lecture Notes in Computer Science, 2337:127–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2337/23370127.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370127.
                                        pdf.

Dash:2002:ELB

[425] Sanjeeb Dash. An exponential lower bound on the length of some classes of branch-and-cut proofs. Lecture Notes in Computer Science, 2337:145– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2337/23370145.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370145.
                                        pdf.

Richard:2002:LIM

[426] Jean-Philippe P. Richard, Ismael R. de Farias, Jr., and George L. Nemhauser. Lifted inequalities for 0-1 mixed integer programming: Basic theory and algorithms. Lecture Notes in Computer Science, 2337:161– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2337/23370161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370161.
                                        pdf.

Aharoni:2002:LS

[427] Ron Aharoni and Tamás Fleiner. On a lemma of scarf. Lecture Notes in Computer Science, 2337:176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2337/23370176.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370176.
pdf.

Guenin:2002:SPS

[428] Bertrand Guenin. A short proof of Seymour’s characterization of the matroids with the max-flow min-cut property. Lecture Notes in Computer Science, 2337:188–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2337/23370188.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370188.
pdf.

Sethuraman:2002:IPA

[429] Jay Sethuraman, Chung-Piaw Teo, and Rakesh V. Vohra. Integer programming and Arrovian social welfare functions. Lecture Notes in Computer Science, 2337:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2337/23370194.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370194.
pdf.

Ravi:2002:ILA

[430] Ramamoorthi Ravi and Amitabh Sinha. Integrated logistics: Approximation algorithms combining facility location and network design. Lecture Notes in Computer Science, 2337:212– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2337/23370212.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370212.
                                        pdf.

Sitters:2002:MLP

[431] René Sitters. The minimum latency problem is NP-hard for weighted trees. Lecture Notes in Computer Science, 2337:230–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2337/23370230.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370230.
                                        pdf.

Sviridenko:2002:IAA

[432] Maxim Sviridenko. An improved approximation algorithm for the metric uncapacitated facility location problem. Lecture Notes in Computer Science, 2337:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2337/23370240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370240.
                                        pdf.

Althaus:2002:PAS

[433] Ernst Althaus and Christian Fink. A polyhedral approach to surface reconstruction from planar contours. Lecture Notes in Computer Science, 2337: 258–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2337/23370258.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370258.
pdf.

Eisenblatter:2002:SRP

[434] Andreas Eisenblätter. The semidefinite relaxation of the k-partition polytope is strong. Lecture Notes in Computer Science, 2337:273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2337/23370273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370273.
pdf.

Farias:2002:PSC

[435] Ismael R. de Farias, Jr. and George L. Nemhauser. A polyhedral study of the cardinality constrained knapsack problem. Lecture Notes in Computer Science, 2337:291–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2337/23370291.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2337/23370291.pdf.

Cai:2002:PMT

[436] Mao-Cheng Cai, Xiaotie Deng, Haodi Feng, Guojun Li, and Guizhen Liu. A PTAS for minimizing total completion time of bounded batch scheduling. Lecture Notes in Computer Science, 2337: 304–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2337/23370304.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370304.
                                        pdf.

Caprara:2002:AST

[437] Alberto Caprara, Andrea Lodi, and Michele Monaci. An approximation scheme for the two-stage, twodimensional bin packing problem. Lecture Notes in Computer Science, 2337: 315–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2337/23370315.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370315.
                                        pdf.

Jansen:2002:PRC

[438] Klaus Jansen and Lorant Porkolab. On preemptive resource constrained scheduling: Polynomial-time approximation schemes. Lecture Notes in Computer Science, 2337:329–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2337/23370329.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370329.
                                        pdf.

Aardal:2002:HEC

[439] Karen Aardal and Arjen K. Lenstra. Hard equality constrained integer knapsacks. Lecture Notes in Computer Science, 2337:350–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2337/23370350.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2337/23370350.pdf.

Barvinok:2002:DVQ

[440] Alexander Barvinok and Tamon Stephen. The distribution of values in the quadratic assignment problem. Lecture Notes in Computer Science, 2337: 367–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2337/23370367.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370367.
pdf.

Klabjan:2002:NSA

[441] Diego Klabjan. A new subadditive approach to integer programming. Lecture Notes in Computer Science, 2337: 384–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2337/23370384.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370384.
pdf.

Calinescu:2002:IAA

[442] Gruia Calinescu, Amit Chakrabarti, Howard Karloff, and Yuval Rabani. Improved approximation algorithms for resource allocation. Lecture Notes in Computer Science, 2337:401–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2337/23370401.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370401.
                                        pdf.

Freund:2002:AAP

[443] Ari Freund and Joseph (Seffi) Naor. Approximating the advertisement placement problem. Lecture Notes in Computer Science, 2337:415–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2337/23370415.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370415.
                                        pdf.

Gandhi:2002:AMR

[444] Rajiv Gandhi, Samir Khuller, Yoo-Ah Kim, and Yung-Chun (Justin) Wan. Algorithms for minimizing response time in broadcast scheduling. Lecture Notes in Computer Science, 2337:425– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2337/23370425.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2337/23370425.
                                        pdf.

Chekuri:2002:BEF

[445] Chandra Chekuri, Anupam Gupta, Amit Kumar, Joseph (Seffi) Naor, and Danny Raz. Building edgefailure resilient networks. Lecture Notes in Computer Science, 2337:439– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2337/23370439.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370439.
pdf.

Shepherd:2002:DMP

[446] Bruce Shepherd and Adrian Vetta. The demand matching problem. Lecture Notes in Computer Science, 2337:457– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2337/23370457.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2337/23370457.
pdf.

Talwar:2002:SSB

[447] Kunal Talwar. The single-sink buy-at-bulk LP has constant integrality gap. Lecture Notes in Computer Science, 2337:475–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2337/23370475.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2337/23370475.
pdf.

Anonymous:2002:AIh

[448] Anonymous. Author index. Lecture Notes in Computer Science, 2337:487– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2337/2337auth.pdf.

DeLoach:2002:MOR

[449] Scott A. DeLoach. Modeling organizational rules in the multi-agent systems engineering methodology. Lecture Notes in Computer Science, 2338: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380001.
pdf.

Chen:2002:AOA

[450] David Chen and Robin Cohen. AERO: An outsourced approach to exception handling in multi-agent systems. Lecture Notes in Computer Science, 2338: 16–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380016.
pdf.

Tran:2002:LABa

[451] Thomas Tran and Robin Cohen. A learning algorithm for buying and selling agents in electronic marketplaces. Lecture Notes in Computer Science, 2338:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2338/23380031.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2338/23380031.
pdf.

Yap:2002:GBP

[452] Peter Yap. Grid-based path-finding. Lecture Notes in Computer Science, 2338:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2338/23380044.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2338/23380044.
pdf.

Kishimoto:2002:TTD

[453] Akihiro Kishimoto and Jonathan Schaeffer. Transposition table driven work scheduling in distributed game-tree search. Lecture Notes in Computer Science, 2338:56–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2338/23380056.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2338/23380056.pdf.

Neufeld:2002:CTA

[454] Eric Neufeld. Clue as a testbed for automated theorem proving. Lecture Notes in Computer Science, 2338:69– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380069.
pdf.

Lashkia:2002:NFM

[455] George V. Lashkia. A noise filtering method for inductive concept learning.

Lecture Notes in Computer Science, 2338:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2338/23380079.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2338/23380079.
pdf.

Silver:2002:TRM

[456] Daniel L. Silver and Robert E. Mercer. The task rehearsal method of life-long learning: Overcoming impoverished data. Lecture Notes in Computer Science, 2338:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380090.
pdf.

Mackworth:2002:RCP

[457] Alan Mackworth. Recycling the cycle of perception: a polemic. Lecture Notes in Computer Science, 2338:102– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380102.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380102.
pdf.

Horsch:2002:GAC

[458] Michael C. Horsch, William S. Havens, and Aditya K. Ghose. Generalized arc consistency with application to MaxCSP. Lecture Notes in Computer Science, 2338:104–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380104.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380104.
pdf.

Huang:2002:TLL

[459] Guan-Shieng Huang, Xiumei Jia, Churn-Jung Liau, and Jia-Huai You. Two-literal logic programs and satisfiability representation of stable models: a comparison. Lecture Notes in Computer Science, 2338:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380119.
pdf.

Kennedy:2002:UCA

[460] Kevin Kennedy and Robert E. Mercer. Using communicative acts to plan the cinematographic structure of animations. Lecture Notes in Computer Science, 2338:132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380132.
pdf.

Ezeife:2002:MIA

[461] Christie I. Ezeife and Yue Su. Mining incremental association rules with generalized FP-tree. Lecture Notes in Computer Science, 2338:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380147.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380147.
                                        pdf.

Ayad:2002:TDT

[462] Hanan Ayad and Mohamed Kamel. Topic discovery from text using aggregation of different clustering methods. Lecture Notes in Computer Science, 2338:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2338/23380161.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380161.
                                        pdf.

Parker:2002:GAC

[463] James R. Parker. Genetic algorithms for continuous problems. Lecture Notes in Computer Science, 2338:176– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380176.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380176.
                                        pdf.

Butz:2002:RCW

[464] Cory J. Butz and Manon J. Sanscartier. On the role of contextual weak independence in probabilistic inference. Lecture Notes in Computer Science, 2338:185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380185.
pdf.

Wong:2002:SCD

[465] S. K. M. Wong, Dan Wu, and Tao Lin. A structural characterization of DAGisomorphic dependency models. Lecture Notes in Computer Science, 2338: 195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380195.
pdf.

Wong:2002:CNR

[466] S. K. M. Wong, Tao Lin, and Dan Wu. Construction of a non-redundant cover for conditional independencies. Lecture Notes in Computer Science, 2338:210– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380210.
pdf.

Breban:2002:UIA

[467] Silvia Breban and Julita Vassileva. Using inter-agent trust relationships for efficient coalition formation. Lecture Notes in Computer Science, 2338:221– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380221.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380221.
                                        pdf.

Fedoruk:2002:UAR

[468] Alan Fedoruk and Ralph Deters. Using agent replication to enhance reliability and availability of multiagent systems. Lecture Notes in Computer Science, 2338:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380237.
                                        pdf.

Frost:2002:ECS

[469] Richard Frost and Pierre Boulos. An efficient compositional semantics for natural-language database queries with arbitrarily-nested quantification and negation. Lecture Notes in Computer Science, 2338:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380252.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380252.
                                        pdf.

Copeck:2002:TSC

[470] Terry Copeck, Nathalie Japkowicz, and Stan Szpakowicz. Text summarization as controlled search. Lecture Notes in Computer Science, 2338:268– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380268.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380268.
pdf.

Plamondon:2002:QFB

[471] Luc Plamondon and Leila Kosseim. QUANTUM: a function-based question answering system. Lecture Notes in Computer Science, 2338:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380281.
pdf.

Chali:2002:GQB

[472] Yllias Chali. Generic and querybased text summarization using lexical cohesion. Lecture Notes in Computer Science, 2338:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380293.
pdf.

Mahdi:2002:LFM

[473] Tamer S. Mahdi and Robert E. Mercer. A lexical functional mapping algorithm. Lecture Notes in Computer Science, 2338:303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380303.
                                        pdf.

Kemke:2002:CAP

[474] Christel Kemke. A constructive approach to parsing with neural networks — the hybrid connectionist parsing method. Lecture Notes in Computer Science, 2338:310–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380310.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380310.
                                        pdf.

Bakus:2002:ETP

[475] Jan Bakus, Mohamed Kamel, and Tom Carey. Extraction of text phrases using hierarchical grammar. Lecture Notes in Computer Science, 2338:319– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380319.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380319.
                                        pdf.

Grewal:2002:EGA

[476] G. Grewal, T. Wilson, and C. Nell. An enhanced genetic algorithm approach to the channel assignment problem in mobile cellular networks. Lecture Notes in Computer Science, 2338: 325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380325.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380325.
pdf.

Karimi:2002:RAB

[477] Kamran Karimi and Howard J. Hamilton. RFCT: An association-based causality miner. Lecture Notes in Computer Science, 2338:334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380334.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380334.
pdf.

Wang:2002:RSD

[478] Fulai Wang and Jim Greer. Retrieval of short documents from discussion forums. Lecture Notes in Computer Science, 2338:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380339.
pdf.

Xiang:2002:ABN

[479] Yang Xiang, Chenwen Ye, and Deborah Ann Stacey. Application of Bayesian networks to shopping assistance. Lecture Notes in Computer Science, 2338:344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380344.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380344.
                                        pdf.

Salle:2002:ECF

[480] Mathias Sallé. Electronic contract framework for contractual agents. Lecture Notes in Computer Science, 2338: 349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380349.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380349.
                                        pdf.

Abi-Aad:2002:UMC

[481] Rony Abi-Aad, Thiruvengadam Radhakrishnan, and Ahmed Seffah. User models: Customizing E-commerce Websites to the context of use. Lecture Notes in Computer Science, 2338: 354–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380354.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380354.
                                        pdf.

Sharifi:2002:SNM

[482] Golha Sharifi, Julita Vassileva, and Ralph Deters. Supporting the needs of mobile home care workers: a case study for saskatoon district health system. Lecture Notes in Computer Science, 2338:356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380356.
pdf.

Shakshuki:2002:MAS

[483] E. Shakshuki and P. Kajonpotisuwan. Multi-agent system architecture for computer-based tutoring systems. Lecture Notes in Computer Science, 2338: 358–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380358.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380358.
pdf.

Renaud:2002:WPE

[484] Alfred I. Renaud. Word prediction evaluation measures with performance benchmarking. Lecture Notes in Computer Science, 2338:361–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2338/23380361.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2338/23380361.
pdf.

Abou-Assaleh:2002:RUP

[485] Tony Abou-Assaleh and Nick Cercone.

Relaxed unification — proposal. Lecture Notes in Computer Science, 2338: 364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380364.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380364.
                                        pdf.

Tran:2002:LABb

[486] Thomas Tran and Robin Cohen. A learning algorithm for buying and selling agents in electronic marketplaces. Lecture Notes in Computer Science, 2338:366–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2338/23380366.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380366.
                                        pdf.

Zapata-Rivera:2002:CKE

[487] Diego Zapata-Rivera. cbCPT: Knowledge engineering support for CPTs in Bayesian networks. Lecture Notes in Computer Science, 2338:368–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2338/23380368.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2338/23380368.
                                        pdf.

Mishra:2002:QLR

[488] Laxmikanta Mishra. Query-less retrieval of interesting postings in a Web-Forum. Lecture Notes in Computer Science, 2338:371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2338/23380371.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2338/23380371.pdf.

Anonymous:2002:AIi

[489] Anonymous. Author index. Lecture Notes in Computer Science, 2338:373– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2338/2338ind.pdf.

Liskov:2002:AC

[490] Moses Liskov and Silvio Micali. Amortized E-cash. Lecture Notes in Computer Science, 2339:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390001.
pdf.

Blaze:2002:OMT

[491] Matt Blaze, John Ioannidis, and Angelos D. Keromytis. Offline micropayments without trusted hardware. Lecture Notes in Computer Science, 2339: 21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390021.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2339/23390021.
                                        pdf.

vanSomeren:2002:PPI

[492] Nicko van Someren. The practical problems of implementing MicroMint. Lecture Notes in Computer Science, 2339:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2339/23390041.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2339/23390041.
                                        pdf.

Frankel:2002:PDR

[493] Yair Frankel. Protecting digital rights. Lecture Notes in Computer Science, 2339:51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2339/23390051.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2339/23390051.
                                        pdf.

Kravitz:2002:ADR

[494] David W. Kravitz. Aspects of digital rights management and the use of hardware security devices. Lecture Notes in Computer Science, 2339:54– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2339/23390054.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2339/23390054.
                                        pdf.

Moskowitz:2002:SNP

[495] Scott Moskowitz. A solution to the Napster phenomenon: Why value cannot be created absent the transfer of subjective data. Lecture Notes in Computer Science, 2339:59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390059.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390059.
pdf.

Sander:2002:GTD

[496] Tomas Sander. Golden times for digital rights management? Lecture Notes in Computer Science, 2339:64– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390064.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390064.
pdf.

Wyant:2002:APK

[497] Jeremy Wyant. Applicability of public key cryptosystems to digital rights management applications. Lecture Notes in Computer Science, 2339:75– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390075.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390075.
pdf.

Kohno:2002:GCP

[498] Tadayoshi Kohno and Mark McGovern. On the global content PMI: Improved copy-protected internet content distribution. Lecture Notes in Computer Science, 2339:79–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390079.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390079.
pdf.

Camp:2002:TCP

[499] L. Jean Camp, Helen Nissenbaum, and Cathleen McGrath. Trust: a collision of paradigms. Lecture Notes in Computer Science, 2339:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390091.
pdf.

Jaulmes:2002:SHG

[500] Éliane Jaulmes and Guillaume Poupard. On the security of homage group authentication protocol. Lecture Notes in Computer Science, 2339:106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390106.
pdf.

Malkhi:2002:ACE

[501] Dahlia Malkhi and Elan Pavlov. Anonymity without ‘cryptography’ (extended abstract). Lecture Notes in Computer Science, 2339:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390117.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390117.
pdf.

Kugler:2002:FTT

[502] Dennis Kügler and Holger Vogt. Fair tracing without trustees. Lecture Notes in Computer Science, 2339:136– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390136.
pdf.

Rahn:2002:WWM

[503] Richard W. Rahn. Why the war on money laundering should be aborted. Lecture Notes in Computer Science, 2339:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2339/23390149.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2339/23390149.
pdf.

Brown:2002:PSI

[504] Daniel R. L. Brown, Robert Gallant, and Scott A. Vanstone. Provably secure implicit certificate schemes. Lecture Notes in Computer Science, 2339: 156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390156.
pdf.

Li:2002:NUI

[505] Ninghui Li and Joan Feigenbaum. Nonmonotonicity, user interfaces, and risk assessment in certificate revocation (position paper). Lecture Notes in Computer Science, 2339:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390166.
pdf.

Jakobsson:2002:MAL

[506] Markus Jakobsson and David Pointcheval. Mutual authentication for low-power mobile devices. Lecture Notes in Computer Science, 2339:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390178.
pdf.

Rubin:2002:LGL

[507] Aviel D. Rubin and Rebecca N. Wright. Off-line generation of limiteduse credit card numbers. Lecture Notes in Computer Science, 2339:196– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390196.
pdf.

Tsiounis:2002:SFC

[508] Yiannis Tsiounis. A security framework for card-based systems. Lecture Notes in Computer Science, 2339:210– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390210.
pdf.

Shamir:2002:SWP

[509] Adi Shamir. SecureClick: a Web payment system with disposable credit card numbers. Lecture Notes in Computer Science, 2339:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390232.
pdf.

Gerck:2002:BEV

[510] Ed Gerck, C. Andrew Neff, Ronald L. Rivest, Aviel D. Rubin, and Moti Yung. The business of electronic voting. Lecture Notes in Computer Science, 2339:

243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2339/23390243.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2339/23390243.
                                        pdf.

DiCrescenzo:2002:PSM

[511] Giovanni Di Crescenzo. Privacy for the stock market. Lecture Notes in Computer Science, 2339:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2339/23390269.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2339/23390269.
                                        pdf.

Golle:2002:SDC

[512] Philippe Golle and Stuart Stubblebine. Secure distributed computing in a commercial environment. Lecture Notes in Computer Science, 2339:289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2339/23390289.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2339/23390289.
                                        pdf.

Naccache:2002:MS

[513] David Naccache, David Pointcheval, and Christophe Tymen. Monotone signatures. Lecture Notes in Computer Science, 2339:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390305.
pdf.

Bellare:2002:PRI

[514] Mihir Bellare, Chanathip Namprempre, David Pointcheval, and Michael Semanko. The power of RSA inversion oracles and the security of chaum’s RSA-based blind signature scheme. Lecture Notes in Computer Science, 2339:319–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2339/23390319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2339/23390319.
pdf.

Markowitch:2002:OFE

[515] Olivier Markowitch and Shahrokh Saeednia. Optimistic fair exchange with transparent signature recovery. Lecture Notes in Computer Science, 2339:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2339/23390339.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2339/23390339.
pdf.

Kikuchi:2002:MSP

[516] Hiroaki Kikuchi. ( M +1)st-price auction protocol. Lecture Notes in Computer Science, 2339:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2339/23390351.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2339/23390351.
                                        pdf.

Baudron:2002:NIP

[517] Olivier Baudron and Jacques Stern. Non-interactive private auctions. Lecture Notes in Computer Science, 2339: 364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2339/23390364.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2339/23390364.
                                        pdf.

Anonymous:2002:AIj

[518] Anonymous. Author index. Lecture Notes in Computer Science, 2339:379– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2339/23390379.pdf.

Hinze:2002:OOS

[519] Thomas Hinze, Uwe Hatnik, and Monika Sturm. An object oriented simulation of real occurring molecular biological processes for DNA computing and its experimental verification. Lecture Notes in Computer Science, 2340: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400001.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2340/23400001.
pdf.

Kashiwamura:2002:TOP

[520] Satoshi Kashiwamura, Masashi Nakatsugawa, Masahito Yamamoto, Toshikazu Shiba, and Azuma Ohuchi. Towards optimization of PCR protocol in DNA computing. Lecture Notes in Computer Science, 2340:14– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400014.
pdf.

Feldkamp:2002:DPC

[521] Udo Feldkamp, Sam Saghafi, Wolfgang Banzhaf, and Hilmar Rauhe. DNASequenceGenerator: a program for the construction of DNA sequences. Lecture Notes in Computer Science, 2340: 23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400023.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400023.
pdf.

vanNoort:2002:DCM

[522] Danny van Noort, Frank-Ulrich Gast, and John S. McCaskill. DNA computing in microreactors. Lecture Notes in Computer Science, 2340:33– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2340/23400033.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400033.
                                        pdf.

Penchovsky:2002:CHT

[523] Robert Penchovsky and John S. Mc-Caskill. Cascadable hybridisation transfer of specific DNA between microreactor selection modules. Lecture Notes in Computer Science, 2340:46– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400046.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400046.
                                        pdf.

Hussini:2002:CPD

[524] Salah Hussini, Lila Kari, and Stavros Konstantinidis. Coding properties of DNA languages. Lecture Notes in Computer Science, 2340:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400057.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400057.
                                        pdf.

Jonoska:2002:BCT

[525] Natasa Jonoska and Masahico Saito. Boundary components of thickened graphs. Lecture Notes in Computer Science, 2340:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2340/23400070.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2340/23400070.pdf.

Sakakibara:2002:PCM

[526] Yasubumi Sakakibara. Population computation and majority inference in test tube. Lecture Notes in Computer Science, 2340:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400082.
pdf.

Wood:2002:DSL

[527] David Harlan Wood, Hong Bi, Steven O. Kimbrough, Dong-Jun Wu, and Junghuei Chen. DNA starts to learn poker. Lecture Notes in Computer Science, 2340:92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400092.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400092.
pdf.

Rose:2002:PMW

[528] John A. Rose, Russell J. Deaton, Masami Hagiya, and Akira Suyama. PNA-mediated whiplash PCR. Lecture Notes in Computer Science, 2340:104– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400104.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400104.
                                        pdf.

Garzon:2002:BCV

[529] Max H. Garzon and Chris Oehmen. Biomolecular computation in virtual test tubes. Lecture Notes in Computer Science, 2340:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400117.
                                        pdf.

Tanaka:2002:DSS

[530] Fumiaki Tanaka, Masashi Nakatsugawa, Masahito Yamamoto, Toshikazu Shiba, and Azuma Ohuchi. Developing support system for sequence design in DNA computing. Lecture Notes in Computer Science, 2340:129– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400129.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400129.
                                        pdf.

Rose:2002:FTA

[531] John A. Rose, Russell J. Deaton, Masami Hagiya, and Akira Suyama. The fidelity of the tag-antitag system. Lecture Notes in Computer Science, 2340:138–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2340/23400138.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2340/23400138.
pdf.

Ruben:2002:PEA

[532] Adam J. Ruben, Stephen J. Freeland, and Laura F. Landweber. PUNCH: An evolutionary algorithm for optimizing bit set selection. Lecture Notes in Computer Science, 2340:150– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400150.
pdf.

Perez-Jimenez:2002:SKP

[533] Mario J. Pérez-Jiménez and Fernando Sancho-Caparrini. Solving knapsack problems in a sticker based model. Lecture Notes in Computer Science, 2340: 161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400161.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400161.
pdf.

Manca:2002:CSD

[534] Vincenzo Manca and Claudio Zandron. A clause string DNA algorithm for SAT. Lecture Notes in Computer Science, 2340:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2340/23400172.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400172.
                                        pdf.

Yoichi:2002:PDC

[535] Takenaka Yoichi and Hashimoto Akihiro. A proposal of DNA computing on beads with application to SAT problems. Lecture Notes in Computer Science, 2340:182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400182.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400182.
                                        pdf.

Head:2002:ASA

[536] Tom Head, Xia Chen, Matthew J. Nichols, Masayuki Yamamura, and Susannah Gal. Aqueous solutions of algorithmic problems: Emphasizing knights on a 3 × 3. Lecture Notes in Computer Science, 2340:191– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400191.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400191.
                                        pdf.

Yamamoto:2002:SSP

[537] Masahito Yamamoto, Nobuo Matsuura, Toshikazu Shiba, Yumi Kawazoe, and Azuma Ohuchi. Solutions of shortest path problems by concentration control. Lecture Notes in Computer Science, 2340:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400203.
pdf.

Yamamura:2002:ARA

[538] Masayuki Yamamura, Yusuke Hiroto, and Taku Matoba. Another realization of aqueous computing with peptide nucleic acid. Lecture Notes in Computer Science, 2340:213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400213.
pdf.

Khodor:2002:ECB

[539] Yevgenia Khodor, Julia Khodor, and Thomas F. Knight, Jr. Experimental confirmation of the basic principles of length-only discrimination. Lecture Notes in Computer Science, 2340:223– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400223.
pdf.

Reif:2002:ECV

[540] John H. Reif, Thomas H. LaBean, Michael Pirrung, Vipul S. Rana, Bo Guo, Carl Kingsford, and Gene S.

Wickham. Experimental construction of very large scale DNA databases with associative search capability. Lecture Notes in Computer Science, 2340:231– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400231.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400231.
                                        pdf.

Simmel:2002:OPD

[541] Friedrich C. Simmel and Bernard Yurke. Operation of a purified DNA nanoactuator. Lecture Notes in Computer Science, 2340:248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400248.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400248.
                                        pdf.

Mitchell:2002:DS

[542] James C. Mitchell and Bernard Yurke. DNA scissors. Lecture Notes in Computer Science, 2340:258–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400258.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400258.
                                        pdf.

Wakabayashi:2002:RIG

[543] Kenichi Wakabayashi and Masayuki Yamamura. A realization of information gate by using enterococcus faecalis pheromone system. Lecture Notes in Computer Science, 2340:269– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400269.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400269.
pdf.

Ehrenfeucht:2002:PMG

[544] Andrzej Ehrenfeucht, Tero Harju, Ion Petre, and Grzegorz Rozenberg. Patterns of micronuclear genes in ciliates. Lecture Notes in Computer Science, 2340:279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400279.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400279.
pdf.

Balan:2002:PCU

[545] M. Sakthi Balan, Kamala Krithivasan, and Y. Sivasubramanyam. Peptide computing — universality and complexity. Lecture Notes in Computer Science, 2340:290–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2340/23400290.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2340/23400290.pdf.

Khodor:2002:PMU

[546] Julia Khodor and David K. Gifford.

Programmed mutagenesis is a universal model of computation. Lecture Notes in Computer Science, 2340:300– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400300.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400300.
                                        pdf.

Uejima:2002:HCC

[547] Hiroki Uejima, Masami Hagiya, and Satoshi Kobayashi. Horn clause computation by self-assembly of DNA molecules. Lecture Notes in Computer Science, 2340:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400308.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400308.
                                        pdf.

Hug:2002:DBP

[548] Hubert Hug and Rainer Schuler. DNA-based parallel computation of simple arithmetic. Lecture Notes in Computer Science, 2340:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400321.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400321.
                                        pdf.

Paun:2002:PSG

[549] Andrei Paun. On P systems with global rules. Lecture Notes in Computer Science, 2340:329–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400329.
pdf.

Margenstern:2002:CMV

[550] Maurice Margenstern, Carlos Martín-Vide, and Gheorghe Paun. Computing with membranes: Variants with an enhanced membrane handling. Lecture Notes in Computer Science, 2340:340– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400340.
pdf.

Baranda:2002:TEI

[551] Angel V. Baranda, Fernando Arroyo, Juan Castellanos, and Rafael Gonzalo. Towards an electronic implementation of membrane computing: a formal description of non-deterministic evolution in transition P systems. Lecture Notes in Computer Science, 2340:350– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2340/23400350.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2340/23400350.
pdf.

Krishna:2002:IDP

[552] Shankara Narayanan Krishna and Raghavan Rama. Insertion-deletion P systems. Lecture Notes in Computer Science, 2340:360–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2340/23400360.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2340/23400360.pdf.

Margenstern:2002:UTV

[553] Maurice Margenstern and Yurii Rogozhin. A universal time-varying distributed H system of degree 1. Lecture Notes in Computer Science, 2340: 371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400371.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400371.
                                        pdf.

David:2002:NGS

[554] N. Gnanamalar David, K. G. Subramanian, and D. Gnanaraj Thomas. A note on graph splicing languages. Lecture Notes in Computer Science, 2340: 381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2340/23400381.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2340/23400381.
                                        pdf.

Anonymous:2002:AIk

[555] Anonymous. Author index. Lecture Notes in Computer Science, 2340:391– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2340/2340indx.pdf.

Fensel:2002:SWE

[556] Dieter Fensel, Christoph Bussler, and Alexander Maedche. Semantic Web enabled Web services. Lecture Notes in Computer Science, 2342:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420001.
pdf.

Kesselman:2002:GGS

[557] Carl Kesselman. The Grid, Grid services and the Semantic Web: Technologies and opportunities. Lecture Notes in Computer Science, 2342:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420003.
pdf.

Carroll:2002:MRG

[558] Jeremy J. Carroll. Matching RDF graphs. Lecture Notes in Computer Science, 2342:5–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2342/23420005.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2342/23420005.pdf.

Patel-Schneider:2002:LSW

[559] Peter F. Patel-Schneider and Dieter Fensel. Layering the Semantic Web: Problems and directions. Lecture Notes in Computer Science, 2342:16– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420016.
pdf.

Kamps:2002:NIS

[560] Jaap Kamps and Maarten Marx. Notions of indistinguishability for Semantic Web languages. Lecture Notes in Computer Science, 2342:30–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420030.
pdf.

Missikoff:2002:UOE

[561] Michele Missikoff, Roberto Navigli, and Paola Velardi. The usable ontology: An environment for building and assessing a domain ontology. Lecture Notes in Computer Science, 2342:39– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420039.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420039.
pdf.

Broekstra:2002:SGA

[562] Jeen Broekstra, Arjohn Kampman, and Frank van Harmelen. Sesame: a generic architecture for storing and querying RDF and RDF schema. Lecture Notes in Computer Science, 2342: 54–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420054.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420054.
pdf.

Auillans:2002:FMT

[563] Pascal Auillans, Patrice Ossona de Mendez, Pierre Rosenstiehl, and Bernard Vatant. A formal model for topic maps. Lecture Notes in Computer Science, 2342:69–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420069.
pdf.

Bernstein:2002:THP

[564] Abraham Bernstein and Mark Klein. Towards high-precision service retrieval. Lecture Notes in Computer Science, 2342:84–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2342/23420084.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2342/23420084.pdf.

Eberhart:2002:AGJ

[565] Andreas Eberhart. Automatic generation of Java/SQL based inference engines from RDF schema and RuleML. Lecture Notes in Computer Science, 2342:102–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2342/23420102.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420102.
                                        pdf.

Amann:2002:OBI

[566] Bernd Amann, Catriel Beeri, Irini Fundulaki, and Michel Scholl. Ontologybased integration of XML Web resources. Lecture Notes in Computer Science, 2342:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2342/23420117.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2342/23420117.pdf.

Magkanaraki:2002:BRS

[567] Aimilia Magkanaraki, Sofia Alexaki, Vassilis Christophides, and Dimitris Plexousakis. Benchmarking RDF schemas for the Semantic Web. Lecture Notes in Computer Science, 2342: 132–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420132.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420132.
pdf.

Patel-Schneider:2002:BSW

[568] Peter F. Patel-Schneider and Jérôme Siméon. Building the Semantic Web on XML. Lecture Notes in Computer Science, 2342:147–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420147.
pdf.

Gil:2002:TIS

[569] Yolanda Gil and Varun Ratnakar. Trusting information sources one citizen at a time. Lecture Notes in Computer Science, 2342:162–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420162.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420162.
pdf.

Horrocks:2002:QSW

[570] Ian Horrocks and Sergio Tessaris. Querying the Semantic Web: a formal approach. Lecture Notes in Computer Science, 2342:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420177.
                                        pdf.

Felfernig:2002:SCW

[571] Alexander Felfernig, Gerhard Friedrich, Dietmar Jannach, and Markus Zanker. Semantic configuration Web services in the CAWICOMS project. Lecture Notes in Computer Science, 2342:192– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420192.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420192.
                                        pdf.

Omelayenko:2002:IVD

[572] Borys Omelayenko. Integrating vocabularies: Discovering and representing vocabulary maps. Lecture Notes in Computer Science, 2342:206– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420206.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420206.
                                        pdf.

Sure:2002:OCO

[573] York Sure, Michael Erdmann, Juergen Angele, Steffen Staab, Rudi Studer, and Dirk Wenke. OntoEdit: Collaborative ontology development for the Semantic Web. Lecture Notes in Computer Science, 2342:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420221.
pdf.

Nejdl:2002:TME

[574] Wolfgang Nejdl, Wolf Siberski, Bernd Simon, and Julien Tane. Towards a modification exchange language for distributed RDF repositories. Lecture Notes in Computer Science, 2342:236– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420236.
pdf.

McDermott:2002:RDQ

[575] Drew McDermott and Dejing Dou. Representing disjunction and quantifiers in RDF. Lecture Notes in Computer Science, 2342:250–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420250.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420250.
pdf.

Berendt:2002:TSW

[576] Bettina Berendt, Andreas Hotho, and Gerd Stumme. Towards Semantic Web mining. Lecture Notes in Computer Science, 2342:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420264.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420264.
                                        pdf.

Peer:2002:BTS

[577] Joachim Peer. Bringing together Semantic Web and Web services. Lecture Notes in Computer Science, 2342: 279–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420279.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420279.
                                        pdf.

Hatala:2002:GVC

[578] Marek Hatala and Griff Richards. Global vs. community metadata standards: Empowering users for knowledge exchange. Lecture Notes in Computer Science, 2342:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420292.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420292.
                                        pdf.

Lassila:2002:TRM

[579] Ora Lassila. Taking the RDF model theory out for a spin. Lecture Notes in Computer Science, 2342:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2342/23420307.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420307.
pdf.

Ankolekar:2002:CES

[580] Anupriya Ankolekar, Frank Huch, and Katia Sycara. Concurrent execution semantics of DAML-S with subtypes. Lecture Notes in Computer Science, 2342:318–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420318.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420318.
pdf.

Paolucci:2002:SMW

[581] Massimo Paolucci, Takahiro Kawamura, Terry R. Payne, and Katia Sycara. Semantic matching of Web services capabilities. Lecture Notes in Computer Science, 2342:333–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420333.
pdf.

Ankolekar:2002:DWS

[582] Anupriya Ankolekar, Mark Burstein, Jerry R. Hobbs, Ora Lassila, David Martin, Drew McDermott, Sheila A. McIlraith, Srini Narayanan, Massimo Paolucci, Terry Payne, and Katia Sycara. DAML-S: Web service description for the Semantic Web. Lecture Notes in Computer Science, 2342:348– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420348.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420348.
                                        pdf.

Sintek:2002:TQI

[583] Michael Sintek and Stefan Decker. TRIPLE — A query, inference, and transformation language for the Semantic Web. Lecture Notes in Computer Science, 2342:364–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420364.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420364.
                                        pdf.

Bergamaschi:2002:DIF

[584] Sonia Bergamaschi, Francesco Guerra, and Maurizio Vincini. A data integration framework for e-commerce product classification. Lecture Notes in Computer Science, 2342:379–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420379.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420379.
                                        pdf.

Antoniou:2002:NRS

[585] Grigoris Antoniou. Nonmonotonic rule systems on top of ontology layers. Lecture Notes in Computer Science, 2342: 394–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420394.
pdf.

Seaborne:2002:RN

[586] Andy Seaborne. An RDF NetAPI. Lecture Notes in Computer Science, 2342: 399–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420399.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420399.
pdf.

Schreiber:2002:MES

[587] Guus Schreiber, Inger I. Blok, Daan Carlier, Wouter P. C. van Gent, Jair Hokstam, and Uri Roos. A miniexperiment in semantic annotation. Lecture Notes in Computer Science, 2342:404–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2342/23420404.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2342/23420404.
pdf.

Brickley:2002:SES

[588] D. Brickley, S. Buswell, B. M. Matthews, L. Miller, D. Reynolds, and M. D. Wilson. SWAD-Europe: Semantic Web advanced development in europe. Lecture Notes in Computer Science, 2342:409–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420409.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420409.
                                        pdf.

Newcomb:2002:PR

[589] Steven R. Newcomb. Preemptive reification. Lecture Notes in Computer Science, 2342:414–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420414.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420414.
                                        pdf.

McBride:2002:FST

[590] Brian McBride. Four steps towards the widespread adoption of a Semantic Web. Lecture Notes in Computer Science, 2342:419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420419.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420419.
                                        pdf.

Miller:2002:TIS

[591] Libby Miller, Andy Seaborne, and Alberto Reggiori. Three implementations of SquishQL, a simple RDF query language. Lecture Notes in Computer Science, 2342:423–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420423.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420423.
pdf.

Li:2002:CWS

[592] Gangmin Li, Victoria Uren, Enrico Motta, Simon Buckingham Shum, and John Domingue. ClaiMaker: Weaving a Semantic Web of research papers. Lecture Notes in Computer Science, 2342:436–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420436.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420436.
pdf.

Missikoff:2002:BEO

[593] Michele Missikoff and Francesco Taglino. Business and enterprise ontology management with SymOntoX. Lecture Notes in Computer Science, 2342: 442–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420442.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420442.
pdf.

Haustein:2002:PSW

[594] Stefan Haustein and Jörg Pleumann. Is participation in the Semantic Web too difficult? Lecture Notes in Computer Science, 2342:448–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420448.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420448.
                                        pdf.

Baclawski:2002:CCS

[595] Kenneth Baclawski, Mieczyslaw M. Kokar, Richard Waldinger, and Paul A. Kogut. Consistency checking of Semantic Web ontologies. Lecture Notes in Computer Science, 2342:454–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420454.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420454.
                                        pdf.

Whiting:2002:WUW

[596] Mark A. Whiting and Nick Cramer. WebTheme ^{T M} : Understanding Web information through visual analytics. Lecture Notes in Computer Science, 2342:460–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2342/23420460.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2342/23420460.
                                        pdf.

Payne:2002:BSA

[597] Terry R. Payne, Rahul Singh, and Katia Sycara. Browsing schedules — an agent-based approach to navigating the Semantic Web. Lecture Notes in Computer Science, 2342:469– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2342/23420469.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2342/23420469.
pdf.

Anonymous:2002:AIl

[598] Anonymous. Author index. Lecture Notes in Computer Science, 2342:475– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2342/2342auth.pdf.

Crowcroft:2002:CIR

[599] Jon Crowcroft. Channel islands in a reflective ocean: Large scale event distribution in heterogeneous networks. Lecture Notes in Computer Science, 2345: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450001.
pdf.

Maxemchuk:2002:RMP

[600] Nicholas F. Maxemchuk. A reliable multicast protocol with delay guarantees. Lecture Notes in Computer Science, 2345:10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450010.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450010.
                                        pdf.

Roy:2002:OQB

[601] Abhishek Roy and Sajal K. Das. Optimizing QoS-based multicast routing in wireless networks: a multi-objective genetic algorithmic approach. Lecture Notes in Computer Science, 2345:28– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450028.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450028.
                                        pdf.

Sargento:2002:ESP

[602] Susana Sargento, Roger Salgado, Miguel Carmo, Victor Marques, Rui Valadas, and Edward Knightly. An experimental study of probingbased admission control for Diff-Serv architectures. Lecture Notes in Computer Science, 2345:49–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450049.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450049.
                                        pdf.

Wydrowski:2002:HPD

[603] Bartek Wydrowski and Moshe Zukerman. High performance DiffServ mechanism for routers and switches: Packet arrival rate based queue management for class based scheduling. Lecture Notes in Computer Science, 2345:62– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450062.
pdf.

Mendes:2002:SAP

[604] Paulo Mendes, Henning Schulzrinne, and Edmundo Monteiro. Session-aware popularity resource allocation for assured differentiated services. Lecture Notes in Computer Science, 2345:74– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450074.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450074.
pdf.

Norros:2002:MPP

[605] Ilkka Norros. Most probable path techniques for Gaussian queueing systems. Lecture Notes in Computer Science, 2345:86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2345/23450086.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2345/23450086.
pdf.

Molnar:2002:QTA

[606] Sándor Molnár, Trang Dinh Dang, and István Maricza. On the queue tail asymptotics for general multifractal traffic. Lecture Notes in Computer Science, 2345:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450105.
pdf.

Boxma:2002:SMC

[607] Onno Boxma, Dee Denteneer, and Jacques Resing. Some models for contention resolution in cable networks. Lecture Notes in Computer Science, 2345:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2345/23450117.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2345/23450117.
pdf.

Itao:2002:ACN

[608] Tomoko Itao, Tetsuya Nakamura, Masato Matsuo, Tatsuya Suda, and Tomonori Aoyama. Adaptive creation of network applications in the jack-in-the-net architecture. Lecture Notes in Computer Science, 2345:129– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450129.
pdf.

Blazevic:2002:APD

[609] Ljubica Blažević, Silvia Giordano, and Jean-Yves Le Boudec. Anchored path discovery in terminode routing. Lecture Notes in Computer Science, 2345: 141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450141.
pdf.

Bao:2002:DTS

[610] Lichun Bao and J. J. Garcia-Luna-Aceves. Distributed transmission scheduling using code-division channelization. Lecture Notes in Computer Science, 2345:154–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2345/23450154.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2345/23450154.pdf.

Seres:2002:TED

[611] Gergely Seres, Árpád Szlávik, János Zátonyi, and József Bíró. Towards efficient decision rules for admission control based on the many sources asymptotics. Lecture Notes in Computer Science, 2345:166–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450166.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450166.
pdf.

Jeske:2002:QEB

[612] Daniel R. Jeske, Behrokh Samadi, Kazem Sohraby, Yung-Terng Wang, and Qinqing Zhang. QoS with an edge-based call admission control in IP networks. Lecture Notes in Computer Science, 2345:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450178.
                                        pdf.

Sun:2002:ACC

[613] Yeali S. Sun, Yung-Cheng Tu, and Meng Chang Chen. Admission control and capacity management for advance reservations with uncertain service duration. Lecture Notes in Computer Science, 2345:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450190.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450190.
                                        pdf.

Alexandraki:2002:PED

[614] Adamantia Alexandraki and Michael Paterakis. Performance evaluation of the deadline credit scheduling algorithm for soft-real-time applications in distributed video-on-demand systems. Lecture Notes in Computer Science, 2345:202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2345/23450202.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450202.
pdf.

Balafoutis:2002:IRG

[615] Elias Balafoutis, Antonis Panagakis, Nikolaos Laoutaris, and Ioannis Stavrakakis. The impact of replacement granularity on video caching. Lecture Notes in Computer Science, 2345: 214–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450214.
pdf.

Dube:2002:UAS

[616] Parijat Dube and Eitan Altman. Utility analysis of simple FEC schemes for VoIP. Lecture Notes in Computer Science, 2345:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450226.
pdf.

Anastasi:2002:PSA

[617] Giuseppe Anastasi, Marco Conti, Enrico Gregori, and Andrea Passarella. A power saving architecture for Web access from mobile computers. Lecture Notes in Computer Science, 2345:240– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450240.
                                        pdf.

Okamoto:2002:RCM

[618] Takuya Okamoto, Tatsuhiko Terai, Go Hasegawa, and Masayuki Murata. A resource/connection management scheme for HTTP proxy servers. Lecture Notes in Computer Science, 2345: 252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450252.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450252.
                                        pdf.

Ohsaki:2002:MBM

[619] Hiroyuki Ohsaki, Mitsushige Morita, and Masayuki Murata. Measurementbased modeling of internet round-trip time dynamics using system identification. Lecture Notes in Computer Science, 2345:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2345/23450264.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2345/23450264.pdf.

Pioro:2002:OLC

[620] Michał Pióro, Gábor Malicskó, and Gábor Fodor. Optimal link capacity dimensioning in proportionally fair networks. Lecture Notes in Computer Science, 2345:277–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450277.
pdf.

Brunato:2002:LBW

[621] Mauro Brunato, Roberto Battiti, and Elio Salvadori. Load balancing in WDM networks through adaptive routing table changes. Lecture Notes in Computer Science, 2345:289–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450289.
pdf.

Puech:2002:MLT

[622] Nicolas Puech, Josué Kuri, and Maurice Gagnaire. Models for the logical topology design problem. Lecture Notes in Computer Science, 2345:301– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450301.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450301.
pdf.

Elbiaze:2002:DSS

[623] Halima Elbiaze, Tijani Chahed, Tülin Atmaca, and Gérard Hébuterne. Dynamic shaping for self-similar traffic using network calculus. Lecture Notes in Computer Science, 2345:314– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450314.
                                        pdf.

Bianchi:2002:ACT

[624] Giuseppe Bianchi, Vincenzo Mancuso, and Giovanni Neglia. Is admissioncontrolled traffic self-similar? Lecture Notes in Computer Science, 2345:327– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450327.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450327.
                                        pdf.

Garroppo:2002:ACA

[625] R. G. Garroppo, S. Giordano, S. Lucetti, and M. Pagano. Analysis of CMPP approach in modeling broadband traffic. Lecture Notes in Computer Science, 2345:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450340.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450340.
                                        pdf.

Atov:2002:MMI

[626] Irena Atov and Richard J. Harris. A mathematical model for IP over ATM. Lecture Notes in Computer Science, 2345:352–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450352.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450352.
pdf.

Magoni:2002:ACI

[627] Damien Magoni and Jean-Jacques Pansiot. Analysis and comparison of internet topology generators. Lecture Notes in Computer Science, 2345:364– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450364.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450364.
pdf.

Chiasserini:2002:EED

[628] Carla-Fabiana Chiasserini, Imrich Chlamtac, Paolo Monti, and Antonio Nucci. Energy efficient design of wireless ad hoc networks. Lecture Notes in Computer Science, 2345:376– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450376.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450376.
pdf.

Jacquet:2002:PMR

[629] Philippe Jacquet, Anis Laouiti, Pascale Minet, and Laurent Viennot. Performance of multipoint relaying in ad hoc mobile routing protocols. Lecture Notes in Computer Science, 2345:387– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450387.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450387.
                                        pdf.

Chou:2002:ALA

[630] Zi-Tsan Chou, Ching-Chi Hsu, and Ferng-Ching Lin. An adaptive locationaware MAC protocol for multichannel multihop ad-hoc networks. Lecture Notes in Computer Science, 2345:399– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450399.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450399.
                                        pdf.

Zussman:2002:CAB

[631] Gil Zussman and Adrian Segall. Capacity assignment in Bluetooth scatternets — analysis and algorithms. Lecture Notes in Computer Science, 2345: 411–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450411.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450411.
                                        pdf.

Shapiro:2002:OBC

[632] Jonathan K. Shapiro, Don Towsley, and Jim Kurose. Optimizationbased congestion control for multicast communications. Lecture Notes in Computer Science, 2345:423–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450423.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450423.
pdf.

Csaszar:2002:SCH

[633] András Császár, Attila Takács, Róbert Szabó, Vlora Rexhepi, and Georgios Karagiannis. Severe congestion handling with resource management in Diffserv on demand. Lecture Notes in Computer Science, 2345:443–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450443.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450443.
pdf.

Deb:2002:RAP

[634] Supratim Deb, Ayalvadi Ganesh, and Peter Key. Resource allocation with persistent and transient flows. Lecture Notes in Computer Science, 2345:455– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450455.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450455.
pdf.

Foh:2002:NSM

[635] Chuan Heng Foh and Moshe Zukerman. A novel and simple MAC protocol for high speed passive optical LANs. Lecture Notes in Computer Science, 2345: 467–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450467.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450467.
                                        pdf.

Cassioli:2002:BTS

[636] Dajana Cassioli, Andrea Detti, Pierpaolo Loreti, Franco Mazzenga, and Francesco Vatalaro. The Bluetooth technology: State of the art and networking aspects. Lecture Notes in Computer Science, 2345:479–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450479.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450479.
                                        pdf.

Berno:2002:TFS

[637] Anna Berno and Nicola Laurenti. Time and frequency synchronization for Hiperlan/2. Lecture Notes in Computer Science, 2345:491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450491.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450491.
                                        pdf.

Blondia:2002:PAF

[638] Chris Blondia, Olga Casals, Llorenç Cerdà, and Gert Willems. Performance analysis of a forwarding scheme for handoff in HAWAII. Lecture Notes in Computer Science, 2345:503– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450503.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450503.
pdf.

Rubinstein:2002:EPN

[639] Marcelo G. Rubinstein, Otto Carlos Muniz Bandeira Duarte, and Guy Pujolle. Evaluating the performance of a network management application based on mobile agents. Lecture Notes in Computer Science, 2345:515– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450515.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450515.
pdf.

Ueno:2002:PEW

[640] Hidetoshi Ueno, Norihiro Ishikawa, Hideharu Suzuki, Hiromitsu Sumino, and Osamu Takahashi. Performance evaluation on WAP and internet protocol over 3G wireless networks. Lecture Notes in Computer Science, 2345: 527–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450527.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450527.
                                        pdf.

Freytes:2002:PEH

[641] Matías Freytes. Performance evaluation of H.263-based video transmission in an experimental ad-hoc wireless LAN system. Lecture Notes in Computer Science, 2345:539–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450539.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450539.
                                        pdf.

Fidler:2002:DSB

[642] Markus Fidler. Differentiated services based priority dropping and its application to layered video streams. Lecture Notes in Computer Science, 2345: 551–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450551.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450551.
                                        pdf.

Goncalves:2002:OFQ

[643] Paulo André da Silva Gonçalves, JoséFerreira de Rezende, Otto Carlos Muniz Bandeira Duarte, and Guy Pujolle. Optimal feedback for quality source-adaptive schemes in multicast multi-layered video environments. Lecture Notes in Computer Science, 2345: 563–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450563.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450563.
pdf.

George:2002:FCD

[644] Laurent George, Dana Marinca, and Pascale Minet. A fibre channel dimensioning for a multimedia system with deterministic QoS. Lecture Notes in Computer Science, 2345:575–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450575.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450575.
pdf.

Pentikousis:2002:REE

[645] Kostas Pentikousis and Hussein Badr. On the resource efficiency of explicit congestion notification. Lecture Notes in Computer Science, 2345:588–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450588.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450588.
pdf.

LoCigno:2002:SST

[646] R. Lo Cigno, G. Procissi, and Mario Gerla. Sender-side TCP modifications: An analytical study. Lecture Notes in Computer Science, 2345:600– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450600.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450600.
                                        pdf.

DeVendictis:2002:MMT

[647] Andrea De Vendictis and Andrea Baiocchi. Modeling a mixed TCP Vegas and TCP Reno scenario. Lecture Notes in Computer Science, 2345:612– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450612.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450612.
                                        pdf.

Misra:2002:PSF

[648] Archan Misra and Teunis J. Ott. Performance sensitivity and fairness of ECN-aware ‘modified TCP’. Lecture Notes in Computer Science, 2345:624– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450624.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450624.
                                        pdf.

Huang:2002:CAC

[649] Qian Huang, Hui Min Chen, King Tim Ko, Sammy Chan, and King Sun Chan. Call admission control for 3G CDMA networks with differentiated QoS. Lecture Notes in Computer Science, 2345: 636–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450636.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450636.
pdf.

Ohta:2002:PEC

[650] Yoshiaki Ohta, Kenji Kawahara, Takeshi Ikenaga, and Yuji Oie. Performance evaluation of channel switching scheme for packet data transmission in radio network controller. Lecture Notes in Computer Science, 2345: 648–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450648.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450648.
pdf.

Hu:2002:ORP

[651] Fei Hu and Neeraj K. Sharma. An optimal reservation-pool approach for guaranteeing the call-level QoS in nextgeneration wireless networks. Lecture Notes in Computer Science, 2345:660– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450660.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450660.
pdf.

Xu:2002:NAC

[652] Zhong Xu, Zhenqiang Ye, Srikanth V. Krishnamurthy, Satish K. Tripathi, and Mart Molle. A new adaptive channel reservation scheme for handoff calls in wireless cellular networks. Lecture Notes in Computer Science, 2345:672– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450672.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450672.
                                        pdf.

deLaunois:2002:CES

[653] Cédric de Launois, Aurélien Bonnet, and Marc Lobelle. Connection of extruded subnets: a solution based on RSIP. Lecture Notes in Computer Science, 2345:685–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450685.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450685.
                                        pdf.

Peng:2002:APP

[654] Tao Peng, Christopher Leckie, and Kotagiri Ramamohanarao. Adjusted probabilistic packet marking for IP traceback. Lecture Notes in Computer Science, 2345:697–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450697.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450697.
                                        pdf.

Sousa:2002:TDD

[655] Pedro Sousa, Paulo Carvalho, and Vasco Freitas. Tuning delay differentiation in IP networks using priority queueing models. Lecture Notes in Computer Science, 2345:709–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450709.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450709.
pdf.

Fu:2002:QCH

[656] Xiaoming Fu, Holger Karl, and Cornelia Kappler. QoS-conditionalized handoff for mobile IPv6. Lecture Notes in Computer Science, 2345:721– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450721.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450721.
pdf.

Dube:2002:LPP

[657] Parijat Dube, Omar Ait-Hellal, and Eitan Altman. On loss probabilities in presence of redundant packets with random drop. Lecture Notes in Computer Science, 2345:731–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450731.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450731.
pdf.

Walraevens:2002:PAG

[658] Joris Walraevens, Bart Steyaert, and Herwig Bruneel. Performance analysis of a GI-G-1 preemptive resume priority buffer. Lecture Notes in Computer Science, 2345:745–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450745.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450745.
                                        pdf.

Wittevrongel:2002:ADT

[659] Sabine Wittevrongel, Herwig Bruneel, and Bart Vinck. Analysis of the discrete-time G^{(G)} /Geom/c queueing model. Lecture Notes in Computer Science, 2345:757–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450757.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450757.
                                        pdf.

Stepanenko:2002:TIQ

[660] Alexander Stepanenko, Costas C. Constantinou, Theodoros N. Arvanitis, and Kevin Baughan. On a theory of interacting queues. Lecture Notes in Computer Science, 2345:769– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450769.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2345/23450769.
pdf.

Fantacci:2002:AMP

[661] Romano Fantacci and Giovanni Giambene. Analysis of a MAC protocol for a time-code air interface in LEO mobile satellite systems. Lecture Notes in Computer Science, 2345:778– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450778.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450778.
pdf.

Zaim:2002:PAL

[662] A. Halim Zaim, Harry G. Perros, and George N. Rouskas. Performance analysis of LEO satellite networks. Lecture Notes in Computer Science, 2345: 790–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450790.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450790.
pdf.

Pietrabissa:2002:GAD

[663] Antonio Pietrabissa and Cristiana Santececca. Gateway architecture for DVB-RCS satellite networks. Lecture Notes in Computer Science, 2345:802– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450802.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2345/23450802.
                                        pdf.

Fadel:2002:CAC

[664] Rima Abi Fadel and Samir Tohmé. Connection admission control CAC and differentiated resources allocation RA in a low Earth orbit LEO satellite constellation. Lecture Notes in Computer Science, 2345:814–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450814.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450814.
                                        pdf.

Fan:2002:DBE

[665] Zhong Fan. Dimensioning bandwidth for elastic traffic. Lecture Notes in Computer Science, 2345:826– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450826.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450826.
                                        pdf.

Kamra:2002:FAB

[666] Abhinav Kamra, Huzur Saran, Sandeep Sen, and Rajeev Shorey. Fair adaptive bandwidth allocation: a rate control based active queue management discipline. Lecture Notes in Computer Science, 2345:838–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2345/23450838.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450838.
pdf.

Heikkinen:2002:DSP

[667] Tiina Heikkinen. Distributed scheduling via pricing in a communication network. Lecture Notes in Computer Science, 2345:850–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450850.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450850.
pdf.

Xu:2002:SSA

[668] Lisong Xu, Harry G. Perros, and George N. Rouskas. A simulation study of access protocols for optical burst-switched ring networks. Lecture Notes in Computer Science, 2345:863– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450863.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450863.
pdf.

Rousseau:2002:CED

[669] Bart Rousseau and Fabrice Poppe. Capacity efficiency of distributed path restoration mechanisms in optical mesh networks. Lecture Notes in Computer Science, 2345:875–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450875.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450875.
                                        pdf.

Baldine:2002:HBO

[670] Ilia Baldine, Laura E. Jackson, and George Rouskas. Helios: a broadcast optical architecture. Lecture Notes in Computer Science, 2345:887– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450887.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450887.
                                        pdf.

Tountopoulos:2002:SNM

[671] V. Tountopoulos, V. Stavroulaki, P. Demestichas, N. Mitrou, and M. Theologou. Service and network management interworking in future wireless systems. Lecture Notes in Computer Science, 2345:899–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450899.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450899.
                                        pdf.

Marias:2002:SDT

[672] Giannis F. Marias and Lazaros Merakos. Scheduling differentiated traffic in multicarrier unlicensed systems. Lecture Notes in Computer Science, 2345: 911–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450911.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450911.
pdf.

Kist:2002:SMC

[673] Alexander A. Kist and Richard J. Harris. A simple model for calculating SIP signalling flows in 3GPP IP multimedia subsystems. Lecture Notes in Computer Science, 2345:924– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450924.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450924.
pdf.

Szeto:2002:DOR

[674] W. Szeto, R. Boutaba, and Y. Iraqi. Dynamic online routing algorithm for MPLS traffic engineering. Lecture Notes in Computer Science, 2345:936– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450936.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450936.
pdf.

Bruni:2002:OCP

[675] C. Bruni, C. Scoglio, and S. Vergari. Optimal capacity provisioning for label switched paths in MPLS networks. Lecture Notes in Computer Science, 2345: 947–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450947.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450947.
                                        pdf.

Iliadis:2002:NCO

[676] Ilias Iliadis and Daniel Bauer. A new class of online minimum-interference routing algorithms. Lecture Notes in Computer Science, 2345:959–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450959.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450959.
                                        pdf.

Tachibana:2002:PAD

[677] Takuji Tachibana and Shoji Kasahara. Performance analysis of dynamic lightpath configuration for WDM asymmetric ring networks. Lecture Notes in Computer Science, 2345:972–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23450972.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23450972.
                                        pdf.

Kouvatsos:2002:QMW

[678] D. D. Kouvatsos, K. Al-Begain, and I. Awan. A queueing model for a wireless GSM/GPRS cell with multiple service classes. Lecture Notes in Computer Science, 2345:984–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450984.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450984.
pdf.

Chan:2002:IMP

[679] Hung Nguyen Chan, Belen Carro Martinez, Rafa Mompo Gomez, and Judith Redoli Granados. Integrated multi-purposed testbed to characterize the performance of internet access over hybrid fiber coaxial access networks. Lecture Notes in Computer Science, 2345:996–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23450996.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23450996.
pdf.

Vishnevsky:2002:LST

[680] Vladimir Vishnevsky and Andrey Lyakhov. 802.11 LANs: Saturation throughput in the presence of noise. Lecture Notes in Computer Science, 2345:1008–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451008.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451008.
pdf.

Karvo:2002:ESB

[681] Jouni Karvo. Efficient simulation of blocking probabilities for multi-layer multicast streams. Lecture Notes in Computer Science, 2345:1020–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451020.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451020.
                                        pdf.

Cui:2002:AMC

[682] Jun-Hong Cui, Jinkyu Kim, Dario Maggiorini, Khaled Boussetta, and Mario Gerla. Aggregated multicast — A comparative study. Lecture Notes in Computer Science, 2345:1032– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451032.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451032.
                                        pdf.

Shim:2002:NCL

[683] Young-Chul Shim and Shin-Kyu Kang. New center location algorithms for shared multicast trees. Lecture Notes in Computer Science, 2345:1045–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451045.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451045.
                                        pdf.

Bonuccelli:2002:MFO

[684] Maurizio A. Bonuccelli and Alessandro Urpi. A multicast FCFS output queued switch without speedup. Lecture Notes in Computer Science, 2345:1057–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451057.
pdf.

Anastasi:2002:FTS

[685] Giuseppe Anastasi, Alberto Bartoli, and Flaminia L. Luccio. Fault-tolerant support for reliable multicast in mobile wireless systems. Lecture Notes in Computer Science, 2345:1069–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451069.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451069.
pdf.

Baldine:2002:JJT

[686] Ilia Baldine, Harry Perros, George N. Rouskas, and Dan Stevenson. Jump-Start: a just-in-time signaling architecture for WDM burst-switched networks. Lecture Notes in Computer Science, 2345:1081–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2345/23451081.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2345/23451081.pdf.

Basagni:2002:DDB

[687] Stefano Basagni, Raffaele Bruno, and Chiara Petrioli. Device discovery in Bluetooth networks: a scatternet perspective. Lecture Notes in Computer Science, 2345:1087–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451087.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451087.
                                        pdf.

Carrozzo:2002:QER

[688] G. Carrozzo, V. Chionsini, S. Giordano, and S. Niccolini. QoS evaluation of real-time applications over a multi-domain DiffServ experimental test-bed. Lecture Notes in Computer Science, 2345:1093–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451093.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451093.
                                        pdf.

Chaouchi:2002:NPB

[689] Hakima Chaouchi and Guy Pujolle. A new policy based management of mobile IP users. Lecture Notes in Computer Science, 2345:1099–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451099.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2345/23451099.
pdf.

Cremonese:2002:FPB

[690] P. Cremonese, M. Esposito, S. Giordano, M. Mondini, S. P. Romano, and G. Ventre. A framework for policy-based management of QoS aware IP networks. Lecture Notes in Computer Science, 2345:1105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451105.
pdf.

DeMarco:2002:SHS

[691] G. De Marco, S. Loreto, G. Sorrentino, and L. Veltri. SIP-H323: a solution for interworking saving existing architecture. Lecture Notes in Computer Science, 2345:1111–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451111.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451111.
pdf.

Duret:2002:HRF

[692] Christian Duret, Francis Rischette, Joël Lattmann, Valéry Laspreses, Pim Van Heuven, Steven Van den Berghe, and Piet Demeester. High router flexibility and performance by combining dedicated lookup hardware (IFT), off the shelf switches and Linux. Lecture Notes in Computer Science, 2345:1117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451117.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451117.
                                        pdf.

Hardjono:2002:GSP

[693] Thomas Hardjono and Hugh Harney. Group security policy management for IP multicast and group security. Lecture Notes in Computer Science, 2345: 1123–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2345/23451123.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451123.
                                        pdf.

Ichikawa:2002:IIR

[694] Yasushi Ichikawa, Kensuke Arakawa, Keisuke Wano, and Yuko Murayama. Issues in internet radio. Lecture Notes in Computer Science, 2345:1129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451129.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451129.
                                        pdf.

Lepe-Aldama:2002:BUC

[695] Oscar-Iván Lepe-Aldama and Jorge García-Vidal. I/O bus usage control in PC-based software routers. Lecture Notes in Computer Science, 2345:1135– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451135.
pdf.

Konorski:2002:MAA

[696] Jerzy Konorski. Multiple access in ad-hoc wireless LANs with noncooperative stations. Lecture Notes in Computer Science, 2345:1141–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451141.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451141.
pdf.

Kos:2002:NGN

[697] Andrej Kos, Janez Bešter, and Peter Homan. Next generation networks and services in Slovenia. Lecture Notes in Computer Science, 2345:1147– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451147.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451147.
pdf.

Liang:2002:MRD

[698] Ben Liang and Zygmunt J. Haas. Minimizing the routing delay in ad hoc networks through route-cache TTL optimization. Lecture Notes in Computer Science, 2345:1153–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451153.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451153.
                                        pdf.

Lima:2002:LRD

[699] Solange Lima, Magda Silva, Paulo Carvalho, Alexandre Santos, and Vasco Freitas. Long-range dependence of internet traffic aggregates. Lecture Notes in Computer Science, 2345:1159– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451159.
                                        pdf.

Lomi:2002:IIS

[700] Valentina Lomi, Gianfranco L. Pierobon, Daniele Tonetto, and Lorenzo Vangelista. Improved initial synchronisation in the presence of frequency offset in UMTS FDD mode. Lecture Notes in Computer Science, 2345:1165– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451165.
                                        pdf.

Mathy:2002:SAH

[701] Laurent Mathy, Roberto Canonico, Steven Simpson, and David Hutchison. Scalable adaptive hierarchical clustering. Lecture Notes in Computer Science, 2345:1172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451172.
pdf.

Nyberg:2002:HAF

[702] Eeva Nyberg and Samuli Aalto. How to achieve fair differentiation. Lecture Notes in Computer Science, 2345:1178– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451178.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451178.
pdf.

Pagani:2002:MBA

[703] Elena Pagani and Gian Paolo Rossi. Measurement-based admission control for dynamic multicast groups in diff-serv networks. Lecture Notes in Computer Science, 2345:1184–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451184.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451184.
pdf.

Papadopoulou:2002:FSN

[704] L.-M. Papadopoulou, V. Stavroulaki, P. Demestichas, and M. Theologou.

A framework to service and network resource management in composite radio environments. Lecture Notes in Computer Science, 2345:1190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451190.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451190.
                                        pdf.

Preuss:2002:JSD

[705] Stephan Preuß. JESA service discovery protocol. Lecture Notes in Computer Science, 2345:1196–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451196.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451196.
                                        pdf.

Raatikainen:2002:PSQ

[706] Pertti Raatikainen, Mika Wikström, and Timo Hämäläinen. Performance simulations of a QoS aware caching method. Lecture Notes in Computer Science, 2345:1202–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451202.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451202.
                                        pdf.

Senouci:2002:CAC

[707] Sidi-Mohammed Senouci, André-Luc Beylot, and Guy Pujolle. Call admission control for multimedia cellular networks using neuro-dynamic programming. Lecture Notes in Computer Science, 2345:1208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451208.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451208.
pdf.

Speer:2002:AAS

[708] Anke Speer, Marcus Schöller, Thomas Fuhrmann, and Martina Zitterbart. Aspects of AMnet signaling. Lecture Notes in Computer Science, 2345:1214– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451214.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451214.
pdf.

Tomarchio:2002:VHE

[709] Orazio Tomarchio, Andrea Calvagna, and Giuseppe Di Modica. Virtual home environment for multimedia services in 3rd generation networks. Lecture Notes in Computer Science, 2345:1221– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451221.
pdf.

Tsolakou:2002:PEE

[710] E. Tsolakou, E. Nikolouzou, and S. Venieris. On providing end-to-end QoS introducing a set of network services in large-scale IP networks. Lecture Notes in Computer Science, 2345:1227– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451227.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451227.
                                        pdf.

Velenis:2002:STP

[711] Dimitris Velenis, Dimitris Kalogeras, and Basil Maglaris. SaTPEP: a TCP performance enhancing proxy for satellite links. Lecture Notes in Computer Science, 2345:1233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451233.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451233.
                                        pdf.

Yeo:2002:OUS

[712] Chai Kiat Yeo, Bu Sung Lee, and Meng Hwa Er. An overlay for ubiquitous streaming over internet. Lecture Notes in Computer Science, 2345:1239– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2345/23451239.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2345/23451239.
                                        pdf.

Zander:2002:MBD

[713] Roland Zander and Johan M. Karlsson. A measurement-based dynamic guard channel scheme for handover prioritization in cellular networks. Lecture Notes in Computer Science, 2345:1245– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2345/23451245.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2345/23451245.
pdf.

Anonymous:2002:AIm

[714] Anonymous. Author index. Lecture Notes in Computer Science, 2345:1253– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2345/2345auth.pdf.

Aaronson:2002:PAA

[715] Jack Aaronson. Is personalization all about technology? Lecture Notes in Computer Science, 2347:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470001.
pdf.

Lieberman:2002:ALB

[716] Henry Lieberman and Hugo Liu. Adaptive linking between text and photos using common sense reasoning. Lecture Notes in Computer Science, 2347:

2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470002.
pdf.

Wahlster:2002:RAI

[717] Wolfgang Wahlster. Resourceadaptive interfaces to hybrid navigation systems. Lecture Notes in Computer Science, 2347:12–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470012.
pdf.

Ardissono:2002:UUA

[718] Liliana Ardissono, Anna Goy, Giovanna Petrone, Marino Segnan, and Pietro Torasso. Ubiquitous user assistance in a tourist information server. Lecture Notes in Computer Science, 2347:14–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470014.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470014.
pdf.

Arjona:2002:AES

[719] J. L. Arjona, R. Corchuelo, A. Ruiz, and M. Toro. Automatic extraction of semantically-meaningful information from the Web. Lecture Notes in Computer Science, 2347:24– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470024.
pdf.

Bailey:2002:TOA

[720] Christopher Bailey, Wendy Hall, David E. Millard, and Mark J. Weal. Towards open adaptive hypermedia. Lecture Notes in Computer Science, 2347:36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2347/23470036.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2347/23470036.
pdf.

Barra:2002:GGA

[721] Maria Barra, Paul Maglio, Alberto Negro, and Vittorio Scarano. GAS: Group adaptive system. Lecture Notes in Computer Science, 2347:47– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470047.
pdf.

Baudisch:2002:TSL

[722] Patrick Baudisch and Lars Brueckner. TV scout: Lowering the entry barrier to personalized TV program recommendation. Lecture Notes in Computer Science, 2347:58–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470058.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470058.
                                        pdf.

Bontcheva:2002:AAR

[723] Kalina Bontcheva. Adaptivity, adaptability, and reading behaviour: Some results from the evaluation of a dynamic hypertext system. Lecture Notes in Computer Science, 2347:69– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470069.
                                        pdf.

Calvi:2002:TGA

[724] Licia Calvi and Alexandra Cristea. Towards generic adaptive systems: Analysis of a case study. Lecture Notes in Computer Science, 2347:79– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470079.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470079.
                                        pdf.

Carro:2002:MDA

[725] Rosa M. Carro, Ana M. Breda, Gladys Castillo, and Antonio L. Bajuelos. A methodology for developing adaptive educational-game environments. Lecture Notes in Computer Science, 2347: 90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470090.
pdf.

Conlan:2002:MMM

[726] Owen Conlan, Vincent Wade, Catherine Bruen, and Mark Gargan. Multimodel, metadata driven approach to adaptive hypermedia services for personalized eLearning. Lecture Notes in Computer Science, 2347:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470100.
pdf.

Console:2002:APB

[727] L. Console, S. Gioria, I. Lombardi, V. Surano, and I. Torre. Adaptation and personalization on board cars: a framework and its application to tourist services. Lecture Notes in Computer Science, 2347:112– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470112.
pdf.

Cristea:2002:AAA

[728] Alexandra Cristea and Lora Aroyo. Adaptive authoring of adaptive educational hypermedia. Lecture Notes in Computer Science, 2347:122–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470122.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470122.
                                        pdf.

Frasincar:2002:HPA

[729] Flavius Frasincar and Geert-Jan Houben. Hypermedia presentation adaptation on the Semantic Web. Lecture Notes in Computer Science, 2347: 133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470133.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470133.
                                        pdf.

Gaudioso:2002:UDM

[730] Elena Gaudioso and Jesus G. Boticario. User data management and usage model acquisition in an adaptive educational collaborative environment. Lecture Notes in Computer Science, 2347: 143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470143.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470143.
                                        pdf.

Gouli:2002:PAA

[731] Evangelia Gouli, Kyparisia Papanikolaou, and Maria Grigoriadou. Personalizing assessment in adaptive educational hypermedia systems. Lecture Notes in Computer Science, 2347:153– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470153.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470153.
pdf.

Gu:2002:VBC

[732] Xiao-Dong Gu, Jinlin Chen, Wei-Ying Ma, and Guo-Liang Chen. Visual based content understanding towards Web adaptation. Lecture Notes in Computer Science, 2347:164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470164.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470164.
pdf.

Henze:2002:KMO

[733] Nicola Henze and Wolfgang Nejdl. Knowledge modeling for open adaptive hypermedia. Lecture Notes in Computer Science, 2347:174–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470174.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470174.
pdf.

Hubscher:2002:ANL

[734] Roland Hübscher and Sadhana Puntambekar. Adaptive navigation for learners in hypermedia is scaffolded navigation. Lecture Notes in Computer Science, 2347:184–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470184.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470184.
                                        pdf.

Jameson:2002:PCC

[735] Anthony Jameson and Eric Schwarzkopf. Pros and cons of controllability: An empirical study. Lecture Notes in Computer Science, 2347:193–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470193.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470193.
                                        pdf.

Kay:2002:PSU

[736] Judy Kay, Bob Kummerfeld, and Piers Lauder. Personis: a server for user models. Lecture Notes in Computer Science, 2347:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470203.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470203.
                                        pdf.

Koch:2002:MRM

[737] Nora Koch and Martin Wirsing. The munich reference model for adaptive hypermedia applications. Lecture Notes in Computer Science, 2347:213– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470213.
pdf.

Koychev:2002:TCU

[738] Ivan Koychev. Tracking changing user interests through priorlearning of context. Lecture Notes in Computer Science, 2347:223–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470223.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470223.
pdf.

Kukulenz:2002:PNP

[739] Dirk Kukulenz. Prediction of navigation profiles in a distributed internet environment through learning of graph distributions. Lecture Notes in Computer Science, 2347:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470233.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470233.
pdf.

Kurhila:2002:DCL

[740] Jaakko Kurhila, Miikka Miettinen, Petri Nokelainen, and Henry Tirri. E DUCO — A collaborative learning environment based on social navigation. Lecture Notes in Computer Science, 2347:242–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470242.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470242.
                                        pdf.

Liu:2002:GGO

[741] Hugo Liu, Henry Lieberman, and Ted Selker. GOOSE: a goal-oriented search engine with commonsense. Lecture Notes in Computer Science, 2347:253– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470253.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470253.
                                        pdf.

Macias:2002:AWS

[742] Mercedes Macías, Julia González, and Fernando Sánchez. On adaptability of Web sites for visually handicapped people. Lecture Notes in Computer Science, 2347:264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470264.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2347/23470264.
pdf.

Martins:2002:FFP

[743] Lucimar C. Martins, Tatiana A. S. Coelho, Simone D. J. Barbosa, Marco A. Casanova, and Carlos J. P. de Lucena. A framework for filtering and packaging hypermedia documents. Lecture Notes in Computer Science, 2347:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470274.
pdf.

Medina-Medina:2002:AEH

[744] Nuria Medina-Medina, Lina García-Cabrera, M. José Rodríguez-Fortiz, and José Parets-Llorca. Adaptation in an evolutionary hypermedia system: Using semantic and petri nets. Lecture Notes in Computer Science, 2347: 284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470284.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470284.
pdf.

Mitrovic:2002:EEO

[745] Antonija Mitrovic and Brent Martin. Evaluating the effects of open student models on learning. Lecture Notes in Computer Science, 2347:296– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470296.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470296.
                                        pdf.

Mizzaro:2002:EPP

[746] Stefano Mizzaro and Carlo Tasso. Ephemeral and persistent personalization in adaptive information access to scholarly publications on the Web. Lecture Notes in Computer Science, 2347: 306–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470306.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470306.
                                        pdf.

Sicilia:2002:FLS

[747] Miguel-Ángel Sicilia, Paloma Díaz, Ignacio Aedo, and Elena García. Fuzzy linguistic summaries in rule-based adaptive hypermedia systems. Lecture Notes in Computer Science, 2347:317– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470317.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470317.
                                        pdf.

Smyth:2002:PNS

[748] Barry Smyth and Paul Cotter. The plight of the navigator: Solving the navigation problem for wireless portals. Lecture Notes in Computer Science, 2347:328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470328.
pdf.

Souto:2002:TAW

[749] Maria A. M. Souto, Regina Verdin, Ricardo Wainer, Milton Madeira, Mariusa Warpechowski, Karine Beschoren, Renata Zanella, Juarez Sagebin Correa, Rosa M. Vicari, and José Palazzo M. de Oliveira. Towards an adaptive Web training environment based on cognitive style of learning: An empirical approach. Lecture Notes in Computer Science, 2347:338– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470338.
pdf.

Sunderam:2002:API

[750] Aditya V. Sunderam. Automated personalization of internet news. Lecture Notes in Computer Science, 2347:348– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470348.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470348.
pdf.

ao:2002:CMP

[751] Silvia Abrah ao, Joan Fons, Magalí González, and Oscar Pastor. Conceptual modeling of personalized Web applications. Lecture Notes in Computer Science, 2347:358–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470358.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470358.
                                        pdf.

Arruabarrena:2002:EAS

[752] Rosa Arruabarrena, Tomás A. Pérez, J. López-Cuadrado, J. Gutiérrez, and J. A. Vadillo. On evaluating adaptive systems for education. Lecture Notes in Computer Science, 2347:363– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470363.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470363.
                                        pdf.

Berka:2002:RID

[753] Tobias Berka, Wernher Behrendt, Erich Gams, and Siegfried Reich. Recommending internet-domains using trails and neural networks. Lecture Notes in Computer Science, 2347:368– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470368.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2347/23470368.
pdf.

Brunstein:2002:LGA

[754] Angela Brunstein, Jacqueline Waniek, Anja Naumann, and Josef F. Krems. Learning grammar with adaptive hypertexts: Reading or searching? Lecture Notes in Computer Science, 2347: 372–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470372.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470372.
pdf.

Carmona:2002:SMW

[755] Cristina Carmona, David Bueno, Eduardo Guzman, and Ricardo Conejo. SIGUE: Making Web courses adaptive. Lecture Notes in Computer Science, 2347:376–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470376.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470376.
pdf.

Coffey:2002:OGA

[756] John W. Coffey. An ontologyguided approach to content adaptation in LEO: a learning environment organizer. Lecture Notes in Computer Science, 2347:380–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2347/23470380.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470380.
                                        pdf.

Czarkowski:2002:SAH

[757] Marek Czarkowski and Judy Kay. A scrutable adaptive hypertext. Lecture Notes in Computer Science, 2347:384– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470384.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470384.
                                        pdf.

DeBra:2002:AMA

[758] Paul De Bra, Ad Aerts, David Smits, and Natalia Stash. AHA! meets AHAM. Lecture Notes in Computer Science, 2347:388–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470388.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470388.
                                        pdf.

Esuli:2002:ART

[759] Andrea Esuli, Antonio Cisternino, Giuliano Pacini, and Maria Simi. Adaptive real time comment generation for sail racing scenarios. Lecture Notes in Computer Science, 2347:392– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470392.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2347/23470392.
pdf.

Grimsley:2002:WBS

[760] Michael Grimsley and Anthony Meehan. A Web-based selling agent that maintains customer loyalty through integrative negotiation. Lecture Notes in Computer Science, 2347:397–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470397.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470397.
pdf.

Healey:2002:ACD

[761] Jennifer Healey, Rafah Hosn, and Stéphane H. Maes. Adaptive content for device independent multimodal browser applications. Lecture Notes in Computer Science, 2347:401– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470401.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470401.
pdf.

Kanawati:2002:BSN

[762] Rushed Kanawati and Maria Malek. Behavioral sequences: a new logcoding scheme for effective prediction of Web user accesses. Lecture Notes in Computer Science, 2347:406– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470406.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470406.
                                        pdf.

Martinovska:2002:FBA

[763] Cveta Martinovska. A fuzzy-based approach to user model refinement in adaptive hypermedia systems. Lecture Notes in Computer Science, 2347:411– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470411.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470411.
                                        pdf.

Masthoff:2002:TAC

[764] Judith Masthoff. Towards an authoring coach for adaptive Webbased instruction. Lecture Notes in Computer Science, 2347:415–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470415.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470415.
                                        pdf.

Montero:2002:GPW

[765] Susana Montero, Ignacio Aedo, and Paloma Díaz. Generation of personalized Web courses using RBAC. Lecture Notes in Computer Science, 2347: 419–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2347/23470419.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470419.
pdf.

Mun:2002:ART

[766] Hyeonjeong Mun, Sooho Ok, and Yongtae Woo. An automatic rating technique based on XML document. Lecture Notes in Computer Science, 2347: 424–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470424.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470424.
pdf.

Ng:2002:UER

[767] Muan Hong Ng, Wendy Hall, Pat Maier, and Ray Armstrong. Using effective reading speed to integrate adaptivity into Web-based learning. Lecture Notes in Computer Science, 2347:428– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470428.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470428.
pdf.

Roselli:2002:NIC

[768] Teresa Roselli, Enrica Gentile, Paola Plantamura, Veronica Rossano, and Vittorio Saponaro. A non-invasive cooperative student model. Lecture Notes in Computer Science, 2347:432– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470432.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470432.
                                        pdf.

Sollenborn:2002:CBF

[769] Mikael Sollenborn and Peter Funk. Category-based filtering in recommender systems for improved performance in dynamic domains. Lecture Notes in Computer Science, 2347:436– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470436.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470436.
                                        pdf.

Spalka:2002:PUD

[770] Adrian Spalka and Hanno Langweg. Protecting the user from the data: Security and privacy aspects of public Web access. Lecture Notes in Computer Science, 2347:440–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470440.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470440.
                                        pdf.

Villanova-Oliver:2002:EPA

[771] Marlène Villanova-Oliver, Jérôme Gensel, and Hervé Martin. Exploiting a progressive access model to information. Lecture Notes in Computer Science, 2347:444–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470444.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470444.
pdf.

Weibelzahl:2002:APK

[772] Stephan Weibelzahl and Gerhard Weber. Adapting to prior knowledge of learners. Lecture Notes in Computer Science, 2347:448–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470448.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470448.
pdf.

Yamada:2002:MPU

[773] Seiji Yamada and Yuki Nakai. Monitoring partial updates in Web pages using relational learning. Lecture Notes in Computer Science, 2347:452– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470452.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470452.
pdf.

Abraham:2002:AWB

[774] David Abraham and Kalina Yacef. Adaptation in the Web-based logic-ITA. Lecture Notes in Computer Science, 2347:456–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2347/23470456.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2347/23470456.pdf.

Avesani:2002:CRC

[775] Paolo Avesani, Paolo Massa, Michele Nori, and Angelo Susi. Collaborative radio community. Lecture Notes in Computer Science, 2347:462–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470462.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470462.
                                        pdf.

Silveira:2002:IIL

[776] Ricardo Azambuja Silveira and Rosa Maria Vicari. Improving interactivity in e-learning systems with multiagent architecture. Lecture Notes in Computer Science, 2347:466–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470466.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470466.
                                        pdf.

Bry:2002:AHM

[777] François Bry and Michael Kraus. Adaptive hypermedia made simple with HTML/XML style sheet selectors. Lecture Notes in Computer Science, 2347:472–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2347/23470472.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2347/23470472.
pdf.

Buendia:2002:FEA

[778] Félix Buendía and Paloma Díaz. A framework for educational adaptive hypermedia applications. Lecture Notes in Computer Science, 2347:476–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470476.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470476.
pdf.

Bueno:2002:MPR

[779] David Bueno, Ricardo Conejo, Cristina Carmona, and Amos A. David. METIORE: a publications reference for the adaptive hypermedia community. Lecture Notes in Computer Science, 2347:480–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470480.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470480.
pdf.

Carrara:2002:UDA

[780] P. Carrara, D. Fogli, G. Fresta, and P. Mussio. User-driven adaptation of e-documents. Lecture Notes in Computer Science, 2347:484–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470484.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470484.
                                        pdf.

Cassel:2002:SIP

[781] Lillian N. Cassel, Ursula Wolz, and Robert E. Beck. Server independent personalization for effective searching in the World Wide Web. Lecture Notes in Computer Science, 2347:488– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470488.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470488.
                                        pdf.

Cawsey:2002:PMP

[782] Alison Cawsey, Euan Dempster, Diana Bental, Daniel Pacey, Howard Williams, Lachlan MacKinnon, and David Marwick. Preventing misleading presentations of XML documents: Some initial proposals. Lecture Notes in Computer Science, 2347: 492–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470492.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470492.
                                        pdf.

Cini:2002:ACE

[783] Alessandra Cini and José Valdeni de Lima. Adaptivity conditions evaluation for the user of hypermedia presentations built with AHA! Lecture Notes in Computer Science, 2347:497– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470497.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470497.
pdf.

FernandezdeArriba:2002:DAW

[784] Marta Fernández de Arriba and José A. López Brugos. Development of adaptive Web sites with usability and accessibility features. Lecture Notes in Computer Science, 2347:501–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470501.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470501.
pdf.

Garcia:2002:ACS

[785] Francisco José García, Fabio Paternò, and Ana Belén Gil. An adaptive ecommerce system definition. Lecture Notes in Computer Science, 2347:505– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470505.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470505.
pdf.

Goel:2002:WSP

[786] Mohit Goel and Sudeshna Sarkar. Web site personalization using user profile information. Lecture Notes in Computer Science, 2347:510–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470510.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470510.
                                        pdf.

McGowan:2002:WDY

[787] J. P. McGowan, Nicholas Kushmerick, and Barry Smyth. Who do you want to be today? Web personae for personalised information access. Lecture Notes in Computer Science, 2347:514– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470514.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470514.
                                        pdf.

Kashihara:2002:ANP

[788] Akihiro Kashihara, Shinobu Hasegawa, and Jun’ichi Toyoda. Adaptive navigation path previewing for learning on the Web. Lecture Notes in Computer Science, 2347:518–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470518.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470518.
                                        pdf.

Kim:2002:CBR

[789] Youngji Kim, Sooho Ok, and Yongtae Woo. A case-based recommender system using implicit rating techniques. Lecture Notes in Computer Science, 2347:522–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470522.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470522.
pdf.

Kraus:2002:IIM

[790] Sarit Kraus, Alexander Kröner, and Lea Tsaban. IMAP — intelligent multimedia authoring tools for electronic publishing. Lecture Notes in Computer Science, 2347:527–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470527.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470527.
pdf.

Lee:2002:HRS

[791] Meehee Lee, Pyungseok Choi, and Yongtae Woo. A hybrid recommender system combining collaborative filtering with neural network. Lecture Notes in Computer Science, 2347:531– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470531.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470531.
pdf.

Martinez-Enriquez:2002:ACW

[792] Ana María Martínez-Enríquez, Dominique Decouchant, Alberto L. Morán, and Jesus Favela. An adaptive cooperative Web authoring environment. Lecture Notes in Computer Science, 2347:535–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470535.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470535.
                                        pdf.

Martens:2002:ALE

[793] Alke Martens and Adelinde M. Uhrmacher. Adapting learner evaluation by plans. Lecture Notes in Computer Science, 2347:539–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470539.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470539.
                                        pdf.

Martin:2002:WWB

[794] Brent Martin and Antonija Mitrovic. WETAS: a Web-based authoring system for constraint-based ITS. Lecture Notes in Computer Science, 2347:543– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470543.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470543.
                                        pdf.

Mitsuhara:2002:OEA

[795] Hiroyuki Mitsuhara, Youji Ochi, and Yoneo Yano. Open-ended adaptive system for facilitating knowledge construction in Web-based exploratory learning. Lecture Notes in Computer Science, 2347:547–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470547.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470547.
pdf.

Ng:2002:AWI

[796] Wilfred Ng and Jerry Yau. Adapting Web interfaces by WHAT. Lecture Notes in Computer Science, 2347:551– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470551.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470551.
pdf.

Paredes:2002:CSI

[797] Pedro Paredes and Pilar Rodriguez. Considering sensing-intuitive dimension to exposition-exemplification in adaptive sequencing. Lecture Notes in Computer Science, 2347:556–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470556.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470556.
pdf.

Redondo:2002:PAP

[798] M. A. Redondo, C. Bravo, M. Ortega, and M. F. Verdejo. PlanEdit: An adaptive problem solving tool for design. Lecture Notes in Computer Science, 2347:560–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2347/23470560.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470560.
                                        pdf.

Rodriguez:2002:DGI

[799] Martín González Rodríguez, Benjamin López Pérez, María del Puerto Paule Ruíz, and Juan Ramón Pérez Pérez. Dynamic generation of interactive dialogs based on intelligent agents. Lecture Notes in Computer Science, 2347:564–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470564.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470564.
                                        pdf.

Signore:2002:SAA

[800] Oreste Signore. A simple architecture for adaptive hypertext. Lecture Notes in Computer Science, 2347:568– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2347/23470568.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470568.
                                        pdf.

Specht:2002:ALE

[801] Marcus Specht, Milos Kravcik, Roland Klemke, Leonid Pesin, and Rüdiger Hüttenhain. Adaptive learning environment for teaching and learning in WINDS. Lecture Notes in Computer Science, 2347:572–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470572.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470572.
pdf.

Toppano:2002:ATI

[802] Elio Toppano. Adaptation through interaction modelling in educational hypermedia. Lecture Notes in Computer Science, 2347:576–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2347/23470576.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2347/23470576.pdf.

Zakaria:2002:HMA

[803] Mohamed Ramzy Zakaria, Adam Moore, Helen Ashman, Craig Stewart, and Tim Brailsford. The hybrid model for adaptive educational hypermedia. Lecture Notes in Computer Science, 2347:580–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470580.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2347/23470580.
                                        pdf.

Dolog:2002:NMA

[804] Peter Dolog and Mária Bieliková. Navigation modelling in adaptive hypermedia. Lecture Notes in Computer Science, 2347:586–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2347/23470586.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2347/23470586.pdf.

Gil:2002:AAM

[805] Ana B. Gil, Francisco García, and Zahia Guessoum. An adaptive agent model for e-commerce architecture. Lecture Notes in Computer Science, 2347:592–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2347/23470592.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2347/23470592.
                                        pdf.

Herder:2002:PAD

[806] Eelco Herder and Betsy van Dijk. Personalized adaptation to device characteristics. Lecture Notes in Computer Science, 2347:598–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2347/23470598.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2347/23470598.pdf.

Torre:2002:UMA

[807] Ilaria Torre. Users modeling for adaptive call centers. Lecture Notes in Computer Science, 2347:603–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470603.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470603.
pdf.

Yong:2002:EPA

[808] Lim Tek Yong and Tang Enya Kong. The exploratory of personal assistants. Lecture Notes in Computer Science, 2347:608–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2347/23470608.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2347/23470608.
pdf.

Anonymous:2002:AIn

[809] Anonymous. Author index. Lecture Notes in Computer Science, 2347:613– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2347/2347auth.pdf.

Brodie:2002:GCI

[810] Michael L. Brodie. The grand challenge in information technology and the illusion of validity. Lecture Notes in Computer Science, 2348:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480001.
pdf.

Jarke:2002:MCKa

[811] Matthias Jarke. Metadata and cooperative knowledge management. Lecture Notes in Computer Science, 2348: 2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480002.
pdf.

Welty:2002:ODC

[812] Christopher Welty. Ontology-driven conceptual modeling. Lecture Notes in Computer Science, 2348:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480003.
pdf.

Jarke:2002:MCKb

[813] Matthias Jarke and Ralf Klamma. Metadata and cooperative knowledge management. Lecture Notes in Computer Science, 2348:4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

bibs/2348/23480004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480004.
pdf.

Yang:2002:WCS

[814] Jian Yang and Mike. P. Papazoglou. Web component: a substrate for Web service reuse and composition. Lecture Notes in Computer Science, 2348:21– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480021.
pdf.

Liu:2002:DWB

[815] Lin Liu and Eric Yu. Designing Webbased systems in social context: a goal and scenario based approach. Lecture Notes in Computer Science, 2348:37– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480037.
pdf.

Mohan:2002:SMB

[816] Rakesh Mohan, Mitchell A. Cohen, and Josef Schiefer. A state machine based approach for a process driven development of Web-applications. Lecture Notes in Computer Science, 2348:52– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2348/23480052.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480052.
                                        pdf.

Vaisman:2002:SDU

[817] Alejandro A. Vaisman, Alberto O. Mendelzon, Walter Ruaro, and Sergio G. Cymerman. Supporting dimension updates in an OLAP server. Lecture Notes in Computer Science, 2348: 67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480067.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480067.
                                        pdf.

Eder:2002:CMT

[818] Johann Eder, Christian Koncilia, and Tadeusz Morzy. The COMET metamodel for temporal data warehouses. Lecture Notes in Computer Science, 2348:83–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2348/23480083.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480083.
                                        pdf.

Liu:2002:ERE

[819] Ping Liu, Jayne Curson, and Peter Dew. Exploring RDF for expertise matching within an organizational memory. Lecture Notes in Computer Science, 2348:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2348/23480100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480100.
pdf.

Smolander:2002:DCS

[820] Kari Smolander and Tero Päivärinta. Describing and communicating software architecture in practice: Observations on stakeholders and rationale. Lecture Notes in Computer Science, 2348:117–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2348/23480117.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2348/23480117.
pdf.

Huisman:2002:IDS

[821] Magda Huisman and Juhani Iivari. The individual deployment of systems development methodologies. Lecture Notes in Computer Science, 2348:134– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480134.
pdf.

Hou:2002:SDO

[822] Daqing Hou, H. James Hoover, and Eleni Stroulia. Supporting the deployment of object-oriented frameworks. Lecture Notes in Computer Science, 2348:151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2348/23480151.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480151.
                                        pdf.

Brasethvik:2002:CMA

[823] Terje Brasethvik and Jon Atle Gulla. A conceptual modeling approach to semantic document retrieval. Lecture Notes in Computer Science, 2348:167– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480167.
                                        pdf.

Stavrakas:2002:MSD

[824] Yannis Stavrakas and Manolis Gergatsoulis. Multidimensional semistructured data: Representing contextdependent information on the Web. Lecture Notes in Computer Science, 2348:183–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2348/23480183.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480183.
                                        pdf.

Papadopoulos:2002:RSR

[825] Nick Papadopoulos and Dimitris Plexousakis. The role of semantic relevance in dynamic user community management and the formulation of recommendations. Lecture Notes in Computer Science, 2348:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480200.
pdf.

Lipson:2002:CWE

[826] Howard F. Lipson, Nancy R. Mead, and Andrew P. Moore. Can we ever build survivable systems from COTS components? Lecture Notes in Computer Science, 2348:216–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480216.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480216.
pdf.

Kim:2002:TDM

[827] Henry M. Kim and Mark S. Fox. Towards a data model for quality management Web services: An ontology of measurement for enterprise modeling. Lecture Notes in Computer Science, 2348:230–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480230.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480230.
pdf.

Norrie:2002:MAR

[828] Moira C. Norrie and Alexios Palinginis. A modelling approach to the realisation of modular information spaces. Lecture Notes in Computer Science, 2348: 245–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480245.
                                        pdf.

Cali:2002:DIU

[829] Andrea Calì, Diego Calvanese, Giuseppe De Giacomo, and Maurizio Lenzerini. Data integration under integrity constraints. Lecture Notes in Computer Science, 2348:262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480262.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480262.
                                        pdf.

Zhang:2002:BXB

[830] Huaxin Zhang and Eleni Stroulia. Babel: An XML-based application integration framework. Lecture Notes in Computer Science, 2348:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480280.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480280.
                                        pdf.

Aoumeur:2002:IRP

[831] Nasreddine Aoumeur and Gunter Saake. Integrating and rapidprototyping UML structural and behavioural diagrams using rewriting logic. Lecture Notes in Computer Science, 2348:296–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2348/23480296.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2348/23480296.pdf.

Benerecetti:2002:VPP

[832] M. Benerecetti, M. Panti, L. Spalazzi, and S. Tacconi. Verification of payment protocols via MultiAgent model checking. Lecture Notes in Computer Science, 2348:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480311.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480311.
pdf.

Gans:2002:SMS

[833] Günter Gans, Gerhard Lakemeyer, Matthias Jarke, and Thomas Vits. SNet: a modeling and simulation environment for agent networks based on i* and ConGolog. Lecture Notes in Computer Science, 2348:328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480328.
pdf.

Paik:2002:UCA

[834] Hye young Paik, Boualem Benatallah, and Rachid Hamadi. Usage-centric adaptation of dynamic E-catalogs. Lecture Notes in Computer Science, 2348: 344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480344.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480344.
                                        pdf.

Lu:2002:RDA

[835] Jianguo Lu. Reengineering of database applications to EJB based architecture. Lecture Notes in Computer Science, 2348:361–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2348/23480361.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480361.
                                        pdf.

Lee:2002:ESS

[836] Sangjun Lee, Dongseop Kwon, and Sukho Lee. Efficient similarity search for time series data based on the minimum distance. Lecture Notes in Computer Science, 2348:377–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480377.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480377.
                                        pdf.

Wilson:2002:HPD

[837] John N. Wilson. A high-performance data structure for mobile information systems. Lecture Notes in Computer Science, 2348:392–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480392.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480392.
pdf.

Leicher:2002:ERV

[838] Andreas Leicher and Felix Bübl. External requirements validation for component-based systems. Lecture Notes in Computer Science, 2348:404– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480404.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480404.
pdf.

Leonardi:2002:UBR

[839] Maria Carmen Leonardi and Julio Cesar Sampaio do Prado Leite. Using business rules in EXtreme requirements. Lecture Notes in Computer Science, 2348:420–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2348/23480420.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2348/23480420.pdf.

Kajko-Mattsson:2002:ECP

[840] Mira Kajko-Mattsson. Evaluating CM3: Problem management. Lecture Notes in Computer Science, 2348:436– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480436.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480436.
                                        pdf.

Berlin:2002:DSM

[841] Jacob Berlin and Amihai Motro. Database schema matching using machine learning with feature selection. Lecture Notes in Computer Science, 2348:452–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2348/23480452.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480452.
                                        pdf.

Gomez:2002:EPC

[842] Cristina Gómez and Antoni Olivé. Evolving partitions in conceptual schemas in the UML. Lecture Notes in Computer Science, 2348:467–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480467.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480467.
                                        pdf.

Brien:2002:SEH

[843] Peter Mc. Brien and Alexandra Poulovassilis. Schema evolution in heterogeneous database architectures, A schema transformation approach. Lecture Notes in Computer Science, 2348: 484–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480484.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480484.
pdf.

Wetzel:2002:SBW

[844] Ingrid Wetzel and Ralf Klischewski. Serviceflow beyond workflow? concepts and architectures for supporting interorganizational service processes. Lecture Notes in Computer Science, 2348: 500–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480500.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480500.
pdf.

Fent:2002:DCE

[845] Alfred Fent, Herbert Reiter, and Burkhard Freitag. Design for change: Evolving workflow specifications in ULTRAflow. Lecture Notes in Computer Science, 2348:516–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480516.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2348/23480516.
                                        pdf.

vanderAalst:2002:AWA

[846] W. M. P. van der Aalst, A. Hirnschall, and H. M. W. Verbeek. An alternative way to analyze workflow graphs. Lecture Notes in Computer Science, 2348: 535–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480535.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480535.
                                        pdf.

Li:2002:AIB

[847] Yingjiu Li, Lingyu Wang, X. Sean Wang, and Sushil Jajodia. Auditing interval-based inference. Lecture Notes in Computer Science, 2348:553– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480553.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480553.
                                        pdf.

Liu:2002:LFX

[848] Mengchi Liu. A logical foundation for XML. Lecture Notes in Computer Science, 2348:568–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480568.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480568.
                                        pdf.

Vdovjak:2002:PSL

[849] Richard Vdovjak and Geert-Jan Houben. Providing the semantic layer for WIS design. Lecture Notes in Computer Science, 2348:584–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480584.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480584.
pdf.

Soderstrom:2002:TFC

[850] Eva Söderström, Birger Andersson, Paul Johannesson, Erik Perjons, and Benkt Wangler. Towards a framework for comparing process modelling languages. Lecture Notes in Computer Science, 2348:600–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480600.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480600.
pdf.

Prakash:2002:GME

[851] N. Prakash and M. P. S. Bhatia. Generic models for engineering methods of diverse domains. Lecture Notes in Computer Science, 2348:612– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480612.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480612.
pdf.

Saeki:2002:RMT

[852] Motoshi Saeki. Role of model transformation in method engineering. Lecture Notes in Computer Science, 2348: 626–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480626.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480626.
                                        pdf.

Dahchour:2002:GRM

[853] Mohamed Dahchour, Alain Pirotte, and Esteban Zimányi. A generic role model for dynamic objects. Lecture Notes in Computer Science, 2348:643– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480643.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480643.
                                        pdf.

Costal:2002:URU

[854] Dolors Costal, Maria-Ribera Sancho, and Ernest Teniente. Understanding redundancy in UML models for object-oriented analysis. Lecture Notes in Computer Science, 2348:659–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480659.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480659.
                                        pdf.

Olive:2002:RGR

[855] Antoni Olivé. Representation of generic relationship types in conceptual modeling. Lecture Notes in Computer Science, 2348:675–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480675.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480675.
pdf.

Adiba:2002:BST

[856] Michel Adiba and José-Luis Zechinelli-Martini. Building spatio-temporal presentations warehouses from heterogeneous multimedia Web servers. Lecture Notes in Computer Science, 2348: 692–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480692.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480692.
pdf.

Arjona:2002:PAB

[857] J. L. Arjona, R. Corchuelo, A. Ruiz, and M. Toro. A practical agentbased method to extract semantic information from the Web. Lecture Notes in Computer Science, 2348:697– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480697.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480697.
pdf.

Bussler:2002:PI

[858] Christoph Bussler. Process inheritance. Lecture Notes in Computer Science, 2348:701–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2348/23480701.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480701.
                                        pdf.

Cai:2002:APR

[859] Zhiming Cai and Eric Yu. Addressing performance requirements using a goal and scenario-oriented approach. Lecture Notes in Computer Science, 2348: 706–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480706.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480706.
                                        pdf.

Combi:2002:QDM

[860] Carlo Combi and Angelo Montanari. Querying data with multiple temporal dimensions. Lecture Notes in Computer Science, 2348:711– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480711.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480711.
                                        pdf.

Domenig:2002:QEI

[861] Ruxandra Domenig and Klaus R. Dittrich. Query explorativeness for integrated search in heterogeneous data sources. Lecture Notes in Computer Science, 2348:715–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480715.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480715.
pdf.

Filha:2002:UNT

[862] Irna M. R. Evangelista Filha, Altigran S. da Silva, Alberto H. F. Laender, and David W. Embley. Using nested tables for representing and querying semistructured Web data. Lecture Notes in Computer Science, 2348:719– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480719.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480719.
pdf.

Genero:2002:DVM

[863] Marcela Genero, Geert Poels, and Mario Piattini. Defining and validating measures for conceptual data model quality. Lecture Notes in Computer Science, 2348:724–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480724.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480724.
pdf.

Grundy:2002:ABM

[864] John Grundy and Wenjing Zou. An architecture for building multi-device thin-client Web user interfaces. Lecture Notes in Computer Science, 2348: 728–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480728.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480728.
                                        pdf.

Harren:2002:FTI

[865] Arne Harren and Heiko Tapken. A framework for tool-independent modeling of data acquisition processes for data warehousing. Lecture Notes in Computer Science, 2348:733–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480733.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480733.
                                        pdf.

Ivan:2002:MCD

[866] Cosmina Ivan, Vasile Dadarlat, and Kalman Pusztai. Managing complexity of designing routing protocols using a middleware approach. Lecture Notes in Computer Science, 2348:737– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480737.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480737.
                                        pdf.

Kang:2002:DIR

[867] Hyunchul Kang and JaeGuk Lim. Deferred incremental refresh of XML materialized views. Lecture Notes in Computer Science, 2348:742–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480742.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480742.
pdf.

Montero:2002:RHD

[868] Susana Montero, Paloma Díaz, and Ignacio Aedo. Requirements for hypermedia development methods: a survey of outstanding methods. Lecture Notes in Computer Science, 2348:747– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480747.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480747.
pdf.

Pontieri:2002:ASC

[869] Luigi Pontieri, Domenico Ursino, and Ester Zumpano. An approach for synergically carrying out intensional and extensional integration of data sources having different formats. Lecture Notes in Computer Science, 2348: 752–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480752.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2348/23480752.
                                        pdf.

Sengupta:2002:DSS

[870] Arijit Sengupta and Mehmet Dalkilic. DSQL — an SQL for structured documents. Lecture Notes in Computer Science, 2348:757–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480757.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480757.
                                        pdf.

Su:2002:CSO

[871] Xiaomeng Su and Lars Ilebrekke. A comparative study of ontology languages and tools. Lecture Notes in Computer Science, 2348:761–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480761.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480761.
                                        pdf.

Sugumaran:2002:LCB

[872] Vijayan Sugumaran and Mohan Tanniru. Life cycle based approach for knowledge management: A knowledge organization case study. Lecture Notes in Computer Science, 2348:766– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480766.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2348/23480766.
pdf.

Sun:2002:PQP

[873] Wenjun Sun and Kevin Lü. Parallel query processing algorithms for semi-structured data. Lecture Notes in Computer Science, 2348:770–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2348/23480770.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2348/23480770.
pdf.

Torchiano:2002:DSI

[874] Marco Torchiano and Giorgio Bruno. Domain-specific instance models in UML. Lecture Notes in Computer Science, 2348:774–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2348/23480774.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2348/23480774.pdf.

Tzitzikas:2002:EFO

[875] Yannis Tzitzikas, Nicolas Spyratos, Panos Constantopoulos, and Anastasia Analyti. Extended faceted ontologies. Lecture Notes in Computer Science, 2348:778–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2348/23480778.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2348/23480778.
pdf.

Vassiliadis:2002:LME

[876] Panos Vassiliadis, Alkis Simitsis, and Spiros Skiadopoulos. On the logical modeling of ETL processes. Lecture Notes in Computer Science, 2348:782– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480782.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480782.
                                        pdf.

Wang:2002:IAS

[877] Minhong Wang and Huaiqing Wang. Intelligent agent supported flexible workflow monitoring system. Lecture Notes in Computer Science, 2348:787– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2348/23480787.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2348/23480787.
                                        pdf.

Zhu:2002:MSS

[878] Kenny Qili Zhu and Andrew E. Santosa. A meeting scheduling system based on open constraint programming. Lecture Notes in Computer Science, 2348:792–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2348/23480792.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2348/23480792.pdf.

Anonymous:2002:AIo

[879] Anonymous. Author index. Lecture Notes in Computer Science, 2348:797– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2348/2348ind.pdf.

Rombach:2002:SQV

[880] H. Dieter Rombach. Software quality versus time-to-market: How to resolve these conflicts? Lecture Notes in Computer Science, 2349:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490001.
pdf.

Terho:2002:MWS

[881] Mikko Terho. Mobile Web services and software quality. Lecture Notes in Computer Science, 2349:2– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490002.
pdf.

Pfleeger:2002:SSI

[882] Shari Lawrence Pfleeger. Solid software: Is it rocket science? Lecture Notes in Computer Science, 2349:7– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490007.
pdf.

Rifkin:2002:PII

[883] Stan Rifkin. Is process improvement irrelevant to produce new era software? Lecture Notes in Computer Science, 2349:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490013.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490013.
pdf.

Dehli:2002:MDB

[884] Einar Dehli. Model-driven business operations. Lecture Notes in Computer Science, 2349:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490017.
pdf.

Anttila:2002:PQS

[885] Juhani Anttila. Product quality in software business connection. Lecture Notes in Computer Science, 2349:25– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490025.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2349/23490025.
pdf.

Watson:2002:BDS

[886] Gregory H. Watson. Breakthrough in delivering software quality: Capability maturity model and six sigma. Lecture Notes in Computer Science, 2349: 36–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490036.
pdf.

Chang:2002:UMA

[887] Wen-Kui Chang, Min-Hsiang Chuang, and Chao-Tung Yang. Using mobile agents for security testing in Web environments. Lecture Notes in Computer Science, 2349:42–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490042.
pdf.

Kanayama:2002:QCT

[888] Toyohiro Kanayama, Hideto Ogasawara, and Hiroshi Kimijima. Quality control techniques for constructing attractive corporate Websites: Usability in relation to the popularity ranking of Websites. Lecture Notes in Computer Science, 2349:53– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490053.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490053.
                                        pdf.

Chang:2002:EPW

[889] Wen-Kui Chang and Shing-Kai Hon. Evaluating the performance of a Web site via queuing theory. Lecture Notes in Computer Science, 2349:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490063.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490063.
                                        pdf.

Kauppinen:2002:LLA

[890] Marjo Kauppinen, Tapani Aaltio, and Sari Kujala. Lessons learned from applying the requirements engineering good practice guide for process improvement. Lecture Notes in Computer Science, 2349:73–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490073.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490073.
                                        pdf.

Nakata:2002:QAA

[891] Masahiro Nakata and Katsuyuki Yasuda. Quality assurance activities for ASP based on SLM in hitachi. Lecture Notes in Computer Science, 2349: 82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490082.
pdf.

Raghavan:2002:ISQ

[892] Gopalakrishna Raghavan. Improving software quality in product families through systematic reengineering. Lecture Notes in Computer Science, 2349: 90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490090.
pdf.

Richardson:2002:SMW

[893] Ita Richardson. SPI models: What characteristics are required for small software development companies? Lecture Notes in Computer Science, 2349: 100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490100.
pdf.

Schneider:2002:EBP

[894] Kurt Schneider. Experience based process improvement. Lecture Notes in Computer Science, 2349:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490114.
                                        pdf.

Ogasawara:2002:HEP

[895] Hideto Ogasawara, Atsushi Yamada, Takumi Kusanagi, and Mikako Arami. How to effectively promote the software process improvement activities in a large-scale organization. Lecture Notes in Computer Science, 2349:124– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490124.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490124.
                                        pdf.

Kitajima:2002:CET

[896] Yoshihiro Kitajima, Hitoshi Fuji, Seiichiro Satou, Hitoshi Ohsugi, Isao Gotou, and Hitoshi Oono. Consideration of EVMS technique application to software development. Lecture Notes in Computer Science, 2349:135– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490135.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490135.
                                        pdf.

Rosendahl:2002:PIR

[897] Esa Rosendahl and Ton Vullinghs. Performing initial risk assessments in software acquisition projects. Lecture Notes in Computer Science, 2349:146– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490146.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490146.
pdf.

Larvet:2002:UDC

[898] Philippe Larvet and Frédérique Vallée. UML developments: Cost estimation from requirements. Lecture Notes in Computer Science, 2349:156–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490156.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490156.
pdf.

Grutter:2002:PSP

[899] Georg Grütter and Stefan Ferber. The personal software process in practice: Experience in two cases over five years. Lecture Notes in Computer Science, 2349:165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490165.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490165.
pdf.

Abrahamsson:2002:PSP

[900] Pekka Abrahamsson and Karlheinz Kautz. Personal software process: Classroom experiences from finland. Lecture Notes in Computer Science, 2349:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2349/23490175.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490175.
                                        pdf.

Gantner:2002:GES

[901] Thomas Gantner and Tobias Häberlein. GARP — the evolution of a software acquisition process model. Lecture Notes in Computer Science, 2349:186– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490186.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490186.
                                        pdf.

Okazaki:2002:CCP

[902] Yasuko Okazaki. Cooperation and competition with partner companies: Practices for quality control through competition among teams. Lecture Notes in Computer Science, 2349:197– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490197.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490197.
                                        pdf.

Olesen:2002:CCD

[903] Robert Olesen and Jørn Johansen. Cooperate or conquer? A Danish survey of the customer-supplier relationship. Lecture Notes in Computer Science, 2349:207–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2349/23490207.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2349/23490207.
pdf.

Fukui:2002:ISC

[904] Shinji Fukui. Introduction of the software configuration management team and defect tracking system for global distributed development. Lecture Notes in Computer Science, 2349: 217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490217.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490217.
pdf.

Goldin:2002:SDB

[905] Leah Goldin and Lilach Rochell. Software development bug tracking: “Tool Isn’t User Friendly” or “User Isn’t Process Friendly”. Lecture Notes in Computer Science, 2349:226–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490226.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490226.
pdf.

Hashino:2002:PMD

[906] Nobuyuki Hashino, Satoshi Kurokawa, Mamoru Wakaki, and Junji Nakasone. I-P-O/multilateral design quality evaluation methods: Process improvements and effects. Lecture Notes in Computer Science, 2349:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490236.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490236.
                                        pdf.

Jaccheri:2002:CCP

[907] Letizia Jaccheri and Marco Torchiano. Classifying COTS products. Lecture Notes in Computer Science, 2349:246– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490246.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490246.
                                        pdf.

Helander:2002:USC

[908] Nina Helander, Pauliina Ulkuniemi, and Veikko Seppänen. Understanding software component markets: The value creation perspective. Lecture Notes in Computer Science, 2349:256– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490256.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490256.
                                        pdf.

Helokunnas:2002:CBC

[909] Tuija Helokunnas and Marko Nyby. Collaboration between a COTS integrator and vendors. Lecture Notes in Computer Science, 2349:267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490267.
pdf.

Shimanaka:2002:CGT

[910] Kazutoshi Shimanaka, Masato Matsumoto, Junji Koga, and Hiroyuki Domae. Creation of a guideline for tailoring development processes using project metrics data. Lecture Notes in Computer Science, 2349:274– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490274.
pdf.

Nawrocki:2002:CCL

[911] Jerzy R. Nawrocki, Bartosz Walter, and Adam Wojciechowski. Comparison of CMM level 2 and eXtreme programming. Lecture Notes in Computer Science, 2349:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490288.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490288.
pdf.

Calero:2002:ESM

[912] Coral Calero, Houari Sahraoui, and Mario Piattini. An empirical study with metrics for object-relational databases. Lecture Notes in Computer Science, 2349:298–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2349/23490298.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2349/23490298.pdf.

Takahashi:2002:EMB

[913] Juichi Takahashi and Yoshiaki Kakuda. Extended model-based testing toward high code coverage rate. Lecture Notes in Computer Science, 2349:310– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490310.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490310.
                                        pdf.

Chan:2002:RRT

[914] Kwok Ping Chan, Tsong Yueh Chen, and Dave Towey. Restricted random testing. Lecture Notes in Computer Science, 2349:321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2349/23490321.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2349/23490321.
                                        pdf.

Nishi:2002:QAT

[915] Yasuharu Nishi. Quality-adaptive testing: a strategy for testing with focusing on where bugs have been detected. Lecture Notes in Computer Science, 2349: 331–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490331.
pdf.

Stark:2002:PRQ

[916] Jacqueline Stark. Peer reviews as a quality management technique in open-source software development projects. Lecture Notes in Computer Science, 2349:340–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490340.
pdf.

Tenhunen:2002:EIA

[917] Vesa Tenhunen and Jorma Sajaniemi. An evaluation of inspection automation tools. Lecture Notes in Computer Science, 2349:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2349/23490351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2349/23490351.
pdf.

Anonymous:2002:AIp

[918] Anonymous. Author index. Lecture Notes in Computer Science, 2349:363– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2349/2349auth.pdf.

Pece:2002:TEC

[919] A. E. C. Pece and A. D. Worrall. Tracking with the EM contour algorithm. Lecture Notes in Computer Science, 2350:3–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500003.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500003.
                                        pdf.

Mittal:2002:MMV

[920] A. Mittal and L. S. Davis. M2Tracker: a multi-view approach to segmenting and tracking people in a cluttered scene using region-based stereo. Lecture Notes in Computer Science, 2350: 18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500018.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500018.
                                        pdf.

Srivastava:2002:AIM

[921] A. Srivastava, X. Liu, and U. Grenander. Analytical image models and their applications. Lecture Notes in Computer Science, 2350:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500037.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500037.
pdf.

Lindeberg:2002:TRV

[922] T. Lindeberg. Time-recursive velocityadapted spatio-temporal scale-space filters. Lecture Notes in Computer Science, 2350:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2350/23500052.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2350/23500052.pdf.

Tell:2002:CAT

[923] D. Tell and S. Carlsson. Combining appearance and topology for wide baseline matching. Lecture Notes in Computer Science, 2350:68– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500068.
pdf.

Tordoff:2002:GSC

[924] B. Tordoff and D. W. Murray. Guided sampling and consensus for motion estimation. Lecture Notes in Computer Science, 2350:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500082.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500082.
                                        pdf.

Geusebroek:2002:FAG

[925] J.-M. Geusebroek, A. W. M. Smeulders, and J. van de Weijer. Fast anisotropic Gauss filtering. Lecture Notes in Computer Science, 2350:99– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500099.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500099.
                                        pdf.

Rivera:2002:ARC

[926] M. Rivera and J. L. Marroquin. Adaptive rest condition potentials: Second order edge-preserving regularization. Lecture Notes in Computer Science, 2350:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2350/23500113.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500113.
                                        pdf.

Mikolajczyk:2002:AII

[927] K. Mikolajczyk and C. Schmid. An affine invariant interest point detector. Lecture Notes in Computer Science, 2350:128–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2350/23500128.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2350/23500128.
pdf.

Kuijper:2002:UME

[928] Arjan Kuijper and Luc Florack. Understanding and modeling the evolution of critical points under Gaussian blurring. Lecture Notes in Computer Science, 2350:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500143.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500143.
pdf.

Koenderink:2002:IPD

[929] J. J. Koenderink and A. J. van Doorn. Image processing done right. Lecture Notes in Computer Science, 2350:158– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500158.
pdf.

Zhu:2002:MDR

[930] Y. Zhu, D. Comaniciu, S. Schwartz, and V. Ramesh. Multimodal data representations with parameterized local structures. Lecture Notes in Computer Science, 2350:173–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500173.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500173.
                                        pdf.

Kuijper:2002:RNG

[931] A. Kuijper and L. Florack. The relevance of non-generic events in scale space models. Lecture Notes in Computer Science, 2350:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500190.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500190.
                                        pdf.

Wang:2002:LCP

[932] B. Wang, K. K. Sung, and T. K. Ng. The localized consistency principle for image matching under nonuniform illumination variation and affine distortion. Lecture Notes in Computer Science, 2350:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500205.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500205.
                                        pdf.

Hadjidemetriou:2002:RSU

[933] E. Hadjidemetriou, M. D. Grossberg, and S. K. Nayar. Resolution selection using generalized entropies of multiresolution histograms. Lecture Notes in Computer Science, 2350:220– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500220.
pdf.

Chen:2002:RCV

[934] H. Chen and P. Meer. Robust computer vision through kernel density estimation. Lecture Notes in Computer Science, 2350:236–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500236.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500236.
pdf.

Chefdhotel:2002:CFM

[935] C. Chefd’hotel, D. Tschumperlé, R. Deriche, and O. Faugeras. Constrained flows of matrix-valued functions: Application to diffusion tensor regularization. Lecture Notes in Computer Science, 2350:251–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500251.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500251.
pdf.

Carcassoni:2002:HFS

[936] M. Carcassoni and E. R. Hancock. A hierarchical framework for spectral correspondence. Lecture Notes in Computer Science, 2350:266–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500266.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500266.
                                        pdf.

Carneiro:2002:PBL

[937] G. Carneiro and A. D. Jepson. Phasebased local features. Lecture Notes in Computer Science, 2350:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500282.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500282.
                                        pdf.

Vasconcelos:2002:WRI

[938] N. Vasconcelos and G. Carneiro. What is the role of independence for visual recognition? Lecture Notes in Computer Science, 2350:297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500297.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500297.
                                        pdf.

Ren:2002:PMS

[939] X. Ren and J. Malik. A probabilistic multi-scale model for contour completion based on image statistics. Lecture Notes in Computer Science, 2350:312– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500312.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500312.
pdf.

Pedersen:2002:TFP

[940] K. S. Pedersen and A. B. Lee. Toward a full probability model of edges in natural images. Lecture Notes in Computer Science, 2350:328–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500328.
pdf.

Malladi:2002:FDS

[941] R. Malladi and I. Ravve. Fast difference schemes for edge enhancing Beltrami flow. Lecture Notes in Computer Science, 2350:343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500343.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500343.
pdf.

Loy:2002:FRS

[942] G. Loy and A. Zelinsky. A fast radial symmetry transform for detecting points of interest. Lecture Notes in Computer Science, 2350:358– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500358.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500358.
                                        pdf.

Felsberg:2002:IFB

[943] M. Felsberg and G. Sommer. Image features based on a new approach to 2D rotation invariant quadrature filters. Lecture Notes in Computer Science, 2350:369–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500369.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500369.
                                        pdf.

Huggins:2002:REM

[944] P. S. Huggins and S. W. Zucker. Representing edge models via local principal component analysis. Lecture Notes in Computer Science, 2350:384– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500384.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500384.
                                        pdf.

Gilboa:2002:RSF

[945] Guy Gilboa, Nir A. Sochen, and Yehoshua Y. Zeevi. Regularized shock filters and complex diffusion. Lecture Notes in Computer Science, 2350:399– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2350/23500399.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500399.
pdf.

Schaffalitzky:2002:MVM

[946] F. Schaffalitzky and A. Zisserman. Multi-view matching for unordered image sets, or “How Do I Organize My Holiday Snaps?”. Lecture Notes in Computer Science, 2350:414– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500414.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500414.
pdf.

Speyer:2002:PEP

[947] G. Speyer and M. Werman. Parameter estimates for a pencil of lines: Bounds and estimators. Lecture Notes in Computer Science, 2350:432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500432.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500432.
pdf.

Vasilescu:2002:MAI

[948] M. A. O. Vasilescu and D. Terzopoulos. Multilinear analysis of image ensembles: TensorFaces. Lecture Notes in Computer Science, 2350:447–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500447.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500447.
                                        pdf.

Goldenberg:2002:DDL

[949] R. Goldenberg, R. Kimmel, E. Rivlin, and M. Rudzsky. ’dynamism of a dog on a leash’ or behavior classification by eigen-decomposition of periodic motions. Lecture Notes in Computer Science, 2350:461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500461.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500461.
                                        pdf.

Fablet:2002:ADT

[950] R. Fablet and M. J. Black. Automatic detection and tracking of human motion with a view-based representation. Lecture Notes in Computer Science, 2350:476–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2350/23500476.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500476.
                                        pdf.

Tarel:2002:URE

[951] J.-P. Tarel, S.-S. Ieng, and P. Charbonnier. Using robust estimation algorithms for tracking explicit curves. Lecture Notes in Computer Science, 2350: 492–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500492.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500492.
pdf.

Swaminathan:2002:MAS

[952] R. Swaminathan, S. B. Kang, R. Szeliski, A. Criminisi, and S. K. Nayar. On the motion and appearance of specularities in image sequences. Lecture Notes in Computer Science, 2350:508– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500508.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500508.
pdf.

Ringer:2002:MHT

[953] M. Ringer and J. Lasenby. Multiple hypothesis tracking for automatic optical motion capture. Lecture Notes in Computer Science, 2350:524– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500524.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500524.
pdf.

Jiang:2002:SAG

[954] Guang Jiang, Hung tat Tsui, Long Quan, and Andrew Zisserman. Single axis geometry by fitting conics. Lecture Notes in Computer Science, 2350: 537–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500537.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500537.
                                        pdf.

Bhat:2002:CPP

[955] K. S. Bhat, S. M. Seitz, J. Popović, and P. K. Khosla. Computing the physical parameters of rigid-body motion from video. Lecture Notes in Computer Science, 2350:551–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500551.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500551.
                                        pdf.

Sminchisescu:2002:BRL

[956] C. Sminchisescu and B. Triggs. Building roadmaps of local minima of visual models. Lecture Notes in Computer Science, 2350:566–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500566.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500566.
                                        pdf.

Wang:2002:GMT

[957] Y. Wang and S.-C. Zhu. A generative method for textured motion: Analysis and synthesis. Lecture Notes in Computer Science, 2350:583–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500583.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500583.
pdf.

Zhao:2002:SRO

[958] W. Zhao and H. S. Sawhney. Is superresolution with optical flow feasible? Lecture Notes in Computer Science, 2350:599–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2350/23500599.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2350/23500599.
pdf.

Weinshall:2002:NVG

[959] D. Weinshall, M.-S. Lee, T. Brodsky, M. Trajkovic, and D. Feldman. New view generation with a bi-centric camera. Lecture Notes in Computer Science, 2350:614–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500614.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500614.
pdf.

Sullivan:2002:RTH

[960] J. Sullivan and S. Carlsson. Recognizing and tracking human action. Lecture Notes in Computer Science, 2350: 629–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2350/23500629.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500629.
                                        pdf.

Vermaak:2002:TIO

[961] J. Vermaak, P. Pérez, M. Gangnet, and A. Blake. Towards improved observation models for visual tracking: Selective adaptation. Lecture Notes in Computer Science, 2350:645– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500645.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500645.
                                        pdf.

Perez:2002:CBP

[962] P. Pérez, C. Hue, J. Vermaak, and M. Gangnet. Color-based probabilistic tracking. Lecture Notes in Computer Science, 2350:661–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2350/23500661.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500661.
                                        pdf.

Corpetti:2002:DMA

[963] T. Corpetti, É. Mémin, and P. Pérez. Dense motion analysis in fluid imagery. Lecture Notes in Computer Science, 2350:676–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2350/23500676.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2350/23500676.
pdf.

Jepson:2002:LMR

[964] A. D. Jepson, D. J. Fleet, and M. J. Black. A layered motion representation with occlusion and compact spatial support. Lecture Notes in Computer Science, 2350:692–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500692.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500692.
pdf.

Brand:2002:ISV

[965] M. Brand. Incremental singular value decomposition of uncertain data with missing values. Lecture Notes in Computer Science, 2350:707–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500707.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500707.
pdf.

Alvarez:2002:SDO

[966] L. Alvarez, R. Deriche, T. Papadopoulo, and J. Sánchez. Symmetrical dense optical flow estimation with occlusions detection. Lecture Notes in Computer Science, 2350:721– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2350/23500721.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2350/23500721.
                                        pdf.

Beal:2002:AVS

[967] Matthew J. Beal, Hagai Attias, and Nebojsa Jojic. Audio-video sensor fusion with probabilistic graphical models. Lecture Notes in Computer Science, 2350:736–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2350/23500736.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2350/23500736.pdf.

Shechtman:2002:IST

[968] E. Shechtman, Y. Caspi, and M. Irani. Increasing space-time resolution in video. Lecture Notes in Computer Science, 2350:753–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2350/23500753.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2350/23500753.pdf.

Sminchisescu:2002:HIS

[969] Cristian Sminchisescu and Bill Triggs. Hyperdynamics importance sampling. Lecture Notes in Computer Science, 2350:769–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2350/23500769.htm;   http:
                                        //link.springer-ny.com/link/service/

series/0558/papers/2350/23500769.
pdf.

Sidenbladh:2002:IPM

[970] H. Sidenbladh, M. J. Black, and L. Sigal. Implicit probabilistic models of human motion for synthesis and tracking. Lecture Notes in Computer Science, 2350:784–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2350/23500784.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2350/23500784.
pdf.

Torresani:2002:STT

[971] L. Torresani and C. Bregler. Spacetime tracking. Lecture Notes in Computer Science, 2350:801–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/2350/23500801.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2350/23500801.pdf.

Anonymous:2002:AIq

[972] Anonymous. Author index. Lecture Notes in Computer Science, 2350:813– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2350/2350auth.pdf.

Gomes:2002:VAR

[973] J. Gomes and A. Mojsilovic. A variational approach to recovering a manifold from sample points. Lecture Notes in Computer Science, 2351:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510003.
pdf.

Jin:2002:VAS

[974] H. Jin and P. Favaro. A variational approach to shape from defocus. Lecture Notes in Computer Science, 2351: 18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510018.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510018.
pdf.

Yu:2002:SGS

[975] Y. Yu and J. T. Chang. Shadow graphs and surface reconstruction. Lecture Notes in Computer Science, 2351:31– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510031.
pdf.

Drbohlav:2002:SRA

[976] O. Drbohlav and R. Šára. Specularities reduce ambiguity of uncalibrated photometric stereo. Lecture Notes in Computer Science, 2351:46– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510046.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510046.
pdf.

Robles-Kelly:2002:PCM

[977] A. Robles-Kelly and E. R. Hancock. Pairwise clustering with matrix factorisation and the EM algorithm. Lecture Notes in Computer Science, 2351:63– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510063.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510063.
pdf.

Rousson:2002:SPL

[978] M. Rousson and N. Paragios. Shape priors for level set representations. Lecture Notes in Computer Science, 2351: 78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510078.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510078.
pdf.

Cremers:2002:NSS

[979] D. Cremers, T. Kohlberger, and C. Schnörr. Nonlinear shape statistics in Mumford–Shah based segmentation. Lecture Notes in Computer Science, 2351:93–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/2351/23510093.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510093.
                                        pdf.

Borenstein:2002:CST

[980] E. Borenstein and S. Ullman. Classspecific, top-down segmentation. Lecture Notes in Computer Science, 2351: 109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510109.
                                        pdf.

Lhuillier:2002:QDR

[981] M. Lhuillier and L. Quan. Quasi-dense reconstruction from image sequence. Lecture Notes in Computer Science, 2351:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2351/23510125.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510125.
                                        pdf.

Geyer:2002:PCF

[982] C. Geyer and K. Daniilidis. Properties of the catadioptric fundamental matrix. Lecture Notes in Computer Science, 2351:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510140.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2351/23510140.
pdf.

Robertson:2002:BAM

[983] D. P. Robertson and R. Cipolla. Building architectural models from many views using map constraints. Lecture Notes in Computer Science, 2351:155– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510155.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510155.
pdf.

Strecha:2002:MSI

[984] C. Strecha and L. Van Gool. Motion — stereo integration for depth estimation. Lecture Notes in Computer Science, 2351:170–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510170.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510170.
pdf.

Cornelis:2002:LDR

[985] K. Cornelis, M. Pollefeys, and L. Van Gool. Lens distortion recovery for accurate sequential structure and motion recovery. Lecture Notes in Computer Science, 2351:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510186.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2351/23510186.
                                        pdf.

Huang:2002:GRC

[986] K. Huang, R. Fossum, and Y. Ma. Generalized rank conditions in multiple view geometry with applications to dynamical scenes. Lecture Notes in Computer Science, 2351:201– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510201.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510201.
                                        pdf.

Ernst:2002:DSM

[987] F. Ernst, P. Wilinski, and K. van Overveld. Dense structure-from-motion: An approach based on segment matching. Lecture Notes in Computer Science, 2351:217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510217.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510217.
                                        pdf.

Maciel:2002:MRO

[988] J. Maciel and J. Costeira. Maximizing rigidity: Optimal matching under scaled-orthography. Lecture Notes in Computer Science, 2351:232– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510232.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2351/23510232.
pdf.

Smelyansky:2002:DIF

[989] V. N. Smelyansky, R. D. Morris, F. O. Kuehnel, D. A. Maluf, and P. Cheeseman. Dramatic improvements to feature based stereo. Lecture Notes in Computer Science, 2351:247– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510247.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510247.
pdf.

Bazin:2002:MCP

[990] P.-L. Bazin and J.-M. Vézien. Motion curves for parametric shape and motion estimation. Lecture Notes in Computer Science, 2351:262–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510262.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510262.
pdf.

Qian:2002:BSC

[991] G. Qian and R. Chellappa. Bayesian self-calibration of a moving camera. Lecture Notes in Computer Science, 2351:277–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2351/23510277.htm;   http:
//link.springer-ny.com/link/service/

                                        series/0558/papers/2351/23510277.
                                        pdf.

Georgescu:2002:BRS

[992] B. Georgescu and Peter Meer. Balanced recovery of 3D structure and camera motion from uncalibrated image sequences. Lecture Notes in Computer Science, 2351:294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510294.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510294.
                                        pdf.

Rother:2002:LMV

[993] C. Rother and S. Carlsson. Linear multi view reconstruction with missing data. Lecture Notes in Computer Science, 2351:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2351/23510309.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2351/23510309.pdf.

Plaenkers:2002:MBS

[994] R. Plaenkers and P. Fua. Modelbased silhouette extraction for accurate people tracking. Lecture Notes in Computer Science, 2351:325–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510325.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2351/23510325.
pdf.

Bartoli:2002:NLO

[995] A. Bartoli. On the non-linear optimization of projective motion using minimal parameters. Lecture Notes in Computer Science, 2351:340– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510340.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510340.
pdf.

Martinec:2002:SMP

[996] D. Martinec and T. Pajdla. Structure from many perspective images with occlusions. Lecture Notes in Computer Science, 2351:355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510355.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510355.
pdf.

Wolf:2002:SSS

[997] L. Wolf and A. Zomet. Sequenceto-sequence self calibration. Lecture Notes in Computer Science, 2351:370– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510370.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510370.
pdf.

Vidal:2002:SPM

[998] R. Vidal and J. Oliensis. Structure from planar motions with small baselines. Lecture Notes in Computer Science, 2351:383–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510383.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510383.
                                        pdf.

Levin:2002:RSV

[999] A. Levin and A. Shashua. Revisiting single-view shape tensors: Theory and applications. Lecture Notes in Computer Science, 2351:399–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510399.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510399.
                                        pdf.

Cobzas:2002:TRU

[1000] D. Cobzas and M. Jagersand. Tracking and rendering using dynamic textures on geometric structure from motion. Lecture Notes in Computer Science, 2351:415–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2351/23510415.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510415.
                                        pdf.

Hartley:2002:SCP

[1001] R. I. Hartley and R. Kaucic. Sensitivity of calibration to principal point position. Lecture Notes in Computer Science, 2351:433–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510433.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510433.
pdf.

Kahl:2002:CCS

[1002] F. Kahl and R. Hartley. Critical curves and surfaces for Euclidean reconstruction. Lecture Notes in Computer Science, 2351:447–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510447.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510447.
pdf.

Jelinek:2002:VSO

[1003] D. Jelinek and C. J. Taylor. View synthesis with occlusion reasoning using quasi-sparse feature correspondences. Lecture Notes in Computer Science, 2351:463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2351/23510463.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2351/23510463.
pdf.

Yang:2002:EGC

[1004] R. Yang and Z. Zhang. Eye gaze correction with stereovision for video-teleconferencing. Lecture Notes in Computer Science, 2351:479–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510479.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510479.
                                        pdf.

Clerc:2002:WBC

[1005] M. Clerc. Wavelet-based correlation for stereopsis. Lecture Notes in Computer Science, 2351:495–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510495.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510495.
                                        pdf.

Sun:2002:SMU

[1006] J. Sun, H.-Y. Shum, and N.-N. Zheng. Stereo matching using belief propagation. Lecture Notes in Computer Science, 2351:510–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510510.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510510.
                                        pdf.

Szeliski:2002:SSP

[1007] R. Szeliski and D. Scharstein. Symmetric sub-pixel stereo matching. Lecture Notes in Computer Science, 2351: 525–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510525.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510525.
pdf.

Werner:2002:NTA

[1008] T. Werner and A. Zisserman. New techniques for automated architectural reconstruction from photographs. Lecture Notes in Computer Science, 2351: 541–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510541.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510541.
pdf.

Zhang:2002:SMS

[1009] Y. Zhang and C. Kambhamettu.

Stereo matching with segmentationbased cooperation. Lecture Notes in Computer Science, 2351:556–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510556.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510556.
pdf.

VandenWyngaerd:2002:CRS

[1010] J. Vanden Wyngaerd and L. Van Gool. Coarse registration of surface patches with local symmetries. Lecture Notes in Computer Science, 2351: 572–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510572.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510572.
pdf.

Sharp:2002:MRS

[1011] G. C. Sharp, S. W. Lee, and D. K. Wehe. Multiview registration of 3D scenes by minimizing error between coordinate frames. Lecture Notes in Computer Science, 2351:587–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510587.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510587.
pdf.

Kamberov:2002:RSR

[1012] G. Kamberov and G. Kamberova. Recovering surfaces from the restoring force. Lecture Notes in Computer Science, 2351:598–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2351/23510598.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2351/23510598.pdf.

Noakes:2002:ISD

[1013] L. Noakes and R. Kozera. Interpolating sporadic data. Lecture Notes in Computer Science, 2351:613–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510613.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510613.
pdf.

Ragheb:2002:HRU

[1014] H. Ragheb and E. R. Hancock.

Highlight removal using shape-from-shading. Lecture Notes in Computer Science, 2351:626–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2351/23510626.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2351/23510626.pdf.

Kazhdan:2002:RSD

[1015] M. Kazhdan, B. Chazelle, D. Dobkin, A. Finkelstein, and T. Funkhouser. A reflective symmetry descriptor. Lecture Notes in Computer Science, 2351:642– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510642.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510642.
pdf.

Liu:2002:GSA

[1016] Y. Liu, R. Collins, and Y. Tsin. Gait sequence analysis using frieze patterns.

Lecture Notes in Computer Science, 2351:657–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2351/23510657.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2351/23510657.
pdf.

Tam:2002:FPM

[1017] R. Tam and W. Heidrich. Featurepreserving medial axis noise removal. Lecture Notes in Computer Science, 2351:672–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2351/23510672.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2351/23510672.
pdf.

Liu:2002:HSM

[1018] C. Liu, H.-Y. Shum, and C. Zhang. Hierarchical shape modeling for automatic face localization. Lecture Notes in Computer Science, 2351:687– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510687.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510687.
pdf.

Slobodan:2002:UDF

[1019] I. Slobodan and P. Fua. Using Dirichlet free form deformation to fit deformable models to noisy 3-D data. Lecture Notes in Computer Science, 2351:704– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510704.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510704.
pdf.

Giblin:2002:TMA

[1020] P. Giblin and B. B. Kimia. Transitions of the 3D medial axis under a oneparameter family of deformations. Lecture Notes in Computer Science, 2351: 718–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510718.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510718.
pdf.

Favaro:2002:LSD

[1021] P. Favaro and S. Soatto. Learning shape from defocus. Lecture Notes in Computer Science, 2351:735– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510735.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510735.
pdf.

Zunic:2002:RMP

[1022] J. Žunić and P. L. Rosin. A rectilinearity measurement for polygons. Lecture Notes in Computer Science, 2351: 746–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2351/23510746.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510746.
                                        pdf.

Savarese:2002:LAR

[1023] S. Savarese and P. Perona. Local analysis for 3D reconstruction of specular surfaces — Part II. Lecture Notes in Computer Science, 2351:759– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510759.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510759.
                                        pdf.

Paragios:2002:MDF

[1024] N. Paragios, M. Rousson, and V. Ramesh. Matching distance functions: a shape-to-area variational approach for global-to-local registration. Lecture Notes in Computer Science, 2351:775–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2351/23510775.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510775.
                                        pdf.

Prados:2002:SSV

[1025] E. Prados, O. Faugeras, and E. Rouy. Shape from shading and viscosity solutions. Lecture Notes in Computer Science, 2351:790–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2351/23510790.

htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2351/23510790.pdf.

Fusiello:2002:MAR

[1026] A. Fusiello, U. Castellani, L. Ronchetti, and V. Murino. Model acquisition by registration of multiple acoustic range views. Lecture Notes in Computer Science, 2351:805–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510805.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510805.
pdf.

Kaminski:2002:GTT

[1027] J. Y. Kaminski and M. Teicher. General trajectory triangulation. Lecture Notes in Computer Science, 2351:823– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510823.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2351/23510823.
pdf.

Pollefeys:2002:SDP

[1028] M. Pollefeys, F. Verbiest, and L. Van Gool. Surviving dominant planes in uncalibrated structure and motion recovery. Lecture Notes in Computer Science, 2351:837– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2351/23510837.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510837.
                                        pdf.

Dick:2002:BEB

[1029] A. R. Dick, P. H. S. Torr, and R. Cipolla. A Bayesian estimation of building shape using MCMC. Lecture Notes in Computer Science, 2351:852– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510852.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510852.
                                        pdf.

Sturm:2002:SMD

[1030] P. Sturm. Structure and motion for dynamic scenes — the case of points moving in planes. Lecture Notes in Computer Science, 2351:867– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510867.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2351/23510867.
                                        pdf.

Irani:2002:WDS

[1031] M. Irani, T. Hassner, and P. Anandan. What does the scene look like from a scene point? Lecture Notes in Computer Science, 2351:883– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2351/23510883.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2351/23510883.
pdf.

Anonymous:2002:AIr

[1032] Anonymous. Author index. Lecture Notes in Computer Science, 2351:899– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2351/2351auth.pdf.

Davies:2002:SSM

[1033] Rhodri H. Davies, Carole J. Twining, Tim F. Cootes, John C. Waterton, and Chris J. Taylor. 3D statistical shape models using direct optimisation of description length. Lecture Notes in Computer Science, 2352:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520003.
pdf.

Donato:2002:ATP

[1034] Gianluca Donato and Serge Belongie. Approximate thin plate spline mappings. Lecture Notes in Computer Science, 2352:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520021.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520021.
pdf.

Soatto:2002:DDM

[1035] Stefano Soatto and Anthony J. Yezzi. DEFORMOTION: Deforming motion, shape average and the joint registration and segmentation of images. Lecture Notes in Computer Science, 2352: 32–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520032.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520032.
pdf.

Duci:2002:RMM

[1036] Alessandro Duci, Anthony J. Yezzi, Sanjoy Mitter, and Stefano Soatto. Region matching with missing parts. Lecture Notes in Computer Science, 2352: 48–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520048.
pdf.

Kolmogorov:2002:WEF

[1037] Vladimir Kolmogorov and Ramin Zabih. What energy functions can be minimized via graph cuts? Lecture Notes in Computer Science, 2352:65– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520065.
pdf.

Kolmogorov:2002:MCS

[1038] Vladimir Kolmogorov and Ramin Zabih. Multi-camera scene reconstruction via graph cuts. Lecture Notes in Computer Science, 2352:82– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520082.
pdf.

Senegas:2002:MCM

[1039] Julien Sénégas. A Markov chain Monte Carlo approach to stereovision. Lecture Notes in Computer Science, 2352: 97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520097.
pdf.

Bhotika:2002:PTO

[1040] Rahul Bhotika, David J. Fleet, and Kiriakos N. Kutulakos. A probabilistic theory of occupancy and emptiness. Lecture Notes in Computer Science, 2352:112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520112.
pdf.

Mathiassen:2002:TSM

[1041] John Reidar Mathiassen, Amund Skavhaug, and Ketil Bø. Texture similarity measure using Kullback–Leibler divergence between gamma distributions. Lecture Notes in Computer Science, 2352:133–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2352/23520133.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2352/23520133.pdf.

Narasimhan:2002:AIO

[1042] Srinivasa G. Narasimhan, Chi Wang, and Shree K. Nayar. All the images of an outdoor scene. Lecture Notes in Computer Science, 2352:148– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520148.
pdf.

Luong:2002:RRV

[1043] Q.-Tuan Luong, Pascal Fua, and Yvan Leclerc. Recovery of reflectances and varying illuminants from multiple views. Lecture Notes in Computer Science, 2352:163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520163.
pdf.

Zalesny:2002:CTD

[1044] Alexey Zalesny, Vittorio Ferrari, Geert Caenen, Dominik Auf der Maur, and Luc Van Gool. Composite texture descriptions. Lecture Notes in Computer Science, 2352:180–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520180.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520180.
pdf.

Nakashima:2002:CII

[1045] Akiko Nakashima, Atsuto Maki, and Kazuhiro Fukui. Constructing illumination image basis from object motion. Lecture Notes in Computer Science, 2352:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520195.
pdf.

Lin:2002:DSS

[1046] Stephen Lin, Yuanzhen Li, Sing Bing Kang, Xin Tong, and Heung-Yeung Shum. Diffuse-specular separation and depth recovery from image sequences. Lecture Notes in Computer Science, 2352:210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520210.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2352/23520210.
                                        pdf.

Forsyth:2002:STB

[1047] D. A. Forsyth. Shape from texture without boundaries. Lecture Notes in Computer Science, 2352:225– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520225.
                                        pdf.

Wu:2002:SMT

[1048] Ying Nian Wu, Song-Chun Zhu, and Cheng en Guo. Statistical modeling of texture sketch. Lecture Notes in Computer Science, 2352:240–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520240.
                                        pdf.

Varma:2002:CIM

[1049] Manik Varma and Andrew Zisserman. Classifying images of materials: Achieving viewpoint and illumination independence. Lecture Notes in Computer Science, 2352:255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520255.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2352/23520255.
pdf.

Wang:2002:EMI

[1050] Yang Wang and Dimitris Samaras. Estimation of multiple illuminants from a single image of arbitrary known geometry. Lecture Notes in Computer Science, 2352:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520272.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520272.
pdf.

Chantler:2002:EIR

[1051] M. Chantler, M. Schmidt, M. Petrou, and G. McGunnigle. The effect of illuminant rotation on texture filters: Lissajous’s ellipses. Lecture Notes in Computer Science, 2352:289– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520289.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520289.
pdf.

Fitzgibbon:2002:AIC

[1052] Andrew Fitzgibbon and Andrew Zisserman. On affine invariant clustering and automatic cast listing in movies. Lecture Notes in Computer Science, 2352: 304–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520304.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520304.
                                        pdf.

Kim:2002:FMR

[1053] Junhwan Kim and Ramin Zabih. Factorial Markov random fields. Lecture Notes in Computer Science, 2352:321– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520321.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520321.
                                        pdf.

Kanatani:2002:ESM

[1054] Kenichi Kanatani. Evaluation and selection of models for motion segmentation. Lecture Notes in Computer Science, 2352:335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2352/23520335.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2352/23520335.pdf.

Tohka:2002:SEV

[1055] Jussi Tohka. Surface extraction from volumetric images using deformable meshes: a comparative study. Lecture Notes in Computer Science, 2352:350– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520350.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520350.
                                        pdf.

Jehan-Besson:2002:DDR

[1056] Stéphanie Jehan-Besson, Michel Barlaud, and Gilles Aubert. DREAM^{2} S: Deformable regions driven by an Eulerian accurate minimization method for image and video segmentation. Lecture Notes in Computer Science, 2352: 365–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520365.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520365.
pdf.

Lo:2002:NFS

[1057] Benny P. L. Lo and Guang-Zhong Yang. Neuro-fuzzy shadow filter. Lecture Notes in Computer Science, 2352: 381–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520381.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520381.
pdf.

Tu:2002:PIR

[1058] Zhuowen Tu and Song-Chun Zhu.

Parsing images into region and curve processes. Lecture Notes in Computer Science, 2352:393–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520393.
pdf.

Freixenet:2002:YAS

[1059] J. Freixenet, X. Muñoz, D. Raba, J. Martí, and X. Cufí. Yet another survey on image segmentation: Region and boundary information integration. Lecture Notes in Computer Science, 2352:408–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520408.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520408.
pdf.

Nicolescu:2002:PGM

[1060] Mircea Nicolescu and Gérard Medioni. Perceptual grouping from motion cues using tensor voting in 4-D. Lecture Notes in Computer Science, 2352:423– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520423.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520423.
pdf.

Taton:2002:DMN

[1061] Benjamin Taton and Jacques-Olivier Lachaud. Deformable model with non-euclidean metrics. Lecture Notes in Computer Science, 2352:438–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520438.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520438.
pdf.

Coughlan:2002:FDS

[1062] James M. Coughlan and Sabino J. Ferreira. Finding deformable shapes using loopy belief propagation. Lecture Notes in Computer Science, 2352:453– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520453.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520453.
pdf.

Hayman:2002:PVA

[1063] Eric Hayman and Jan-Olof Eklundh. Probabilistic and voting approaches to cue integration for figureground segmentation. Lecture Notes in Computer Science, 2352:469–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520469.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520469.
pdf.

Wexler:2002:BEL

[1064] Y. Wexler, A. Fitzgibbon, and A. Zisserman. Bayesian estimation of layers from multiple images. Lecture Notes in Computer Science, 2352:487– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520487.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520487.
pdf.

Han:2002:SAS

[1065] Feng Han, Zhouwen Tu, and Song-Chun Zhu. A stochastic algorithm for 3D scene segmentation and reconstruction. Lecture Notes in Computer Science, 2352:502–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520502.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520502.
pdf.

Yu:2002:NGV

[1066] Zeyun Yu and Chandrajit Bajaj. Normalized gradient vector diffusion and image segmentation. Lecture Notes in Computer Science, 2352:517–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520517.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520517.
pdf.

Belongie:2002:SPI

[1067] Serge Belongie, Charless Fowlkes, Fan Chung, and Jitendra Malik. Spectral partitioning with indefinite kernels using the nyström extension. Lecture Notes in Computer Science, 2352:531– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520531.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520531.
pdf.

Harville:2002:FHL

[1068] Michael Harville. A framework for high-level feedback to adaptive, per-pixel, mixture-of-Gaussian background models. Lecture Notes in Computer Science, 2352:543–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520543.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520543.
pdf.

Comaniciu:2002:MSP

[1069] Dorin Comaniciu, Visvanathan Ramesh, and Alessio Del Bue. Multivariate saddle point detection for statistical clustering. Lecture Notes in Computer Science, 2352:561–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520561.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520561.
pdf.

Hermes:2002:PDC

[1070] Lothar Hermes, Thomas Zöller, and Joachim M. Buhmann. Parametric distributional clustering for image segmentation. Lecture Notes in Computer Science, 2352:577–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520577.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2352/23520577.
                                        pdf.

Fisher:2002:PMI

[1071] John W. Fisher and Trevor Darrell. Probabalistic models and informative subspaces for audiovisual correspondence. Lecture Notes in Computer Science, 2352:592–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520592.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520592.
                                        pdf.

August:2002:VFN

[1072] Jonas August. Volterra filtering of noisy images of curves. Lecture Notes in Computer Science, 2352:604–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520604.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520604.
                                        pdf.

Rivera:2002:ISF

[1073] Mariano Rivera and James Gee. Image segmentation by flexible models based on robust regularized networks. Lecture Notes in Computer Science, 2352: 621–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520621.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2352/23520621.
pdf.

Levin:2002:PCA

[1074] Anat Levin and Amnon Shashua.

Principal component analysis over continuous subspaces and intersection of half-spaces. Lecture Notes in Computer Science, 2352:635–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520635.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520635.
pdf.

Lazebnik:2002:PTP

[1075] Svetlana Lazebnik, Amit Sethi, Cordelia Schmid, David J. Kriegman, Jean Ponce, and Martial Hebert. On pencils of tangent planes and the recognition of smooth 3D shapes from silhouettes. Lecture Notes in Computer Science, 2352:651–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520651.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520651.
pdf.

Mori:2002:EHB

[1076] Greg Mori and Jitendra Malik. Estimating human body configurations using shape context matching. Lecture Notes in Computer Science, 2352:666– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520666.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520666.
                                        pdf.

Zhou:2002:PHR

[1077] Shaohua Zhou and Rama Chellappa. Probabilistic human recognition from video. Lecture Notes in Computer Science, 2352:681–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520681.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520681.
                                        pdf.

David:2002:SSP

[1078] Philip David, Daniel DeMenthon, Ramani Duraiswami, and Hanan Samet. SoftPOSIT: Simultaneous pose and correspondence determination. Lecture Notes in Computer Science, 2352:698– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520698.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520698.
                                        pdf.

Giannopoulos:2002:PMW

[1079] Panos Giannopoulos and Remco C. Veltkamp. A pseudo-metric for weighted point sets. Lecture Notes in Computer Science, 2352:715–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2352/23520715.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520715.
pdf.

Sebastian:2002:SBI

[1080] Thomas B. Sebastian, Philip N.

Klein, and Benjamin B. Kimia. Shock-based indexing into large shape databases. Lecture Notes in Computer Science, 2352:731–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520731.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520731.
pdf.

Avidan:2002:EST

[1081] Shai Avidan. EigenSegments: a spatio-temporal decomposition of an ensemble of images. Lecture Notes in Computer Science, 2352:747–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520747.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520747.
pdf.

Shokoufandeh:2002:RMQ

[1082] Ali Shokoufandeh, Sven Dickinson, Clas Jönsson, Lars Bretzner, and Tony Lindeberg. On the representation and matching of qualitative shape at multiple scales. Lecture Notes in Computer Science, 2352:759– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520759.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520759.
                                        pdf.

Mahamud:2002:CSD

[1083] Shyjan Mahamud, Martial Hebert, and John Lafferty. Combining simple discriminators for object discrimination. Lecture Notes in Computer Science, 2352:776–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2352/23520776.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520776.
                                        pdf.

Hillenbrand:2002:PSO

[1084] Ulrich Hillenbrand and Gerd Hirzinger. Probabilistic search for object segmentation and recognition. Lecture Notes in Computer Science, 2352:791– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520791.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520791.
                                        pdf.

Gerard:2002:RTI

[1085] Olivier Gérard, Thomas Deschamps, Myriam Greff, and Laurent D. Cohen. Real-time interactive path extraction with on-the-fly adaptation of the external forces. Lecture Notes in Computer Science, 2352:807–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520807.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520807.
pdf.

Torsello:2002:MET

[1086] Andrea Torsello and Edwin R. Hancock. Matching and embedding through edit-union of trees. Lecture Notes in Computer Science, 2352:822– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520822.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520822.
pdf.

Breuel:2002:CSS

[1087] Thomas M. Breuel. A comparison of search strategies for geometric branch and bound algorithms. Lecture Notes in Computer Science, 2352:837– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2352/23520837.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2352/23520837.
pdf.

Shakhnarovich:2002:FRL

[1088] Gregory Shakhnarovich, John W.

Fisher, and Trevor Darrell. Face recognition from long-term observations. Lecture Notes in Computer Science, 2352:851–??, 2002. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2352/23520851.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520851.
                                        pdf.

Zickler:2002:HSE

[1089] Todd Zickler, Peter N. Belhumeur, and David J. Kriegman. Helmholtz stereopsis: Exploiting reciprocity for surface reconstruction. Lecture Notes in Computer Science, 2352:869–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520869.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520869.
                                        pdf.

Buehler:2002:MSS

[1090] Chris Buehler, Steven J. Gortler, Michael F. Cohen, and Leonard McMillan. Minimal surfaces for stereo. Lecture Notes in Computer Science, 2352: 885–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2352/23520885.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2352/23520885.
                                        pdf.

Sara:2002:FLU

[1091] Radim Šára. Finding the largest unambiguous component of stereo matching. Lecture Notes in Computer Science, 2352:900–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2352/23520900.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2352/23520900.
pdf.

Anonymous:2002:AIs

[1092] Anonymous. Author index. Lecture Notes in Computer Science, 2352:915– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2352/2352auth.pdf.

Romdhani:2002:FIF

[1093] S. Romdhani, V. Blanz, and T. Vetter. Face identification by fitting a 3D morphable model using linear shape and texture error functions. Lecture Notes in Computer Science, 2353:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530003.
pdf.

Barla:2002:HKO

[1094] A. Barla, F. Odone, and A. Verri.

Hausdorff kernel for 3D object acquisition and detection. Lecture Notes in Computer Science, 2353:20– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530020.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2353/23530020.
pdf.

Everingham:2002:EIS

[1095] M. Everingham, H. Muller, and B. Thomas. Evaluating image segmentation algorithms using the Pareto front. Lecture Notes in Computer Science, 2353:34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2353/23530034.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2353/23530034.pdf.

Vogel:2002:PCO

[1096] J. Vogel and B. Schiele. On performance characterization and optimization for image retrieval. Lecture Notes in Computer Science, 2353:49– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530049.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530049.
pdf.

Li:2002:SLM

[1097] S. Z. Li, L. Zhu, Z. Zhang, A. Blake, H. Zhang, and H. Shum. Statistical learning of multi-view face detection. Lecture Notes in Computer Science, 2353:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530067.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2353/23530067.
pdf.

Adams:2002:DTL

[1098] N. J. Adams and C. K. I. Williams.

Dynamic trees: Learning to model outdoor scenes. Lecture Notes in Computer Science, 2353:82–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530082.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530082.
pdf.

Duygulu:2002:ORM

[1099] P. Duygulu, K. Barnard, J. F. G.

de Freitas, and D. A. Forsyth. Object recognition as machine translation: Learning a lexicon for a fixed image vocabulary. Lecture Notes in Computer Science, 2353:97–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530097.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530097.
pdf.

Agarwal:2002:LSR

[1100] S. Agarwal and D. Roth. Learning a sparse representation for object detection. Lecture Notes in Computer Science, 2353:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530113.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530113.
                                        pdf.

Manning:2002:SSC

[1101] R. Manning and C. Dyer. Stratified self calibration from screwtransform manifolds. Lecture Notes in Computer Science, 2353:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530131.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530131.
                                        pdf.

Fisher:2002:SOR

[1102] R. B. Fisher. Self-organization of randomly placed sensors. Lecture Notes in Computer Science, 2353:146– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530146.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530146.
                                        pdf.

Zhang:2002:CCO

[1103] Z. Zhang. Camera calibration with one-dimensional objects. Lecture Notes in Computer Science, 2353:161–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530161.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530161.
                                        pdf.

Deutscher:2002:ACC

[1104] J. Deutscher, M. Isard, and J. Mac-Cormick. Automatic camera calibration from a single Manhattan image. Lecture Notes in Computer Science, 2353:175–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2353/23530175.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2353/23530175.
pdf.

Grossberg:2002:WCK

[1105] M. D. Grossberg and S. K. Nayar.

What can be known about the radiometric response from images? Lecture Notes in Computer Science, 2353:189– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530189.
pdf.

Zhou:2002:EID

[1106] W. Zhou and C. Kambhamettu. Estimation of illuminant direction and intensity of multiple light sources. Lecture Notes in Computer Science, 2353: 206–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530206.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530206.
pdf.

Wilczkowiak:2002:MUG

[1107] M. Wilczkowiak, E. Boyer, and P. Sturm. 3D modelling using geometric constraints: a parallelepiped based approach. Lecture Notes in Computer Science, 2353:221–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530221.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530221.
pdf.

Barreto:2002:GPC

[1108] J. P. Barreto and H. Araujo. Geometric properties of central catadioptric line images. Lecture Notes in Computer Science, 2353:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530237.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530237.
pdf.

Gurdjos:2002:AWL

[1109] P. Gurdjos, A. Crouzil, and R. Payrissat. Another way of looking at plane-based calibration: The centre circle constraint. Lecture Notes in Computer Science, 2353:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530252.
pdf.

Mitran:2002:ASR

[1110] M. Mitran and F. P. Ferrie. Active surface reconstruction using the gradient strategy. Lecture Notes in Computer Science, 2353:267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530267.
pdf.

Ansar:2002:LPE

[1111] A. Ansar and K. Daniilidis. Linear pose estimation from points or lines. Lecture Notes in Computer Science, 2353: 282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530282.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530282.
pdf.

Kam:2002:VBD

[1112] A. H. Kam, W. Lu, and W.- Y. Yau. A video-based drowning detection system. Lecture Notes in Computer Science, 2353:297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530297.
pdf.

Flandin:2002:VDF

[1113] G. Flandin and F. Chaumette. Visual data fusion for objects localization by active vision. Lecture Notes in Computer Science, 2353:312–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530312.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530312.
                                        pdf.

DeCarlo:2002:TRT

[1114] D. DeCarlo. Towards real-time cue integration by using partial results. Lecture Notes in Computer Science, 2353: 327–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530327.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530327.
                                        pdf.

Javed:2002:TOC

[1115] O. Javed and M. Shah. Tracking and object classification for automated surveillance. Lecture Notes in Computer Science, 2353:343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530343.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530343.
                                        pdf.

Schweitzer:2002:VFT

[1116] H. Schweitzer, J. W. Bell, and F. Wu.

Very fast template matching. Lecture Notes in Computer Science, 2353:358– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530358.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530358.
pdf.

Siebel:2002:FMT

[1117] N. T. Siebel and S. Maybank. Fusion of multiple tracking algorithms for robust people tracking. Lecture Notes in Computer Science, 2353:373– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530373.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530373.
pdf.

Aner:2002:VST

[1118] A. Aner and J. R. Kender. Video summaries through mosaic-based shot and scene clustering. Lecture Notes in Computer Science, 2353:388–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530388.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530388.
pdf.

Liu:2002:OAS

[1119] T. Liu and J. R. Kender. Optimization algorithms for the selection of key frame sequences of variable length. Lecture Notes in Computer Science, 2353: 403–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530403.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530403.
                                        pdf.

Granger:2002:MSE

[1120] S. Granger and X. Pennec. Multiscale EM-ICP: a fast and robust approach for surface registration. Lecture Notes in Computer Science, 2353:418– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530418.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530418.
                                        pdf.

Malis:2002:UAM

[1121] E. Malis. An unified approach to model-based and model-free visual servoing. Lecture Notes in Computer Science, 2353:433–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2353/23530433.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2353/23530433.pdf.

Mindru:2002:CIT

[1122] F. Mindru, T. Moons, and L. Van Gool. Comparing intensity transformations and their invariants in the context of color pattern recognition. Lecture Notes in Computer Science, 2353: 448–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530448.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530448.
pdf.

Greenspan:2002:PFS

[1123] H. Greenspan, J. Goldberger, and A. Mayer. A probabilistic framework for spatio-temporal video representation & indexing. Lecture Notes in Computer Science, 2353:461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530461.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530461.
pdf.

Kosecka:2002:VC

[1124] J. Košecká and W. Zhang. Video compass. Lecture Notes in Computer Science, 2353:476–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2353/23530476.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2353/23530476.pdf.

Schweitzer:2002:CCP

[1125] H. Schweitzer. Computing contentplots for video. Lecture Notes in Computer Science, 2353:491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530491.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530491.
                                        pdf.

Osareh:2002:CLD

[1126] A. Osareh, M. Mirmehdi, B. Thomas, and R. Markham. Classification and localisation of diabeticrelated eye disease. Lecture Notes in Computer Science, 2353:502–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530502.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530502.
                                        pdf.

Rogers:2002:RAS

[1127] M. Rogers and J. Graham. Robust active shape model search. Lecture Notes in Computer Science, 2353:517– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530517.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530517.
                                        pdf.

Richard:2002:NIR

[1128] F. Richard and L. Cohen. A new image registration technique with free boundary constraints: Application to mammography. Lecture Notes in Computer Science, 2353:531–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530531.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530531.
pdf.

Vemuri:2002:RAI

[1129] B. C. Vemuri, Y. Chen, and Z. Wang.

Registration assisted image smoothing and segmentation. Lecture Notes in Computer Science, 2353:546–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530546.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530546.
pdf.

Marroquin:2002:AEB

[1130] J. L. Marroquin, B. C. Vemuri, S. Botello, and F. Calderon. An accurate and efficient Bayesian method for automatic segmentation of brain MRI. Lecture Notes in Computer Science, 2353:560–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2353/23530560.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2353/23530560.
pdf.

Yezzi:2002:PAT

[1131] A. Yezzi and J. L. Prince. A PDE approach for thickness, correspondence, and gridding of annular tissues. Lecture Notes in Computer Science, 2353: 575–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530575.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530575.
                                        pdf.

Gao:2002:SCM

[1132] X. Gao, V. Ramesh, and T. Boult. Statistical characterization of morphological operator sequences. Lecture Notes in Computer Science, 2353:590– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530590.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530590.
                                        pdf.

Dornaika:2002:IRF

[1133] F. Dornaika and J. Elder. Image registration for foveated omnidirectional sensing. Lecture Notes in Computer Science, 2353:606–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530606.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530606.
                                        pdf.

Cootes:2002:AMS

[1134] T. F. Cootes, N. Thacker, and C. J. Taylor. Automatic model selection by modelling the distribution of residuals. Lecture Notes in Computer Science, 2353:621–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530621.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530621.
pdf.

Nayar:2002:APM

[1135] S. K. Nayar and S. G. Narasimhan.

Assorted pixels: Multi-sampled imaging with structural models. Lecture Notes in Computer Science, 2353:636– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530636.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530636.
pdf.

DelaTorre:2002:RPC

[1136] F. De la Torre and M. J. Black. Robust parameterized component analysis. Lecture Notes in Computer Science, 2353:653–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2353/23530653.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2353/23530653.
pdf.

Ng:2002:LIV

[1137] J. Ng and S. Gong. Learning intrinsic video content using Levenshtein distance in graph partitioning. Lecture Notes in Computer Science, 2353:670– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530670.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530670.
                                        pdf.

Yang:2002:TTC

[1138] M.-H. Yang, D. Roth, and N. Ahuja. A tale of two classifiers: SNoW vs. SVM in visual recognition. Lecture Notes in Computer Science, 2353:685– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530685.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530685.
                                        pdf.

Ronfard:2002:LPP

[1139] R. Ronfard, C. Schmid, and B. Triggs. Learning to parse pictures of people. Lecture Notes in Computer Science, 2353:700–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2353/23530700.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530700.
                                        pdf.

Pal:2002:LMT

[1140] C. Pal, B. J. Frey, and N. Jojic. Learning montages of transformed latent images as representations of objects that change in appearance. Lecture Notes in Computer Science, 2353:715– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530715.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530715.
pdf.

Kruger:2002:EBF

[1141] V. Krüger and S. Zhou. Exemplarbased face recognition from video. Lecture Notes in Computer Science, 2353: 732–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530732.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530732.
pdf.

Wieghardt:2002:LTO

[1142] J. Wieghardt, R. P. Würtz, and C. von der Malsburg. Learning the topology of object views. Lecture Notes in Computer Science, 2353:747– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530747.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530747.
pdf.

Skocaj:2002:RPA

[1143] D. Skočaj, H. Bischof, and A. Leonardis. A robust PCA algorithm for building representations from panoramic images. Lecture Notes in Computer Science, 2353:761–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2353/23530761.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2353/23530761.pdf.

Shental:2002:ALR

[1144] N. Shental, T. Hertz, D. Weinshall, and M. Pavel. Adjustment learning and relevant component analysis. Lecture Notes in Computer Science, 2353:776– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530776.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530776.
                                        pdf.

Zhu:2002:WT

[1145] S.-C. Zhu, C. e. Guo, Y. Wu, and Y. Wang. What are textons? Lecture Notes in Computer Science, 2353: 793–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2353/23530793.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2353/23530793.
                                        pdf.

Pont:2002:BTC

[1146] S. C. Pont and J. J. Koenderink.

Bidirectional texture contrast function. Lecture Notes in Computer Science, 2353:808–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2353/23530808.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2353/23530808.
pdf.

Finlayson:2002:RSI

[1147] G. D. Finlayson, S. D. Hordley, and M. S. Drew. Removing shadows from images. Lecture Notes in Computer Science, 2353:823–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2353/23530823.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2353/23530823.
pdf.

Anonymous:2002:AIt

[1148] Anonymous. Author index. Lecture Notes in Computer Science, 2353:837– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2353/2353auth.pdf.

Lucks:2002:SAB

[1149] Stefan Lucks. The saturation attack — A bait for twofish. Lecture Notes in Computer Science, 2355:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550001.
pdf.

Biham:2002:LCR

[1150] Eli Biham, Orr Dunkelman, and Nathan Keller. Linear cryptanalysis of reduced round serpent. Lecture Notes in Computer Science, 2355:16– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550016.
pdf.

Fluhrer:2002:CMB

[1151] Scott R. Fluhrer. Cryptanalysis of the mercy block cipher. Lecture Notes in Computer Science, 2355:28– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550028.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550028.
pdf.

Rijmen:2002:PCP

[1152] Vincent Rijmen, Bart Van Rompay, Bart Preneel, and Joos Vandewalle. Producing collisions for PANAMA. Lecture Notes in Computer Science, 2355:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

0558/bibs/2355/23550037.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2355/23550037.
pdf.

Debaert:2002:RRI

[1153] Christophe Debaert and Henri Gilbert.

The RIPEMD and RIPEMD improved variants of MD4 are not collision free. Lecture Notes in Computer Science, 2355:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2355/23550052.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2355/23550052.
pdf.

Tarannikov:2002:NCR

[1154] Yuriy Tarannikov. New constructions of resilient Boolean functions with maximal nonlinearity. Lecture Notes in Computer Science, 2355:66– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550066.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550066.
pdf.

Alkassar:2002:OSS

[1155] Ammar Alkassar, Alexander Geraldy, Birgit Pfitzmann, and Ahmad-Reza Sadeghi. Optimized self-synchronizing mode of operation. Lecture Notes in Computer Science, 2355:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2355/23550078.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2355/23550078.
                                        pdf.

Gligor:2002:FEA

[1156] Virgil D. Gligor and Pompiliu Donescu. Fast encryption and authentication: XCBC encryption and XECB authentication modes. Lecture Notes in Computer Science, 2355:92–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2355/23550092.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2355/23550092.
                                        pdf.

Buonanno:2002:IUE

[1157] Enrico Buonanno, Jonathan Katz, and Moti Yung. Incremental unforgeable encryption. Lecture Notes in Computer Science, 2355:109–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2355/23550109.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2355/23550109.
                                        pdf.

Stay:2002:ZAR

[1158] Michael Stay. ZIP attacks with reduced known plaintext. Lecture Notes in Computer Science, 2355:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2355/23550125.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2355/23550125.
pdf.

Fluhrer:2002:CSP

[1159] Scott R. Fluhrer. Cryptanalysis of the SEAL 3.0 pseudorandom function family. Lecture Notes in Computer Science, 2355:135–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550135.
pdf.

Jakimovski:2002:CS

[1160] Goce Jakimovski and Ljupčo Kocarev. Cryptanalysis of SBLH. Lecture Notes in Computer Science, 2355: 144–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550144.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550144.
pdf.

Mantin:2002:PAB

[1161] Itsik Mantin and Adi Shamir. A practical attack on broadcast RC4. Lecture Notes in Computer Science, 2355: 152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550152.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550152.
pdf.

Barreto:2002:IQA

[1162] Paulo S. L. M. Barreto, Vincent Rijmen, Jorge Nakahara Jr., Bart Preneel, Joos Vandewalle, and Hae Y. Kim. Improved S QUARE attacks against reduced-round H IEROCRYPT. Lecture Notes in Computer Science, 2355: 165–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2355/23550165.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2355/23550165.
                                        pdf.

Biham:2002:DCQ

[1163] Eli Biham, Vladimir Furman, Michal Misztal, and Vincent Rijmen. Differential cryptanalysis of Q. Lecture Notes in Computer Science, 2355:174– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2355/23550174.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2355/23550174.
                                        pdf.

Furman:2002:DCN

[1164] Vladimir Furman. Differential cryptanalysis of nimbus. Lecture Notes in Computer Science, 2355:187–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2355/23550187.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2355/23550187.
                                        pdf.

Mihaljevic:2002:FCA

[1165] Miodrag J. Mihaljević, Marc P. C. Fossorier, and Hideki Imai. Fast correlation attack algorithm with list decoding and an application. Lecture Notes in Computer Science, 2355:196– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550196.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550196.
pdf.

Crowley:2002:BLS

[1166] Paul Crowley and Stefan Lucks. Bias in the LEVIATHAN stream cipher. Lecture Notes in Computer Science, 2355: 211–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550211.
pdf.

Bleichenbacher:2002:AS

[1167] Daniel Bleichenbacher and Willi Meier.

Analysis of SSC2. Lecture Notes in Computer Science, 2355:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550219.
pdf.

Iwata:2002:RSS

[1168] Tetsu Iwata, Tomonobu Yoshino, Tomohiro Yuasa, and Kaoru Kurosawa. Round security and superpseudorandomness of MISTY type structure. Lecture Notes in Computer Science, 2355:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2355/23550233.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2355/23550233.pdf.

Gilbert:2002:NRP

[1169] Henri Gilbert and Marine Minier. New results on the pseudorandomness of some blockcipher constructions. Lecture Notes in Computer Science, 2355: 248–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550248.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550248.
pdf.

Preneel:2002:NEA

[1170] Bart Preneel. NESSIE: a European approach to evaluate cryptographic algorithms. Lecture Notes in Computer Science, 2355:267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550267.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550267.
pdf.

Blunden:2002:RKA

[1171] Mark Blunden and Adrian Escott.

Related key attacks on reduced round KASUMI. Lecture Notes in Computer Science, 2355:277–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550277.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550277.
pdf.

Kanda:2002:SCA

[1172] Masayuki Kanda and Tsutomu Matsumoto. Security of Camellia against truncated differential cryptanalysis. Lecture Notes in Computer Science, 2355:286–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2355/23550286.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2355/23550286.
pdf.

Hong:2002:IDC

[1173] Deukjo Hong, Jaechul Sung, Shiho Moriai, SangjinLee, and Jongin Lim. Impossible differential cryptanalysis of zodiac. Lecture Notes in Computer Science, 2355:300–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2355/23550300.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2355/23550300.
pdf.

Shimoyama:2002:BCS

[1174] Takeshi Shimoyama, Hitoshi Yanami, Kazuhiro Yokoyama, Masahiko Takenaka, Kouichi Itoh, Jun Yajima, Naoya Torii, and Hidema Tanaka. The block cipher SC2000. Lecture Notes in Computer Science, 2355:312– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2355/23550312.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2355/23550312.
                                        pdf.

Granboulan:2002:FDC

[1175] Louis Granboulan. Flaws in differential cryptanalysis of skipjack. Lecture Notes in Computer Science, 2355:328– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2355/23550328.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2355/23550328.
                                        pdf.

Lipmaa:2002:EAC

[1176] Helger Lipmaa and Shiho Moriai. Efficient algorithms for computing differential properties of addition. Lecture Notes in Computer Science, 2355:336– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2355/23550336.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2355/23550336.
                                        pdf.

Anonymous:2002:AIu

[1177] Anonymous. Author index. Lecture Notes in Computer Science, 2355:351– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2355/2355auth.pdf.

Berendt:2002:DCW

[1178] Bettina Berendt. Detail and context in Web usage mining: Coarsening and visualizing sequences. Lecture Notes in Computer Science, 2356:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2356/23560001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2356/23560001.
pdf.

Geyer-Schulz:2002:CPI

[1179] Andreas Geyer-Schulz, Michael Hahsler, and Maximillian Jahn. A customer purchase incidence model applied to recommender services. Lecture Notes in Computer Science, 2356: 25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2356/23560025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2356/23560025.
pdf.

Huang:2002:CMC

[1180] Joshua Zhexue Huang, Michael Ng, Wai-Ki Ching, Joe Ng, and David Cheung. A cube model and cluster analysis for Web access sessions. Lecture Notes in Computer Science, 2356:48– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2356/23560048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2356/23560048.
pdf.

Nanopoulos:2002:EWL

[1181] Alexandros Nanopoulos, Dimitrios Katsaros, and Yannis Manolopoulos. Exploiting Web log mining for Web cache enhancement. Lecture Notes in Computer Science, 2356:68– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2356/23560068.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2356/23560068.
pdf.

Punin:2002:LLM

[1182] John R. Punin, Mukkai S. Krishnamoorthy, and Mohammed J. Zaki. LOGML: Log markup language for Web usage mining. Lecture Notes in Computer Science, 2356:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2356/23560088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2356/23560088.
pdf.

Shahabi:2002:FEA

[1183] Cyrus Shahabi and Farnoush Banaei-Kashani. A framework for efficient and anonymous Web usage mining based on client-side tracking. Lecture Notes in Computer Science, 2356:113– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2356/23560113.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2356/23560113.
pdf.

Tan:2002:MIA

[1184] Pang-Ning Tan and Vipin Kumar. Mining indirect associations in Web data. Lecture Notes in Computer Science, 2356:145–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2356/23560145.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2356/23560145.
pdf.

Anonymous:2002:AIv

[1185] Anonymous. Author index. Lecture Notes in Computer Science, 2356:167– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2356/2356auth.pdf.

Michalopoulos:2002:EBP

[1186] Demetrios Michalopoulos and Chih-Kang Hu. An error back-propagation artificial neural networks application in automatic car license plate recognition. Lecture Notes in Computer Science, 2358:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2358/23580001.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580001.
                                        pdf.

Levinski:2002:UAN

[1187] Oleg Levinski. Use of artificial neural networks for buffet loads prediction. Lecture Notes in Computer Science, 2358:9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2358/23580009.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580009.
                                        pdf.

Kim:2002:CCR

[1188] In-Cheol Kim and Sung-Il Chien. Computational cost reduction by selective attention for fast speaker adaptation in multilayer perceptron. Lecture Notes in Computer Science, 2358:17– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580017.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580017.
                                        pdf.

Yim:2002:CNN

[1189] Juliana Yim. A comparison of neural networks with time series models for forecasting returns on a stock market index. Lecture Notes in Computer Science, 2358:25–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580025.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580025.
pdf.

Walsh:2002:ADF

[1190] Dane Walsh and Christian Omlin. Automatic detection of film orientation with support vector machines. Lecture Notes in Computer Science, 2358:36– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580036.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580036.
pdf.

Quan:2002:GAV

[1191] Vu Hai Quan, Pham Nam Trung, Nguyen Duc Hoang Ha, Lam Tri Tin, Hoang Kiem, and An H. Nguyen. A generic approach for the Vietnamese handwritten and speech recognition problems. Lecture Notes in Computer Science, 2358:47–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580047.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580047.
pdf.

Lee:2002:EAF

[1192] Bae-Ho Lee, Kwang-Hee Kim, Yonggwan Won, and Jiseung Nam. Efficient and automatic faces detection based on skin-tone and neural network model. Lecture Notes in Computer Science, 2358:57–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580057.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580057.
pdf.

Kim:2002:EIS

[1193] Jong-Bae Kim and Hang-Joon Ki. Efficient image segmentation based on wavelet and watersheds for video objects extraction. Lecture Notes in Computer Science, 2358:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580067.
pdf.

Skliarova:2002:FBI

[1194] Iouliia Skliarova and António B. Ferrari. FPGA-based implementation of genetic algorithm for the traveling salesman problem and its industrial application. Lecture Notes in Computer Science, 2358:77–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580077.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580077.
pdf.

Nedjah:2002:MAC

[1195] Nadia Nedjah and Luiza de Macedo Mourelle.

Minimal addition chain for efficient modular exponentiation using genetic algorithms. Lecture Notes in Computer Science, 2358:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580088.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580088.
pdf.

Sklyarov:2002:EAS

[1196] Valery Sklyarov. An evolutionary algorithm for the synthesis of RAM-based FSMs. Lecture Notes in Computer Science, 2358:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580108.
pdf.

Chau:2002:GAD

[1197] K. W. Chau and F. Albermani. Genetic algorithms for design of liquid retaining structure. Lecture Notes in Computer Science, 2358:119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580119.
pdf.

Vollebregt:2002:MCA

[1198] Arjen Vollebregt, Daan Hannessen, Henk Hesselink, and Jelle Beetstra. Modelling crew assistants with multiagent systems in fighter aircraft. Lecture Notes in Computer Science, 2358: 129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580129.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580129.
pdf.

Thawonmas:2002:LHD

[1199] Ruck Thawonmas, Junichiro Hirayama, and Fumiaki Takeda. Learning from human decision-making behaviors — an application to RoboCup software agents. Lecture Notes in Computer Science, 2358:136–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580136.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580136.
pdf.

Ashfield:2002:DDD

[1200] Bruce Ashfield, Dwight Deugo, Franz Oppacher, and Tony White. Distributed deadlock detection in mobile agent systems. Lecture Notes in Computer Science, 2358:146–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580146.htm;   http:/
/link.springer-ny.com/link/service/

series/0558/papers/2358/23580146.
pdf.

Zudor:2002:ABA

[1201] Elisabeth Ilie Zudor and László Monostori. An agent-based approach for production control incorporating environmental and life-cycle issues, together with sensitivity analysis. Lecture Notes in Computer Science, 2358: 157–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580157.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580157.
pdf.

Randall:2002:FRI

[1202] Marcus Randall. Feasibility restoration for iterative meta-heuristics search algorithms. Lecture Notes in Computer Science, 2358:168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580168.
pdf.

Nyakoe:2002:OPP

[1203] George Nyauma Nyakoe, Makoto Ohki, Suichiro Tabuchi, and Masaaki Ohkita. Optimization of pulse pattern for a multi-robot sonar system using genetic algorithm. Lecture Notes in Computer Science, 2358:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580179.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580179.
                                        pdf.

Braendler:2002:SPS

[1204] David Braendler and Tim Hentdlass. The suitability of particle swarm optimisation for training neural hardware. Lecture Notes in Computer Science, 2358:190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580190.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580190.
                                        pdf.

Laumanns:2002:EMO

[1205] Nando Laumanns, Marco Laumanns, and Hartmut Kitterer. Evolutionary multi-objective integer programming for the design of adaptive cruise control systems. Lecture Notes in Computer Science, 2358:200–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580200.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580200.
                                        pdf.

Hendtlass:2002:MES

[1206] Tim Hendtlass and Gerrard Murray. The macronet element: a substitute for the conventional neuron. Lecture Notes in Computer Science, 2358:211– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580211.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580211.
pdf.

Dunn:2002:GAO

[1207] S. Dunn and S. Peucker. Genetic algorithm optimisation of mathematical models using distributed computing. Lecture Notes in Computer Science, 2358:220–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2358/23580220.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2358/23580220.
pdf.

Crispin:2002:GAO

[1208] Alan Crispin, Paul Clay, Gaynor Taylor, Robert Hackney, Tom Bayes, and David Reedman. Genetic algorithm optimisation of part placement using a connection-based coding method. Lecture Notes in Computer Science, 2358: 232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580232.
pdf.

Salami:2002:FEA

[1209] Mehrdad Salami and Tim Hendtlass.

A fast evolutionary algorithm for image compression in hardware. Lecture Notes in Computer Science, 2358:241– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580241.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580241.
                                        pdf.

Daoudi:2002:ASR

[1210] Khalid Daoudi. Automatic speech recognition: The new millennium. Lecture Notes in Computer Science, 2358: 253–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580253.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580253.
                                        pdf.

Loh:2002:AML

[1211] Gabriel H. Loh and Dana S. Henry. Applying machine learning for ensemble branch predictors. Lecture Notes in Computer Science, 2358:264– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580264.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580264.
                                        pdf.

Murtagh:2002:CCT

[1212] Niall Murtagh. A customizable configuration tool for design of multi-part products. Lecture Notes in Computer Science, 2358:275–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580275.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580275.
pdf.

Han:2002:PPW

[1213] Fengling Han, Xinghuo Yu, Yong Feng, and Huifeng Dong. Phaseto-phase wave parameters measurement of distribution lines based on BP networks. Lecture Notes in Computer Science, 2358:284–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580284.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580284.
pdf.

Debnath:2002:LCC

[1214] Rameswar Debnath and Haruhisa Takahashi. Learning capability: Classical RBF network vs. SVM with Gaussian kernel. Lecture Notes in Computer Science, 2358:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580293.
pdf.

Wang:2002:TBM

[1215] Dianhui Wang, Tharam Dillon, and Elizabeth Chang. Trading off between misclassification, recognition and generalization in data mining with continuous features. Lecture Notes in Computer Science, 2358:303–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580303.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580303.
                                        pdf.

Briscoe:2002:INM

[1216] Garry Briscoe. Interacting neural modules. Lecture Notes in Computer Science, 2358:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2358/23580314.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2358/23580314.pdf.

Zhou:2002:AVN

[1217] Zhi-Hua Zhou, Yuan Jiang, Xu-Ri Yin, and Shi-Fu Chen. The application of visualization and neural network techniques in a power transformer condition monitoring system. Lecture Notes in Computer Science, 2358:325– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580325.
                                        pdf.

Debenham:2002:EIE

[1218] John Debenham. Entrepreneurial intervention in an electronic market place. Lecture Notes in Computer Science, 2358:335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2358/23580335.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2358/23580335.pdf.

Natarajan:2002:IAD

[1219] Vikram Natarajan and Angela Goh. Intelligent auto-downloading of images. Lecture Notes in Computer Science, 2358:346–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2358/23580346.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2358/23580346.
pdf.

Chalidabhongse:2002:IFA

[1220] Junalux Chalidabhongse, Wirat Chinnan, Pichet Wechasaethnon, and Arpakorn Tantisirithanakorn. Intelligent facilitation agent for online Webbased group discussion system. Lecture Notes in Computer Science, 2358: 356–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580356.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580356.
pdf.

Ku:2002:TAR

[1221] Sebon Ku, Bogju Lee, and Dongman Lee. TWIMC: An anonymous recipient E-mail system. Lecture Notes in Computer Science, 2358:363–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580363.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580363.
                                        pdf.

Saba:2002:MSA

[1222] Walid S. Saba. Mental states of autonomous agents in competitive and cooperative settings. Lecture Notes in Computer Science, 2358:373–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580373.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580373.
                                        pdf.

Rossetto:2002:ESA

[1223] Silvana Rossetto, Flávio M. Varejão, and Thomas W. Rauber. An expert system application for improving results in a handwritten form recognition system. Lecture Notes in Computer Science, 2358:383–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580383.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580383.
                                        pdf.

Chau:2002:KBS

[1224] K. W. Chau and M. Anson. A knowledge-based system for construction site level facilities layout. Lecture Notes in Computer Science, 2358:393– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580393.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580393.
pdf.

Lee:2002:DSS

[1225] Dongkon Lee, Jaeyong Lee, and K. H. Lee. A decision-support system to improve damage survivability of submarine. Lecture Notes in Computer Science, 2358:403–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580403.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580403.
pdf.

Santos:2002:VES

[1226] Jorge Santos, Zita Vale, and Carlos Ramos. On the verification of an expert system: Practical issues. Lecture Notes in Computer Science, 2358:414– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580414.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580414.
pdf.

Mitrpanont:2002:DRB

[1227] J. L. Mitrpanont and T. Plengpung.

DOWNSIZINGX : a rule-based system for downsizing the corporation’s computer systems. Lecture Notes in Computer Science, 2358:425–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580425.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580425.
                                        pdf.

Hewahi:2002:CAS

[1228] Nabil M. Hewahi and H. Ahmad. Credit apportionment scheme for rulebased systems: Implementation and comparative study. Lecture Notes in Computer Science, 2358:435–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580435.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580435.
                                        pdf.

Tian:2002:AWC

[1229] Wen Tian, Ben Choi, and Vir V. Phoha. An adaptive Web cache access predictor using neural network. Lecture Notes in Computer Science, 2358: 450–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580450.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580450.
                                        pdf.

Matsuo:2002:DBR

[1230] Tokuro Matsuo and Takayuki Ito. A designated bid reverse auction for agent-based electronic commerce. Lecture Notes in Computer Science, 2358: 460–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580460.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580460.
pdf.

Phang:2002:DFU

[1231] K. K. Phang, S. H. Lim, Mashkuri Hj.

Yaacob, and T. C. Ling. Design of a fuzzy usage parameter controller for Diffserv and MPLS. Lecture Notes in Computer Science, 2358:470–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580470.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580470.
pdf.

Galitsky:2002:TER

[1232] Boris Galitsky. A tool for extension and restructuring natural language question answering domains. Lecture Notes in Computer Science, 2358:482– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580482.
pdf.

Jung:2002:ERI

[1233] Sung-Won Jung, Kyung-Hee Sung, Tae-Won Park, and Hyuk chul Kwon.

Effective retrieval of information in tables on the internet. Lecture Notes in Computer Science, 2358:493–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580493.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580493.
pdf.

Salami:2002:FES

[1234] Mehrdad Salami and Tim Hendtlass. A fitness estimation strategy for genetic algorithms. Lecture Notes in Computer Science, 2358:502– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580502.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580502.
pdf.

Runqiang:2002:DSM

[1235] Bian Runqiang, Phoebe Chen, Kevin Burrage, Jim Hanan, Peter Room, and John Belward. Derivation of L-system models from measurements of biological branching structures using genetic algorithms. Lecture Notes in Computer Science, 2358:514–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580514.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580514.
pdf.

Darwen:2002:ESB

[1236] Paul J. Darwen. Evolving a schedule with batching, precedence constraints, and sequence-dependent setup times: Crossover needs building blocks. Lecture Notes in Computer Science, 2358: 525–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580525.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580525.
pdf.

Mitrpanont:2002:DFE

[1237] J. L. Mitrpanont and Surasit Kiwprasopsak. The development of the feature extraction algorithms for Thai handwritten character recognition system. Lecture Notes in Computer Science, 2358:536–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2358/23580536.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2358/23580536.
pdf.

Ito:2002:RPW

[1238] Teruaki Ito. Route planning wizard:

Basic concept and its implementation. Lecture Notes in Computer Science, 2358:547–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2358/23580547.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2358/23580547.
pdf.

Zhang:2002:DIC

[1239] HaiYi Zhang, JianDong, and Barbro Back. The design and implementation of color matching system based on back propagation. Lecture Notes in Computer Science, 2358:557– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580557.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580557.
                                        pdf.

Mitra:2002:COP

[1240] Debasis Mitra and Walter P. Bond. Component-oriented programming as an AI-planning problem. Lecture Notes in Computer Science, 2358:567– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580567.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580567.
                                        pdf.

Mouhoub:2002:DCI

[1241] Malek Mouhoub and Jonathan Yip. Dynamic CSPs for interval-based temporal reasoning. Lecture Notes in Computer Science, 2358:575–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580575.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580575.
                                        pdf.

Lee:2002:EPM

[1242] Sangjun Lee, Dongseop Kwon, and Sukho Lee. Efficient pattern matching of time series data. Lecture Notes in Computer Science, 2358:586– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580586.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580586.
pdf.

Lavagna:2002:MAD

[1243] Michelle Lavagna and Amalia Ercoli Finzi. A multi-attribute decisionmaking approach toward space system design automation through a fuzzy logic-based analytic hierarchical process. Lecture Notes in Computer Science, 2358:596–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580596.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580596.
pdf.

Kravis:2002:CBS

[1244] Simon Kravis and Rosemary Irrgang.

A case based system for oil and gas well design. Lecture Notes in Computer Science, 2358:607–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580607.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2358/23580607.
                                        pdf.

Angus:2002:ACO

[1245] Daniel Angus and Tim Hendtlass. Ant colony optimisation applied to a dynamically changing problem. Lecture Notes in Computer Science, 2358:618– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580618.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580618.
                                        pdf.

Salim:2002:GII

[1246] M. D. Salim, Tim Strauss, and Michael Emch. A GIS-integrated intelligent system for optimization of asset management for maintenance of roads and bridges. Lecture Notes in Computer Science, 2358:628–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580628.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580628.
                                        pdf.

MonteirodeFarias:2002:UAS

[1247] Oscar Luiz Monteiro de Farias and Sueli Bandeira Teixeira Mendes. A unified approach for spatial object modelling and map analysis based on 2nd order many-sorted language. Lecture Notes in Computer Science, 2358:638– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2358/23580638.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580638.
pdf.

Viharos:2002:TAA

[1248] Zs. J. Viharos, L. Monostori, and T. Vincze. Training and application of artificial neural networks with incomplete data. Lecture Notes in Computer Science, 2358:649–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580649.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580649.
pdf.

Heeren:2002:MAR

[1249] Frank Heeren and Wilfried Sihn. Message analysis for the recommendation of contact persons within defined subject fields. Lecture Notes in Computer Science, 2358:660–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580660.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580660.
pdf.

Chau:2002:IKP

[1250] K. W. Chau, Chuntian Cheng, Y. S.

Li, C. W. Li, and O. Wai. An intelligent knowledge processing system on hydrodynamics and water quality modeling. Lecture Notes in Computer Science, 2358:670–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580670.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580670.
                                        pdf.

Chachoua:2002:UMI

[1251] M. Chachoua and D. Pacholczyk. Uncertainty management and informational relevance. Lecture Notes in Computer Science, 2358:680–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580680.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580680.
                                        pdf.

Kang:2002:PGR

[1252] Mi young Kang, Su ho Park, Ae sun Yoon, and Hyuk chul Kwon. Potential governing relationship and a Korean grammar checker using partial parsing. Lecture Notes in Computer Science, 2358:692–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580692.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580692.
                                        pdf.

Shin:2002:LHC

[1253] Jungpil Shin. On-line handwriting character recognition using stroke information. Lecture Notes in Computer Science, 2358:703–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2358/23580703.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2358/23580703.pdf.

Kim:2002:FDI

[1254] Jong-Bae Kim, Su-Woong Jung, and Hang-Joon Kim. Face detection by integrating multiresolution-based watersheds and a skin-color model. Lecture Notes in Computer Science, 2358:715– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580715.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580715.
pdf.

Okuno:2002:SIH

[1255] Hiroshi G. Okuno, Kazuhiro Nakadai, and Hiroaki Kitano. Social interaction of humanoid RobotBased on audio-visual tracking. Lecture Notes in Computer Science, 2358:725–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580725.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580725.
pdf.

Kim:2002:HCM

[1256] Jinyoung Kim, Joohun Lee, and Seungho Choi. Hybrid confidence measure for domain-specific keyword spotting. Lecture Notes in Computer Science, 2358:736–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580736.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580736.
                                        pdf.

Wotawa:2002:MBD

[1257] Franz Wotawa, Markus Stumptner, and Wolfgang Mayer. Model-based debugging or how to diagnose programs automatically. Lecture Notes in Computer Science, 2358:746–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580746.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580746.
                                        pdf.

Hiratsuka:2002:MBD

[1258] Satoshi Hiratsuka and Akira Fusaoka. On a model-based design verificationfor combinatorial Boolean networks. Lecture Notes in Computer Science, 2358: 758–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580758.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580758.
                                        pdf.

Nedjah:2002:OAP

[1259] Nadia Nedjah and Luiza de Macedo Mourelle. Optimal adaptive pattern matching.

Lecture Notes in Computer Science, 2358:768–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2358/23580768.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2358/23580768.
pdf.

Xia:2002:AAC

[1260] Mao Xia, Chen Bin, Zhu Gang, and Muta Itsuya. Analysis of affective characteristics and evaluationof harmonious feeling of image based on 1/f fluctuation theory. Lecture Notes in Computer Science, 2358:780–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580780.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580780.
pdf.

White:2002:CIP

[1261] Tony White, Bernard Pagurek, and Dwight Deugo. Collective intelligence and priority routing in networks. Lecture Notes in Computer Science, 2358: 790–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2358/23580790.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2358/23580790.
pdf.

Wernstedt:2002:ABA

[1262] Fredrik Wernstedt and Paul Davidsson. An agent-based approach to monitoring and control of district heating systems. Lecture Notes in Computer Science, 2358:801–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580801.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580801.
                                        pdf.

Bratko:2002:UML

[1263] Ivan Bratko and Dorian Šuc. Using machine learning to understand operator’s skill. Lecture Notes in Computer Science, 2358:812–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580812.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580812.
                                        pdf.

Islam:2002:RLC

[1264] Fakhrul Islam, Baikunth Nath, and Joarder Kamruzzaman. Reactive load control of parallel transformer operations using neural networks. Lecture Notes in Computer Science, 2358:824– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2358/23580824.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2358/23580824.
                                        pdf.

Anonymous:2002:AIw

[1265] Anonymous. Author index. Lecture Notes in Computer Science, 2358:831– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2358/2358ind.pdf.

Deniz:2002:ILA

[1266] O. Déniz, M. Castrillón, J. Lorenzo, and M. Hernández. An incremental learning algorithm for face recognition. Lecture Notes in Computer Science, 2359:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2359/23590001.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2359/23590001.
pdf.

Yi:2002:FRB

[1267] Juneho Yi, Jongsun Kim, Jongmoo Choi, Junghyun Han, and Eunseok Lee. Face recognition based on ICA combined with FLD. Lecture Notes in Computer Science, 2359:10– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2359/23590010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2359/23590010.
pdf.

Tistarelli:2002:UII

[1268] Massimo Tistarelli, Andrea Lagorio, and Enrico Grosso. Understanding iconic image-based face biometrics. Lecture Notes in Computer Science, 2359:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2359/23590019.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2359/23590019.
                                        pdf.

Marcialis:2002:FLP

[1269] Gian Luca Marcialis and Fabio Roli. Fusion of LDA and PCA for face verification. Lecture Notes in Computer Science, 2359:30–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2359/23590030.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2359/23590030.
                                        pdf.

Nilsson:2002:CFA

[1270] Kenneth Nilsson and Josef Bigun. Complex filters applied to fingerprint images detecting prominent symmetry points used for alignment. Lecture Notes in Computer Science, 2359:39– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2359/23590039.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2359/23590039.
                                        pdf.

Ross:2002:FMU

[1271] Arun Ross, James Reisman, and Anil Jain. Fingerprint matching using feature space correlation. Lecture Notes in Computer Science, 2359:48– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2359/23590048.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2359/23590048.
pdf.

Bolle:2002:FMC

[1272] Ruud M. Bolle, Andrew W. Senior, Nalini K. Ratha, and Sharath Pankanti. Fingerprint minutiae: a constructive definition. Lecture Notes in Computer Science, 2359:58–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2359/23590058.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2359/23590058.
pdf.

Kompanets:2002:PES

[1273] Leonid Kompanets, Janusz Bobulski, and Roman Wyrzykowski. Pseudoentropy similarity for human biometrics. Lecture Notes in Computer Science, 2359:67–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2359/23590067.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2359/23590067.pdf.

Valchuk:2002:MCP

[1274] Tetiana Valchuk, Roman Wyrzykowski, and Leonid Kompanets. Mental characteristics of person as basic biometrics. Lecture Notes in Computer Science, 2359:78–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2359/23590078.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2359/23590078.
                                        pdf.

Santana:2002:DFF

[1275] M. Castrillón Santana, J. Lorenzo Navarro, J. Cabrera Gámez, F. M. Hernández Tejera, and J. Méndez Rodríguez. Detection of frontal faces in video streams. Lecture Notes in Computer Science, 2359:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2359/23590091.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2359/23590091.
                                        pdf.

Kirchberg:2002:GMO

[1276] Klaus J. Kirchberg, Oliver Jesorsky, and Robert W. Frischholz. Genetic model optimization for Hausdorff distance-based face localization. Lecture Notes in Computer Science, 2359: 103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2359/23590103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2359/23590103.
                                        pdf.

Sahbi:2002:CFF

[1277] Hichem Sahbi and Nozha Boujemaa. Coarse to fine face detection based on skin color adaption. Lecture Notes in Computer Science, 2359:112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2359/23590112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2359/23590112.
pdf.

Sahbi:2002:RFR

[1278] Hichem Sahbi and Nozha Boujemaa.

Robust face recognition using dynamic space warping. Lecture Notes in Computer Science, 2359:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2359/23590121.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2359/23590121.
pdf.

Cappelli:2002:SCF

[1279] Raffaele Cappelli, Dario Maio, and Davide Maltoni. Subspace classification for face recognition. Lecture Notes in Computer Science, 2359:133– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2359/23590133.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2359/23590133.
pdf.

Lee:2002:GAR

[1280] L. Lee and W. E. L. Grimson. Gait appearance for recognition. Lecture Notes in Computer Science, 2359:143– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2359/23590143.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2359/23590143.
                                        pdf.

BenAbdelkader:2002:VIE

[1281] Chiraz BenAbdelkader, Ross Cutler, and Larry Davis. View-invariant estimation of height and stride for gait recognition. Lecture Notes in Computer Science, 2359:155–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2359/23590155.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2359/23590155.
                                        pdf.

Kawamoto:2002:ILS

[1282] Masato Kawamoto, Takayuki Hamamoto, and Seiichiro Hangai. Improvement of on-line signature verification system robust to intersession variability. Lecture Notes in Computer Science, 2359:168– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2359/23590168.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2359/23590168.
                                        pdf.

Gonzalez-Rodriguez:2002:BIF

[1283] J. Gonzalez-Rodriguez, J. Fiérrez-Aguilar, J. Ortega-Garcia, and J. J. Lucena-Molina. Biometric identification in forensic cases according to the Bayesian approach. Lecture Notes in Computer Science, 2359:177– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2359/23590177.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2359/23590177.
pdf.

Thomaz:2002:NQC

[1284] Carlos E. Thomaz, Duncan F. Gillies, and Raul Q. Feitosa. A new quadratic classifier applied to biometric recognition. Lecture Notes in Computer Science, 2359:186–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2359/23590186.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2359/23590186.
pdf.

Anonymous:2002:AIx

[1285] Anonymous. Author index. Lecture Notes in Computer Science, 2359:197– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2359/2359indx.pdf.

vanderAalst:2002:MWF

[1286] Wil M. P. van der Aalst. Making work flow: On the application of Petri nets to business process management. Lecture Notes in Computer Science, 2360: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2360/23600001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600001.
                                        pdf.

Desel:2002:MVT

[1287] Jörg Desel. Model validation — A theoretical issue? Lecture Notes in Computer Science, 2360:23–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2360/23600023.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600023.
                                        pdf.

Hayes:2002:RTR

[1288] Ian J. Hayes. The real-time refinement calculus: a foundation for machine-independent realtime programming. Lecture Notes in Computer Science, 2360:44–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2360/23600044.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600044.
                                        pdf.

Lakos:2002:COO

[1289] Charles Lakos. The challenge of object orientation for the analysis of concurrent systems. Lecture Notes in Computer Science, 2360:59–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2360/23600059.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2360/23600059.
pdf.

Thiagarajan:2002:ACC

[1290] P. S. Thiagarajan. Abstract cyclic communicating processes: a logical view. Lecture Notes in Computer Science, 2360:68–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2360/23600068.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2360/23600068.
pdf.

Yakovlev:2002:CTG

[1291] Alex Yakovlev. Is the die cast for the token game? Lecture Notes in Computer Science, 2360:70–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2360/23600070.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2360/23600070.
pdf.

Cortadella:2002:QSS

[1292] Jordi Cortadella, Alex Kondratyev, Luciano Lavagno, Claudio Passerone, and Yosinori Watanabe. Quasistatic scheduling of independent tasksfor reactive systems. Lecture Notes in Computer Science, 2360:80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2360/23600080.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2360/23600080.
                                        pdf.

Couvreur:2002:DDD

[1293] Jean-Michel Couvreur, Emmanuelle Encrenaz, Emmanuel Paviot-Adet, Denis Poitrenaud, and Pierre-André Wacrenier. Data decision diagrams for Petri net analysis. Lecture Notes in Computer Science, 2360:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2360/23600101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600101.
                                        pdf.

Dehnert:2002:NCC

[1294] Juliane Dehnert. Non-controllable choice robustness expressing the controllability of workflow processes. Lecture Notes in Computer Science, 2360: 121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2360/23600121.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600121.
                                        pdf.

DiMarzoSerugendo:2002:RTS

[1295] Giovanna Di Marzo Serugendo, Dino Mandrioli, Didier Buchs, and Nicolas Guelfi. Real-time synchronised Petri nets. Lecture Notes in Computer Science, 2360:142–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2360/23600142.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2360/23600142.
pdf.

Fleischhack:2002:CFP

[1296] Hans Fleischhack and Christian Stehno. Computing a finite prefix of a time Petri net. Lecture Notes in Computer Science, 2360:163– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2360/23600163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2360/23600163.
pdf.

Gordon:2002:VRW

[1297] Steven Gordon, Lars Michael Kristensen, and Jonathan Billington. Verification of a revised WAP wireless transaction protocol. Lecture Notes in Computer Science, 2360:182–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2360/23600182.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2360/23600182.
pdf.

Jiao:2002:CLP

[1298] Li Jiao, To-Yat Cheung, and Weiming Lu. Characterizing liveness of Petri nets in terms of siphons. Lecture Notes in Computer Science, 2360: 203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2360/23600203.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600203.
                                        pdf.

Kindler:2002:PNS

[1299] Ekkart Kindler. Petri nets, situations, and automata. Lecture Notes in Computer Science, 2360:217–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2360/23600217.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600217.
                                        pdf.

Lautenbach:2002:REM

[1300] Kurt Lautenbach. Reproducibility of the empty marking. Lecture Notes in Computer Science, 2360:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2360/23600237.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600237.
                                        pdf.

LeNy:2002:MAM

[1301] Louis-Marie Le Ny and Bruno Tuffin. Modeling and analysis of multi-class threshold-based queues with hysteresis using stochastic petri nets. Lecture Notes in Computer Science, 2360:254– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2360/23600254.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2360/23600254.
pdf.

Liu:2002:TISb

[1302] Lin Liu and Jonathan Billington.

Tackling the infinite state space of a multimedia control protocol service specification. Lecture Notes in Computer Science, 2360:273–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2360/23600273.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2360/23600273.
pdf.

Lorentsen:2002:MFF

[1303] Louise Lorentsen, Antti-Pekka Tuovinen, and Jianli Xu. Modelling of features and feature interactions in Nokia mobile phones using coloured Petri nets. Lecture Notes in Computer Science, 2360:294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2360/23600294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2360/23600294.
pdf.

Mailund:2002:AIS

[1304] Thomas Mailund. Analysing infinitestate systems by combining equivalence reduction and the sweep-line method. Lecture Notes in Computer Science, 2360:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/2360/23600314.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600314.
                                        pdf.

Nielsen:2002:RES

[1305] Mogens Nielsen and P. S. Thiagarajan. Regular event structures and finite Petri nets: The conflict-free case. Lecture Notes in Computer Science, 2360: 335–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2360/23600335.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600335.
                                        pdf.

Ouyang:2002:FSS

[1306] Chun Ouyang, Lars Michael Kristensen, and Jonathan Billington. A formal service specification for the internet open trading protocol. Lecture Notes in Computer Science, 2360:352– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2360/23600352.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600352.
                                        pdf.

Peuker:2002:TRD

[1307] Sibylle Peuker. Transition refinement for deriving a distributed minimum weight spanning tree algorithm. Lecture Notes in Computer Science, 2360: 374–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2360/23600374.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2360/23600374.
pdf.

Stork:2002:TCP

[1308] David G. Stork and Rob van Glabbeek.

Token-controlled place refinement in hierarchical petri nets with application to active document workflow. Lecture Notes in Computer Science, 2360:394– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2360/23600394.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2360/23600394.
pdf.

Valero:2002:TTS

[1309] Valentín Valero, Juan José Pardo, and Fernando Cuartero. Translating TPAL specifications into timedarc Petri nets. Lecture Notes in Computer Science, 2360:414–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2360/23600414.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2360/23600414.
pdf.

Makela:2002:MMR

[1310] Marko Mäkelä. Maria: Modular reachability analyser for algebraic system nets. Lecture Notes in Computer Science, 2360:434–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2360/23600434.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2360/23600434.
                                        pdf.

Anonymous:2002:AIy

[1311] Anonymous. Author index. Lecture Notes in Computer Science, 2360:445– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2360/2360ind.pdf.

Boasson:2002:ESU

[1312] Maarten Boasson. Embedded systems unsuitable for object orientation. Lecture Notes in Computer Science, 2361: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610001.
                                        pdf.

Jazayeri:2002:ASE

[1313] Mehdi Jazayeri. On architectural stability and evolution. Lecture Notes in Computer Science, 2361:13–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610013.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610013.
                                        pdf.

Doudou:2002:EFD

[1314] Assia Doudou, Benoît Garbinato, and Rachid Guerraoui. Encapsulating failure detection: From crash to Byzantine failures. Lecture Notes in Computer Science, 2361:24–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2361/23610024.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2361/23610024.
pdf.

Ferscha:2002:CBP

[1315] Alois Ferscha. Contextware: Bridging physical and virtual worlds. Lecture Notes in Computer Science, 2361:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2361/23610051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2361/23610051.
pdf.

Chatzigeorgiou:2002:EPP

[1316] Alexander Chatzigeorgiou and George Stephanides. Evaluating performance and power of object-oriented vs. procedural programming in embedded processors. Lecture Notes in Computer Science, 2361:65–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2361/23610065.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2361/23610065.
pdf.

Lobato:2002:OIM

[1317] Jose Manuel Pérez Lobato and Eva Martín Lobo. OMC-INTEGRAL memory management. Lecture Notes in Computer Science, 2361:76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610076.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610076.
                                        pdf.

Ward:2002:LIC

[1318] Michael Ward and Neil C. Audsley. Language issues of compiling Ada to hardware. Lecture Notes in Computer Science, 2361:88–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610088.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610088.
                                        pdf.

Hoe:2002:SDR

[1319] Adrian Hoe. Software development reengineering — an experience report. Lecture Notes in Computer Science, 2361:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2361/23610100.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610100.
                                        pdf.

Ortiz:2002:DCS

[1320] Francisco J. Ortiz, Alejandro S.

Martínez, Barbara Álvarez, Andres Iborra, and José M. Fernández. Development of a control system for teleoperated robots using UML and Ada95. Lecture Notes in Computer Science, 2361:113–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2361/23610113.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2361/23610113.
pdf.

Binder:2002:USJ

[1321] Walter Binder and Balázs Lichtl.

Using a secure Java micro-kernel on embedded devices for the reliable execution of dynamically uploaded applications. Lecture Notes in Computer Science, 2361:125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2361/23610125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2361/23610125.
pdf.

Rivas:2002:PAI

[1322] Mario Aldea Rivas and Michael González Harbour. A POSIX-Ada interface for application-defined scheduling. Lecture Notes in Computer Science, 2361:136– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2361/23610136.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2361/23610136.
                                        pdf.

Amey:2002:CLI

[1323] Peter Amey. Closing the loop: The influence of code analysis on design. Lecture Notes in Computer Science, 2361: 151–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610151.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610151.
                                        pdf.

Parkinson:2002:HIS

[1324] Paul Parkinson and Franco Gasperoni. High-integrity systems development for integrated modular avionics using VxWorks and GNAT. Lecture Notes in Computer Science, 2361:163– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610163.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610163.
                                        pdf.

Miranda:2002:HUG

[1325] Javier Miranda, Francisco Guerra, Ernestina Martel, José Martín, and Alexis González. How to use GNAT to efficiently preprocess new Ada sentences. Lecture Notes in Computer Science, 2361:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610179.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2361/23610179.
pdf.

Dewar:2002:EUV

[1326] Robert Dewar, Olivier Hainque, Dirk Craeynest, and Philippe Waroquiers. Exposing uninitialized variables: Strengthening and extending run-time checks in Ada. Lecture Notes in Computer Science, 2361:193– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2361/23610193.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2361/23610193.
pdf.

Lamm:2002:ADC

[1327] Ehud Lamm. Adding design by contract to the Ada language. Lecture Notes in Computer Science, 2361:205– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2361/23610205.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2361/23610205.
pdf.

Chen:2002:SDA

[1328] Zhenqiang Chen, Baowen Xu, Jianjun Zhao, and Hongji Yang. Static dependency analysis for concurrent Ada 95 programs. Lecture Notes in Computer Science, 2361:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2361/23610219.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610219.
                                        pdf.

Czarnecki:2002:DPA

[1329] Krzysztof Czarnecki, Michael Himsolt, Ernst Richter, Falk Vieweg, and Alfred Rosskopf. DataFAN: a practical approach to data flow analysis for Ada 95. Lecture Notes in Computer Science, 2361:231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610231.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610231.
                                        pdf.

Yu:2002:PTC

[1330] Yuen T. Yu and Man F. Lau. Prioritization of test cases in MUMCUT test sets: An empirical study. Lecture Notes in Computer Science, 2361:245– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610245.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610245.
                                        pdf.

Rybin:2002:ADB

[1331] Sergey Rybin and Alfred Strohmeier. About the difficulties of building a pretty-printer for Ada. Lecture Notes in Computer Science, 2361:257–268, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2361/23610257.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2361/23610257.
pdf.

Martel:2002:TDP

[1332] Ernestina Martel, Francisco Guerra, and Javier Miranda. A tailorable distributed programming environment. Lecture Notes in Computer Science, 2361:269–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2361/23610269.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2361/23610269.
pdf.

Medina:2002:MSA

[1333] Julio L. Medina, J. Javier Gutiérrez, José M. Drake, and Michael González Harbour. Modeling and schedulability analysis of hard real-time distributed systems based on Ada components. Lecture Notes in Computer Science, 2361:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2361/23610282.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2361/23610282.
pdf.

Pinho:2002:TER

[1334] Luís Miguel Pinho and Francisco Vasques. Transparent environment for replicated Ravenscar applications. Lecture Notes in Computer Science, 2361: 297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610297.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610297.
                                        pdf.

Patino-Martinez:2002:CCT

[1335] Marta Patiño-Martínez, Ricardo Jiménez-Peris, Jörg Kienzle, and Sergio Arévalo. Concurrency control in transactional drago. Lecture Notes in Computer Science, 2361:309–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610309.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610309.
                                        pdf.

Minguet:2002:ABI

[1336] Agustín Espinosa Minguet, Ana García Fornes, and Alfons Crespo i Lorente. An Ada binding to the IEEE 1003.1q (POSIX Tracing) Standard. Lecture Notes in Computer Science, 2361:321– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2361/23610321.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2361/23610321.
                                        pdf.

Erdmann:2002:GAD

[1337] Michael Erdmann. GNAT Ada database development environment. Lecture Notes in Computer Science, 2361:334–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2361/23610334.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2361/23610334.
pdf.

Rosen:2002:AIL

[1338] Jean-Pierre Rosen. Ada, interfaces and the listener paradigm. Lecture Notes in Computer Science, 2361:344– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2361/23610344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2361/23610344.
pdf.

Alonso:2002:UOO

[1339] Alejandro Alonso, Roberto López, Tullio Vardanega, and Juan Antonio de la Puente. Using object orientation in high integrity applications: a case study. Lecture Notes in Computer Science, 2361:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2361/23610357.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2361/23610357.
pdf.

Anonymous:2002:AIz

[1340] Anonymous. Author index. Lecture Notes in Computer Science, 2361:367– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        papers/2361/2361indx.pdf.

vandenBesselaar:2002:IDC

[1341] Peter van den Besselaar, Makoto Tanabe, and Toru Ishida. Introduction: Digital cities research and open issues. Lecture Notes in Computer Science, 2362:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2362/23620001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2362/23620001.
                                        pdf.

Wellman:2002:LBG

[1342] Barry Wellman. Little boxes, glocalization, and networked individualism. Lecture Notes in Computer Science, 2362:10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2362/23620010.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2362/23620010.
                                        pdf.

Gumpert:2002:PPS

[1343] Gary Gumpert and Susan Drucker. Privacy, predictability or serendipity and digital cities. Lecture Notes in Computer Science, 2362:26–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2362/23620026.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2362/23620026.
pdf.

Lelis:2002:RFE

[1344] Stelios Lelis, Petros Kavassalis, Jakka Sairamesh, Seif Haridi, Fredrik Holmgren, Mahmoud Rafea, and Antonis Hatzistamatiou. Regularities in the formation and evolution of information cities. Lecture Notes in Computer Science, 2362:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620041.
pdf.

Kryssanov:2002:CSA

[1345] Victor V. Kryssanov, Masayuki Okabe, Koh Kakusho, and Michihiko Minoh. Communication of social agents and the digital city — A semiotic perspective. Lecture Notes in Computer Science, 2362:56–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620056.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620056.
pdf.

Schuler:2002:DCD

[1346] Doug Schuler. Digital cities and digital citizens. Lecture Notes in Computer Science, 2362:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2362/23620071.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2362/23620071.
                                        pdf.

Day:2002:DDC

[1347] Peter Day. Designing democratic community networks: Involving communities through civil participation. Lecture Notes in Computer Science, 2362: 86–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2362/23620086.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2362/23620086.
                                        pdf.

Gotzl:2002:TDC

[1348] Ingrid Götzl. TeleCities — digital cities network. Lecture Notes in Computer Science, 2362:101–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2362/23620101.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2362/23620101.
                                        pdf.

Pinkett:2002:CEC

[1349] Randal D. Pinkett. The camfield estates-MIT creating community connections project: Strategies for active participation in a low- to moderateincome community. Lecture Notes in Computer Science, 2362:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2362/23620110.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620110.
pdf.

Hanson:2002:CWL

[1350] Karrie J. Hanson and Gerald M.

Karam. Community Websites as a local communication network: “Directory Westfield”, an experience report. Lecture Notes in Computer Science, 2362: 125–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620125.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620125.
pdf.

McQuillan:2002:EIA

[1351] Helen McQuillan. Ennis information age town: Virtuality rooted in reality. Lecture Notes in Computer Science, 2362:139–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2362/23620139.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2362/23620139.
pdf.

Harada:2002:FSD

[1352] Kikuko Harada and Hiroshi Hoshino.

Feasibility study of digital community through virtual enterprise network. Lecture Notes in Computer Science, 2362:152–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/

                                        0558/bibs/2362/23620152.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2362/23620152.
                                        pdf.

Rehfeld:2002:KBE

[1353] Dieter Rehfeld and Ileana Hamburg. Knowledge-based economic services supported by digital experiments. Lecture Notes in Computer Science, 2362: 164–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2362/23620164.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2362/23620164.
                                        pdf.

Venkatesh:2002:CND

[1354] Murali Venkatesh and Donghee Shin. Community network development: a dialectical view. Lecture Notes in Computer Science, 2362:177–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2362/23620177.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2362/23620177.
                                        pdf.

Ranerup:2002:CUC

[1355] Agneta Ranerup. The complexity of using commercial forces to counteract the digital divide: a case study of the TUC of sweden. Lecture Notes in Computer Science, 2362:191– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2362/23620191.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620191.
pdf.

Cheng:2002:LLS

[1356] Lili Cheng, Shelly Farnham, and Linda Stone. Lessons learned: Social interaction in virtual environments. Lecture Notes in Computer Science, 2362: 203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620203.
pdf.

Rommes:2002:WAE

[1357] Els Rommes. Worlds apart: Exclusionprocesses in DDS. Lecture Notes in Computer Science, 2362:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620219.
pdf.

Hiramatsu:2002:LAM

[1358] Kaoru Hiramatsu. Log analysis of map-based Web page search on digital city kyoto. Lecture Notes in Computer Science, 2362:233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620233.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2362/23620233.
                                        pdf.

Ishida:2002:CDP

[1359] Toru Ishida, Hiroshi Ishiguro, and Hideyuki Nakanishi. Connecting digital and physical cities. Lecture Notes in Computer Science, 2362:246–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2362/23620246.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2362/23620246.
                                        pdf.

Borner:2002:TWA

[1360] Katy Börner. Twin worlds: Augmenting, evaluating, and studying three-dimensional digital cities and their evolving communities. Lecture Notes in Computer Science, 2362:257– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2362/23620257.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2362/23620257.
                                        pdf.

Fukuhara:2002:CCC

[1361] Tomohiro Fukuhara, Ken’ichi Matsumura, Shintaro Azechi, Nobuhiko Fujihara, Kazunori Terada, Koji Yamashita, and Toyoaki Nishida. Creating city community consanguinity: Use of public opinion channel in digital cities. Lecture Notes in Computer Science, 2362:270–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620270.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620270.
pdf.

Akahani:2002:ABC

[1362] Jun ichi Akahani, Kaoru Hiramatsu, Yoshikazu Furukawa, and Kiyoshi Kogure. Agent-based coordination of regional information services. Lecture Notes in Computer Science, 2362:283– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620283.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620283.
pdf.

Okada:2002:RDE

[1363] Masaya Okada, Hiroyuki Tarumi, Tetsuhiko Yoshimura, Kazuyuki Moriya, and Tetsuro Sakai. Realization of digital environmental education – A future style of environmental education in dynamically changing virtual environment –. Lecture Notes in Computer Science, 2362:292–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620292.
pdf.

Tanaka:2002:DPC

[1364] Hiroya Tanaka, Masatoshi Arikawa, and Ryosuke Shibasaki. A 3-D photo collage system for spatial navigations. Lecture Notes in Computer Science, 2362:305–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620305.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620305.
pdf.

Goto:2002:SMP

[1365] Koichi Goto and Yahiko Kambayashi. Study on mobile passenger support systems for public transportation using multi-channel data dissemination. Lecture Notes in Computer Science, 2362: 317–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620317.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620317.
pdf.

Tsuji:2002:SIS

[1366] Hiroshi Tsuji, Masato Terada, Yuki Kadowaki, Masaaki Tanizaki, and Shigeru Shimada. Spatial information sharing for mobile phones. Lecture Notes in Computer Science, 2362:331– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620331.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620331.
pdf.

Burg:2002:AWA

[1367] Bernard Burg. Agents in the world of active Web-services. Lecture Notes in Computer Science, 2362:343–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620343.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620343.
pdf.

Koizumi:2002:TDB

[1368] Satoshi Koizumi, Guiming Dai, and Hiroshi Ishiguro. Town digitizing for building an image-based cyber space. Lecture Notes in Computer Science, 2362:357–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2362/23620357.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2362/23620357.
pdf.

Noda:2002:LDR

[1369] Itsuki Noda, Tomoichi Takahashi, Shuji Morita, Tetsuhiko Koto, and Satoshi Tadokoro. Language design for rescue agents. Lecture Notes in Computer Science, 2362:371–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2362/23620371.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2362/23620371.
pdf.

Goel:2002:UP

[1370] Aradhana Goel. Urban pilot. Lecture Notes in Computer Science, 2362:384– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2362/23620384.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2362/23620384.
                                        pdf.

Anonymous:2002:AIba

[1371] Anonymous. Author index. Lecture Notes in Computer Science, 2362:399– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2362/2362auth.pdf.

Bradshaw:2002:LAH

[1372] Jeffrey M. Bradshaw. Living with agents: From human-agent teamwork to cognitive prostheses. Lecture Notes in Computer Science, 2363:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630001.
                                        pdf.

Stock:2002:CH

[1373] Oliviero Stock. Computational humor. Lecture Notes in Computer Science, 2363:2–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/

0558/bibs/2363/23630002.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2363/23630002.
pdf.

Mulder:2002:LSI

[1374] Fred Mulder, Jos Rikers, and Piet Henderikx. Large-scale introduction of E-learning at universities throughout europe. Lecture Notes in Computer Science, 2363:4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630004.
pdf.

Koriche:2002:REL

[1375] Frédéric Koriche. A roadmap of epistemic logics for learning agents. Lecture Notes in Computer Science, 2363: 6–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630006.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630006.
pdf.

Vicari:2002:ABA

[1376] Rosa Maria Vicari. ITS, agents, BDI, and affection: Trying to make a plan come together. Lecture Notes in Computer Science, 2363:8– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630008.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630008.
                                        pdf.

Hayashi:2002:IGG

[1377] Yusuke Hayashi, Hiroyuki Tsumoto, Mitsuru Ikeda, and Riichiro Mizoguchi. An intellectual genealogy graph affording a fine prospect of organizational learning. Lecture Notes in Computer Science, 2363:10–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630010.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630010.
                                        pdf.

Pinkwart:2002:GOM

[1378] Niels Pinkwart, H. Ulrich Hoppe, Lars Bollen, and Eva Fuhlrott. Grouporiented modelling tools with heterogeneous semantics. Lecture Notes in Computer Science, 2363:21–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630021.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630021.
                                        pdf.

Reffay:2002:SNA

[1379] Christophe Reffay and Thierry Chanier. Social network analysis used for modelling collaboration in distance learning groups. Lecture Notes in Computer Science, 2363:31–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630031.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630031.
pdf.

daNobrega:2002:SRM

[1380] Germana M. da Nóbrega, Stefano A.

Cerri, and Jean Sallantin. On the social rational mirror: Learning Ecommerce in a Web-served learning environment. Lecture Notes in Computer Science, 2363:41–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630041.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630041.
pdf.

Lopez:2002:IMA

[1381] Natalia López, Manuel Núñez, Ismael Rodríguez, and Fernando Rubio. Including malicious agents into a collaborative learning environment. Lecture Notes in Computer Science, 2363:51– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630051.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630051.
pdf.

Rasseneur:2002:LVA

[1382] Dorothée Rasseneur, Elisabeth Delozanne, Pierre Jacoboni, and Brigitte Grugeon. Learning with virtual agents: Competition and cooperation in AMICO. Lecture Notes in Computer Science, 2363:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630061.
                                        pdf.

Martens:2002:ATP

[1383] Alke Martens and Adelinde M. Uhrmacher. Adaptive tutoring processes and mental plans. Lecture Notes in Computer Science, 2363:71–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630071.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630071.
                                        pdf.

Devedzic:2002:APP

[1384] Vladan Devedzic and Andreas Harrer. Architectural patterns in pedagogical agents. Lecture Notes in Computer Science, 2363:81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630081.
                                        pdf.

TrojahndosSantos:2002:DPA

[1385] Cássia Trojahn dos Santos, Rejane Frozza, Alessandra Dhamer, and Luciano Paschoal Gaspary. DÓRIS — pedagogical agent in intelligent tutoring systems. Lecture Notes in Computer Science, 2363:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630091.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630091.
pdf.

Silveira:2002:DDI

[1386] Ricardo Azambuja Silveira and Rosa Maria Vicari. Developing distributed intelligent learning environment with JADE — Java agents for distance education framework. Lecture Notes in Computer Science, 2363:105–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630105.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630105.
pdf.

Prentzas:2002:WBI

[1387] Jim Prentzas, Ioannis Hatzilygeroudis, and John Garofalakis. A Web-based intelligent tutoring system using hybrid rules as its representational basis. Lecture Notes in Computer Science, 2363: 119–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630119.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630119.
pdf.

Kurhila:2002:UVS

[1388] Jaakko Kurhila, Matti Lattu, and Anu Pietilä. Using vector-space model in adaptive hypermedia for learning. Lecture Notes in Computer Science, 2363: 129–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630129.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630129.
                                        pdf.

Verdejo:2002:SSD

[1389] M. F. Verdejo, B. Barros, T. Read, and M. Rodriguez-Artacho. A system for the specification and development of an environment for distributed CSCL scenarios. Lecture Notes in Computer Science, 2363:139– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630139.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630139.
                                        pdf.

Raad:2002:MAH

[1390] Hussein Raad and Bernard Causse. Modelling of an adaptive hypermedia system based on active rules. Lecture Notes in Computer Science, 2363:149– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630149.
                                        pdf.

VanLehn:2002:AWA

[1391] Kurt VanLehn, Pamela W. Jordan, Carolyn P. Rosé, Dumisizwe Bhembe, Michael Böttner, Andy Gaydos, Maxim Makatchev, Umarani Pappuswamy, Michael Ringenberg, Antonio Roque, Stephanie Siler, and Ramesh Srivastava. The architecture of Why2-Atlas: a coach for qualitative physics essay writing. Lecture Notes in Computer Science, 2363:158– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630158.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630158.
pdf.

Moreno:2002:IMT

[1392] L. Moreno, C. González, V. Muñoz, J. Estévez, R. Aguilar, J. Sánchez, J. Sigut, and J. Piñeiro. Integrating multimedia technology, knowledge based system and speech processing for the diagnostic and treatment of developmental dyslexia. Lecture Notes in Computer Science, 2363: 168–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630168.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630168.
pdf.

Nicaud:2002:AEN

[1393] Jean-François Nicaud, Denis Bouhineau, and Thomas Huguet. The aplusixeditor: a new kind of software for the learning of algebra. Lecture Notes in Computer Science, 2363:178–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630178.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630178.
                                        pdf.

Rouane:2002:LLK

[1394] Khalid Rouane, Claude Frasson, and Marc Kaltenbach. LKC: Learning by knowledge construction. Lecture Notes in Computer Science, 2363:188– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630188.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630188.
                                        pdf.

El-Sheikh:2002:GIT

[1395] Eman El-Sheikh and Jon Sticklen. Generating intelligent tutoring systems from reusable components and knowledge-based systems. Lecture Notes in Computer Science, 2363:199– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630199.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630199.
                                        pdf.

Herin:2002:ERC

[1396] Daniéle Hérin, Michel Sala, and Pierre Pompidor. Evaluating and revising courses from Web resources educational. Lecture Notes in Computer Science, 2363:208–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2363/23630208.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2363/23630208.pdf.

Kashihara:2002:APC

[1397] Akihiro Kashihara, Kunitaka Kumei, Shinobu Hasegawa, and Junichi Toyoda. Adaptive postviewer for constructive learning in hyperspace. Lecture Notes in Computer Science, 2363:219– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630219.
pdf.

Brusilovsky:2002:AVC

[1398] Peter Brusilovsky and Hoah-Der Su.

Adaptive visualization component of a distributed Web-based adaptive educational system. Lecture Notes in Computer Science, 2363:229–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630229.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630229.
pdf.

Kurhila:2002:DPR

[1399] Jaakko Kurhila, Miikka Miettinen, Petri Nokelainen, and Henry Tirri. Dynamic profiling in a real-time collaborative learning environment. Lecture Notes in Computer Science, 2363:239– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630239.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630239.
pdf.

Rebai:2002:DGI

[1400] Issam Rebaï and Brigitte de La Passardiére. Dynamic generation of an interface for the capture of educational metadata. Lecture Notes in Computer Science, 2363:249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630249.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630249.
pdf.

Trausan-Matu:2002:OCP

[1401] Stefan Trausan-Matu, Daniele Maraschi, and Stefano Cerri. Ontology-centered personalized presentation of knowledge extracted from the Web. Lecture Notes in Computer Science, 2363:259– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630259.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630259.
pdf.

Aroyo:2002:OSW

[1402] Lora Aroyo, Darina Dicheva, and Alexandra Cristea. Ontological support for Web courseware authoring. Lecture Notes in Computer Science, 2363:270–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2363/23630270.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2363/23630270.
pdf.

Meyer:2002:HEL

[1403] Cécile Meyer. Hypermedia environment for learning concepts based on inter-domain analogies as an educational strategy. Lecture Notes in Computer Science, 2363:281–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630281.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630281.
pdf.

Moebus:2002:TSD

[1404] Claus Moebus, Bernd Albers, Stefan Hartmann, Heinz-Juergen Thole, and Jochen Zurborg. Towards a specification of distributed and intelligent Web based training systems. Lecture Notes in Computer Science, 2363:291– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630291.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2363/23630291.
                                        pdf.

Geoffroy:2002:VAW

[1405] Frédéric Geoffroy, Esma Aimeur, and Denis Gillet. A virtual assistant for Web-based training in engineering education. Lecture Notes in Computer Science, 2363:301–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630301.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630301.
                                        pdf.

Woo:2002:WBT

[1406] Chong woo Woo, Jin woo Choi, and Martha Evens. Web-based ITS for training system managers on the computer intrusion. Lecture Notes in Computer Science, 2363:311–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630311.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630311.
                                        pdf.

Mostow:2002:RDT

[1407] Jack Mostow, Greg Aist, Joseph Beck, Raghuvee Chalasani, Andrew Cuneo, Peng Jia, and Krishna Kadaru. A la recherche du temps perdu, or as time goes by: Where does the time go in a reading tutor that listens? Lecture Notes in Computer Science, 2363:320– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630320.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630320.
pdf.

Mathan:2002:EAC

[1408] Santosh A. Mathan and Kenneth R.

Koedinger. An empirical assessment of comprehension fostering features in an intelligent tutoring system. Lecture Notes in Computer Science, 2363:330– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630330.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630330.
pdf.

Aleven:2002:PTT

[1409] Vincent Aleven, Octav Popescu, and Kenneth Koedinger. Pilot-testing a tutorial dialogue system that supports self-explanation. Lecture Notes in Computer Science, 2363:344–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630344.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630344.
pdf.

MacLaren:2002:WWD

[1410] Benjamin MacLaren and Kenneth Koedinger. When and why does mastery learning work: Instructional experiments with ACT-R “SimStudents”. Lecture Notes in Computer Science, 2363:355–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2363/23630355.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630355.
                                        pdf.

VanLehn:2002:MIT

[1411] Kurt VanLehn, Collin Lynch, Linwood Taylor, Anders Weinstein, Robert Shelby, Kay Schulze, Don Treacy, and Mary Wintersgill. Minimally invasive tutoring of complex physics problem solving. Lecture Notes in Computer Science, 2363:367–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630367.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630367.
                                        pdf.

Suraweera:2002:KCB

[1412] Pramuditha Suraweera and Antonija Mitrovic. KERMIT: a constraint-based tutor for database modeling. Lecture Notes in Computer Science, 2363:377– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630377.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630377.
                                        pdf.

Martin:2002:APG

[1413] Brent Martin and Antonija Mitrovic. Automatic problem generation in constraint-based tutors. Lecture Notes in Computer Science, 2363:388– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630388.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630388.
pdf.

Bourdeau:2002:COE

[1414] Jacqueline Bourdeau and Riichiro Mizoguchi. Collaborative ontological engineering of instructional design knowledge for an ITS authoring environment. Lecture Notes in Computer Science, 2363:399–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630399.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630399.
pdf.

Paraskakis:2002:TTC

[1415] Iraklis Paraskakis. TeLoDe: Towards creating an intelligent computer algebra system. Lecture Notes in Computer Science, 2363:410–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630410.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630410.
pdf.

Lesta:2002:ITA

[1416] Leanna Lesta and Kalina Yacef.

An intelligent teaching assistant system for logic. Lecture Notes in Computer Science, 2363:421–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630421.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630421.
pdf.

Dimitrova:2002:DIG

[1417] Vania Dimitrova, Paul Brna, and John Self. The design and implementation of a graphical communication medium for interactive open learner modelling. Lecture Notes in Computer Science, 2363: 432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630432.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630432.
pdf.

Zapata-Rivera:2002:EVG

[1418] Diego Zapata-Rivera and Jim E. Greer. Exploring various guidance mechanisms to support interaction with inspectable learner models. Lecture Notes in Computer Science, 2363:442– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630442.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630442.
pdf.

Hartley:2002:SLO

[1419] Danita Hartley and Antonija Mitrovic. Supporting learning by opening the student model. Lecture Notes in Computer Science, 2363:453–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630453.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630453.
pdf.

Liew:2002:RAS

[1420] Chun Wai Liew and Donald E.

Smith. Reasoning about systems of physics equations. Lecture Notes in Computer Science, 2363:463–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630463.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630463.
pdf.

Richards:2002:PHB

[1421] Robert A. Richards. Principle hierarchy based intelligent tutoring system for common cockpit helicopter training. Lecture Notes in Computer Science, 2363:473–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630473.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630473.
pdf.

Aka:2002:AOD

[1422] Michel Aka and Claude Frasson. ASIMIL: Overview of a distance learning flight-training system. Lecture Notes in Computer Science, 2363:484– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630484.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630484.
                                        pdf.

Ludtke:2002:CMA

[1423] Andreas Lüdtke, Claus Möbus, and Heinz-Jürgen Thole. Cognitive modelling approach to diagnose oversimplification in simulation-based training. Lecture Notes in Computer Science, 2363:496–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2363/23630496.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2363/23630496.pdf.

Duquesnoy:2002:STP

[1424] Laurent Duquesnoy, Jean-Luc Berger, Patrick Prévôt, and Françoise Sandoz-Guermond. SIMPA: a training platform in work station including computing tutors. Lecture Notes in Computer Science, 2363:507–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630507.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2363/23630507.
pdf.

Joab:2002:CAE

[1425] Michelle Joab, Odette Auzende, Michel Futtersack, Brigitte Bonnet, and Patrice Le Leydour. Computer aided evaluation of trainee skills on a simulator network. Lecture Notes in Computer Science, 2363:521–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630521.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630521.
pdf.

Minko:2002:QAA

[1426] Anton Minko, Guy Gouardères, and Orest Popov. Qualitative assessment on aeronautical training with cognitive agents. Lecture Notes in Computer Science, 2363:531–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630531.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630531.
pdf.

Rickel:2002:CDT

[1427] Jeff Rickel, Neal Lesh, Charles Rich, Candace L. Sidner, and Abigail Gertner. Collaborative discourse theory as a foundation for tutorial dialogue. Lecture Notes in Computer Science, 2363: 542–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630542.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630542.
                                        pdf.

Rose:2002:HLU

[1428] Carolyn P. Rosé, Dumisizwe Bhembe, Antonio Roque, Stephanie Siler, Ramesh Srivastava, and Kurt Van-Lehn. A hybrid language understanding approach for robust selection of tutoring goals. Lecture Notes in Computer Science, 2363:552–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630552.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630552.
                                        pdf.

Malatesta:2002:BSA

[1429] Kalliopi-Irini Malatesta, Peter Wiemer-Hastings, and Judy Robertson. Beyond the short answer question with research methods tutor. Lecture Notes in Computer Science, 2363:562–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630562.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630562.
                                        pdf.

Zinn:2002:TPA

[1430] Claus Zinn, Johanna D. Moore, and Mark G. Core. A 3-tier planning architecture for managing tutorial dialogue. Lecture Notes in Computer Science, 2363:574–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630574.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630574.
pdf.

Ashley:2002:TCB

[1431] Kevin D. Ashley, Ravi Desai, and John M. Levine. Teaching casebased argumentation concepts using dialectic arguments vs. didactic explanations. Lecture Notes in Computer Science, 2363:585–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630585.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630585.
pdf.

Heffernan:2002:ITS

[1432] Neil T. Heffernan and Kenneth R.

Koedinger. An intelligent tutoring system incorporating a model of an experienced human tutor. Lecture Notes in Computer Science, 2363:596– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630596.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630596.
pdf.

Chen:2002:IAA

[1433] Weiqin Chen and Barbara Wasson. An instructional assistant agent for distributed collaborative learning. Lecture Notes in Computer Science, 2363: 609–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630609.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630609.
pdf.

Tedesco:2002:CGP

[1434] Patricia Azevedo Tedesco. Conflicts in group planning: Results from the experimental study of MArCo. Lecture Notes in Computer Science, 2363:619– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630619.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630619.
pdf.

George:2002:AAA

[1435] Sébastien George and Pascal Leroux. An approach to automatic analysis of learners’ social behavior during computer-mediated synchronous conversations. Lecture Notes in Computer Science, 2363:630–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630630.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630630.
pdf.

Katz:2002:GBP

[1436] Sandra Katz and David Allbritton. Going beyond the problem given: How human tutors use post-practice discussions to support transfer. Lecture Notes in Computer Science, 2363:641– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630641.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630641.
pdf.

Michel:2002:CLL

[1437] Johan Michel and Jérôme Lehuen.

Conception of a language learning environment based on the communicative and actional approaches. Lecture Notes in Computer Science, 2363:651– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630651.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630651.
pdf.

Kim:2002:DAP

[1438] Jihie Kim and Yolanda Gil. Deriving acquisition principles from tutoring principles. Lecture Notes in Computer Science, 2363:661–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630661.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630661.
pdf.

Bravo:2002:CDL

[1439] Crescencio Bravo, Miguel A. Redondo, Manuel Ortega, and M. Felisa Verdejo. Collaborative discovery learning of model design. Lecture Notes in Computer Science, 2363:671– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630671.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630671.
                                        pdf.

Woolf:2002:GPI

[1440] Beverly Park Woolf, John Reid, Neil Stillings, Merle Bruno, Dan Murray, Paula Reese, Alan Peterfreund, and Kenneth Rath. A general platform for inquiry learning. Lecture Notes in Computer Science, 2363:681– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630681.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630681.
                                        pdf.

Bunt:2002:AEE

[1441] Andrea Bunt and Cristina Conati. Assessing effective exploration in open learning environments using Bayesian networks. Lecture Notes in Computer Science, 2363:698–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630698.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2363/23630698.
pdf.

Tchetagni:2002:HRE

[1442] Joséphine M. P. Tchétagni and Roger Nkambou. Hierarchical representation and evaluation of the student in an intelligent tutoring system. Lecture Notes in Computer Science, 2363:708– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630708.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630708.
pdf.

Aimeur:2002:CML

[1443] Esma Aïmeur, Gilles Brassard, Hugo Dufort, and Sébastien Gambs. CLARISSE: a machine learning tool to initialize student models. Lecture Notes in Computer Science, 2363:718–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630718.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630718.
pdf.

Kelly:2002:ILC

[1444] Declan Kelly and Brendan Tangney.

Incorporating learning characteristics into an intelligent tutor. Lecture Notes in Computer Science, 2363:729– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630729.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630729.
                                        pdf.

Guzman:2002:SEM

[1445] Eduardo Guzmán and Ricardo Conejo. Simultaneous evaluation of multiple topics in SIETTE. Lecture Notes in Computer Science, 2363:739–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630739.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630739.
                                        pdf.

Murray:2002:TMM

[1446] Tom Murray and Ivon Arroyo. Toward measuring and maintaining the zone of proximal development in adaptive instructional systems. Lecture Notes in Computer Science, 2363:749– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630749.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630749.
                                        pdf.

Luckin:2002:GKM

[1447] Rosemary Luckin and Louise Hammerton. Getting to know me: Helping learners understand their own learning needs through metacognitive scaffolding. Lecture Notes in Computer Science, 2363:759–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2363/23630759.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630759.
pdf.

Dessus:2002:UPA

[1448] Philippe Dessus and Benoît Lemaire.

Using production to assess learning: An ILE that fosters selfregulated learning. Lecture Notes in Computer Science, 2363:772–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630772.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630772.
pdf.

Guin-Duclosson:2002:AIH

[1449] Nathalie Guin-Duclosson, Stéphanie Jean-Daubias, and Sandra Nogry. The ambre ILE: How to use case-based reasoning to teach methods. Lecture Notes in Computer Science, 2363:782– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630782.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630782.
pdf.

Kumar:2002:MBR

[1450] Amruth N. Kumar. Model-based reasoning for domain modeling in a Web-based intelligent tutoring system to help students learn to debug C++ programs. Lecture Notes in Computer Science, 2363:792–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630792.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630792.
                                        pdf.

Budenbender:2002:UCA

[1451] Jochen Büdenbender, Adrian Frischauf, Georgi Goguadze, Erica Melis, Paul Libbrecht, and Carsten Ullrich. Using computer algebra systems as cognitive tools. Lecture Notes in Computer Science, 2363:802–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630802.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630802.
                                        pdf.

Georgouli:2002:DMI

[1452] Katerina Georgouli. The design of a ‘motivating’ intelligent assessment system. Lecture Notes in Computer Science, 2363:811–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630811.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630811.
                                        pdf.

Person:2002:HCA

[1453] Natalie Person and Arthur C. Graesser. Human or computer? AutoTutor in a bystander Turing test. Lecture Notes in Computer Science, 2363:821– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630821.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630821.
pdf.

VanLabeke:2002:RDW

[1454] Nicolas Van Labeke and Shaaron Ainsworth. Representational decisions when learning population dynamics with an instructional simulation. Lecture Notes in Computer Science, 2363: 831–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630831.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630831.
pdf.

Mbala:2002:IAT

[1455] Aloys Mbala, Christophe Reffay, and Thierry Chanier. Integration of automatic tools for displaying interaction data in computer environments for distance learning. Lecture Notes in Computer Science, 2363:841–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630841.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630841.
pdf.

Beck:2002:DDE

[1456] Joseph E. Beck. Directing development effort with simulated students. Lecture Notes in Computer Science, 2363:851– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630851.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630851.
                                        pdf.

Nakano:2002:EIL

[1457] Akira Nakano, Tsukasa Hirashima, and Akira Takeuchi. An evaluation of intelligent learning environment for problem posing. Lecture Notes in Computer Science, 2363:861– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630861.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630861.
                                        pdf.

Ainsworth:2002:UED

[1458] Shaaron Ainsworth, David Clarke, and Robert Gaizauskas. Using edit distance algorithms to compare alternative approaches to ITS authoring. Lecture Notes in Computer Science, 2363:873– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630873.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630873.
                                        pdf.

Ainsworth:2002:ICR

[1459] Shaaron Ainsworth and Shirley Grimshaw.

Are ITSs created with the REDEEM authoring tool more effective than “Dumb” couresware? Lecture Notes in Computer Science, 2363:883–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630883.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630883.
pdf.

Gerdt:2002:SXE

[1460] Petri Gerdt, Piet Kommers, Jarkko Suhonen, and Erkki Sutinen. StoryML: An XML extension for woven stories. Lecture Notes in Computer Science, 2363:893–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2363/23630893.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2363/23630893.
pdf.

Rizzo:2002:AHC

[1461] Paola Rizzo, Erin Shaw, and W. Lewis Johnson. An agent that helps children to author rhetorically-structured digital puppet presentations. Lecture Notes in Computer Science, 2363:903– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630903.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630903.
pdf.

Prada:2002:YCU

[1462] Rui Prada, Ana Paiva, Isabel Machado, and Catarina Gouveia. “You Cannot Use My Broom! I’m the Witch, You’re the Prince”: Collaboration in a virtual dramatic game. Lecture Notes in Computer Science, 2363:913– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630913.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630913.
                                        pdf.

Robertson:2002:FCS

[1463] Judy Robertson and Peter Wiemer-Hastings. Feedback on children’s stories via multiple interface agents. Lecture Notes in Computer Science, 2363: 923–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630923.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630923.
                                        pdf.

deVicente:2002:IDS

[1464] Angel de Vicente and Helen Pain. Informing the detection of the students’ motivational state: An empirical study. Lecture Notes in Computer Science, 2363:933–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630933.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2363/23630933.
pdf.

Conati:2002:MSE

[1465] Cristina Conati and Xiaoming Zhou.

Modeling students’ emotions from cognitive appraisal in educational games. Lecture Notes in Computer Science, 2363:944–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2363/23630944.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2363/23630944.
pdf.

Kort:2002:AMI

[1466] Barry Kort and Rob Reilly. An affective module for an intelligent tutoring system. Lecture Notes in Computer Science, 2363:955–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630955.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630955.
pdf.

Moreno:2002:PCP

[1467] Kristen N. Moreno, Bianca Klettke, Kiran Nibbaragandla, and Arthur C. Graesser. Perceived characteristics and pedagogical efficacy of animated conversational agents. Lecture Notes in Computer Science, 2363:963–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630963.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630963.
                                        pdf.

Graesser:2002:LAE

[1468] Arthur C. Graesser, Xiangen Hu, Natalie K. Person, Craig Stewart, Joe Toth, G. Tanner Jackson, Suresh Susarla, and Matthew Ventura. Learning about the ethical treatment of human subjects in experiments on a Web facility with a conversational agent and ITS components. Lecture Notes in Computer Science, 2363:972– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630972.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630972.
                                        pdf.

Masthoff:2002:DEN

[1469] Judith Masthoff. Design and evaluation of a navigation agent with a mixed locus of control. Lecture Notes in Computer Science, 2363:982– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2363/23630982.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2363/23630982.
                                        pdf.

Aist:2002:AHP

[1470] Gregory Aist, Barry Kort, Rob Reilly, Jack Mostow, and Rosalind Picard. Adding human-provided emotional scaffolding to an automated reading tutor that listens increases student persistence. Lecture Notes in Computer Science, 2363:992–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630992.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630992.
pdf.

Bennane:2002:ARL

[1471] Abdellah Bennane. An approach of reinforcement learning use in tutoring systems. Lecture Notes in Computer Science, 2363:993–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630993.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630993.
pdf.

Deuff:2002:MAT

[1472] Dominique Deuff, Nicole Devoldère, Béatrice Foucault, Isabelle Chanclou, Delphine Mantsos, Cédric Gégout, and Michael Picaud. An MPEG-4 authoring tool dedicated to teachers. Lecture Notes in Computer Science, 2363:994– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630994.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630994.
pdf.

duBoulay:2002:RRB

[1473] Benedict du Boulay and Rosemary Luckin. Resource reuse and broadband user modelling in ie-TV. Lecture Notes in Computer Science, 2363:995– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630995.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630995.
pdf.

Faivre:2002:IAE

[1474] Jessica Faivre, Roger Nkambou, and Claude Frasson. Integrating adaptive emotional agents in ITS. Lecture Notes in Computer Science, 2363:996– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630996.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630996.
pdf.

Giroire:2002:TDA

[1475] Hélène Giroire, Françoise Le Calvez, Jacques Duma, Gérard Tisseau, and Marie Urtasun. Targeted detection: Application to error detection in a pedagogical system. Lecture Notes in Computer Science, 2363:998–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630998.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630998.
pdf.

Harrop:2002:DEL

[1476] Andrew G. Harrop. The design of EN-CAL, a learning environment utilising multiple linked representations. Lecture Notes in Computer Science, 2363: 999–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23630999.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23630999.
pdf.

Inaba:2002:SIA

[1477] Akiko Inaba, Ryoji Ohkubo, Mitsuru Ikeda, Riichiro Mizoguchi, and Jun’ichi Toyoda. Supporting interaction analysis for collaborative learning. Lecture Notes in Computer Science, 2363:1000– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23631000.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23631000.
pdf.

Lloyd:2002:SEN

[1478] Anna Lloyd. A support environment for novice Web searchers. Lecture Notes in Computer Science, 2363:1002– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23631002.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23631002.
pdf.

Mavrikis:2002:TMA

[1479] Manolis Mavrikis and John Lee. Towards more affective tutoring systems. Lecture Notes in Computer Science, 2363:1003–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2363/23631003.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2363/23631003.
pdf.

Moulin:2002:MLU

[1480] Claude Moulin, Sylvain Giroux, Antonio Pintus, and Raffaella Sanna. Mobile lessons using geo-referenced data in e-learning. Lecture Notes in Computer Science, 2363:1004–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23631004.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23631004.
pdf.

Neau:2002:SSA

[1481] Nadège Neau and Alain Derycke. Shared scientific argumentation for lab work in physics: An application of ethnomethodology in the conception process. Lecture Notes in Computer Science, 2363:1005–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23631005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23631005.
pdf.

Rosselle:2002:SSA

[1482] Marilyne Rosselle, Pierre Gillois, Josette Morinet, and François Kohler. Student’s setup to access a WEDS. Lecture Notes in Computer Science, 2363:1006–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2363/23631006.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2363/23631006.
pdf.

Rubens:2002:LEW

[1483] M. Rubens, P. Trigano, and D. Lenne.

Learner evaluation in Web-based learning environments. Lecture Notes in Computer Science, 2363:1007–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23631007.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23631007.
pdf.

Saad:2002:CKB

[1484] Ashraf Saad and A. R. M. Zaghloul. A constructivist knowledge base architecture for adaptive intelligent tutoring systems. Lecture Notes in Computer Science, 2363:1008–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23631008.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23631008.
pdf.

Salles:2002:QMI

[1485] Paulo Salles, Bert Bredeweg, Symone Araujo, and Walter Neto. Qualitative models of interactions between two populations. Lecture Notes in Computer Science, 2363:1009–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23631009.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23631009.
pdf.

Seridi-Bouchelaghem:2002:ADL

[1486] Hassina Seridi-Bouchelaghem and Mokhtar Sellami. An adaptive distance learning architecture. Lecture Notes in Computer Science, 2363:1010– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23631010.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23631010.
pdf.

Spagnol:2002:MAR

[1487] J. P. Spagnol. Modelisation and automation of reasoning in geometry. the ARGOS system: a learning companion for high-school pupils. Lecture Notes in Computer Science, 2363:1011– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23631011.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23631011.
pdf.

Underwood:2002:ADS

[1488] Joshua Underwood and Rosemary Luckin. Adequate decision support systems must also be good learning environments. Lecture Notes in Computer Science, 2363:1012–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2363/23631012.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2363/23631012.
pdf.

Anonymous:2002:AIbb

[1489] Anonymous. Author index. Lecture Notes in Computer Science, 2363:1013– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2363/2363auth.pdf.

Ghosh:2002:MSB

[1490] Joydeep Ghosh. Multiclassifier systems: Back to the future. Lecture Notes in Computer Science, 2364:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640001.
pdf.

Zhu:2002:SVM

[1491] Ji Zhu and Trevor Hastie. Support vector machines, kernel logistic regression and boosting. Lecture Notes in Computer Science, 2364:16– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640016.
pdf.

Raudys:2002:MCS

[1492] Sarunas Raudys. Multiple classification systems in the context of feature extraction and selection. Lecture Notes in Computer Science, 2364:27– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640027.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640027.
pdf.

Windeatt:2002:BTE

[1493] Terry Windeatt and Gholamreza Ardeshir. Boosted tree ensembles for solving multiclass problems. Lecture Notes in Computer Science, 2364:42– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640042.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640042.
pdf.

Chawla:2002:DPS

[1494] N. V. Chawla, L. O. Hall, K. W. Bowyer, T. E. Moore, Jr., and W. P. Kegelmeyer. Distributed pasting of small votes. Lecture Notes in Computer Science, 2364:52–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640052.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640052.
pdf.

Skurichina:2002:BBN

[1495] Marina Skurichina, Liudmila I. Kuncheva, and Robert P. W. Duin. Bagging and boosting for the nearest mean classifier: Effects of sample size on diversity and accuracy. Lecture Notes in Computer Science, 2364:62– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640062.
pdf.

Caprile:2002:HHP

[1496] Bruno Caprile, Cesare Furlanello, and Stefano Merler. Highlighting hard patterns via AdaBoost weights evolution. Lecture Notes in Computer Science, 2364:72–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640072.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640072.
pdf.

Kuncheva:2002:UDT

[1497] Ludmila I. Kuncheva and Christopher J. Whitaker. Using diversity with three variants of boosting: Aggressive, conservative, and inverse. Lecture Notes in Computer Science, 2364: 81–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640081.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640081.
                                        pdf.

Yang:2002:MNN

[1498] Shuang Yang, Antony Browne, and Philip D. Picton. Multistage neural network ensembles. Lecture Notes in Computer Science, 2364:91–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640091.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640091.
                                        pdf.

Cohen:2002:FBS

[1499] Shimon Cohen and Nathan Intrator. Forward and backward selection in regression hybrid network. Lecture Notes in Computer Science, 2364:98– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640098.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640098.
                                        pdf.

Sharkey:2002:TMS

[1500] Amanda J. C. Sharkey. Types of multinet system. Lecture Notes in Computer Science, 2364:108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640108.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640108.
pdf.

Mola:2002:DAF

[1501] Francesco Mola and Roberta Siciliano. Discriminant analysis and factorial multiple splits in recursive partitioning for data mining. Lecture Notes in Computer Science, 2364:118– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640118.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640118.
pdf.

Ruta:2002:NMC

[1502] Dymitr Ruta and Bogdan Gabrys. New measure of classifier dependency in multiple classifier systems. Lecture Notes in Computer Science, 2364:127– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640127.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640127.
pdf.

Pekalska:2002:DCP

[1503] Elzbieta Pekalska, Robert P. W. Duin, and Marina Skurichina. A discussion on the classifier projection space for classifier combining. Lecture Notes in Computer Science, 2364:137–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640137.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640137.
                                        pdf.

Windridge:2002:GAT

[1504] D. Windridge and J. Kittler. On the general application of the tomographic classifier fusion methodology. Lecture Notes in Computer Science, 2364:149– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640149.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640149.
                                        pdf.

Altincay:2002:PPC

[1505] Hakan Altinçay and Mübeccel Demirekler. Post-processing of classifier outputs in multiple classifier systems. Lecture Notes in Computer Science, 2364:159– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640159.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640159.
                                        pdf.

Sirlantzis:2002:TMC

[1506] K. Sirlantzis, S. Hoque, and M. C. Fairhurst. Trainable multiple classifier schemes for handwritten character recognition. Lecture Notes in Computer Science, 2364:169–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640169.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640169.
pdf.

Gunter:2002:GCE

[1507] Simon Günter and Horst Bunke. Generating classifier ensembles from multiple prototypes and its application to handwriting recognition. Lecture Notes in Computer Science, 2364:179– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640179.
pdf.

Morgan:2002:AFS

[1508] Joseph T. Morgan, Alex Henneguelle, Melba M. Crawford, Joydeep Ghosh, and Amy Neuenschwander. Adaptive feature spaces for land cover classification with limited ground truth data. Lecture Notes in Computer Science, 2364:189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640189.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640189.
pdf.

Dzeroski:2002:SMR

[1509] Saso Dzeroski and Bernard Zenko. Stacking with multi-response model trees. Lecture Notes in Computer Science, 2364:201–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2364/23640201.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2364/23640201.pdf.

Lai:2002:COC

[1510] Carmen Lai, David M. J. Tax, Robert P. W. Duin, Elzbieta Pekalska, and Pavel Paclík. On combining one-class classifiers for image database retrieval. Lecture Notes in Computer Science, 2364:212–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640212.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640212.
                                        pdf.

Valentini:2002:BVA

[1511] Giorgio Valentini and Thomas G. Dietterich. Bias-variance analysis and ensembles of SVM. Lecture Notes in Computer Science, 2364:222–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640222.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640222.
                                        pdf.

Roli:2002:ECF

[1512] Fabio Roli, Sarunas Raudys, and Gian Luca Marcialis. An experimental comparison of fixed and trained fusion rules for crisp classifier outputs. Lecture Notes in Computer Science, 2364: 232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640232.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640232.
pdf.

Janeliunas:2002:RBB

[1513] Arunas Janeliunas and Sarunas Raudys.

Reduction of the boasting bias of linear experts. Lecture Notes in Computer Science, 2364:242–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640242.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640242.
pdf.

Roli:2002:ALO

[1514] Fabio Roli and Giorgio Fumera.

Analysis of linear and order statistics combiners for fusion of imbalanced classifiers. Lecture Notes in Computer Science, 2364:252–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640252.
pdf.

Masulli:2002:BCE

[1515] Francesco Masulli, Matteo Pardo, Giorgio Sberveglieri, and Giorgio Valentini. Boosting and classification of electronic nose data. Lecture Notes in Computer Science, 2364:262– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640262.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640262.
                                        pdf.

Schettini:2002:CBC

[1516] R. Schettini, C. Brambilla, and C. Cusano. Content-based classification of digital photos. Lecture Notes in Computer Science, 2364:272–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640272.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640272.
                                        pdf.

Minguillon:2002:CCV

[1517] Julià Minguillón, Anne Rosemary Tate, Carles Arús, and John R. Griffiths. Classifier combination for in vivo magnetic resonance spectra of brain tumours. Lecture Notes in Computer Science, 2364:282–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640282.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2364/23640282.
pdf.

Benfenati:2002:CCP

[1518] Emilio Benfenati, Paolo Mazzatorta, Daniel Neagu, and Giuseppina Gini. Combining classifiers of pesticides toxicity through a neurofuzzy approach. Lecture Notes in Computer Science, 2364:293–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640293.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640293.
pdf.

Cordella:2002:MES

[1519] L. P. Cordella, M. De Santo, G. Percannella, C. Sansone, and M. Vento. A multi-expert system for movie segmentation. Lecture Notes in Computer Science, 2364:304–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2364/23640304.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2364/23640304.
pdf.

Kittler:2002:DLF

[1520] J. Kittler, M. Ballette, J. Czyz, F. Roli, and L. Vandendorpe. Decision level fusion of intramodal personal identity verification experts. Lecture Notes in Computer Science, 2364:314–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640314.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640314.
                                        pdf.

Roli:2002:ECC

[1521] Fabio Roli, Josef Kittler, Giorgio Fumera, and Daniele Muntoni. An experimental comparison of classifier fusion rules for multimodal personal identity verification systems. Lecture Notes in Computer Science, 2364:325– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2364/23640325.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2364/23640325.
                                        pdf.

Anonymous:2002:AIbc

[1522] Anonymous. Author index. Lecture Notes in Computer Science, 2364:337– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2364/2364indx.pdf.

Biham:2002:NRB

[1523] Eli Biham, Orr Dunkelman, and Nathan Keller. New results on boomerang and rectangle attacks. Lecture Notes in Computer Science, 2365: 1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2365/23650001.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2365/23650001.
pdf.

Borisov:2002:MD

[1524] Nikita Borisov, Monica Chew, Rob Johnson, and David Wagner. Multiplicative differentials. Lecture Notes in Computer Science, 2365:17–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2365/23650017.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2365/23650017.
pdf.

Yanami:2002:DLC

[1525] Hitoshi Yanami, Takeshi Shimoyama, and Orr Dunkelman. Differential and linear cryptanalysis of a reduced-round SC2000. Lecture Notes in Computer Science, 2365:34–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2365/23650034.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2365/23650034.
pdf.

Moon:2002:IDC

[1526] Dukjae Moon, Kyungdeok Hwang, Wonil Lee, Sangjin Lee, and Jongin Lim. Impossible differential cryptanalysis of reduced round XTEA and TEA. Lecture Notes in Computer Science, 2365:49–61, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2365/23650049.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2365/23650049.
                                        pdf;   http://www.iacr.org/archive/
                                        fse2002/23650050/23650050.pdf.

Kuhn:2002:ICM

[1527] Ulrich Kühn. Improved cryptanalysis of MISTY1. Lecture Notes in Computer Science, 2365:61–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2365/23650061.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2365/23650061.
                                        pdf.

Shimoyama:2002:MLC

[1528] Takeshi Shimoyama, Masahiko Takenaka, and Takeshi Koshiba. Multiple linear cryptanalysis of a reduced round RC6. Lecture Notes in Computer Science, 2365:76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2365/23650076.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2365/23650076.
                                        pdf.

Yeom:2002:SCA

[1529] Yongjin Yeom, Sangwoo Park, and Iljun Kim. On the security of CAMEL-LIA against the square attack. Lecture Notes in Computer Science, 2365:89– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2365/23650089.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2365/23650089.
pdf.

Hwang:2002:SAR

[1530] Kyungdeok Hwang, Wonil Lee, Sungjae Lee, Sangjin Lee, and Jongin Lim. Saturation attacks on reduced round skipjack. Lecture Notes in Computer Science, 2365:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2365/23650100.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2365/23650100.
pdf.

Knudsen:2002:IC

[1531] Lars Knudsen and David Wagner. Integral cryptanalysis. Lecture Notes in Computer Science, 2365:112–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2365/23650112.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2365/23650112.
pdf.

Shirai:2002:IUB

[1532] Taizo Shirai, Shoji Kanamaru, and George Abe. Improved upper bounds of differential and linear characteristic probability for camellia. Lecture Notes in Computer Science, 2365:128– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2365/23650128.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2365/23650128.
                                        pdf.

Wernsdorf:2002:RFR

[1533] Ralph Wernsdorf. The round functions of RIJNDAEL generate the alternating group. Lecture Notes in Computer Science, 2365:143–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2365/23650143.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2365/23650143.
                                        pdf.

Iwata:2002:NCP

[1534] Tetsu Iwata, Tomonobu Yoshino, and Kaoru Kurosawa. Non-cryptographic primitive for pseudorandom permutation. Lecture Notes in Computer Science, 2365:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2365/23650149.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2365/23650149.pdf.

Driscoll:2002:BER

[1535] Kevin Driscoll. BeepBeep: Embedded real-time encryption. Lecture Notes in Computer Science, 2365:164– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2365/23650164.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2365/23650164.
pdf.

Watanabe:2002:NKG

[1536] Dai Watanabe, Soichi Furuya, Hirotaka Yoshida, Kazuo Takaragi, and Bart Preneel. A new keystream generator MUGI. Lecture Notes in Computer Science, 2365:179–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2365/23650179.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2365/23650179.
pdf.

Halevi:2002:SSE

[1537] Shai Halevi, Don Coppersmith, and Charanjit Jutla. Scream: a softwareefficient stream cipher. Lecture Notes in Computer Science, 2365:195–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2365/23650195.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2365/23650195.
pdf.

Ekdahl:2002:DAS

[1538] Patrik Ekdahl and Thomas Johansson. Distinguishing attacks on SOBER-t16 and t32. Lecture Notes in Computer Science, 2365:210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2365/23650210.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2365/23650210.
                                        pdf.

Dichtl:2002:LPS

[1539] Markus Dichtl and Marcus Schafheutle. Linearity properties of the SOBERt32 key loading. Lecture Notes in Computer Science, 2365:225–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2365/23650225.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2365/23650225.
                                        pdf.

Saarinen:2002:TMT

[1540] Markku-Juhani Olavi Saarinen. A time-memory tradeoff attack against LILI-128. Lecture Notes in Computer Science, 2365:231–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2365/23650231.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2365/23650231.pdf.

Jaulmes:2002:SRC

[1541] Éliane Jaulmes, Antoine Joux, and Frédéric Valette. On the security of randomized CBC-MAC beyond the birthday paradox limit: a new construction. Lecture Notes in Computer Science, 2365:237–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2365/23650237.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2365/23650237.
pdf.

Han:2002:CMV

[1542] Daewan Han, Sangwoo Park, and Seongtaek Chee. Cryptanalysis of the modified version of the hash function proposed at PKC’98. Lecture Notes in Computer Science, 2365:252– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2365/23650252.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2365/23650252.
pdf.

Kelsey:2002:CIL

[1543] John Kelsey. Compression and information leakage of plaintext. Lecture Notes in Computer Science, 2365:263– 276, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2365/23650263.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2365/23650263.
pdf.

Anonymous:2002:AIbd

[1544] Anonymous. Author index. Lecture Notes in Computer Science, 2365:277– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2365/2365auth.pdf.

Rousset:2002:KRI

[1545] Marie-Christine Rousset. Knowledge representation for information integration. Lecture Notes in Computer Science, 2366:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660001.
pdf.

Sycara:2002:IIA

[1546] Katia Sycara. Infrastructure and interoperability for agent-mediated services. Lecture Notes in Computer Science, 2366:4–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2366/23660004.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2366/23660004.
pdf.

Lallich:2002:ICR

[1547] Stéphane Lallich, Fabrice Muhlenbach, and Djamel A. Zighed. Improving classification by removing or relabeling mislabeled instances. Lecture Notes in Computer Science, 2366:5– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660005.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660005.
pdf.

Maloof:2002:ILP

[1548] Marcus A. Maloof and Ryszard S. Michalski. Incremental learning with partial instance memory. Lecture Notes in Computer Science, 2366:16– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660016.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660016.
pdf.

Slezak:2002:KBA

[1549] Dominik Ślȩzak, Piotr Synak, Alicja Wieczorkowska, and Jakub Wróblewski. KDD-based approach to musical instrument sound recognition. Lecture Notes in Computer Science, 2366:28–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2366/23660028.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2366/23660028.
pdf.

Breimer:2002:LSA

[1550] Eric Breimer and Mark Goldberg.

Learning significant alignments: An alternative to normalized local alignment. Lecture Notes in Computer Science, 2366:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2366/23660037.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2366/23660037.pdf.

Palopoli:2002:PDL

[1551] Luigi Palopoli, Giorgio Terracina, and Domenico Ursino. A plausibility description logics for reasoning with information sources having different formats and structures. Lecture Notes in Computer Science, 2366:46– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660046.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660046.
                                        pdf.

Zhong:2002:ROW

[1552] Ning Zhong and Norichika Hayazaki. Roles of ontologies for Web intelligence. Lecture Notes in Computer Science, 2366:55–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2366/23660055.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660055.
                                        pdf.

Ras:2002:HSI

[1553] Zbigniew W. Raś and Agnieszka Dardzińska. Handling semantic inconsistencies in distributed knowledge systems using ontologies. Lecture Notes in Computer Science, 2366:66– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660066.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660066.
                                        pdf.

Eklund:2002:SOI

[1554] Peter Eklund and Richard Cole. Structured ontology and information retrieval for email search and discovery. Lecture Notes in Computer Science, 2366:75–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2366/23660075.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2366/23660075.
pdf.

McClean:2002:CCH

[1555] Sally McClean, Bryan Scotney, and Fiona Palmer. Conceptual clustering of heterogeneous sequences via schema mapping. Lecture Notes in Computer Science, 2366:85–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660085.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660085.
pdf.

Esposito:2002:GCS

[1556] Roberto Esposito and Lorenza Saitta.

Is a greedy covering strategy an extreme boosting? Lecture Notes in Computer Science, 2366:94–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660094.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660094.
pdf.

Sakurai:2002:AKD

[1557] Shigeaki Sakurai, Yumi Ichimura, Akihiro Suyama, and Ryohei Orihara.

Acquisition of a knowledge dictionary from training examples including multiple values. Lecture Notes in Computer Science, 2366:103–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660103.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660103.
                                        pdf.

Klopotek:2002:MBN

[1558] Mieczysław A. Kłopotek. Mining Bayesian network structure for large sets of variables. Lecture Notes in Computer Science, 2366:114–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660114.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660114.
                                        pdf.

Merzbacher:2002:AGT

[1559] Matthew Merzbacher. Automatic generation of trivia questions. Lecture Notes in Computer Science, 2366:123– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660123.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660123.
                                        pdf.

Cholvy:2002:AQA

[1560] Laurence Cholvy and Christophe Garion. Answering queries addressed to several databases: a query evaluator which implements a majority merging approach. Lecture Notes in Computer Science, 2366:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660131.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660131.
pdf.

Fanizzi:2002:MGU

[1561] Nicola Fanizzi and Stefano Ferilli. Minimal generalizations under OI-implication. Lecture Notes in Computer Science, 2366:140–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660140.
pdf.

DiSciascio:2002:SSI

[1562] E. Di Sciascio, F. M. Donini, and M. Mongiello. I-search: a system for intelligent information search on the Web. Lecture Notes in Computer Science, 2366:149–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660149.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660149.
pdf.

Lalmas:2002:FVK

[1563] Mounia Lalmas and Thomas Roelleke. Four-valued knowledge augmentation for representing structured documents. Lecture Notes in Computer Science, 2366:158–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2366/23660158.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660158.
                                        pdf.

Kroupa:2002:WIA

[1564] Tomas Kroupa, Petr Berka, and Tomas Kocka. WISECON — an intelligent assistant for buying computers on the internet. Lecture Notes in Computer Science, 2366:167–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660167.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660167.
                                        pdf.

Sniezynski:2002:BSL

[1565] Bartłomiej Śnieżyński. Basic semantics of the logic of plausible reasoning. Lecture Notes in Computer Science, 2366: 176–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660176.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660176.
                                        pdf.

Ruan:2002:LBR

[1566] Chun Ruan, Vijay Varadharajan, and Yan Zhang. Logic-based reasoning on delegatable authorizations. Lecture Notes in Computer Science, 2366:185– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660185.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660185.
pdf.

Bosc:2002:MSF

[1567] Patrick Bosc and Olivier Pivert. Mixing selections and foreign key joins in queries against possibilistic databases. Lecture Notes in Computer Science, 2366:194–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2366/23660194.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2366/23660194.
pdf.

Kashikawa:2002:AMF

[1568] Shingo Kashikawa, Shogo Ogura, Isamu Shioya, and Takao Miura. Aggregates as meta functions. Lecture Notes in Computer Science, 2366:203– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660203.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660203.
pdf.

Rodriguez:2002:KBA

[1569] M. Andrea Rodríguez and Marcela Varas. A knowledge-based approach to querying heterogeneous databases. Lecture Notes in Computer Science, 2366: 213–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660213.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660213.
                                        pdf.

Danilowicz:2002:UUP

[1570] Czesław Daniłowicz and Huy Cuong Nguyen. Using user profiles in intelligent information retrieval. Lecture Notes in Computer Science, 2366:223– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660223.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660223.
                                        pdf.

Elomaa:2002:PRA

[1571] Tapio Elomaa. Partition-refining algorithms for learning finite state automata. Lecture Notes in Computer Science, 2366:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2366/23660232.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2366/23660232.pdf.

Laporte:2002:CFI

[1572] Marc Laporte, Noël Novelli, Rosine Cicchetti, and Lotfi Lakhal. Computing full and iceberg datacubes using partitions. Lecture Notes in Computer Science, 2366:244–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660244.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660244.
pdf.

Alphonse:2002:DAD

[1573] Erick Alphonse and Stan Matwin. A dynamic approach to dimensionality reduction in relational learning. Lecture Notes in Computer Science, 2366: 255–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660255.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660255.
pdf.

Dubois:2002:IDT

[1574] Vincent Dubois and Mohamed Quafafou.

Incremental and dynamic text mining. Lecture Notes in Computer Science, 2366:265–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2366/23660265.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2366/23660265.
pdf.

Greco:2002:LFI

[1575] Gianluigi Greco, Sergio Greco, and Ester Zumpano. A logic framework for the integration of databases. Lecture Notes in Computer Science, 2366:274– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660274.
pdf.

Lu:2002:IAL

[1576] James J. Lu, Neil V. Murray, Heydar Radjavi, Erik Rosenthal, and Peter Rosenthal. Inference for annotated logics over distributive lattices. Lecture Notes in Computer Science, 2366: 285–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660285.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660285.
pdf.

Costantini:2002:DLP

[1577] Stefania Costantini, Ramón P. Otero, Alessandro Provetti, and Tran C. Son. Disjunctive logic programs with inheritance revisited. Lecture Notes in Computer Science, 2366:294–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660294.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660294.
pdf.

Pietarinen:2002:KCA

[1578] Ahti Pietarinen. Knowledge constructions for artificial intelligence. Lecture Notes in Computer Science, 2366:303– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660303.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660303.
pdf.

Djelouah:2002:VRK

[1579] R. Djelouah, B. Duval, and S. Loiseau.

Validation and reparation of knowledge bases. Lecture Notes in Computer Science, 2366:312–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2366/23660312.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2366/23660312.pdf.

Tsumoto:2002:ADD

[1580] Shusaku Tsumoto. Automated discovery of decision rule chains using rough sets and medical diagnostic model. Lecture Notes in Computer Science, 2366: 321–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660321.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660321.
pdf.

Riff:2002:IPO

[1581] María-Cristina Riff and Xavier Bonnaire. Inheriting parents operators: a new dynamic strategy for improving evolutionary algorithms. Lecture Notes in Computer Science, 2366:333– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660333.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660333.
pdf.

Ishikawa:2002:IWR

[1582] Hiroshi Ishikawa, Toshiyuki Nakajima, Tokuyo Mizuhara, Shohei Yokoyama, Junya Nakayama, Manabu Ohta, and Kaoru Katayama. An intelligent Web recommendation system: a Web usage mining approach. Lecture Notes in Computer Science, 2366:342–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660342.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660342.
pdf.

Benel:2002:PSS

[1583] Aurélien Bénel, Sylvie Calabretto, Andréa Iacovella, and Jean-Marie Pinon. Porphyry 2001: Semantics for scholarly publications retrieval. Lecture Notes in Computer Science, 2366: 351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660351.
pdf.

Thanapandi:2002:PIP

[1584] Chitra Malini Thanapandi, Aranya Walairacht, Thanapandi Periasamy, and Shigeyuki Ohara. Preprocessor to improve performance of GA in determining bending process for sheet metal industry. Lecture Notes in Computer Science, 2366:362–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660362.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660362.
pdf.

Giordana:2002:RDC

[1585] Attilio Giordana, Paolo Terenziani, and Marco Botta. Recognizing and discovering complex events in sequences. Lecture Notes in Computer Science, 2366:374–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2366/23660374.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2366/23660374.
pdf.

Kryszkiewicz:2002:WAG

[1586] Marzena Kryszkiewicz and Marcin Gajek. Why to apply generalized disjunction-free generators representation of frequent patterns? Lecture Notes in Computer Science, 2366:383– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660383.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2366/23660383.
                                        pdf.

Malerba:2002:TLV

[1587] Donato Malerba, Annalisa Appice, Michelangelo Ceci, and Marianna Monopoli. Trading-off local versus global effects of regression nodes in model trees. Lecture Notes in Computer Science, 2366:393–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660393.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660393.
                                        pdf.

Hwang:2002:EIA

[1588] Byungyeon Hwang, Euichan Kim, and Bogju Lee. An efficient intelligent agent system for automated recommendation in electronic commerce. Lecture Notes in Computer Science, 2366:403– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660403.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660403.
                                        pdf.

Bertino:2002:MXD

[1589] Elisa Bertino, Giovanna Guerrini, and Marco Mesiti. Matching an XML document against a set of DTDs. Lecture Notes in Computer Science, 2366:412– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2366/23660412.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660412.
pdf.

Bentayeb:2002:DTM

[1590] Fadila Bentayeb and Jérôme Darmont. Decision tree modeling with relational views. Lecture Notes in Computer Science, 2366:423–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660423.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660423.
pdf.

Harms:2002:DSA

[1591] Sherri K. Harms, Jitender Deogun, and Tsegaye Tadesse. Discovering sequential association rules with constraints and time lags in multiple sequences. Lecture Notes in Computer Science, 2366:432–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660432.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660432.
pdf.

Greco:2002:MAR

[1592] Salvatore Greco, Roman Slowinski, and Jerzy Stefanowski. Mining association rules in preference-ordered data. Lecture Notes in Computer Science, 2366: 442–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660442.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660442.
                                        pdf.

Bruha:2002:UAV

[1593] Ivan Bruha. Unknown attribute values processing by meta-learner. Lecture Notes in Computer Science, 2366: 451–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660451.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660451.
                                        pdf.

Ryu:2002:IBC

[1594] Yeonseung Ryu, Kyoungwoon Cho, Youjip Won, and Kern Koh. Intelligent buffer cache management in multimedia data retrieval. Lecture Notes in Computer Science, 2366:462– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660462.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660462.
                                        pdf.

Badia:2002:LFS

[1595] Antonio Badia. A logical formalization of semistructured data models. Lecture Notes in Computer Science, 2366: 472–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660472.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2366/23660472.
pdf.

Camacho:2002:STP

[1596] David Camacho, José M. Molina, Daniel Borrajo, and Ricardo Aler. Solving travel problems by integrating WEB information with planning. Lecture Notes in Computer Science, 2366: 482–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660482.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660482.
pdf.

Aberg:2002:FGT

[1597] Cécile Aberg and Nahid Shahmehri. A framework for generating task specific information extraction systems. Lecture Notes in Computer Science, 2366: 491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660491.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660491.
pdf.

Cali:2002:FFR

[1598] Andrea Calì, Diego Calvanese, Giuseppe De Giacomo, and Maurizio Lenzerini. A formal framework for reasoning on UML class diagrams. Lecture Notes in Computer Science, 2366:503– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2366/23660503.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660503.
                                        pdf.

Deeb:2002:BAD

[1599] Kevin Deeb and Steven Lewis. A biological approach to the development of computer autoimmune systems. Lecture Notes in Computer Science, 2366: 514–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660514.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660514.
                                        pdf.

Malerba:2002:ALA

[1600] Donato Malerba, Floriana Esposito, and Oronzo Altamura. Adaptive layout analysis of document images. Lecture Notes in Computer Science, 2366: 526–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660526.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660526.
                                        pdf.

DeCarolis:2002:NBP

[1601] Berardina De Carolis and Francesca A. Lisi. A NLG-based presentation method for supporting KDD end-users. Lecture Notes in Computer Science, 2366:535–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2366/23660535.htm;   http:

//link.springer-ny.com/link/service/
series/0558/papers/2366/23660535.
pdf.

Simpson:2002:TSS

[1602] R. M. Simpson and T. L. McCluskey. A tool supported structured method for planning domain acquisition. Lecture Notes in Computer Science, 2366:544– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660544.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660544.
pdf.

Ross:2002:PA

[1603] Robert Ross, V. S. Subrahmanian, and John Grant. Probabilistic aggregates. Lecture Notes in Computer Science, 2366:553–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660553.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660553.
pdf.

DeMarchi:2002:SUD

[1604] Fabien De Marchi, Stéphane Lopes, and Jean-Marc Petit. Samples for understanding data-semantics in relations. Lecture Notes in Computer Science, 2366:565–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2366/23660565.
htm;   http://link.springer-ny.

                                        com/link/service/series/0558/papers/
                                        2366/23660565.pdf.

Esposito:2002:CMS

[1605] Floriana Esposito, Stefano Ferilli, Nicola Fanizzi, Teresa Maria Altomare Basile, and Nicola Di Mauro. Cooperation of multiple strategies for automated learning in complex environments. Lecture Notes in Computer Science, 2366:574–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660574.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660574.
                                        pdf.

Kim:2002:CFU

[1606] Eunju Kim, Wooju Kim, and Yillbyung Lee. Classifier fusion using local confidence. Lecture Notes in Computer Science, 2366:583–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2366/23660583.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2366/23660583.
                                        pdf.

Tsymbal:2002:FSE

[1607] Alexey Tsymbal, Seppo Puuronen, and David Patterson. Feature selection for ensembles of simple Bayesian classifiers. Lecture Notes in Computer Science, 2366:592–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/2366/23660592.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660592.
pdf.

Inatani:2002:DSS

[1608] Shutaro Inatani and Einoshin Suzuki.

Data squashing for speeding up boosting-based outlier detection. Lecture Notes in Computer Science, 2366: 601–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2366/23660601.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2366/23660601.
pdf.

Anonymous:2002:AIbe

[1609] Anonymous. Author index. Lecture Notes in Computer Science, 2366:613– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2366/2366auth.pdf.

Cai:2002:ENS

[1610] Xing Cai and Glenn Terje Lines. Enabling numerical and software technologies for studying the electrical activity in human heart. Lecture Notes in Computer Science, 2367:3– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670003.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670003.
pdf.

Cebral:2002:PPS

[1611] J. Cebral, R. Löhner, P. L. Choyke, and P. J. Yim. Parallel patient-specific computational haemodynamics. Lecture Notes in Computer Science, 2367: 18–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670018.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670018.
                                        pdf.

Dongarra:2002:HPC

[1612] Jack Dongarra. High performance computing, computational grid, and numerical libraries. Lecture Notes in Computer Science, 2367:35–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670035.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670035.
                                        pdf.

vonLaszewski:2002:GCE

[1613] Gregor von Laszewski. Grid computing: Enabling a vision for collaborative research. Lecture Notes in Computer Science, 2367:37–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670037.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670037.
                                        pdf.

Mirza:2002:HWM

[1614] Jamshed Mirza. HPC — what might the future hold? Lecture Notes in Computer Science, 2367:53– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670053.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670053.
pdf.

Nieminen:2002:MPM

[1615] R. M. Nieminen. Multi-physics and multi-scale modelling of materials processing. Lecture Notes in Computer Science, 2367:55–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670055.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670055.
pdf.

Reid:2002:CAF

[1616] John Reid. Co-array Fortran for full and sparse matrices. Lecture Notes in Computer Science, 2367:61– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670061.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670061.
pdf.

Sibani:2002:MLG

[1617] Paolo Sibani and J. C. Schön. Measuring the local geometry of valleys in complex energy landscapes by exhaustive exploration: The lid method. Lecture Notes in Computer Science, 2367: 62–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670062.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670062.
pdf.

Waananen:2002:OAP

[1618] A. Wäänänen, M. Ellert, A. Konstantinov, B. Kónya, and O. Smirnova. An overview of an architecture proposal for a high energy physics grid. Lecture Notes in Computer Science, 2367: 76–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670076.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670076.
pdf.

Ashrafi:2002:DMA

[1619] Mafruz Zaman Ashrafi, David Taniar, and Kate A. Smith. A data mining architecture for clustered environments. Lecture Notes in Computer Science, 2367:89–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2367/23670089.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2367/23670089.
pdf.

Dhaene:2002:AFR

[1620] Tom Dhaene. Automated fitting and rational modeling algorithm for EMbased S-parameter data. Lecture Notes in Computer Science, 2367:99– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670099.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670099.
pdf.

Liu:2002:PHP

[1621] Zhen Liu and Minyi Guo. A proposal of high performance data mining system. Lecture Notes in Computer Science, 2367:106–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670106.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670106.
pdf.

Szeto:2002:QPR

[1622] Kwok Yip Szeto and Rui Jiang. A quasi-parallel realization of the investment frontier in computer resource allocation using simple genetic algorithm on a single computer. Lecture Notes in Computer Science, 2367:116– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670116.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670116.
pdf.

Talia:2002:PKD

[1623] Domenico Talia. Parallelism in knowledge discovery techniques. Lecture Notes in Computer Science, 2367:127– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670127.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670127.
                                        pdf.

Balle:2002:NAP

[1624] Susanne M. Balle, Bevin R. Brett, Chih-Ping Chen, and David LaFrance-Linden. A new approach to parallel debugger architecture. Lecture Notes in Computer Science, 2367:139– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670139.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670139.
                                        pdf.

DiMartino:2002:AAC

[1625] Beniamino Di Martino. ALCOR — an algorithmic concept recognition tool to support high level parallel program development. Lecture Notes in Computer Science, 2367:150– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670150.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670150.
                                        pdf.

Huttunen:2002:MCC

[1626] Pentti Huttunen, Jouni Ikonen, and Jari Porras. MPIT — communication/computation paradigm for networks of SMP workstations. Lecture Notes in Computer Science, 2367:160– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670160.
pdf.

Pilz:2002:COT

[1627] Norbert A. Pilz and Kenneth Adamson. Code optimization techniques of data-intensive tasks onto statically scheduled architectures: Optimal performance on the TigerSharc. Lecture Notes in Computer Science, 2367:172– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670172.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670172.
pdf.

Baiardi:2002:PLP

[1628] Fabrizio Baiardi, Paolo Mori, and Laura Ricci. PIT: a library for the parallelization of irregular problems. Lecture Notes in Computer Science, 2367: 185–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670185.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2367/23670185.
                                        pdf.

Chung:2002:PIR

[1629] Yoojin Chung. Parallel information retrieval with query expansion. Lecture Notes in Computer Science, 2367:195– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670195.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670195.
                                        pdf.

Guo:2002:RCC

[1630] Minyi Guo, Zhen Liu, Chengfei Liu, and Li Li. Reducing communication cost for parallelizing irregular scientific codes. Lecture Notes in Computer Science, 2367:203–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670203.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670203.
                                        pdf.

Lee:2002:IPC

[1631] Nung Kion Lee, David Taniar, J. Wenny Rahayu, and Mafruz Zaman Ashrafi. Implementation of parallel collection equi-join using MPI. Lecture Notes in Computer Science, 2367:217– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670217.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2367/23670217.
pdf.

Vermin:2002:PEP

[1632] Willem Vermin. Practical experiences in parallelizing existent computer programs. Lecture Notes in Computer Science, 2367:227–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670227.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670227.
pdf.

Aversa:2002:EDO

[1633] Rocco Aversa, Beniamino Di Martino, Thomas Fahringer, and Salvatore Venticinque. On the evaluation of the distributed objects and mobile agents programming models for a distributed optimization application. Lecture Notes in Computer Science, 2367: 233–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670233.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670233.
pdf.

Bawa:2002:PTC

[1634] Seema Bawa and G. K. Sharma. A parallel transitive closure computation algorithm for VLSI test generation. Lecture Notes in Computer Science, 2367: 243–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2367/23670243.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670243.
                                        pdf.

Ha:2002:SEF

[1635] Keum-Sook Ha, Eun-Kyung Ryu, and Kee-Young Yoo. Space-efficient first race detection in shared memory programs with nested parallelism. Lecture Notes in Computer Science, 2367:253– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670253.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670253.
                                        pdf.

Ryu:2002:PMF

[1636] Eun-Kyung Ryu, Keum-Sook Ha, and Kee-Young Yoo. A practical method for on-the-fly data race detection. Lecture Notes in Computer Science, 2367: 264–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670264.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670264.
                                        pdf.

Song:2002:PMA

[1637] Samuel Moon-Ho Song, Gunho Lee, Sunghyun Kim, Manhee Lee, Hyeokman Kim, and Dong-Sik Jang. Parallelisms in MPEG and its applications to 3-D visualization. Lecture Notes in Computer Science, 2367:274– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670274.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670274.
pdf.

Andersen:2002:RFI

[1638] Bjarne S. Andersen, John A. Gunnels, Fred Gustavson, and Jerzy Wasniewski. A recursive formulation of the inversion of symmetric positive definite matrices in packed storage data format. Lecture Notes in Computer Science, 2367:287–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670287.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670287.
pdf.

Jonsson:2002:PTS

[1639] Isak Jonsson and Bo Kågström. Parallel two-sided sylvester-type matrix equation solvers for SMP systems using recursive blocking. Lecture Notes in Computer Science, 2367:297–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670297.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670297.
pdf.

Kowarschik:2002:POM

[1640] Markus Kowarschik, Ulrich Rüde, Nils Thürey, and Christian Weiß. Performance optimization of 3D multigrid on hierarchical memory architectures. Lecture Notes in Computer Science, 2367:307–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2367/23670307.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2367/23670307.
pdf.

Adlerborn:2002:PBA

[1641] Björn Adlerborn, Krister Dackland, and Bo Kågström. Parallel and blocked algorithms for reduction of a regular matrix pair to hessenberg-triangular and generalized Schur forms. Lecture Notes in Computer Science, 2367:319– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670319.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670319.
pdf.

Benner:2002:ESR

[1642] Peter Benner, Rafael Mayo, Enrique S. Quintana-Ortí, and Gregorio Quintana-Ortí. Enhanced services for remote model reduction of largescale dense linear systems. Lecture Notes in Computer Science, 2367:329– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670329.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670329.
pdf.

Hakula:2002:HFI

[1643] Harri Hakula, Juha Ruokolainen, and Jouni Malinen. HUTI: Framework for iterative solvers. Lecture Notes in Computer Science, 2367:339–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670339.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670339.
pdf.

Chen:2002:BFD

[1644] Hsin-Chu Chen. A block Fourier decomposition method. Lecture Notes in Computer Science, 2367:351–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670351.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670351.
pdf.

Ku:2002:NPA

[1645] Kyo-Min Ku, Kyeoung-Ju Ha, Hyun-Sung Kim, and Kee-Young Yoo. New parallel architecture for modular multiplication and squaring based on cellular automata. Lecture Notes in Computer Science, 2367:359–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670359.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670359.
pdf.

Lensu:2002:PIT

[1646] Anssi Lensu and Pasi Koikkalainen. A parallel implementation of the treestructured self-organizing map. Lecture Notes in Computer Science, 2367: 370–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670370.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670370.
                                        pdf.

Takahashi:2002:BAP

[1647] Daisuke Takahashi. A blocking algorithm for parallel 1-D FFT on sharedmemory parallel computers. Lecture Notes in Computer Science, 2367:380– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670380.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670380.
                                        pdf.

Beletskyy:2002:TPL

[1648] Volodymyr Beletskyy. A technique for parallel loop execution. Lecture Notes in Computer Science, 2367:393– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670393.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670393.
                                        pdf.

Bonnaire:2002:SAD

[1649] Xavier Bonnaire and María-Cristina Riff. A self-adaptable distributed evolutionary algorithm to tackle space planning problems. Lecture Notes in Computer Science, 2367:403–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670403.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670403.
pdf.

Garcia:2002:EPS

[1650] G. Cerruela García, E. López Espinosa, I. Luque Ruiz, and M. A. Gómez-Nieto. Efficient parallel solution to calculate all cycles in graphs. Lecture Notes in Computer Science, 2367:411– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670411.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670411.
pdf.

Banino:2002:SSM

[1651] C. Banino, O. Beaumont, A. Legrand, and Y. Robert. Scheduling strategies for master-slave tasking on heterogeneous processor grids. Lecture Notes in Computer Science, 2367:423– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670423.htm;   http:/
/link.springer-ny.com/link/service/

                                        series/0558/papers/2367/23670423.
                                        pdf.

Elster:2002:HPC

[1652] Anne C. Elster. High-performance computing: Past, present, and future. Lecture Notes in Computer Science, 2367:433–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2367/23670433.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670433.
                                        pdf.

Sack:2002:FMB

[1653] Paul Sack and Anne C. Elster. Fast MPI broadcasts through reliable multicasting. Lecture Notes in Computer Science, 2367:445–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670445.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670445.
                                        pdf.

Steinbeck:2002:FBD

[1654] Timm M. Steinbeck, Volker Lindenstruth, Dieter Röhrich, Anders Strand Vestbo, and Arne Wiebalck. A framework for building distributed data flow chains in clusters. Lecture Notes in Computer Science, 2367:454–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670454.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2367/23670454.
pdf.

Andersson:2002:PIP

[1655] Ulf Andersson and Fredrik Hedman. Performance of an IBM pwr4 node for the GEMS TD codes and parallacs. Lecture Notes in Computer Science, 2367:467–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670467.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670467.
pdf.

Mourino:2002:CBS

[1656] José Carlos Mouriño, Patricia González, María J. Martín, and Ramón Doallo. A cluster-based solution for a high performance air quality simulation. Lecture Notes in Computer Science, 2367: 476–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670476.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670476.
pdf.

Nikolova:2002:CCP

[1657] Kirilka Nikolova, Sou Pei You, and Masahiro Sowa. Compiler-controlled parallelism-independent scheduling for parallel and distributed systems. Lecture Notes in Computer Science, 2367: 484–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/2367/23670484.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670484.
                                        pdf.

Qiao:2002:OPA

[1658] Xiangzhen Qiao. Optimization of parallel algorithms on cluster of SMP’s. Lecture Notes in Computer Science, 2367:494–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2367/23670494.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670494.
                                        pdf.

Fard:2002:RBL

[1659] Nasser S. Fard and Indra Gunawan. Reliability bounds for large multistage interconnection networks. Lecture Notes in Computer Science, 2367: 507–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670507.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670507.
                                        pdf.

Galyuk:2002:GTD

[1660] Y. P. Galyuk, V. P. Memnonov, S. E. Zhuravleva, and V. I. Zolotarev. Grid technology with dynamic load balancing for Monte Carlo simulations. Lecture Notes in Computer Science, 2367: 515–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670515.htm;   http:/

/link.springer-ny.com/link/service/
series/0558/papers/2367/23670515.
pdf.

Goodyer:2002:PGB

[1661] Christopher Goodyer, Jason Wood, and Martin Berzins. A parallel grid based PSE for EHL problems. Lecture Notes in Computer Science, 2367:521– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670521.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670521.
pdf.

Salehnamadi:2002:SBR

[1662] Mohammad R. Salehnamadi and Mehdi N. Fesharaki. A “Single-Box” re-routing architecture for a 3stage rearrangeable CLOS interconnection networks. Lecture Notes in Computer Science, 2367:531–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670531.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670531.
pdf.

Huttunen:2002:ELB

[1663] Pentti Huttunen, Jouni Ikonen, and Jari Porras. Enhancing load balancing in a data-parallel GSM network simulation through applicationspecific information. Lecture Notes in Computer Science, 2367:542–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670542.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670542.
                                        pdf.

Diamessis:2002:ATD

[1664] Peter Diamessis, William Kerney, Scott B. Baden, and Keiko Nomura. Automated tracking of 3-D overturn patches in direct numerical simulation of stratified homogeneous turbulence. Lecture Notes in Computer Science, 2367:557–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670557.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670557.
                                        pdf.

Heikonen:2002:ILB

[1665] Jussi Heikonen and Kalle Eerola. Improving load balance in a weather code: Asynchronous output in HIRLAM with MPI. Lecture Notes in Computer Science, 2367:567–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2367/23670567.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2367/23670567.
                                        pdf.

Lhomme:2002:PSP

[1666] Frédéric Lhommé, Delphine Wolfersberger, Stéphane Vialle, and Nicolas Fressengeas. Parallel simulation of photorefractive material for the design of all-optical components. Lecture Notes in Computer Science, 2367:578– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670578.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670578.
pdf.

Raghavan:2002:SSM

[1667] P. Raghavan, M. A. James, J. C.

Newman, and B. R. Seshadri. Scalable sparse matrix techniques for modeling crack growth. Lecture Notes in Computer Science, 2367:588–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2367/23670588.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2367/23670588.
pdf.

Suviola:2002:PLB

[1668] Tomi Suviola. Parallelization of a lattice Boltzmann suspension flow solver. Lecture Notes in Computer Science, 2367:603–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2367/23670603.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2367/23670603.
pdf.

Anonymous:2002:AIbf

[1669] Anonymous. Author index. Lecture Notes in Computer Science, 2367:611– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2367/2367auth.pdf.

Hagerup:2002:EQ

[1670] Torben Hagerup and Rajeev Raman. An efficient quasidictionary. Lecture Notes in Computer Science, 2368:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680001.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680001.
                                        pdf.

Mannila:2002:CPD

[1671] Heikki Mannila. Combining pattern discovery and probabilistic modeling in data mining. Lecture Notes in Computer Science, 2368:19–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680019.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680019.
                                        pdf.

Alstrup:2002:TSE

[1672] Stephen Alstrup, Gerth Stølting Brodal, Inge Li Gørtz, and Theis Rauhe. Time and space efficient multi-method dispatching. Lecture Notes in Computer Science, 2368:20– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2368/23680020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680020.
pdf.

Augustine:2002:LTA

[1673] John E. Augustine and Steven S.

Seiden. Linear time approximation schemes for vehicle scheduling. Lecture Notes in Computer Science, 2368: 30–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680030.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680030.
pdf.

Hepner:2002:MML

[1674] Clint Hepner and Cliff Stein. Minimizing makespan for the lazy bureaucrat problem. Lecture Notes in Computer Science, 2368:40–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680040.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680040.
pdf.

Mastrolilli:2002:PSM

[1675] Monaldo Mastrolilli. A PTAS for the single machine scheduling problem with controllable processing times. Lecture Notes in Computer Science, 2368: 51–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680051.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680051.
                                        pdf.

Eidenbenz:2002:OIR

[1676] Stephan Eidenbenz. Optimum inapproximability results for finding minimum hidden guard sets in polygons and terrains. Lecture Notes in Computer Science, 2368:60–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680060.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680060.
                                        pdf.

Goswami:2002:SRS

[1677] Partha P. Goswami, Sandip Das, and Subhas C. Nandy. Simplex range searching and k nearest neighbors of a line segment in 2D. Lecture Notes in Computer Science, 2368:69– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680069.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680069.
                                        pdf.

Levcopoulos:2002:AAC

[1678] Christos Levcopoulos, Andrzej Lingas, and Joseph S. B. Mitchell. Adaptive algorithms for constructing convex hulls and triangulations of polygonal chains. Lecture Notes in Computer Science, 2368:80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680080.
pdf.

Lev-Tov:2002:EAA

[1679] Nissan Lev-Tov and David Peleg.

Exact algorithms and approximation schemes for base station placement problems. Lecture Notes in Computer Science, 2368:90–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680090.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680090.
pdf.

Qin:2002:FAL

[1680] Zhongping Qin and Binhai Zhu. A factor-2 approximation for labeling points with maximum sliding labels. Lecture Notes in Computer Science, 2368:100–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2368/23680100.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2368/23680100.
pdf.

Roy:2002:OAS

[1681] Sasanka Roy, Partha P. Goswami, Sandip Das, and Subhas C. Nandy. Optimal algorithm for a special pointlabeling problem. Lecture Notes in Computer Science, 2368:110–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680110.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680110.
                                        pdf.

Sanders:2002:RAA

[1682] Peter Sanders and Berthold Vöcking. Random arc allocation and applications. Lecture Notes in Computer Science, 2368:121–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2368/23680121.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2368/23680121.pdf.

Sharir:2002:NGP

[1683] Micha Sharir and Shakhar Smorodinsky. On neighbors in geometric permutations. Lecture Notes in Computer Science, 2368:131–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680131.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680131.
                                        pdf.

Agnarsson:2002:PGI

[1684] Geir Agnarsson, Peter Damaschke, and Magnús M. Halldórsson. Powers of geometric intersection graphs and dispersion algorithms. Lecture Notes in Computer Science, 2368:140– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680140.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680140.
pdf.

Alber:2002:EDR

[1685] Jochen Alber, Michael R. Fellows, and Rolf Niedermeier. Efficient data reduction for DOMINATING SET: a linear problem kernel for the planar case. Lecture Notes in Computer Science, 2368: 150–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680150.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680150.
pdf.

Broersma:2002:PGC

[1686] Hajo Broersma, Fedor V. Fomin, Jan Kratochvíl, and Gerhard J. Woeginger. Planar graph coloring with forbidden subgraphs: Why trees and paths are dangerous. Lecture Notes in Computer Science, 2368:160–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680160.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680160.
pdf.

Chlebik:2002:AHS

[1687] Miroslav Chlebík and Janka Chlebíková.

Approximation hardness of the Steiner tree problem on graphs. Lecture Notes in Computer Science, 2368:170– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680170.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680170.
                                        pdf.

Ellis:2002:DSP

[1688] J. Ellis, H. Fan, and Michael R. Fellows. The dominating set problem is fixed parameter tractable for graphs of bounded genus. Lecture Notes in Computer Science, 2368:180– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680180.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680180.
                                        pdf.

Gabow:2002:DVM

[1689] Harold N. Gabow and Seth Pettie. The dynamic vertex minimum problem and its application to clusteringtype approximation algorithms. Lecture Notes in Computer Science, 2368: 190–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680190.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680190.
                                        pdf.

Golynski:2002:PTA

[1690] Alexander Golynski and Joseph D. Horton. A polynomial time algorithm to find the minimum cycle basis of a regular matroid. Lecture Notes in Computer Science, 2368:200– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680200.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680200.
pdf.

Konemann:2002:AAE

[1691] Jochen Könemann, Yanjun Li, Ojas Parekh, and Amitabh Sinha. Approximation algorithms for edge-dilation k-center problems. Lecture Notes in Computer Science, 2368:210–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680210.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680210.
pdf.

Ramnath:2002:FFA

[1692] Sarnath Ramnath. Forewarned is fore-armed: Dynamic digraph connectivity with lookahead speeds up a static clustering algorithm. Lecture Notes in Computer Science, 2368:220– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680220.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680220.
pdf.

Shamir:2002:IAR

[1693] Ron Shamir and Dekel Tsur. Improved algorithms for the random cluster graph model. Lecture Notes in Computer Science, 2368:230–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680230.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680230.
                                        pdf.

Skulrattanakulchai:2002:LVC

[1694] San Skulrattanakulchai. Δ-list vertex coloring in linear time. Lecture Notes in Computer Science, 2368:240– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680240.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680240.
                                        pdf.

Demaine:2002:RLD

[1695] Erik D. Demaine, Alejandro López-Ortiz, and J. Ian Munro. Robot localization without depth perception. Lecture Notes in Computer Science, 2368: 249–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680249.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680249.
                                        pdf.

Lopez-Ortiz:2002:OPH

[1696] Alejandro López-Ortiz and Sven Schuierer. Online parallel heuristics and robot searching under the competitive framework. Lecture Notes in Computer Science, 2368:260–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680260.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680260.
pdf.

Sztainberg:2002:AHF

[1697] Marcelo O. Sztainberg, Esther M.

Arkin, Michael A. Bender, and Joseph S. B. Mitchell. Analysis of heuristics for the freeze-tag problem. Lecture Notes in Computer Science, 2368:270– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680270.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680270.
pdf.

Arkin:2002:AMT

[1698] Esther M. Arkin, Refael Hassin, Shlomi Rubinstein, and Maxim Sviridenko. Approximations for maximum transportation problem with permutable supply vector and other capacitated star packing problems. Lecture Notes in Computer Science, 2368:280–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680280.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680280.
pdf.

Azar:2002:ANA

[1699] Yossi Azar, Leah Epstein, Yossi Richter, and Gerhard J. Woeginger. All-norm approximation algorithms. Lecture Notes in Computer Science, 2368:288–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2368/23680288.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680288.
                                        pdf.

Bazgan:2002:ADI

[1700] Cristina Bazgan, W. Fernandez de la Vega, and Marek Karpinski. Approximability of dense instances of NEAR-EST CODEWORD problem. Lecture Notes in Computer Science, 2368:298– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680298.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680298.
                                        pdf.

Anand:2002:CCR

[1701] R. Sai Anand, Thomas Erlebach, Alexander Hall, and Stamatis Stefanakos. Call control with k rejections. Lecture Notes in Computer Science, 2368:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2368/23680308.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680308.
                                        pdf.

Even:2002:NDP

[1702] Guy Even, Guy Kortsarz, and Wolfgang Slany. On network design problems: Fixed cost flows and the covering Steiner problem. Lecture Notes in Computer Science, 2368:318–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680318.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680318.
pdf.

Frederiksen:2002:PB

[1703] Jens S. Frederiksen and Kim S.

Larsen. Packet bundling. Lecture Notes in Computer Science, 2368:328– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680328.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680328.
pdf.

Puri:2002:AMC

[1704] Anuj Puri and Stavros Tripakis. Algorithms for the multi-constrained routing problem. Lecture Notes in Computer Science, 2368:338–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680338.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680338.
pdf.

Karkkainen:2002:CTG

[1705] Juha Kärkkäinen. Computing the threshold for q-gram filters. Lecture Notes in Computer Science, 2368:348– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680348.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680348.
                                        pdf.

Peer:2002:GPI

[1706] Itsik Pe’er, Ron Shamir, and Roded Sharan. On the generality of phylogenies from incomplete directed characters. Lecture Notes in Computer Science, 2368:358–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680358.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680358.
                                        pdf.

Albert:2002:SF

[1707] M. H. Albert and M. D. Atkinson. Sorting with a forklift. Lecture Notes in Computer Science, 2368:368– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680368.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680368.
                                        pdf.

Bodlaender:2002:TDS

[1708] Hans L. Bodlaender and Fedor V.

Fomin. Tree decompositions with small cost. Lecture Notes in Computer Science, 2368:378–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/2368/23680378.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
2368/23680378.pdf.

Bodlaender:2002:CTM

[1709] Hans L. Bodlaender and Udi Rotics.

Computing the treewidth and the minimum fill-in with the modular decomposition. Lecture Notes in Computer Science, 2368:388–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680388.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680388.
pdf.

Katajainen:2002:PTA

[1710] Jyrki Katajainen and Jeppe Nejsum Madsen. Performance tuning an algorithm for compressing relational tables. Lecture Notes in Computer Science, 2368:398–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680398.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680398.
pdf.

Katajainen:2002:RPA

[1711] Jyrki Katajainen and Tomi A. Pasanen. A randomized in-place algorithm for positioning the k^{th} element in a multiset. Lecture Notes in Computer Science, 2368:408–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680408.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680408.
                                        pdf.

Lai:2002:PRL

[1712] Tony W. Lai. Paging on a RAM with limited resources. Lecture Notes in Computer Science, 2368:418–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680418.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680418.
                                        pdf.

DalPalu:2002:OAF

[1713] A. Dal Palú, E. Pontelli, and D. Ranjan. An optimal algorithm for finding NCA on pure pointer machines. Lecture Notes in Computer Science, 2368: 428–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2368/23680428.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2368/23680428.
                                        pdf.

Soisalon-Soininen:2002:ACB

[1714] Eljas Soisalon-Soininen and Peter Widmayer. Amortized complexity of bulk updates in AVL-trees. Lecture Notes in Computer Science, 2368:439– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2368/23680439.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2368/23680439.
pdf.

Anonymous:2002:AIbg

[1715] Anonymous. Author index. Lecture Notes in Computer Science, 2368:449– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
papers/2368/2368auth.pdf.

Bhargava:2002:GCG

[1716] Manjul Bhargava. Gauss composition and generalizations. Lecture Notes in Computer Science, 2369:1– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690001.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690001.
pdf.

Coates:2002:ECC

[1717] John Coates. Elliptic curves – the crossroads of theory and computation. Lecture Notes in Computer Science, 2369:9–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2369/23690009.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2369/23690009.
pdf.

Joux:2002:WTP

[1718] Antoine Joux. The Weil and Tate pairings as building blocks for public key cryptosystems. Lecture Notes in Computer Science, 2369:20–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690020.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690020.
pdf.

Poonen:2002:UEC

[1719] Bjorn Poonen. Using elliptic curves of rank one towards the undecidability of hilbert’s tenth problem over rings of algebraic integers. Lecture Notes in Computer Science, 2369:33– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690033.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690033.
pdf.

Satoh:2002:APC

[1720] Takakazu Satoh. On p-adic point counting algorithms for elliptic curves over finite fields. Lecture Notes in Computer Science, 2369:43–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690043.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690043.
pdf.

Bosma:2002:AEN

[1721] Wieb Bosma and Bart de Smit.

On arithmetically equivalent number fields of small degree. Lecture Notes in Computer Science, 2369:67– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690067.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690067.
pdf.

Cohen:2002:SDC

[1722] Henri Cohen, Francisco Diaz y Diaz, and Michel Olivier. A survey of discriminant counting. Lecture Notes in Computer Science, 2369:80–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690080.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690080.
pdf.

Everest:2002:HRM

[1723] Graham Everest, Peter Rogers, and Thomas Ward. A higher-rank Mersenne problem. Lecture Notes in Computer Science, 2369:95–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/2369/23690095.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690095.
                                        pdf.

Fukuda:2002:ASM

[1724] Takashi Fukuda and Keiichi Komatsu. An application of Siegel modular functions to Kronecker’s limit formula. Lecture Notes in Computer Science, 2369: 108–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2369/23690108.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690108.
                                        pdf.

Jacobson:2002:CAN

[1725] Michael J. Jacobson, Jr. and Alfred J. van der Poorten. Computational aspects of NUCOMP. Lecture Notes in Computer Science, 2369:120– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2369/23690120.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690120.
                                        pdf.

Louboutin:2002:ECC

[1726] Stéphane R. Louboutin. Efficient computation of class numbers of real Abelian number fields. Lecture Notes in Computer Science, 2369:134–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/

bibs/2369/23690134.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690134.
pdf.

Vollmer:2002:ABA

[1727] Ulrich Vollmer. An accelerated buchmann algorithm for regulator computation in real quadratic fields. Lecture Notes in Computer Science, 2369:148– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690148.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690148.
pdf.

Auer:2002:SGC

[1728] Roland Auer and Jaap Top. Some genus 3 curves with many points. Lecture Notes in Computer Science, 2369: 163–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690163.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690163.
pdf.

Bruin:2002:TGG

[1729] Nils Bruin and Noam D. Elkies. Trinomials ax^{7} + bx + c and ax^{8} + bx + c with Galois groups of order 168 and 8 · 168. Lecture Notes in Computer Science, 2369:172–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690172.htm;   http:/

                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690172.
                                        pdf.

Gonzalez-Jimenez:2002:CMJ

[1730] Enrique González-Jiménez, Josep González, and Jordi Guàrdia. Computations on modular Jacobian surfaces. Lecture Notes in Computer Science, 2369:189–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/bibs/2369/23690189.htm;   http:
                                        //link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690189.
                                        pdf.

Kresch:2002:IPP

[1731] Andrew Kresch and Yuri Tschinkel. Integral points on punctured Abelian surfaces. Lecture Notes in Computer Science, 2369:198–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2369/23690198.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2369/23690198.pdf.

Shaska:2002:GCS

[1732] Tony Shaska. Genus 2 curves with (3, 3)-split Jacobian and large automorphism group. Lecture Notes in Computer Science, 2369:205–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2369/23690205.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2369/23690205.
pdf.

Verrill:2002:TMS

[1733] Helena A. Verrill. Transportable modular symbols and the intersection pairing. Lecture Notes in Computer Science, 2369:219–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690219.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690219.
pdf.

Couveignes:2002:AMC

[1734] Jean-Marc Couveignes and Thierry Henocq. Action of modular correspondences around CM points. Lecture Notes in Computer Science, 2369:234– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690234.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690234.
pdf.

Elkies:2002:COA

[1735] Noam D. Elkies. Curves Dy^{2} = x^{3} — x of odd analytic rank. Lecture Notes in Computer Science, 2369:244– 251, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690244.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690244.
pdf.

Enge:2002:CIC

[1736] Andreas Enge and François Morain. Comparing invariants for class fields of imaginary quadratic fields. Lecture Notes in Computer Science, 2369:252– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2369/23690252.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690252.
                                        pdf.

Stein:2002:DEC

[1737] William A. Stein and Mark Watkins. A database of elliptic curves — first report. Lecture Notes in Computer Science, 2369:267–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2369/23690267.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2369/23690267.pdf.

Fouquet:2002:IVS

[1738] Mireille Fouquet and François Morain. Isogeny volcanoes and the SEA algorithm. Lecture Notes in Computer Science, 2369:276–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2369/23690276.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2369/23690276.pdf.

Kim:2002:FEC

[1739] Hae Young Kim, Jung Youl Park, Jung Hee Cheon, Je Hong Park, Jae Heon Kim, and Sang Geun Hahn. Fast elliptic curve point counting using Gaussian normal basis. Lecture Notes in Computer Science, 2369:292– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690292.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690292.
pdf.

Denef:2002:EKA

[1740] Jan Denef and Frederik Vercauteren.

An extension of Kedlaya’s algorithm to Artin–Schreier curves in characteristic 2. Lecture Notes in Computer Science, 2369:308–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690308.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690308.
pdf.

Galbraith:2002:ITP

[1741] Steven D. Galbraith, Keith Harrison, and David Soldera. Implementing the tate pairing. Lecture Notes in Computer Science, 2369:324–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690324.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690324.
pdf.

Pomerance:2002:SOC

[1742] Carl Pomerance and Igor E. Shparlinski. Smooth orders and cryptographic applications. Lecture Notes in Computer Science, 2369:338–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2369/23690338.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690338.
                                        pdf.

Shparlinski:2002:CRA

[1743] Igor E. Shparlinski and Ron Steinfeld. Chinese remaindering for algebraic numbers in a hidden field. Lecture Notes in Computer Science, 2369: 349–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2369/23690349.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690349.
                                        pdf.

Hess:2002:ACW

[1744] Florian Hess. An algorithm for computing Weierstrass points. Lecture Notes in Computer Science, 2369:357– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2369/23690357.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690357.
                                        pdf.

Li:2002:NOT

[1745] Wen-Ching W. Li, Hiren Maharaj, Henning Stichtenoth, and Noam D. Elkies. New optimal tame towers of function fields over small finite fields. Lecture Notes in Computer Science, 2369:372–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2369/23690372.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2369/23690372.
pdf.

vanderPoorten:2002:PCF

[1746] Alfred J. van der Poorten and Xuan Chuong Tran. Periodic continued fractions in elliptic function fields. Lecture Notes in Computer Science, 2369: 390–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690390.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690390.
pdf.

Holden:2002:FPT

[1747] Joshua Holden. Fixed points and twocycles of the discrete logarithm. Lecture Notes in Computer Science, 2369: 405–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690405.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690405.
pdf.

Horwitz:2002:RCD

[1748] Jeremy Horwitz and Ramarathnam Venkatesan. Random Cayley digraphs and the discrete logarithm. Lecture Notes in Computer Science, 2369:416– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2369/23690416.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690416.
                                        pdf.

Joux:2002:FFS

[1749] Antoine Joux and Reynald Lercier. The function field sieve is quite special. Lecture Notes in Computer Science, 2369: 431–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2369/23690431.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690431.
                                        pdf.

Leyland:2002:MTL

[1750] Paul Leyland, Arjen Lenstra, Bruce Dodson, Alec Muffett, and Sam Wagstaff. MPQS with three large primes. Lecture Notes in Computer Science, 2369:446–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

                                        link.springer-ny.com/link/service/
                                        series/0558/bibs/2369/23690446.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        2369/23690446.pdf.

Matsuo:2002:IBS

[1751] Kazuto Matsuo, Jinhui Chao, and Shigeo Tsujii. An improved baby step giant step algorithm for point counting of hyperelliptic curves over finite fields. Lecture Notes in Computer Science, 2369:461–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://link.

springer-ny.com/link/service/series/
0558/bibs/2369/23690461.htm;   http:
//link.springer-ny.com/link/service/
series/0558/papers/2369/23690461.
pdf.

Ebinger:2002:FEC

[1752] Peter Ebinger and Edlyn Teske. Factoring N = pq with the elliptic curve method. Lecture Notes in Computer Science, 2369:475–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690475.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690475.
pdf.

Steel:2002:NSC

[1753] Allan Steel. A new scheme for computing with algebraically closed fields. Lecture Notes in Computer Science, 2369: 491–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2369/23690491.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2369/23690491.
pdf.

Rojas:2002:ACR

[1754] J. Maurice Rojas. Additive complexity and roots of polynomials over number fields and √-adic fields. Lecture

Notes in Computer Science, 2369:506– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2369/23690506.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2369/23690506.
                                        pdf.

Anonymous:2002:AIbh

[1755] Anonymous. Author index. Lecture Notes in Computer Science, 2369:517– ??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        papers/2369/2369auth.pdf.